A Complete Bibliography of Publications in Conservation Biology: 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

23 November 2024 Version 1.07

Title word cross-reference

1990 [Ano90f]. 1991 [Ano91g]. 1992 [WF99b]. 1993 [Ano93b]. 1996 [Ano96b]. 1999 [Ano99d].

2 [Dun97].

A. [GWH^{+}99]. Abalone [Wal99b]. Abandonadas [CC99a, CUM96]. Abandoned [CUM96, CC99a]. abiertos [ST96]. Abiotic [CFL96]. Abi’otica [CFL96]. Absence [CZN99, SRR95, SGW90, Str99, SRV96]. Absent [SSTL92]. Abulón [Wal99b]. Abundance [BM98, BGB96, BSR97, CW97a, DH98, DSB95b, GS93, KSO^{+}99, KS97, Mar98, MP97, MH93, PDFG96, VI95, Yah96]. Abundances [RS91a]. Abundancia [BSR97, CW97a, KSO^{+}99, KS97, BGB96, DSB95b, MH93, MKM^{+}96]. Abundant [GB95b, MLD97]. Abundante [MLD97]. abundavies [GB95b]. Acacia [RW99]. Acantilados [KL97]. acaulis [DeM93]. Accelerating [MW93]. Acceptance [WP97]. acceso [PZSG96]. Access [Gra99, PZSG96]. Accidental [MPC99]. acción [Boz93]. Acciones [MDW99]. Accumulation [GWM^{+}97, SL93]. Acedo [Mac91]. Aceleración [MW93]. Aceptación [WP97]. acerca [Car93]. Acercamiento [Pri96b]. Achievable [Blo99a]. Acid [SGK95]. Acid-Damaged [SGK95]. acidificación [FC93]. Acidification [FC93]. Acinonyx [Mer94]. Acknowledgments [Ano98b, Ano99h]. Acorn [LS96]. acrocnema [BBM94b]. Across [BL95a, KASR99, ML94, PWC^{+}98]. Across-Country [BL95a]. Act [WF99b, DME95, GDK99, GD99a, MWN^{+}94, O’C92, PD97, RDE96, RA94, Roh91, TSHG95, Tho94b, Wap98, Win98]. Acta [DME95, GDK99, GD99a, PD97, RDE96, TSHG95]. Actidudes [JM97]. Action [Boz93, KS94a]. Actions [MDW99]. Actitudes [MH99a, RK93]. actividad [KC95]. actividades [BT94]. Activism [Dud95, Hau95]. Activities [Hun96a]. Activity [BHKH96, GS93, KC95]. actuales [BL95a]. Actualizaceón [WGRT96]. acuáticas [CD95, KHP95]. acuerdo [KO93]. acumulación [SL93, GWM^{+}97]. Ad [Pre94]. adamsii [Nor91]. Adaptabilidad [SHMV97]. Adaptation [All93b]. Addition [RH93]. Additions [HM98]. Addressing [HK94]. adecuación [EDFM96, Tho94a, Tsc92]. adelante [Pre94]. Adenophora [GWH^{+}99]. Adequacy [EDFM96]. Adequately [NG92]. adición [RH93]. Adjacent [GMSP98, MKA96, NMGS94, TW92b]. Adjustment [Cha95]. ADN [BHKH96, DMMG96, GSB99, LBR99, OFH94, RFDB94, Rob95c, WLAH92]. adopción [PC93]. Advertencia [HCBH99]. Advocacy [CFH^{+}95, JB95, Kar93, MBJ^{+}97, Mat97a, McC96, Nos96a, Sal95, SF96]. adyacente [NMGS94]. aestivalis [LDP95]. afectado [FC93]. Afectan [FM96, BGTB93]. Affairs [Tem93a]. Affect [BE94, FM96]. Affected [FC93]. Affecting [MG97]. áfidos [Tsc92]. Africa [ACB^{+}99, ACB^{+}99, CC99a, Hac99, HHF^{+}97, HHWB90, IS96, KC93, Lan96, Lin95, LNA95, LCPM97, Mad99, Pom93, RS92, Sam90, Sim96, SF97]. African [Nos98, AL93, BH93, CKCM95, Cow99, CC96, Cre97, CCM97, EG96, EHBE97, FCR^{+}98a, GAC^{+}95, HD99, Lin95, LM93, MBJ^{+}97, McN95, SS98, SF97, TWMM97, VC99]. Africana [Hac99, TWMM97, AL93]. Africanas [HD99]. africano [AL93, GAC^{+}95]. Africanos [CC96, VC99, BH93, Cow99, LM93]. Afrotropical [KMW95]. Afrotropicales [KMW95]. after [ELS93, Gar97, MFR97, Pet99, Ren99]. Again [Bis92, Dal93]. Against [BV97]. Agalleras [MPC99]. Age [DSB95b, PK94]. Age-Structured [PK94]. Aged [CB98, FESS94]. Agency [MBJ^{+}97, TB95]. Agenda [Bal99, MEE93, SBB^{+}94]. Agent [AR97]. aggregata [HP95]. Aggregation [BDSD98]. Aggression [GKA97]. Agícola [War94a]. Agotamiento [HP94]. Agrarias [Sab96]. Agreement [Kau95a]. agrícola [Haa95]. agrícolas [FGG^{+}95]. agricultura [ART94]. Agricultural [CC99a, Haa95, TFHD98, War94a]. Agriculturales [CC99a]. Agriculture [BH97, Gei90, KWR95, Orr91a, Sie96, Sti93, Wue94]. Agroecosystem [VP97]. agroforestales [Thi95]. Agroforestry [Sal93]. Agroforests [Thi95]. Agua [GLFZ96, MLP^{+}97, RR99b, VT99, Ang95, BG95b, BT94]. aguas [BL95b, EBV95, CBLV94]. Aguirre [Car94]. Agulhas [LCPM97, LCPM97]. Aid [SFM96a]. Aimed [RS92]. Aimophila [LDP95]. Ain’t [Rob93a]. Aïr [Ing90]. Airborne [FP95]. aire [FP95]. aislada [DP96, Was96]. Aislado [TCO^{+}96]. aislamiento [Mil95b, DBCM95, New96]. al [BC97b, Cre97, Dia95, Doa95, Duf93c, DME95, Fim94, GK96, JFJ96, ML94, MEE94, NF98, RDN94, Sal92, SO94, TCK^{+}95, WW98a]. Alamo [HNC97]. Alaska [DBH^{+}99]. albens [PB94b]. albicaulis [TCK^{+}95]. albus [FBFO96]. alcance [FC93]. Alcanzar [GG99]. Alcorn [RS93b]. aldea [LLP95]. Alectoris [KASR99]. alejamiento [RS95b]. Alélica [SMV96]. alelo [LRT93]. Alelos [LCA99, RL96]. Algae [MKM^{+}96]. algas [MKM^{+}96]. Algonquin [FT96]. Alien [BS92, HRCTS99, HS95, Mau97, Sou90, TW92b, UCB92]. Aliens [Ric98]. Alkaloid [HS98b]. Alkaloid-Rich [HS98b]. allá [DW93]. Allegheny [BY96]. Allele [LRT93, LC98b, LCA99, RL96, WT90]. Alleles [Mil95a, SW94]. Allelic [SMV96]. Allendorf [WW98a]. Alligator [RSMB99]. Allocating [ML90]. Allocation [CKB98]. Allometric [SD94]. Allopolyploid [RA94]. Allotetraploid [PSB98]. Allouata [Sto96]. Allozyme [CAH97, CSC^{+}94, GZWR97, SSTL92]. Allozymic [WDWG94]. Alluvial [LSC^{+}96]. alométrica [SD94]. Alondras [YS96]. Along [AB90, FESS94, Fra94, JR99, Med93, VI95]. Alouatta [Sto96]. Alozimas [CAH97]. Alpinismo [KL97]. Already [CBM^{+}97, SMS97]. alrededores [LLP95]. alta [GNH94]. Altamente [FCM99, WS97]. Altas [McK99, KB94]. Alteraciones [RBPB96, BC94]. alterada [Bak94]. alteradas [BT94]. Alterado [Pea97]. Alteration [RBPB96]. Alterations [BC94]. Altered [Bak94, BT94, Pea97]. Alternate [PNS99]. Alternativas [GK99, KMW95]. Alternative [Cob99, GK99, JJ99, KMW95, MB92b, SKOO98]. Alternatives [Sie96, SGA98]. Alternos [PNS99]. Altiplano [BB97]. Altitudinal [HY93]. altos [HCB94]. altura [CLM^{+}95]. aluvial [LSC^{+}96]. Alvarez [Cua97]. Alwyn [Hur94]. Amazon [AMGJ95, ART94, Ano99e, CS98, DS95, DPOO94, Dra94, NMV^{+}98, SG98, VJSU98]. Amazonas [AMGJ95, DPOO94]. Amazonia [DOSP94, FL97, PGR^{+}94, CUM96, DS95, Pea97]. Amazonian [BER97b, CUM96, DOSP94, FF95, LFM^{+}98, Pea97, Per94, PT95, PGR^{+}94, SB95b, VAPCO98]. Amazonians [RS93a]. Amazónica [RS93a]. Amazónicas [PT95]. Amazónicos [BER97b]. Ambiental [Sym96, Fol94, HP95, KB94, PSGS94, Tho94a, You94a]. Ambientales [Yaf97, Kau93b]. Ambiente [ACB^{+}99, GNH94, HDE95, KK94]. ambientes [CRHSS96, NGN96, Tsc92]. Amblema [MLP^{+}97]. Ambystoma [MPB96, TLG^{+}98]. Amenaza [GWM^{+}97, HRCTS99, McK99, Ree99, BGF95]. Amenazada [GWH^{+}99, HP99, KB99, Wei99, YBZ99, BY96, GHB95, Gre96, KMW95]. Amenazadas [BPC97, HB96, RDE96, BSW^{+}96, DDM93, Rei95]. Amenazado [LSA99]. Amenazados [MBD^{+}99, NR97, BSW^{+}96]. amenazamiento [Fri93]. Amenazas [SM99, Gre96, MAR96, RBMM97]. America [BSW^{+}96, Fle94a, GL93b, Hou95, KBRB96, Lan96, Mef92, RH97b, Rod93b, RRM93, Vid93, WW98b, BGTB93, GL93b, KBRB96, New95, RR99b, RRM93, SPP96, RH97b]. American [BGTB93, CHK98, DD97, GB97, Hun98, LCSC95, Meh94, MLP^{+}97, NGN96, New95, PWC^{+}98, PSSB95, RR99b, SPP96, SY94, UCB92, WTWS98]. Americana [Can96, GSB99, SY94, BHSO92, CRHSS96, SY94]. Americanas [MLP^{+}97]. Americano [GB97, NGN96]. americanus [LCSC95]. Ammophila [BPS95]. amnitagei [HE91]. Among [UA94, Aco99, FGG^{+}95, HC99, HGNK93, LV94, Loc99]. amortiguadoras [MVH96]. Amount [Bar95]. Amphibian [BWS94, DSB95b, HS97, HG97a, KB97, Lip98, Pea97, PC94, PFSG97, RN96, RB95a]. Amphibians [DH98, DC98, FS96, LSS95, Wel90, Wym90]. Amplia [GWH^{+}99]. Amsinckia [PNH93]. Amur [HV97]. anádromos [WH95]. Anadromous [Moy94, WH95]. Analisis [HBA93, DMMG96, GMMR^{+}96, MDW99, CSBC95, FDV94, HW95, KGS93, LP96, MSHW95, PT95, RHS94, SB95c, Tay95, TM96, WMW93, WGRT96, HC96, AL93, BL95a, BGK^{+}97, Bus96, CBAT97, Gre96, KLR97, LMNV92, MHO97, NGN96, PZSG96, RLR97, RFDB94, RHS94, SDC^{+}93, SAH^{+}95]. Analogous [Win98]. Analyses [BL95a, CBAT97, Gre96, Los95, PZSG96, PL95, RFDB94, TFHD98]. Analysis [Ano90o, AL93, BDSD98, BGK^{+}97, Bus96, CSBC95, CAWH91, DMMG96, EV98, FF95, FDV94, GMMR^{+}96, Gil96, HBA93, HH98, HS97, HC96, HW95, KGS93, KLR97, LMNV92, LP96, Mag91, MHO97, Men90, MDW99, MSHW95, MFW90, NGN96, PT95, RLR97, RB95a, RHS94, Sha90, SAM98, Son96, SDC^{+}93, SB95c, Str94, SAH^{+}95, Tay95, TM96, Tho97, WMW93]. Analytical [Jen95]. Anatidae [Gre96]. Ancestors [Ano99c, Bri95a]. ancho [RDN94]. Ancient [Ber95b]. Andean [BKO^{+}95, MP97]. Anderson [Göt92]. Andrew [Ogd92]. Anfibio [KB97]. Anfibios [Pea97, RN96, BWS94, DSB95b, FS96, PC94]. angostos [RDN94]. angustifolia [SAS95]. anidación [JBB96]. Animal [ACB^{+}99, BEG^{+}94, Swa98]. Animales [RRS96, ACB^{+}99, BEG^{+}94, DMMG96]. Animals [Ano91h, Ano99c, Bri95a, DCL98, FT91, RRS96, VMMK92, WMKF98, War96]. Announcement [Bru90]. Announcements [Ano90b, Ano90c, Ano91a, Ano91b, Ano91c, Ano92a, Ano92b, Ano93a, Ano94a]. Annual [Ano90a, MBH96, Tem92b, Ano99a]. Años [MBBF95, KALG94]. Answer [MH91]. Ant [Hol95]. Antelopes [KMW95]. Antes [MFR97, RGSW96]. Anthologies [Pis94]. Anthracnose [HE97]. Anthropogenic [FH97, FC93]. Antillanos [LBR99]. Antillean [LBR99]. antílopes [KMW95]. antiquus [Ber95b]. Antonio [KALG94]. antropogénica [FC93]. Antropogénicos [FH97]. Ants [HG97b, MB96]. anual [MBH96]. Anunsens [Sch93b]. Anurans [Lip99]. Any [FNN92]. Apalaches [GJH96, PEH93, HGK^{+}96, JHHB99]. Aparente [MNU94]. aparentemente [SRR95]. Aphantopus [ST96]. Aphids [Tsc92]. Apiaceae [KMR94]. Aplicación [AM99, GD99a, HW95, IS96, TB95, BN97, Wit97a, SB95c].

Aplicaciones [Cla95, Doa95]. Aplicada [DHW99, WKS^{+}94]. Aplicando [MWCP94]. Apostrophe [Car98]. apoyo [LSC^{+}96]. Appalachia [SB95c]. Appalachian [DD97, DM92, GJH96, JHHB99, PEH93]. Appalachians [ELS93, HGK^{+}96, Pet99, RFG^{+}98]. Apparent [MNU94]. Apparently [SRR95]. Appeal [And96]. Appear [ST96]. Application [BN97, Cla95, GD99a, HW95, IS96, TB95, TL94, Wit97a]. Applications [Doa95]. Applied [DHW99]. Applying [MWCP94, SB95c]. Approach [BC97b, Cla99, DLP^{+}92, Fit94, GD99a, Gin97, HH95b, HS90, LB95, MFW90, Nos90b, OD98, PS98, Pri96b, SM99, Sti93, Wal99a, WKS^{+}94, YG92]. Approaches [BJ95, Gen94, KB94, MMT^{+}98]. Appropriateness [BHCB99, VM99]. Aproximación [BC97b, GD99a]. Aptitud [BHCB99]. Aptostichus [RF97]. Aquarium [SB95a]. Aquatic [LM95, Vid93, WOA98, Wal96]. aquatico [LM95]. Arabuko [FMF95]. Arabuko-Sokoke [FMF95]. arachnifera [CAH97]. arachnoids [Str91]. Arbitrary [Sid98]. Arbol [HWB96]. arboles [HCB94, BB97, HC99]. arborea [BE94]. Arborícolas [GK99]. arbustivas [CLM^{+}95]. Arbusto [YB96]. Arbustos [YS96]. Arca [BMLW96]. Archie [Fra95b]. Archipelago [RTNK99]. Archipelagoes [Bur95]. archipelagos [Bur95]. Archipiélago [RTNK99]. Architecture [Orr93a, Orr97, Orr99a]. Arctic [FS92]. arctos [TSSB95]. ardilla [War94c]. Area [BFS98, Dia95, DG92, DP96, DF96, ES94, Hei96a, HWG99, Pow91, VHM94, WF99a, DP96, DF96, FGG^{+}95, TCK^{+}95, AFF^{+}97]. Areas [AM99, BHSO92, Bei93, BN98b, GWW93b, GN92, Hei95a, HWG99, Ing90, Ker97, KMW95, LKM90, LP98, MMT^{+}98, NMGS94, NG92, PNS99, SAQMGQ^{+}99, Sim96, VRV^{+}94, WGM^{+}96, WMH94, CLM^{+}95, JFJ96, SAQMGQ^{+}99, Bei93, BGTB93, GWW93b, KMW95, NMGS94, Sim96, VRV^{+}94, WMH94]. Arena [LC99]. arenaria [BPS95]. arenoso [GNH94, MM92]. Arg [RFM97]. Argentina [VMMK92]. Argentine [Hol95, HG97b]. argumento [Mer94]. argus [BGK^{+}97]. Argyroxiphium [RFM97]. Arid [CBLV94, RW99]. árida [BT94]. áridas [CBLV94]. Aristida [WDWG94]. Arizona [BB93, BB99, SMHK98]. Ark [BMLW96, Mor96, Sch93c]. Arrecife [MKA96]. Arrecifes [JRP99]. arrendajo [BLD^{+}95]. Arrogance [Mef92, Sta95]. Arrogancia [Sta95]. arroyo [HGK^{+}96]. Arroyos [JHHB99, LNPC^{+}95]. Art [Ral92, RR99c, Sam96a]. Arthropod [KCE^{+}93, RF97]. Articles [Ano96d]. Artificial [Ang94, BH97, HH98, YM96, Yah96]. Arts [Orr91a, Orr92b]. asentamientos [DOSP94]. Ash [LF97, Pet99]. así [Yu94]. Asia [BPC97]. Asian [HS96b, OK96]. Asiática [OK96]. Asking [Bra94]. asociacines [BEG^{+}94]. Asociación [BT94]. Asociaciones [RDH97]. Asociadas [Ber97a, BY96]. Asociado [RFM97]. asociados [GL95]. Aspen [CB98, MCO98]. Assemblages [JHHB99, KCE^{+}93, LNPC^{+}95, NLS93, OBY98, SS98]. Assess [LB95, PRH98]. Assessing [BM92, BSW^{+}96, Gre92, Kuc92, ML91, Mas91, MSC96, RBPB96, RHK97, SGA98, SG93, WMB^{+}97, WA95]. Assessment [Ber90, DBH^{+}99, GFA90, OB93, PC93, RGH^{+}99, RB95a, SM99, SEC98, STM95, TB95, TB98, TW92a, Vog95]. Assessments [PL94]. Assigning [DHJG91]. Associated [BY96, Ber97a, GL95, RFM97]. Associates [TNN^{+}99]. Association [BT94]. Associations [BEG^{+}94, RDH97]. Assumptions [Ham94, How94, Sie96]. Asteraceae [FCR^{+}98b, RFM97, VIO98]. Astronomy [War96]. Asymmetry [Tsu98]. ataques [LV94]. ater [HH95a]. Atlantic [BAR^{+}93, Goe97, Saf93, WW98b]. Atlántico [Goe97]. Atlases [CV95]. Atmosférico [CT99]. Atmospheric [CT99]. atrás [Pre94]. Atributos [Ang95]. Attack [LV94]. Attempts [Dia96]. Attitude [Bal95]. Attitudes [CPG94, JM97, MH99a, RK93]. Attraction [SP90]. Attributes [Ang95]. Attwater [PS96, PS96]. Augmentation [MS92]. Augmenting [PC93]. August [Ano90f, Ano91g, Ano93b]. aurocapillus [PF99]. Aurora [SGK95, KB98]. Ausencia [CZN99, Str99, SRR95]. Australia [MHL97, BC97a, DHE97, HG97a, Hob96, HM98, LP96, LF97, MB96, NHA95, NR97, OBY98, RS91a, SFM96b]. Australia/Papua [MHL97]. Australian [AR97, BHCB99, DHE97, Lau91, LMS96, LF97, NHA95, ST94, WF99b]. Australiana [LMS96]. Australianas [BHCB99]. australianos [ST94]. Author [Ano99b]. Autocorrelated [Swa98]. Autocorrelation [Koe98]. Availability [LLP95]. available [Ano90n]. Ave [SPR97, WER^{+}97]. Aves [ADFT97, BHS99, BC97a, BB99, BSR97, BPC97, BPO99, CS97, DH97, DHW99, FCM99, Goe97, GSB97, GBAR97, HNC97, MPC99, Ree99, SW99a, Sto97, SS99b, TA99, VM99, VTM99, War94a, Win96a, WF99a, WGRT96, BSC95, BSW^{+}96, BGTB93, BGB96, CUM96, CLM^{+}95, CD95, CPN95, DLK^{+}95, DTFP95, FESS94, FH95, FEM95, FGG^{+}95, Haa95, HVM96, KHP95, KS94b, MVH96, MW93, RH93, RS95b, SPP96, SK93, TM96, VRV^{+}94, WGM^{+}96, WDSA94]. Avian [FESS94, Gat94, HMO98, HNC97, Kar90a, KS94b, PS99, Pat94a, RFG^{+}98, Ree99, WDSA94, Win96a, WGRT96]. Aviares [PS99]. avícola [ART94]. Avifauna [ART94, Loc99, Ren99, Rob99]. Avifaunal [DRP98]. avistamiento [SDC^{+}93]. Avoiding [KM97]. Awareness [Dun97, Mef94, Vid94]. ayuda [SFM96a]. Ayudando [Rab95a]. Azucena [GWH97].

B [AKG98, BHKH96, LP98, LSS95, RH97a]. Baby [SF96, VL91]. Bachman [LDP95, LDP95]. Back [Ano97a, Ano99c, RS95b]. Backed [CA92, HHR99, MKH98, Mil95b]. Backgarden [Che93]. Backward [Pre94]. bactracios [DF96]. Badlands [MBT90]. Bahamas [CLL91]. Bahía [BG95b, SB99]. Baitworm [SB99]. Baja [REDH98, ADM94, KGM^{+}97]. Bajas [GW97, MBD^{+}99, FGG^{+}95, TCO^{+}96]. Bajo [BR99, NM97, RDE96, SFM99, ART94, DBCM95, FS95, LLP95]. Bajos [EKL^{+}99, MPB96, HCB94]. Balance [LWMG93, Mat97b]. Balanced [HS90]. Balancing [MH97]. Baler [Orr94d]. Ballena [GD99a]. Ballenas [GDK99]. Baltic [KLR97]. Báltico [KLR97]. Ban [DP92]. Bananas [Ewe91]. bancarios [WSB92]. bancos [Ham94]. Bandas [Hil95]. Bandicoot [Rob95c]. Bandicoots [SMGB91]. Bank [WSB92]. Banks [Ham94, SGR^{+}98]. Banksia [BBSK96]. Baptisia [CAH97]. Baram [Duf93a]. Bargains [SFCM92]. Barred [Rob95c]. Barro [Kar90a, Rob99]. Basada [BC97b, CAH97, DHW99, LBR99]. Basadas [GMMR^{+}96, KASR99, MDW99, KTA94]. Basado [AA97, JRP99, HW95, LNPC^{+}95]. Basados [SM97]. Base [GMF99, KRG^{+}99, PNS99, RDN94]. Based [AA97, CAH97, DHW99, GMMR^{+}96, GMF99, KASR99, KTA94, KRG^{+}99, LNPC^{+}95, MDW99, PS98, SM97, WDR^{+}98, BC97b]. Bases [PNS99]. Basics [Qui95]. Basin [Fra94, MQ99, BW91, DRP98, GL93a, NTT94, SJB97]. Basis [PEP98, RTOB94]. Bass [KB98]. Basura [MB99]. Bat [ZG99]. Bathwater [SF96, VL91]. Bats [CB98, FCR^{+}98a, PP96, RHCB93, SVC96]. Bay [BG95b, EK98, SB99]. Be [AL93, Fim94, LNA95, MG90, PB94a, RS92, RS93c, Wil92]. Bear [MWM^{+}99, MS92, MR91, MHB^{+}96, PZSG96, Pri96b, TSSB95, WTWS98]. Bearded [Bus96]. Bears [MW98, MHWP96, PWC^{+}98]. Beating [Ano97a]. Beattie [Gol97b]. Beaucarnea [CRGGFT97]. Beauty [Gra98]. Beaver [Har94]. Becoming [Ano99n, Mor95]. Bedfellows [BG95a]. Been [Ano99p]. Beetle [BR96, LCSC95, NLS93]. Beetles [MKH98, NABE95, PC92, RCM97]. before [MFR97]. Beginning [Dal93, Mef99b]. Behavior [AMA^{+}94, DSM90, GB97, Geh96, Hei96a, JBB96, KB97, KLF98, Ree99, Sch98, WSE97]. Behavioral [SSS94, YM97]. Behind [HSD93]. Believing [Rob93a]. Belize [BAR94, PH94]. Bellotero [LS96]. Belowground [PBBA90]. Benchmark [HCH^{+}97]. Benchmarks [Hun96a]. Beneficios [DSGS93]. Benefits [CKB98, DSGS93]. Benevolence [Ebe95]. bennetti [RH97a]. Bensoniella [SSTL92]. Benthic [Aco99, EK98, GW97]. Benthos [WN98b]. Béntica [Aco99]. Bentónicos [GW97]. Bernardino [MBBF95, MBBF95]. Best [PB94a]. Better [BJ95, EV98, Max90, Nor93]. Between [Avi95, Cla95, FYP96, HT95, Ran96, Rym94, BTRM99, BPO99, HD99, KQK93, LWMG93, NMGS94, OLKD98, PQL99, RR99c, Ris99, SB97, ST96, TSSB95, TFHD98, TLG^{+}98, VW99a, WMKF98]. Bewick [KW96]. Beyond [Dia96, DW93, LT92, MBKR92, Neh94, Sch93c, RG99]. Biased [GMG94]. bicolor [KA92]. bicornis [OFH94, AMW90, OFH94, SF97]. bieti [SS95]. Big [AL93, Sad95]. Bighorn [Ber90, Ber99, FBS95, SG93]. Bill [FSCL96]. Billion [Dun97]. Biodiversidad [DBH^{+}99, FM96, Gol99b, MB96, RS93a, RR99a, Ris95, Sab96, SAQMGQ^{+}99, SHM97, WMB^{+}97, Win96a, BL95a, CBCC93, DPOO94, DW93, FYP96, HDE95, KTA94, KK94, OSEC95, OR94, OB93, Rei93, Rob93b, SBL96, SB95c, TTW^{+}94, BL95b]. Biodiversity [All93a, Ang93, Ang94, Ano98d, BL95a, BW92, Baw95b, BS98, BO99, BDR^{+}98, BL95b, Bra98, BABK98, BCP95, CG95, Che93, CBCC93, Con95, Cro95, Cue96, DPOO94, Del99, DG93, DW93, Dud92, DBH^{+}99, FM96, FYP96, Gol99b, Gra99, Gur92, Hei93, Hei96b, Hob96, HDE95, HG97b, Hun93, IN96, Jai93, JOB96, KBM98, KTA94, KK94, Lyd96, Mag92b, MB96, MQ99, MR98a, MMT^{+}98, MT99, MH97, MW90, MD96, Nab95, NKA98, Nos90b, OSEC95, OR94, OB93, PC92, PG91, Pee96, Per93, Phi94, Pom93, Pri92, RTS90, RS93a, RR99a, Rei93, Rex93, Ris95, Ris99, Rob93b, Rod93a, Sab96, SAQMGQ^{+}99, Sar96, Sch93c, SBL96, SC97b, SHM97, SB95c, TTW^{+}94, Wal92, WMB^{+}97, Win96a]. Biogeografia [Ver93, HB96]. Biogeographic [LC98a, RT91]. Biogeography [Jor92, Ver93, HB96]. Bioindicator [PC98]. Biologia [BO96, Sto96, DGG^{+}99, Wit97a, Cla95, HLAS96, SK96, SNP94, NQH^{+}96, ST95]. Biológica [CO99, GMF99, LC99, KO93, PSGS94, TG93, Wal95]. Biological [BAR94, Blo99b, BMN92, CNP^{+}97, DB98, DKOW94, GMF99, Hab96, HSD93, IU92, Kar90b, KS94a, KRG^{+}99, LC99, Mag92a, MS92, Nos90a, O’C92, Orr91a, Roh91, RHS94, RDH97, SHM91, STM95, Sem98, SFM99, SSS94, Str94, Wal95]. Biologically [OD98, RB97b]. Biológicas [Hab96, RDH97, SFM99, RHS94]. Biológicos [KRG^{+}99]. Biologist [All97, Str98, VP97]. Biologists [CDD96, GWW93a, Mag91, Nos97, Pov94, Put97, Ste98, Tem92a, Tro94]. Biólogos [CDD96]. Biology [Ano94b, ALF92, BO96, BB90, CPG94, CO99, Cla95, Cob90, CMB96, Cro92, DGG^{+}99, DMD^{+}92, DDM^{+}92, Ela96, Fle96, FNN92, Gru96, Hag92, Hag95, HSS94, HLAS96, Jac90, Jac93, JVM95, JD98, Kar90b, KO93, Mag96, Mat97a, McC96, Mei99, Mur90, MD96, Nos94, NQH^{+}96, Nos96a, Orr99b, PSGS94, Ran96, RP99, SK96, Sch96, SF96, ST95, Spe95, Sto96, SNP94, TG93, TL94, TB96, Uhl98a, War94b, Wes91, Wit97a, Wit97b, You95, Ano90i, Ano90j, Ano90k, Ano90l, Ano91i, Ano98a, Sto95, Tem92c, Ehr93, Mef98a, Nos93, Tem93c, WM91]. biomasa [Rob95b]. Biomass [Rob95b]. Bioplanning [Bab99]. biopolítica [War94c]. Biopolitics [War94c]. Bioregional [Dud95]. Biosfera [FC99, Kau93b]. Biosphere [Ehr91, FC99, GM91, Kau93b, Sch91, Tho91, Wil91, Avi94]. Biotic [LNPC^{+}95, MCR94, MR98b, TNN^{+}99]. biótica [LNPC^{+}95, MCR94]. Bird [BHS99, BGB96, BSR97, BPO99, CS97, Dav96, FH95, GBAR97, Has95, Hut95c, KALG94, Mar98, MCO98, NY91, RHSM94, RH93, RR99d, RS95b, RBMA98, SJ98, Sto97, Thi92, Thi95, TM96, Tho93, WRSP97, WER^{+}97]. Bird-Habitat [FH95]. Bird-Nest [Has95]. Birds [Ama93, ADFT97, BSC95, BSW^{+}96, BC97a, Bla91, BB99, BPC97, CUM96, CLM^{+}95, CPN95, DLK^{+}95, DTFP95, FS91, FGG^{+}95, Goe97, GSB97, Haa95, HVM96, HY93, JS93, Kat92, KHS98, KR95, Koe98, LV91, MVH96, MH93, MW93, Meh94, New91, RS91a, RDN94, SLNP95, SMHK98, SGW90, SS99b, TA99, Tsc92, VM99, VHM94, VTM99, VRV^{+}94, War94a, WGM^{+}96, WF99a, Yah95]. Birth [Bab99, TL94]. Birth-Death [TL94]. Bison [BC99, Gei90, HP94, MBT90, MM94, PSSB95, HP94, HP94, MBT90]. bisonte [MM94]. Bisontes [BC99]. Black [AMA^{+}94, AMW90, BCGL93, BCG94, Ber97a, BC98, BDSD98, EKL^{+}99, GW98, GR98, MAR96, OFH94, PZSG96, PGW97, Pri99, RK93, RTOB94, SB97, SF97]. Black-Faced [GR98]. Black-Footed [EKL^{+}99, GW98, RK93, RTOB94]. Blackbird [GD99b]. Blackbuck [Jha93]. Blanca [CRHSS96, HOF^{+}96]. Blancas [FBFO96]. Blanco [BRHD97, KL97]. Blanding [CDV93]. blandingii [CDV93]. Blind [OE95]. Blindness [LRT93]. Blister [TCK^{+}95]. Blood [AMA^{+}94, SGW90]. Blooming [SNP94]. Blue [Ash97, ES94]. Bluefin [Saf93]. Blumstein [Wil93b]. Blunt [Ehr94b]. Boarding [Mor96]. Body [GB95a, GB96a, GB96b, RHCB93, WMKF98]. Bolín [EE97]. Boloria [BBM94b]. bonasus [HP94]. Book [Ano90d, Ano90e, Ano91d, Ano91e, Ano91f, Ano96e, Sam96b]. Books [Ano90f, Ano91g, Ano93b, Ano96a, Ano96b, Ano97b, Ano99d, MH97]. Borde [SB95c, Fra94, Pat94a]. borde-hacia-el [Fra94]. Bordering [Sal93]. Borders [AB90, RG99]. Bordes [JR99, SPR97, Mil95b]. Boreal [Car93, NLS93, Nie97, Sto97]. Boreales [Nie97, Car93, NLS93]. Bornean [VI95]. Borneo [Rab95a]. Boscosas [AM99]. boscoso [KGD96]. boscosos [Haa95]. Bosque [BRHD97, CFL96, Goe97, HWB96, HC99, JHHB99, JR99, Ren99, RLD99, Sto97, TCO^{+}96, TWMM97, VTM99, YB96, AMGJ95, HVM96, KK94, LMNV92, LDP95, MVH96, MW93, Mil95b, NHA95, PEH93, RDN94, RHSM94, RH93, SGCJ94, ST96, Vog95]. Bosques [BB97, BPO99, CC99a, CC99b, DH97, DHE97, GFR^{+}99, HL99, HV97, LF97, Lug97, MBBF95, NM97, Nie97, RJBB96, WW99, WER^{+}97, ADM94, BSC95, BEG^{+}94, BHM94, Bun95, CPN95, FESS94, Fra94, FEM95, HH95a, KB94, LLP95, Lug94, MPB96, MWCP94, PB94b, SB94, Tho93, WDSA94, You94a]. Boston [DP96]. Botanical [PS99]. Botánicas [PS99]. Botella [BGK^{+}97, GSB99, LCA99, RFM97, RL96]. Bottleneck [BCL^{+}98, GSB99, Hed96, LCA99, PPR^{+}91, RFM97]. Bottlenecked [LC98b]. Bottlenecks [BGK^{+}97, Leb91, RL96]. Bottom [Kai98, PCM98, SGR^{+}98]. Bottom-Fishing [Kai98]. Boundaries [Mef98c, SC97b, SMPN99]. Boundary [FT96]. bouvieri [GL95]. Bovine [PSSB95]. Box [PB94b, PSB98]. Brachyteles [Str91]. Brandegee [SNP94]. Brasil [Goe97]. Brasilena [DS95]. Brasilera [DOSP94]. Brazil [FF95, Goe97, SB95b, SS99b]. Brazilian [CS98, DS95, DOSP94, NMV^{+}98, Tur93, VJSU98]. Bread [Day91]. Breda [CM94]. Breeding [All93b, BBHM95, BMLW96, BHS99, BGTB93, BSR97, BBCR99, CBAT97, DeM93, DC98, DSB95b, Fle94b, FNC98, GKA97, GW91, GC97, Hug91, KHM99, KR95, MH91, Mil95a, Put97, RLD99, RL91, Rym94, RJL95, RJL99, Sem98, Sim96, ST95, SDB^{+}96, SDB^{+}97, Tem94, TM96]. Breeds [HR93]. Bretaña [FGG^{+}95]. brevifolia [BHS95]. Brian [Bla93]. Briofitas [PBB99]. Britain [DHBG97, FGG^{+}95, TJ99, UCB92]. Británica [Wal99b, Ber95b, BGK^{+}97, Bun95, WPSZ96]. Británicas [GW97, WGM^{+}96]. British [Ber95b, BGK^{+}97, Bun95, GW97, HS96a, SRV96, Wal99b, WPSZ96, WGM^{+}96]. Britten [Ehr95, SO94]. Broadening [Bis94a, HM98]. Broadleaved [Bra94]. Brokering [RR99c]. Brood [SW99b]. Brook [HGK^{+}96]. Brown [GMSP98, PWC^{+}98, TSSB95, TB93, WTWS98, Dud97, FFH96, TB93]. Brown-Headed [TB93, TB93]. Brownlash [Ano97a]. Browsing [MS93]. brozal [GNH94]. Brucellosis [BC99, MM94]. Brucellosis-Exposed [BC99]. Brucelosis [BC99, MM94]. Brussard [Ehr95]. Bryophyte [PBB99]. Bryozoan [GSAZ98]. Buceo [HRV^{+}99]. Buckmoth [LRD^{+}96]. Buena [Hig97]. Buffalo [Gei90, HS96b, HVG91]. Buffer [BG95b, MVH96, Sal93, Sem98]. Bufo [DHBG97, SFM96b]. Búho [LNVM94, TG93]. Build [CZN99, Rob95b]. Build-up [Rob95b]. Building [Kau95a]. Bull [LAF93, RM97]. bullata [GHB95]. Bulldozers [Gur92]. Bullfrogs [KB98, LDSH99]. Bunnies [Gur92]. búo [LMNV92]. Burden [LN96]. Burn [Ano99e, TCK^{+}95]. Burning [GLFM98]. Burying [LCSC95]. Bush [KA92, WER^{+}97]. Bush-Tanager [WER^{+}97]. Bushland [BC97a]. Business [Ehr92a, Orr92a]. Butte [DS92]. Butterflies [Fle95, SSEM94, ST96, Wei99]. Butterfly [BBM94b, BBM94c, BGK^{+}97, HMPK96, Sch98, SO94, WRM^{+}91, Ste93]. Buxbaum [SNP94]. Buzzwords [BO99, Gol99b]. Bycatch [MPC99].

c [AA97]. Caballos [Sym96]. cabeza [HBA93, RWES93]. cabeziblancas [SB94]. Cable [Nos98]. Cacería [MN99, WSB92]. Cachorro [EE97]. Cactaceae [NCE96]. Cactácuas [HB96]. Cacteceae [NCE96]. Cacti [HB95, HB96, PP96]. Cactus [PP96]. cadaverina [AKG98]. Cafetaleras [GBAR97]. cahuama [JBB96]. Caiman [TV99, TV99]. calamita [DHBG97]. Calcáreos [FS97]. Calcareous [FS97]. calentamiento [SVC96]. calidad [RCHP94]. California [ADM94, AA97, Bla94, BSR97, BTG93, BPS95, CK98, CD95, DF96, Dud99, FS96, FFL95, Fri93, GK96, GKA97, HM98, Hol95, HG97b, Koe98, LDSH99, LEMC95, Mil95a, MW90, Moy94, MR98b, PV98, Rei94, REDH98, SDC^{+}93]. Californian [MBBF95]. Californianos [MBBF95]. Californica [AA97, DS92]. californicus [Mil95b]. caliza [MNU94]. Call [Ano95a]. Callings [Orr99b]. callos [SB94]. Cambio [SM97, DPOO94, MN96]. Cambios [MBBF95, WMB^{+}97, LS96, SPP96, SBL96, BGB96, CW97a, DDM93, FC99, LCA99, Ren99, RL96]. Cameroon [MN99]. Camerún [MN99]. Caminos [RJBB96]. Campaign [Tem93a]. campestre [Mil95b]. Campo [WSE97, DS92]. Can [Nos90a, PM92, RS93c, SDT93]. Canada [BRHD97, DHW99, Gei90, KL96, KL97, NLS93, Sal99, Scu99, Wal99b, KL96, NLS93, Wal99b]. Canadian [Fra97]. Canadienses [Fra97]. canales [Sha93]. Cane [SFM96b]. cangrejo [FC93, GK96, LEMC95]. Cangrejos [RCMR97]. Canguro [EKL^{+}99, GP97]. canguros [PK94]. Canis [LR91, LRT93]. Cañón [HWG99]. Canopy [FCR^{+}98a, WRM^{+}91]. Cantoras [SW99a, BGTB93, SPP96]. Canyon [BTG93, HWG99]. Capacidad [BTRM99, PP96, Fol94]. Capacitación [CDD96]. Capacity [BTRM99, Fol94, PP96]. Cape [RS92]. Capital [CD92]. Capricious [Sid98]. Captive [All93b, BBHM95, BMLW96, BLF^{+}93, BMR^{+}92, BBCR99, BR99, Bus96, CRJ^{+}94, CBAT97, DMMG96, Fle94b, GW91, GC97, Hug91, KHM99, LR91, LRT93, MJ97, MH91, PC93, RFDB94, SG98, SDB^{+}96, SDB^{+}97, VL91]. Captive-Raised [SG98]. Captive-Reared [MJ97]. Captive-Rearing [PC93]. Captura [MPC99]. Carabidae [NLS93]. carábidos [NLS93]. Características [TTJ^{+}99, BY96, LNPC^{+}95, KS97, RDH97]. Caras [Yaf99].

Carbon [CT99, PP93]. Carbon-Offset [PP93]. Carbono [CT99]. Carcasses [Gar97]. cardo [HT95]. Care [Dra96]. Careers [Orr99b]. Cares [CGG^{+}97]. Caretta [BAR^{+}93, JBB96, MB99]. Carga [BTRM99, PP96, Fol94, LN96]. Caribbean [Aco99, HRV^{+}99, HSGY98, Rob95b]. Caribe [Aco99, HRV^{+}99, Rob95b]. Caribou [WPSZ96]. caribues [WPSZ96]. Caring [HM93, Rob93b, You93]. Carnada [SB99]. carneros [FBS95]. Carnivora [TSP97]. Carnivore [BTRM99, CCR96, CPC96a, KL96, KBRB96, NQH^{+}96, PC96, WR96]. Carnivores [Bro96, CC96, RH96, SCAB91, WPR96]. Carnivoros [NQH^{+}96, RH96, BTRM99, CCR96, KL96, KBRB96, PC96, WPR96, WR96, Bro96, CC96]. Carnosas [HC99]. carolensis [GD99b]. Carolina [BG95b, EBV95, Fra94]. Carpaneto [SDB^{+}97]. Carpintero [LS96, HBA93, RWES93]. Carr [Fra95b]. carreteras [You94a]. Carros [Wei99]. Carry [DM90a]. Carrying [BTRM99, Fol94, PP96]. Cars [Wei99]. Cary [HC96]. cascada [GWW93b]. Cascading [GWW93b]. Case [BG95b, Can96, CC99a, DHJG91, Deh96, DME95, EE97, FCR^{+}98b, GSAZ98, Hed95, LS96, LOT^{+}99, MNU94, MHH98, Mer94, OB96, PPR^{+}91, PC94, PB95, Sch92, UCB92, VT99, Wei92, WDR^{+}98]. Cash [Ano99e]. Caso [AHMM97, CC99a, EE97, LOT^{+}99, BG95b, Can96, Deh96, MNU94, OB96, PC94, PB95, VT99]. Castor [Har94, Har94]. Casualty* [BK99]. Catalina [RG95]. catastrófica [LMS96]. Catastrophes [You99]. Catastrophic [AR97, LMS96, SB97]. Categories [ML91]. Categorizacíon [LP96]. Catfish [TW92a]. Cats [SSS94]. Cattle [Ano94b, BD92, HS96b]. Caughley [HLAS96]. causa [CC96]. Causality [Bra96]. causas [BGTB93, Car93, SPP96]. Cause [FSCL96, Mil94]. Causes [BGTB93, Car93, SPP96]. Cautionary [GB95a, HCBH99, Hes94]. cautiva [Bus96, LRT93]. Cautivas [BR99, BLF^{+}93, RFDB94]. Cautiverio [BMLW96, BBCR99, CBAT97, KHM99, MJ97, DMMG96, PC93, SDB^{+}96]. cautivos [CRJ^{+}94]. Cave [KF95, RHCB93]. cavernas [RHCB93]. Caviar [BDSD98]. Caviar-Producing [BDSD98]. Cay [CLL91]. Cayos [BSC95]. Caza [ARRK97, KLR97, CLM^{+}95, PZSG96, BER97b, LA97, TA99]. Cedar [KL97]. Cedro [KL97]. Ceguera [LRT93]. censos [RC94]. Censuses [RC94, VW98]. Censusing [SB95a]. Centaurea [FCR^{+}98b]. Center [BKO^{+}95]. Centers [FS92, Pom93, WA95]. Central [FS96, GBAR97, Kat92, LSM99, Tho93, BGB96, BGB96, BS92, DPOO94, FL97, Fis94, FB96, GBAR97, HHF^{+}97, LP98, LNPC^{+}95, MA98, Nos98]. Centrales [FB96]. centro [LNPC^{+}95, Tho93]. centro-oeste [LNPC^{+}95]. centros [WA95]. Century [Bla94, TCO^{+}96]. Cercano [GWH97, SS99b]. Cercocarpus [RG95]. Cercocebus [KO91]. Cerdeña [VS95]. Cereus [SNP94]. cernícalo [CJ93]. Cetaceans [HWG99]. Cetaceos [HWG99]. cetros [MW93]. Challenge [DME95, Kau93a, Mec95]. Challenges [Cue96, Sou90]. chalumnae [HFS98]. Chance [Mor94]. Change [DPOO94, FSCL96, LS96, MB92a, MN96, MBBF95, PBBA90, Rod92, REDH98, RT91, Sad95, SPP96, SM97, SS94, Swa98, SJB97, WMB^{+}97].

Changed [GG90a]. Changes [Bal95, BGB96, CW97a, DDM93, FC99, Lit93, LCA99, Mar98, PRH98, RS91a, Ren99, RL96, Rob99, SBL96, WT90]. Changing [DSB^{+}95a, Geh96]. Channel [Dud99, Sha93, TW92a]. chaparral [BLD^{+}95]. Character [GQ99]. Characteristics [BY96, Hea94, KR95, NPM98, OLKD98, RDH97]. Characters [Cro92, DMD^{+}92, DDM^{+}92, Way92]. Charadrius [RRM93]. Chatham [AMA^{+}94]. Checkerspot [Wei99]. Checklists [DCL98]. cheeked [ES94]. Cheese [YH91]. Cheetah [CSD98, Mer94, Mil96, O’B94]. Cheetahs [Hed96, LWC95]. Chelonia [KB99]. Chelydra [PDC96]. Chiapas [Med94, WGO95]. Chicken [BCL^{+}98, PS96, SMS97]. Chihuahuan [HB95, HB96]. Chihuahuense [HB96]. Children [Kio91]. Chile [JFJ96, WDSA94]. Chiles [TNN^{+}99]. Chimpanzees [Gol97a]. China [GZWR97]. Chinchinero [WER^{+}97]. Chinkara [RH97a]. Chinook [BB98, Fis94, Hea94, HH94b, HHH95, NTT94, RLR97]. Chiroptera [FT91]. Chiropteran [HK94]. Choice [Ker97]. Choosing [RS95a]. Chorlitejos [PC93]. chorlitos [RRM93]. Chronic [Cre97]. Chubs [DDM93, DDM93]. Chukar [KASR99, KASR99]. Cichlid [KO93, Rei93, SWK^{+}97]. Cichlids [CKCM95, OO93]. Cicindelidae [PC92]. Cíclidos [SWK^{+}97, KO93, OO93, Rei93]. ciclónicas [PERC96]. Ciencia [BO96, MHWP96]. científicas [BGF95]. Científico [WP97]. científicos [LN96]. cinco [GAC^{+}95]. Cinegéticas [TA99]. cingulatum [MPB96]. cinturón [NHA95]. Circumboreal [Bus92]. Circundante [HNC97]. Citations [SK99]. Citrus [PBKM95]. Civil [HHF^{+}97]. Clapper [FFL95]. clareo [MPB96]. Clarifying [Gol99a]. Clark [MS99a]. clarki [GL95]. Clarkia [MBH96]. Claros [DH97, ST96]. clases [PL95]. Clasificación [GD99a, Tay95, OSEC95]. Classes [PL95]. Classification [GD99a, Kei98, OSEC95, Tay95]. Classifications [MGM^{+}98]. clausura [RH93]. claves [WH95]. Clear [NLS93]. Clear-Cut [NLS93]. Clearcut [Ash97, KDG98, Str96]. Clearcuts [AP99, MCO98]. Clearcutting [DM92, Duf93b, Fra97, JFH93, KGD96, Pet99, WN98a]. Cleared [WGO95]. Clearings [ST96]. Clemmys [Gra95]. Clethrionomys [Mil95b]. Cliff [CK98]. Cliffs [KL97]. Clima [BWS97]. Climate [BWS97, PBBA90, PC94, RT91]. Climatic [Rod92, RS93c, SS94]. climáticas [PC94]. climáticos [RS93c]. Climbing [CK98, KL97]. Close [GWH97]. Closed [RH93]. Closest [BJ96]. Clover [HVG91]. Coast [BPS95, Mef92]. Coastal [BSR97, CD95, MPC99, NPM98, RF97, Sha93]. Cobertura [VTM99, PL94, PL95]. cociente [TB95]. Cockaded [HBA93, RWES93, SLS92]. Code [Col92]. Coelacanth [HFS98]. Coespecificos [RMCM97]. Coevolutionary [AS94, Dud99]. coevolutivas [AS94]. Coexistence [HD99]. Coexistencia [HD99]. Coffee [GBAR97, MT99]. Cola [WSE97]. colapso [DF96]. colecciones [BGF95]. Colecta [AM99, SB99]. Colectados [Gar97]. Colémbolos [Deh96]. Coleoptera [NLS93, PC92]. colibríes [RCHP94]. Colinas [LC99]. Colisiones [BH96]. Collaboration [Med98]. Collapse [DF96, FAP95, LC98a]. Collected [Gar97]. Collecting [GL94, KQK93]. Collection [AM99]. Collections [BGF95]. College [Dun99]. Colleges [Orr91d]. Collembola [Deh96]. Collision [RS93a]. Collisions [BH96]. Colombia [Kat92]. Colombian [ART94]. Colonialism [Baw92]. colonias [RS95b]. Colonies [RS95b]. Colonizacíon [BTRM99, DBCM95, KL97, Tho94a, Fra97]. Colonizadores [FB96]. Colonization [BTRM99, DBCM95, Fra97, SP90, Tho94a]. Colonizing [FB96]. Colorado [GLFZ96, Kar90a, Rob99, Bro96, Car93, GLFZ96]. Columbia [Wal99b, Ber95b, Bun95, LAF93, MGM^{+}98, Wal99b, WPSZ96]. columbiae [CG98]. Columbian [DD97]. Columnar [PP96]. Columnares [PP96]. Comando [HM96]. Comando-y-control [HM96]. Come [Sie96]. Comentarios [HLAS96]. Comercial [SB99, ADM94, Ber95b, WSB92]. comercio [Yu94]. Comfortable [BG95a]. Coming [Sou90]. Command [HM96]. Comment [VL91]. Commentaries [Ano96e]. Comments [Ano98b, GM91, HLAS96, IU92, MR98a, OG96, PH92, Sch91, Tho91, Wil91]. Commercial [ADM94, Ber95b, Mat94, SB99]. Committees [KM98]. Commodity [MWCP94]. Common [Alt90, BL99, Gur92, HCB94, HH94c, PDC96, Pov94, RS93b, RS93a, YB96]. Commons [Kay97, San99]. Communal [PVZJCI90]. Communities [ACB^{+}99, BT94, CK98, CD95, FS92, GSAZ98, GNH94, HNC97, HRV^{+}99, HGNK93, Hut95c, KB94, LFM^{+}98, Mun95, OSEC95, PTTN^{+}96, RFG^{+}98, RN96, SEC98, WP97, WDSA94, WH95]. Community [BGB96, EK98, Hac99, NBH^{+}98, RF99]. como [BSW^{+}96, Ber97a, BHM94, CM97, DW93, Gol99b, GK99, GS99, Hil95, KMW95, LF97, MVH96, MP99, NMGS94, OB96, PD97, PNS99, PBB99, Rei95, RCM97, SFM96a, WH95]. comparación [KB94, CFL96, TA97, WGM^{+}96, YB96]. Comparaciones [BL99, GWH^{+}99]. Comparative [SMHK98, YB96]. Comparing [BH97]. Comparison [CFL96, FNC98, KB94, KR98, SDT93, TA97, WN98a, WGM^{+}96]. Comparisons [BL99, GWH^{+}99]. compatibilidad [Cur93]. Compatibility [Cur93]. competencia [CC96]. Competition [CC96, KW96, Kin92]. complementarias [WGM^{+}96]. Complementary [WGM^{+}96]. Completamente [Ash97]. Complex [KF95, Leb98]. Comportamiento [GB97, WSE97, JBB96]. Composición [HNC97, KSO^{+}99, Ren99]. Composition [HNC97, HE97, Hut95c, KSO^{+}99, NBH^{+}98, Ren99]. Comprehensive [BN98b]. Comprensión [PC96, BJ95]. Computer [Rol95]. Comú [WER^{+}97, YB96]. común [HCB94, RS93a]. Comunes [BL99]. comunidad [BGB96]. Comunidades [ACB^{+}99, Hac99, HNC97, HRV^{+}99, HGNK93, WP97, WDSA94, BT94, GNH94, KB94, LNPC^{+}95, OSEC95, PTTN^{+}96, Sab96, WH95, CD95, RN96]. Comunitaria [RF99]. con [BY96, Ber97a, BB93, BH96, CRHSS96, CM93, CC96, DS95, DPOO94, DSB95b, Gre96, GL95, Hac99, HDE95, LRT93, LDP95, MPC99, OO93, PKH97, PZSG96, PL95, RVK^{+}94, RL96, RFM97, Rob95c, RMCM97, RS93c, SHM97, Sym96, VM99, VW96]. concecuencias [CLM^{+}95, PB95, Was96, WLAH92]. concentración [WGM^{+}96]. Concept [BFS98, CM97, DLP^{+}92, LN90, Pai95, UKG^{+}96]. Concepto [CM97]. Conceptos [RHS94, PNS99]. Concepts [CCM99, Cro92, PNS99, RHS94]. Conceptual [And91, Aus98]. Conceptualización [MH99b]. Conceptualizing [MH99b]. concequencias [BLF^{+}93]. Concern [DHW99]. Concerning [Car93]. Concerns [Est98, MS92]. concervacionistas [CLM^{+}95]. Conclusiones [PKH97]. Conclusions [PKH97, War94b]. Concordance [GQ99]. Condado [DS92]. Condición [BR99, LSM99]. Condiciones [MJ97]. Condition [YG92]. Condor [Mil95a, SDC^{+}93]. Conducen [MHB^{+}96]. Conducta [KB97, Ree99]. conectarse [RS93c]. Conectividad [TA97]. Conference [NY91]. Conferences [Ano91n]. Confiabilidad [MDW99, Tay95]. Confidence [Ree96]. Configuración [VTM99]. Configuration [RS92, VTM99]. Conflict [NMGS94, SGCJ94]. Conflicto [SGCJ94, NMGS94, RS93a]. conflictos [FYP96]. Conflicts [FYP96, TFHD98]. confluentus [RM97]. Congener [GWH^{+}99, GP97, YB96]. Congenere [GP97, GWH^{+}99, YB96]. Congo [Sto96, WSB92]. Congruence [BL95a]. congruencia [BL95a]. Conifer [Bil91, Hut95c, MBBF95]. Conífera [MBBF95]. coníferas [Ben93]. Conifers [Ben93]. Conjunto [MHB^{+}96]. Conmoción [War94a]. Connection [Avi95, PBBA90, Ran96, Wes91]. Connectivity [BN98a, DSB^{+}95a, HMO98, TA97]. connotaciones [Bun95]. Conocimiento [JM97, GL93a, HDE95]. Consecuencias [SFLC95, Bur95, ST95, RFM97]. Consensus [UA94]. Consequence [SS94]. Consequences [Ano98d, BW91, BLF^{+}93, GG90a, PCM98, RFM97, SHM91, SKOO98, SFLC95, SJB97, Was96, WLAH92]. Conservacion [BO96, Aco99, ABB^{+}97, AFF^{+}97, BTRM99, BN97, Bri96, CM97, CDD96, CO99, CCR96, DGG^{+}99, DSGS93, GSB97, Hac99, HB96, KL96, KBRB96, Ker97, LA97, LBR99, MRN99, MS97a, MB99, MA96, MDW99, MBD^{+}99, NQH^{+}96, PS99, PNS99, PDS96, Pri96b, RTNK99, RH96, RRS96, RW99, Sab96, SM99, SAQMGQ^{+}99, SHM97, SG99, Sto97, Sto96, SF97, TTJ^{+}99, WPR96, Wit97a, YS96, NQH^{+}96, GB95b, ADM94, BL95a, BSW^{+}96, Ben93, BC94, BLF^{+}93, Bro96, Bun95, Can96, CMMLRR95, Cla95, DS95, DP96, FBFO96, Gin94, GMG94, HP94, HLAS96, Hes94, HD96, KO93, KSC^{+}96, KMM94, LRT93, Lan95, LA95, LP96, Lin95, LCSC95, LNPC^{+}95, Med94, Mer94, MSC96, OSEC95, OFH94, OO93, PT95, PGR^{+}94, PDC96, PSGS94, RS93a, SK96, SFM96a, SFCM92, ST95, SK93, SB94]. conservación [TG93, Thi95, TTW^{+}94, Ver93, VD94, Wal95, WKS^{+}94, WMH94, You94b, BG95b, Ham94, Med93, BH93, CM93, FM94, Hac99, MLP^{+}97, PC96, PB94b, RSR96, RR99a, WR96]. conservacional [WA95]. conservacionista [FDV94]. Conservacionistas [Hab96]. Conservacionsitas [RB97a]. conservando [Rei93]. Conservar [Lam97, KGS93, Rob95c, WGM^{+}96]. Conservation [Aco99, Alc93, All97, ABB^{+}97, Ano90g, Ano90i, Ano90j, Ano90k, Ano90l, Ano91h, Ano91i, Ano91m, Ano94b, Ano97c, Ano98d, Ano98a, Ano99f, Ano99j, Ano99m, ALF92, AFF^{+}97, ADM94, AMW90, Avi95, Bal96, BL95a, BHSO92, BO96, BS98, BSW^{+}96, Ben93, BC94, BC98, BTRM99, BB90, BN97, Bir93, BWH90, BDR^{+}98, BH93, BLF^{+}93, Boz93, Bri96, Bri95a, BM95, Bro96, Bun95, Bur90, BG95b, CP94, Cal90a, Cal90b, CM97, CCM99, CDD96, Can96, CPG94, CO99, CMMLRR95, CV95, CCR96, CPC96a, Cla95, Cob90, CLM^{+}95, CMB96, CDV93, CDB99, CM93, CEPR91, CSD98, Cro92, CAB^{+}98, DS95, DHJG91, Dav96, Dea95, DGG^{+}99, DC98, DMD^{+}92, DDM^{+}92, DP96, Edu91, Ehr92a, Ehr92b, EV98, Ela96, Eri99, FM94, Fai96, FF95, FDV94, Fit94, Fle90, Fle96]. Conservation [FWB^{+}99, Fle94b, FS91, Fra92, FBFO96, FNN92, Fri92, GWW93a, GB96a, Gin94, GMG94, Gin97, Gin99, GFP98, G96, GSB97, GW98, GB95b, Göt92, GQ99, GLFM98, Gru96, GG90b, Hab96, Hac99, Hag92, Hag95, HMO98, HR93, Ham94, HJ99, HHH96, HHF^{+}97, HP94, Har93, Hau95, Hea94, HSS94, HLAS96, HE91, HB96, Hes94, HD96, HH92, HAD^{+}97, HS90, Hun94, Hut95b, Jac90, Jac93, JVM95, JD98, JS98, JG91, JS93, JB99, Jop94, Jor92, KASR99, Kar90b, KO93, Kau95a, KGS93, KL96, Kei98, KBRB96, Ker97, Ker96, KSC^{+}96, KLF98, Kio91, KM98, KF95, Kra95, KCE^{+}93, KMM94, KRL98, LRT93, Lam97, Lan95, LAF93, Leo90, LA95, LT92, LKM90, LA97, LP96, Lin95, LCSC95, LC98a, LBR99, LNPC^{+}95, MBJ^{+}97, ML90, Mag91, Mag96, MRN99, Man94]. Conservation [Mar92, MS97a, Mas91, Mat97a, Mat99, MR91, MB99, McC96, McL93, MJ96, MBD95, MQ99, Med94, Med93, Mef98b, Mef99a, Mef99b, MM96, Mei99, MB95, Mer94, MA96, MDW99, MBD^{+}99, MR98a, MMT^{+}98, MT99, MAS^{+}97, MSC96, MLP^{+}97, Mun95, Mur90, MD96, NF93, Nab95, NABE95, Nos94, NQH^{+}96, Nos96a, Nos96b, Nos97, O’B94, OPRCB95, OFH94, OO93, Orr90a, Orr90d, Orr91c, Orr91b, Orr95, Orr99b, Pae99, Tru94, PS98, Par95, PS99, PC92, PC98, PB94a, Per94, PT95, PNS99, PEP98, PGR^{+}94, PDC96, PH94, Pov94, PSGS94, PDS96, PC96, Pri96b, PB94b, PSB98, Put93, Put97, PAG^{+}99, Qui95, RB97a, Ran96, RTNK99, RH96, RSR96, RS93a, RR99a, RSM90, Rei94, RS91b, RR99c, RO99, RRS96, RP99, RW99, Roj92, SK96, Sab96, Sab98, SJ98, SM99, Sal90, Sam90]. Conservation [Sam96b, SAQMGQ^{+}99, Sar96, Sch96, SCAB91, Sch92, SF96, Sib94, Sid98, SFM96a, SFCM92, Sim94, Sim95, SPD^{+}98, ST95, SW94, SK93, SHM97, SK99, SG99, Spe95, Ste98, Sto97, Sto96, Sto95, Str91, SB94, Str98, SF97, TG93, TD96, TFHD98, Tem92a, Tem92c, TNN^{+}99, Thi95, TB98, TL94, TD94, TB96, Tri98, Tro94, TTW^{+}94, TTJ^{+}99, Uhl98a, VP97, Ver93, Vid93, VIO98, VD94, WW98a, WTWS98, Wal99a, Wal96, WT90, War94b, WLG^{+}91, WPR96, WR96, Wes91, Wik90, WDR^{+}98, WKS^{+}94, WA95, WL97, Wit97a, Wit97b, WMH94, YS96, You93, You94b, You95, DSGS93, Boz93, Ehr93, Mef98a, Nos93, Tem93c, WM91]. Conservation-Minded [Hun94]. Conservationists [Hun96b, Pea92]. Conservatism [Orr95]. Conservative [GFA90]. Conserve [DW93, Rob95c]. Conserved [NE94]. Conserving [Alt94, BMN92, Bus92, CG95, Day91, HH94c, Mag92a, MSM99, MWW99, OSEC95, OD98, Phi93, Rei93, SS98, Wal95, WGM^{+}96]. Consevación [KASR99]. Consider [Sta96]. consideradas [BHM94]. Considerations [ABM91, WDWG94, Wap94, YM97]. Consolidation [Str96]. Conspecific [SP90]. Conspecifics [RMCM97]. Constituency [Orr93b]. Constraints [Ber97a, Tho99]. Construcción [LC99]. Constructing [LC99]. Construction [CKB98]. Construir [CZN99]. consumo [MWCP94]. Consumption [Cob97b, Cob99, JJ99]. Contact [Bis93, TSSB95]. contagiosas [Hes94]. Contagious [Hes94]. contaminación [CBCC93, OR94]. Contemporary [DM90b, JDEB98]. Contiguous [DSB^{+}95a, PBG^{+}93]. Continental [RLD99]. continente [SK93]. Continued [Hei95b]. Continues [Hut95a]. Continuing [NY91]. Contra [BLS99, HRCTS99, MFR97, KB99]. contracciones [FGG^{+}95]. Contractions [FGG^{+}95]. Contraste [PS99]. Contrasting [BHL^{+}92, PS99]. Contribución [HP99, RJBB96]. Contribution [HP99, MCO98, RJBB96, SM91]. Control [Ehr94b, GB95b, Hab96, HM96, MEE93, Mef94, WRSP97]. Controlled [GLFM98]. Controlling [Hab96, PGFD95]. Controversy [Hut95a, Lym94, O’B94, Sch93a, WM91]. Convenience [Nos96b]. Conveniencia [VM99]. Convention [Baw95b]. Conversation [Ano99g, Pai95]. Conversion [KTA94, KTA94]. Cook [RHSM94, SGW90]. Cookbook [Ano99j]. cooperación [Can96]. Cooperation [Can96]. Coral [Cob97b, Cob99, Guz91, HRV^{+}99, JJ99, MKM^{+}96, MKA96, MBW^{+}98, Rob95a, SB95a]. Coral-reef [MKA96]. corales [MKM^{+}96]. Coralino [MKA96]. Cordillera [Kat92]. Core [BN98b, BFS98]. Coreana [HOF^{+}96]. cornamenta [FBS95]. corporal [RHCB93]. corredes [Bei93]. Corredores [BC97a, DHE97, MP99, TA97, DBCM95, Haa95, Hil95, MVH96, RDN94, Hes94, SFCM92, ST96]. correlación [Fim94]. Correlaciones [CRHSS96, KQK93, Pea97]. Correlates [ARR90, CRHSS96, Fim94, FH97, Lau91, Pea97]. Correlation [ES94]. Correlations [KQK93]. Correlativos [FH97]. corridas [HHH95]. Corridor [DBCM95, Har92, RDN94]. Corridors [Ano97e, Bei93, BN98a, BC97a, DHE97, Haa95, Hes94, Hil95, IU92, LN93, MVH96, MP99, RDN94, SFCM92, SMHK98, ST96, TA97, TW92b, Yah95]. Corta [BB99]. Cosecha [BRHD97, PKH97, OK96, GMG94]. cosechados [NGN96]. Cost [SM99]. Cost-Effective [SM99]. Costa [BBL93, BWBM99, Boz93, CC99b, Guz91, Sto96, BSR97, BJW95, BPS95, CC99b, FS95, HDE95, Hun94, PCG91, Sto96]. costas [KQK93]. costera [Sha93]. Costeras [MPC99, CD95, MSC96, SK93]. Costo [SM99]. Costo-Efectiva [SM99]. Costos [MRN99, Fle94a]. Costs [BCGL93, Fle94a, MRN99, McC95, WC98]. cotorras [WKS^{+}94]. Cotton [CRJ^{+}94]. Cotton-Top [CRJ^{+}94]. Cottonwood [SAS95]. Cougars [Bei93]. Countries [ABM91, Bal99, Baw95b, Gra99, SK96]. Country [BL95a]. Counts [RM97]. County [DS92]. Cousteau [Gra98]. Cover [BB93, HH98, VTM99]. Coverage [PL94, PL95]. Cowbird [TB93, TB93]. Cowbirds [HH95a]. Cows [Nos94, Wei99]. Crab [RCMR97]. Cracks [SS94]. Crafting [CCR96]. Crane [Can96, CRHSS96, GSB99, HOF^{+}96]. Crater [MAR96, PPR^{+}91, MAR96]. Crayfish [FC93, GK96, GKA97, GW97, KF95, LEMC95]. Creando [PNH93]. Creating [PNH93]. Creatively [LG97]. Crecimiento [CC99b, FBS95, PC93, Rob95b]. Crested [KO91]. Cría [VM99, DMMG96, SDB^{+}96]. criados [DMMG96]. crianza [PC93, RJL95]. Crías [MB99, JBB96]. Cricket [KA92]. Crisis [Hag92, HK94, SBB^{+}94]. Criteria [Bot94, GN92, Kei98, LRD^{+}96, PT95]. Criterio [RDE96, PT95]. Criterion [Dud92, Göt92, RDE96]. Criterios [LRD^{+}96]. Crítica [PD97]. Critical [KBM98]. críticas [TSHG95]. Criticisms [TSHG95]. Críticos [NM97]. Critique [Gru90, Pae99, PD97, Pet99]. Crocodile [Tho99]. Crocodilian [Tur93]. Crocodilians [Tho99]. Cronartium [TCK^{+}95]. Crop [Alt94, Ano98d, BKO^{+}95, KEE91, NT98]. Crop-Weed [KEE91]. Crops [TTJ^{+}99]. Cross [FT96, PC93]. Cross-Boundary [FT96]. Cross-Fostering [PC93]. Crossbills [Ben93]. Crossing [DH97]. Crowned [TB93, SB94]. Cruce [DH97]. Crucial [SRR95]. cruciales [SRR95]. Crumbling [Win96a]. Cruza [HWB96]. cruzada [PC93]. Cryan [LRD^{+}96]. Cryptic [RSMB99]. Cuan [Pat94a, AL93]. Cuándo [LA95]. Cuantitativa [GD99a, CRJ^{+}94, PGR^{+}94]. cuantitativo [HMPK96]. Cuatro [GJH96]. Cuba [Edu91]. Cubierta [BB93]. Cuello [GSB99, LCA99, RFM97]. Cuellos [BGK^{+}97, RL96]. cuenca [Fra94, GL93a]. Cuernos [RB97a]. Cuerpos [Gar97]. cuidari [Rob93b]. Cultivation [ART94]. Cultura [KBRB96]. Cultural [BAR94, BM97, Dud97, Mag92a]. Culture [HB94, KBRB96]. Cumberland [HE97]. cumplimiento [LWMG93]. Curacao [PP96]. Current [BL95a, CCM99, Pae99]. Curricula [SK96]. curriculum [SK96]. Curves [Dia95]. Cut [NLS93]. Cutthroat [GL95]. Cutting [HHR99]. Cyclonic [PERC96]. Cyrtaucheniidae [RF97]. Cytochrome [RH97a].

D. [YB96]. Dacus [Tsu98]. Daily [Day91]. Daisy [DeM93, YBZ99]. Dakota [MBT90]. Damage [FL97, NT98]. Damaged [SGK95]. Dams [HSGY98]. Dances [Sym96]. Dangers [Nab95]. Danube [LT92, Pri91]. Danzas [Sym96]. Darling [Dud93]. Darwin [GG97, YJG^{+}96]. Data [BBGB92, BCG94, CZN99, CG98, GDK99, Gre92, KSC^{+}96, KRG^{+}99, Los95, LC98b, MHB^{+}96, PRH98, PC98, RHK97, SGA98, SRR95, SDC^{+}93, Str99, TSHG95, WGRT96]. Date [Cor96]. Datos [CZN99, GDK99, KRG^{+}99, MHB^{+}96, Str99, WGRT96, KSC^{+}96, SRR95, SDC^{+}93, TSHG95, TM96]. Daviesia [YB96]. Dayak [LLP95]. dc [ART94, SL93]. Deals [RR99c]. Death [TL94]. Debate [Cle93, HM98]. deben [AL93]. deberían [LNA95]. debida [Cow99]. debidas [BGTB93]. Debido [EE97, WMB^{+}97, Was96]. Debris [Fra97, MB99]. Debt [Cow99, GL93b]. Debts [MLD97]. Decades [Sou90]. decaryi [RSR96]. Decay [Jam94]. Deception [Wil94]. Deciduos [GFR^{+}99, SB94]. Deciduous [AF98, Bra94, GFR^{+}99, SB94]. Decision [LSC^{+}96, RS95a, SFM99]. Decision-Making [RS95a]. Decision-Support [LSC^{+}96]. Decisiones [SFM99, LSC^{+}96, RS95a, DH97]. Decisions [DH97, MHH98]. declinación [Ber95b, BJ95, FS96, LMS96, RHCB93, ST94, WLAH92, BWS94, PC94]. declinaciones [BGTB93, FGG^{+}95, LEMC95, Sim96, You94b]. Decline [AR97, Ber95b, BBL93, BR99, BJ95, Cow99, FS96, Gel99, GD99b, KW96, LMS96, LEMC95, Lip98, MW90, Moy94, RHCB93, ST94, VT99, WLAH92, WCD97]. Declined [Zei92]. Declines [Ano98d, BWS94, BGTB93, EB99, FGG^{+}95, HG97a, Lip99, NY91, PC94, PFSG97, RB97b, Sim96, Str99]. Declining [Agu94, HHH96]. Declive [Cow99, VT99, GD99b]. Decreased [FP95]. Deducciones [Rob95c]. Deep [BL95b, BCP95, JOB96]. Deep-Sea [BL95b, BCP95, JOB96]. Deer [AF98, KLF98, MS93, Son96]. Defaunation [DM90b]. Defensa [JM97]. Defense [JM97, LMS97]. Defenses [BV97, VB99]. Defensibility [PT95]. Define [GQ99, Jor93]. Defined [CFL96]. Definición [PD97, DGG^{+}99]. Definidos [CFL96]. Definiendo [RDN94]. Defining [DGG^{+}99, PAG^{+}99, RDN94, WDR^{+}98]. Definition [PD97]. Definitions [Gen94]. Deforestación [BPO99, Cow99, FP95, GL93b, BSC95]. Deforestaciones [BPC97]. Deforestation [BSC95, BPC97, BPO99, Cow99, DG92, FP95, GL93b, Tom92]. Degradación [Sab96, Doa95, You94a]. Degradation [Doa95, Sab96, Sab98, You94a]. Degree [DBH^{+}99]. Dehorned [RB97a]. Dehorning [BCG94, LM94]. del [Aco99, ACB^{+}99, ABB^{+}97, Ama93, AMGJ95, AFF^{+}97, AL93, Ash97, Bak94, BY96, BWS97, BHCB99, BB97, Ber97a, BC99, BB93, BLF^{+}93, Boz93, BBSK96, BHM94, BLD^{+}95, BPC97, Bro96, BRHD97, BJ95, CJ93, CW97a, CSBC95, CRGGFT97, CUM96, CCOO^{+}96, Cla95, CBCC93, CD95, Cow99, Cua97, CBAT97, Cur93, DOSP94, DeM93, Doa95, DS92, DLK^{+}95, DMMG96, DSGS93, DSB95b, EE97, EDFM96, EKL^{+}99, EBV95, FESS94, FS95, Fim94, FH97, Fle94a, Fol94, FT96, FB97, FC93, Fra94, FEM95, GB97, GK96, GMMR^{+}96, Gar97, GD99a, GAC^{+}95, GLFZ96, Goe97, GWW93b, GF97, GK99, Gre96, GD99b, GL95, Gru94, HVM96, HBA93, HWB96, HNC97, HRV^{+}99, HM97, HHH95, HDE95, HM96, HV97, HW95, IS96, JM97, JHHB99, KGS93, KL97, KA92, KHP95, KS94b, KSO^{+}99, KTA94]. del [KRG^{+}99, LMNV92, LNVM94, LDSH99, LV94, LEMC95, LF97, Lin95, LDP95, LSC^{+}96, Loc99, LCSC95, LBR99, LCA99, LM95, MWM^{+}99, Mad99, MHO97, MS97a, MJ97, MW93, MM94, Med93, MN96, MHB^{+}96, MWCP94, MSHW95, Mor99, MSC96, NGN96, New95, NHA95, NE94, OO93, OR94, PKH97, PT95, PEH93, PERC96, PF99, PC94, PL95, PC96, Pri96b, QLDV96, RDE96, RF99, RLR97, RK93, Rei93, Ren99, RDN94, Ris95, Rob95b, RHSM94, RH93, Rob95c, RMCM97, RHS94, RCM97, SWK^{+}97, SHMV97, SB99, ST94, Sim96, SK93, Sta95, SDC^{+}93, Sun96, SMPN99, SF97, Sym96, TG93, Tay95, TM96, Tho93, TV99, TB95, TB93, Tsc92, TTJ^{+}99, VS95, VHM94, VRV^{+}94, Vog95, Wal99b, War94a, WPSZ96, War94c, Was96, WLAH92, WP97, WDSA94, WRSP97, Yaf99, YB96]. Delay [All93b]. Delayed [CDV93]. deletereo [LRT93]. Deleterious [LRT93]. Deletions [HM98]. del [OO93]. Delineation [Sem98]. Delisting [Bot94, RDE96]. Delta [GLFZ96, LT92, Pri91]. deltoides [SAS95]. Demands [SBB^{+}94]. Demilitarized [HOF^{+}96]. Demografía [WSE97, DLK^{+}95, MFR97, BBCS99, Mad99, SGMP99]. demográfico [BLD^{+}95, PZSG96, PK94, RRM93]. demográficos [LS96]. Demográificas [MAR96]. Demographic [BCL^{+}98, BLD^{+}95, CSD98, GG90a, Leb98, LS96, MAR96, PZSG96, PK94, RRM93, SGA98, VW98]. Demographics [BBCS99, CDV93, MC95a, SO94]. Demography [Avi95, DLK^{+}95, MW98, Mad99, MFR97, Mat94, OG96, Ran96, SGMP99, Str91, TD96, WSE97]. Dendrocopos [CA92]. Denial [OE95]. Densidad [NMGS94, RMCM97, SD94, WA95, GS99]. Densities [SD94]. Density [GFA90, GS99, KHS98, MG97, NMGS94, PV98, RMCM97, SCAB91, TJ99, WA95]. dentro [DP96, Mor99, RBPB96]. Departamento [JM97]. Department [JM97]. Dependence [GFA90]. Dependencia [PSGS94]. Dependency [PSGS94]. Dependent [NABE95, SPR97, WP97]. Dependiente [SPR97]. Dependientes [WP97]. Depletion [Ano90g, HP94]. Deportiva [LA97]. Deposición [Wei99]. Deposition [Wei99]. Depredación [SW99a, YS96]. Depredadores [CS97, MP99]. Depredation [YM96]. Depresión [EKL^{+}99, Mer94, MGB94, PSGS94, HP95, KHM99]. Depression [EKL^{+}99, FNC98, HP95, KHM99, LR91, Mer94, PH96, PSGS94]. depressus [WOA98]. derechos [DME95]. Dermatemys [PH94]. Derrame [Gar97]. desafio [Mec95, DME95]. desaparecidos [GL93a]. desapareciendo [YJG^{+}96]. desaparición [Car93, Ash97]. Desarrollo [HCBH99, SM99, Cla95, CBLV94, FEM95, HVM96, KMM94, OK96, Pre94, SK96, VS95, CAH97, Har94, RDE96, RH96]. Descornados [RB97a]. Described [PDSB98]. desde [DP96]. Desenlistar [RDE96]. Desert [BW91, Ber97a, BWH90, HB95, HB96, LV94, QLDV96, RMCM97, SMPN99, SJB97]. Desert-dwelling [BWH90]. Deserts [CV95]. Desierto [Ber97a, HB96, QLDV96, RMCM97, SMPN99, LV94]. Design [ABM91, De 95, Har92, HWG99, Hug91, LNVM94, LN93, MH91, MWCP94, Orr92b, PT95, PB95, RRS96, Sch98, SB95c]. Designing [BMLW96, KRG^{+}99]. Designs [IU92]. Desktop [Hun96b]. Desmilitarizada [HOF^{+}96]. Desmoronamiento [Win96a]. Después [Gar97, MFR97, Ren99]. Destiny [Web99]. destrucción [OR94]. Destruction [Bra96, OR94]. detección [LWMG93, LS96]. Detect [RM97, Str99]. Detectar [Str99]. Detected [NBH^{+}98]. Detection [LWMG93, LS96, SW94]. Determinación [HM97, PNS99]. Determinando [Bei93]. Determination [HM97, Leb98, SP90]. Determining [Bei93, PNS99]. Deters [GKA97]. detritívoras [GWW93b]. detritívoras/fitoplanctívoras [GWW93b]. Detritivorous [GWW93b]. Detritivorous/Phytoplanktivorous [GWW93b]. detritos [FP95]. Detritus [Fra97]. Deuda [Cow99, GL93b]. Deudas [MLD97]. Developing [Baw95b, CAH97, Gra99, Pre94, RDE96, SK96]. Development [AKG98, BAR94, CBLV94, CD92, Edu91, Fit94, FEM95, HVM96, Har94, KMM94, Mar93, MBW^{+}98, OK96, PBKM95, Pel92, RH96, SM99, VS95, Wil94, You93]. Developmental [Cla95, HCBH99]. Diagnosing [Kob99, VD94]. Diagnosticando [VD94]. Diamondback [RCMR97]. Diceros [OFH94, AMW90, OFH94, SF97]. Die-Offs [You94b]. Dietary [MB99]. Dietas [MB99]. Diferenciacion [SF97, HCB94]. Diferenciadas [WS97]. Diferencias [RWES93]. Diferentes [MHB^{+}96, SPR97]. Differences [RWES93, Rym94]. Different [Ber90, CB98, FESS94, Goo94, Lym94, MHB^{+}96, SPR97, WLDGGZ98]. Different-Aged [CB98, FESS94]. Different-sized [Ber90]. Differential [Hel98]. Differentiated [WS97]. Differentiation [HCB94, SF97]. Dilemma [Cob90, Goo95]. Dilución [MB99]. Dilution [MB99]. diluvialis [ST95]. Dimension [SK96]. dimensiones [SK96, WP96]. Dimensionless [WP96]. Dimensions [JD98]. Dimmick [Wap98]. Dinámica [BB97, VPW97, GMG94, JFJ96, KA92, QAM96, SY94]. dinámico [LMNV92]. Ding [Dud93]. Dioecious [Bil91]. Dioxide [CT99]. Dióxido [CT99]. Dipterocarpacea [MGB94]. Dipterocarpaceae [MGB94]. Direcciones [HLAS96]. Direct [EV98, KB94]. directas [KB94]. Directions [HLAS96, JVM95]. Disappearance [Ash97, Car93]. Disciplinary [Sch96]. Discipline [Orr93d]. Disciplines [Orr93d]. Disciplines* [HJ99]. Disciplines/The [Orr93d]. Discovered [PSSB95]. Disease [AS94, AR97, Cun96, Gin94, HG97a, Hes94, LMS96, LMS97, Sim95]. Disease-Mediated [Sim95]. Diseñando [BMLW96]. Diseño [RRS96, De 95, MWCP94, PT95, PB95, SB95c, HWG99, KRG^{+}99, LNVM94]. Disequilibrium [BBGB92, BW91]. Disjunct [HMW98]. disminución [FP95, BR99]. Disminuciones [Str99, Gel99]. Dispersal [Aco99, Haa95, Hil95, PM92, RH93, RHK97, RHK99, Sch98, Sou99, SB94, ST96, TA97]. Dispersers [CC95]. Dispersión [TA97, Haa95, PM92, ST96, Aco99, Sou99, SB94]. dispersoras [RH93]. dispersores [Hil95]. Disponibilidad [LLP95]. Distances [RS95b]. Distancia [RS95b]. Distanciamiento [PQL99]. Distant [WCD97]. Distinct [Wap98, YJG^{+}96, PD97]. Distinctiveness [WOA98, ZDM98]. distinta [FESS94]. Distribución [AFF^{+}97, CW97a, GWH^{+}99, HRCTS99, HWG99, RDH97, BEG^{+}94, BGB96, EDFM96, FGG^{+}95, Gre96, KHP95, NHA95, PM92, VHM94]. Distribucional [SM97]. Distribuciones [BHS99]. Distribution [BEG^{+}94, Bla91, BWH90, BGB96, BE94, CW97a, CDB99, DH98, EB99, Gre96, GG90b, HS98b, HRCTS99, Hol95, HWG99, KHP95, MA98, MG97, NHA95, PM92, RDH97, SAM98, SFM96b, TSP97, VHM94, HB95]. Distributional [SM97]. Distributions [BHS99, EDFM96, HY93]. Disturbaciones [SPR97]. Disturbance [ACB^{+}99, Fim94, HH92, Kai98, LC99, ML94, NHA95, PC94, RFG^{+}98, RS95b, SPR97, WBKH97, WN98a]. Disturbance- [Fim94]. Disturbance-Dependent [SPR97]. Disturbances [SB97]. Ditch [Cra94]. Divergence [Leb91]. Divergent [TSSB95]. Diverse [LM95]. Diversidad [BBCS99, BRHD97, CAH97, Cow99, GWH^{+}99, HWB96, HRCTS99, HGNK93, KASR99, LC99, NM97, Pea97, PBB99, RCM97, SMV96, BGB96, CB95, DS95, Deh96, HGK^{+}96, HD96, How94, KMW95, KS94b, LM93, MCR94, Sun96, Thi95, Wal95, WGM^{+}96, Bri96, GHB95, GJH96, GWH97, LSM99, Med94, PBB99, TBL^{+}97]. diversificado [ML94]. Diversity [Ang94, ABM91, BAR94, BBCS99, BGB96, Bri96, BKO^{+}95, BMN92, BRHD97, CB95, CNP^{+}97, CAH97, Cow99, DS95, Deh96, DB98, DKOW94, DM90b, ELS93, GWH^{+}99, GHB95, GJH96, GWH97, HBD90, HWB96, HGK^{+}96, HD96, HRCTS99, HH92, HGNK93, How94, KASR99, KMW95, KS94a, KS94b, LSM99, LB95, LC99, LM93, Mag92a, Med94, MS99b, MCO98, MA98, MCR94, NM97, Orr91a, PWC^{+}98, Pea97, PBB99, PSB98, RS92, RCM97, Sal91, SMV96, SSS94, SS95, Sun96, Thi92, Thi95, TBL^{+}97, VL91, Wal95, WDWG94, WGM^{+}96]. Diversity-Science [KS94a]. diverso [LM95]. Dividing [RDN94, Yah95]. Diving [GSAZ98, HRV^{+}99]. divisores [RDN94]. Dja [MN99]. DNA [AMW90, Avi95, BHKH96, DMMG96, GSB99, Gol97a, Gra95, LBR99, MAS^{+}97, NTT94, OFH94, PDSB98, Ran96, RFDB94, Rob95c, TSSB95, WTWS98, WLAH92, YAM^{+}98]. Do [BN98a, DM92, Jor93, O’C92, Roh91, Tho90]. doctrina [Cur93]. Doctrine [Cur93]. Document [GD99b]. Documentar [GD99b]. Does [Ang94, BE94]. Doesn’t [O’C92, Roh91]. Dog [GAC^{+}95, Mad99, MCR94]. Dogs [CC96, Cre97, CCM97, EG96, EHBE97, MG97, Sta98, VW99a, VC99]. Dogwood [HE97]. Dolphin [PDSB98]. Domain [KB94]. Domestic [Bri95a]. Domestication [BKO^{+}95]. Dominance [BC98]. dominio [KB94]. Donde [LNA95]. Don’t [BCP96]. dorado [PC94]. dorsalis [Tsu98]. Dragonfly [SS98]. Drainages [LAF93]. draweth [Ste93]. Drawing [HS95, War94b]. drosophila [BLF^{+}93, BLF^{+}93, BMR^{+}92]. Drought [BB99, SS94]. Drought-Induced [SS94]. Dry [CG95, GFR^{+}99]. Dudley [BM97]. Due [BGTB93, CD98, MB92a, Was96]. Dugong [MHL97]. Dulce [MLP^{+}97, RR99b, VT99, Ang95, BG95b]. Dulceacuícola [RBMM97]. Duncan [MS99c]. Dune [Bar96, BR96, RF97]. Duración [BB99]. Durango [FC99]. Durante [DH97, GL93a]. duras [Tho93]. Duration [BB99]. During [PBBA90, DH97, Tsu98]. Dutch [BD92, SRV96]. Dwelling [Bun95, FEM95, RS93a, BWH90]. Dying [BBA^{+}96, Nos96c]. Dynamic [LMNV92]. Dynamics [BB97, CA92, GMG94, JFJ96, KA92, QAM96, SÅS98, SY94, SB97, SS98, Tho90, VPW97, Woo99].

E.U.A [GF97]. E.U.A. [BHM94, PD97]. Eacham [ZDM98]. eachamensis [ZDM98]. Early [CPN95, HVM96, Hun98, Lit93, LM94]. Early-Successional [CPN95]. Earth [Che99, Mor95, OD98]. East [ACB^{+}99, CKCM95, CC99a, Cre97, RW99]. Eastern [BRHD97, KL97, MS93, Rob95c, HHH96, HH96, KGS93, KMR94, New91, OBY98]. Eating [GHP97]. Ebenus [HE91]. Echinocereus [NCE96]. Eco [Pat91]. Eco-redevelopment [Pat91]. Ecocentrism [Wil93b]. ecología [CUM96, HH95b, LCSC95]. Ecológica [BTRM99, CM97, CRGGFT97, Hig97, BL95b, FR93, Fim94, JFJ96, RHS94, WMH94]. Ecological [Ang95, ARR90, Bar96, BTRM99, BHL^{+}92, BL95b, Bra94, CM97, Cam95, CRGGFT97, CRHSS96, DSGS93, FR93, Fim94, Fle94a, FT91, GG99, GW97, HB90, Hig97, JFJ96, Jam94, KASR99, KO98, KMM94, Lau91, LN93, Nos90a, Orr92b, OE95, RS93c, RHS94, SDT93, SKOO98, Sti93, TB95, VS95, VMMK92, VC99, Wal92, WMH94]. Ecologically [Dun99, LF97]. Ecológicamente [LF97]. Ecológicas [GG99, VC99, CRHSS96]. Ecológico [GW97, KASR99, KMM94, TB95, VS95]. ecológicos [Ang95, DSGS93, Fle94a, RS93c]. Ecology [Ano97c, BKC^{+}98, BHS95, CUM96, Fro92, HH94a, HE91, HH95b, LSV^{+}96, LCSC95, Mat95, WDR^{+}98, Wil96, WL97]. Ecology-Based [WDR^{+}98]. Econometric [CMB96]. Economía [STA^{+}99, Lan96]. Economic [BM92, Ehr92b, FT91, Jam94, Mar93, MA98, RH96, Tho99, TV99, VS95]. Economically [PH94]. Económico [RH96, VS95]. Económicos [TV99]. Economics [BKC^{+}98, Cau93, HB90, Lan96, Orr91c, RMJ92, STA^{+}99]. Economies [OPRCB95]. Economy [Pow91]. Ecoregions [OD98]. Ecosistema [BC99, Gol99b, TTJ^{+}99, Gru94, LM95, Ris95]. Ecosistemas [BC97b, Gru97, JM97, NR97, RBPB96, Sta95, Yaf99, BSW^{+}96, GAC^{+}95, KS94b, SBL96, Wal95]. Ecosystem [Alp95, Aug95, Bar96, BSW^{+}96, BC99, Blo99a, BC97b, CAWH91, Cla99, CMB96, Die94, Gol99a, Gru94, Gru97, HCH^{+}97, JM97, LB95, LM95, Nae98, Nor93, NR97, Pel95, PBBA90, Pow91, RGH^{+}99, Ris95, Ris99, Rob95a, Sal92, Sal98, SHM91, SGR^{+}98, SBL96, Sta95, SHS^{+}97, SJB97, TTJ^{+}99, Wal95, Wal99a, Yaf99, Bru91, CAWH91, MA91, Pat91, RT91]. Ecosystems [BSW^{+}96, CEPR91, GAC^{+}95, Gol99b, Hol93, Hun96a, Jor93, KS94b, Ogd92, RBPB96, Run95, Sal91]. ecotipos [WPSZ96]. Ecotopia [BJW95]. Ecotourism [Gia93, KHP95, BDR^{+}98]. ecotourismo [KHP95]. Ecotypes [WPSZ96]. Ecuador [Pea97]. Ecuatorial [TWMM97]. Ecuatoriana [Pea97]. edad [DSB95b, FESS94]. edades [PK94]. Edge [ER98, Fra94, Gat94, HH98, KHS98, KDG98, Mil95b, Pat94a, SB95c]. Edge-Related [KDG98]. Edge-to-Interior [Fra94]. Edges [DH98, DSB^{+}95a, JR99, SPR97, WLDGGZ98]. Edit [Tem93c]. Editor [Ano98b, Ano99h, Tem92d]. Editorials [Ano96e]. Educating [Geh96, Orr93b]. Education [Bal96, CPG94, Eri99, Fle90, Jac90, Kio91, Mat97b, MJ96, Orr90a, Orr90d, Orr91b, Orr92a, Orr92b, Orr99b, PAG^{+}99, RSM90, Str95, Tro94, UKG^{+}96, Vid94]. Educational [Orr98]. EEUU [MBW96, AS94, DF96, Fri93]. Efectiva [SM99]. Efectivas [CCR96]. Efectividad [BHS99, CS97]. efectivo [HHH95, NE94, RWES93, RJL95, WP96]. Efecto [GNH94, RF99, Wal99b, BBSK96, DBCM95, Fra95a, JBB96, KHP95, LNVM94, MPB96, Mil95b, MGB94, Pat94a, PERC96, RHCB93, TM96, BRHD97, GB97, GK96, NLS93, OK96]. Efectos [FEM95, HV97, LOT^{+}99, MB99, MH99b, PEH93, BSC95, DOSP94, DLK^{+}95, FS95, FP95, HVM96, PL94, TCK^{+}95, VHM94, ACB^{+}99, BB97, CT99, GSB99, GWW93b, GK99, HWB96, HRV^{+}99, JHHB99, KL97, KC95, KGD96, KSO^{+}99, MKM^{+}96, Mor99, OO93, RKLS99, RLD99, SW99a, SB99, SBL96, Sun96, SMPN99, VTM99, VPW97, WRSP97, WSE97]. Effect [BBHM95, Bil91, BBSK96, CLL91, Fra95a, GK96, GK99, GNH94, HWB96, MKM^{+}96, MKA96, OK96, Pat94a, RF99, RHCB93]. Effective [BBGB92, Cha95, CCR96, HH94b, HHH95, NE94, Orr94a, RWES93, RB95b, RL91, Rym94, RJL95, RJL99, SM99, SM91, VW98, WP96]. Effectively [RO99]. Effectiveness [BHS99, CS97, PS98]. Effectos [LDSH99]. Effects [ACB^{+}99, AKG98, AMA^{+}94, AF98, BSC95, BH97, BB97, BRHD97, BCP95, CK98, CPG94, CT99, DOSP94, DH98, DLK^{+}95, DBCM95, EK98, ER98, FS95, FL97, FMF95, FP95, FEM95, Fur95, GB97, Gat94, GLFZ96, GSB99, GWW93b, HVM96, HH98, Has95, HRV^{+}99, HE97, HSGY98, HG97b, HV97, JFH93, JBB96, JHHB99, KL97, KSSM95, KC95, KHS98, KB98, KGD96, KO98, KHP95, KSO^{+}99, LNVM94, LFM^{+}98, LDSH99, Leb93, LDP95, LOT^{+}99, LP98, MP97, MB99, MH99b, MPB96, Mil95b, Mor99, MGB94, NLS93, OO93, PGFD95, PEH93, PERC96, PL94, PSB98, Rob95a, RKLS99, RR99d, Rod92, RLD99, RL91, SLNP95, SW99a, SGR^{+}98, SB99, SBL96, SPD^{+}98, SG93, SB95b, Sun96, SMPN99, TCK^{+}95, VHM94, VTM99, VPW97, Wal99b, WN98b, WRSP97, WSE97, Yos94]. Efficacy [LM94]. Efficiency [Sto97]. Effort [BL95a]. Efforts [Cra94, Wap94]. Eficiencia [Sto97]. Egg [KB97]. Eggs [BHKH96, KB97]. eight [Gil96]. Eighty [KALG94]. Ejemplo [Win96a, LS96]. Elaeagnus [SAS95]. Elasticidad [MDW99, SFM96a]. Elasticity [MDW99, SFM96a]. Eld [Son96]. electroforética [HP94]. Electrophoresis [SS95]. Electrophoretic [HP94]. elefante [AL93]. Elefantes [HD99, WW99]. elegans [HWB96]. Elephant [AL93, WW99]. Elephants [Cau93, HD99]. Elevados [CT99]. Elevated [CT99]. Elevation [CLM^{+}95, SG93, TLG^{+}98]. Eliot [Sti97]. Elliot [Duf93b]. Embryonic [AKG98]. Embryos [GFP98]. Emergency [DKOW94]. Emphasis [DS95, IN96]. Empirical [Ber90, LC98b]. Empowerment [Kio91]. Emptying [ROT^{+}99]. Emydoidea [CDV93]. enanos [MM92]. Encouraging [BWBM99]. Encubierta [Gel99]. Endangered [AMA^{+}94, BHSO92, Ber95a, Bro96, BBCR99, BJ96, Can96, CCOO^{+}96, Cho95, CBLV94, CBAT97, CKB98, DDM93, DM90a, DS92, DMMG96, FFL95, GMMR^{+}96, GWH^{+}99, GDK99, GWM^{+}97, GWH97, GB95b, GR98, HMPK96, HSS94, HB95, HB96, HP99, JDEB98, KB99, KO91, KMR94, LSA99, LCSC95, LBR99, LM95, ML90, MS92, MFR97, Men90, Mor99, Nab94, NM95, OO93, PSH99, PNH93, PD97, PS96, PK94, QLDV96, RDE96, RA94, RG99, RH97a, Rei95, RS91b, RHCB93, RHSM94, Roy95b, Saf93, STM95, Scu99, SHMV97, SMGB91, STA^{+}99, SPD^{+}98, SDB^{+}96, Str91, SAH^{+}95, Tho94b, WLG^{+}91, WMW93, WC98, WF99b, YJG^{+}96, YBZ99, ZDM98, DME95, GD99a, MWN^{+}94, O’C92, Roh91, TSHG95, Wap98, Win98]. Endangerment [CB95, Fri93]. Endangerments [Ree99]. Endemic [Ama93, BV97, CRGGFT97, Dud99, EE97, GZWR97, Goo95, HE91, HT95, OO93, RKLS99, SSTL92, SAH^{+}95, VB99, Wik90, WS97]. Endémica [CRGGFT97, WS97, SAH^{+}95]. endémico [HT95]. Endémicos [EE97, Ama93, OO93]. Endemism [CB95, Deh96, FCR^{+}98b, Ker97, Lin95]. Endemismo [Ker97, CB95, Deh96, Lin95]. endocria [HP95, Mer94, PSGS94, RVK^{+}94, Fra95a]. Endocrinología [BTRM99]. Endocrinology [BTRM99]. endogamia [MGB94]. Enemies [BLS99]. Enemigos [BLS99]. énfasis [DS95]. enfermedades [Cun96, Hes94, LMS96, AS94]. Enfocada [TTJ^{+}99]. Enfoque [JRP99, HH95b]. Enforcement [AM99, LWMG93]. Enfrentados [MLD97]. engelmannii [NCE96]. Engineered [KEE91]. England [Bra98, Vog95]. Enhancement [HBA93, MWCP94]. Enough [Bal95, CBM^{+}97, SMS97]. Ensamblajes [JHHB99]. Entender [BTRM99]. Entradas [RHCB93]. Entrances [RHCB93]. Entre [Fra96, PQL99, Aco99, BL95a, BTRM99, BL99, BPO99, BT94, Cla95, CD95, DeM93, DSB95b, FS97, FYP96, HC99, HD99, HGNK93, HT95, KQK93, LWMG93, LV94, NMGS94, SGCJ94, ST96]. Entrenamiento [BHCB99]. Environment [HDE95, KK94, MFW90, Web99]. Environment-metapopulation [MFW90]. Environmental [Bis94b, Duf94, Fol94, Gen94, HP95, Kar90b, Kau93b, KB94, MHH98, PSGS94, Pri91, REDH98, Sab98, Sym96, Tho94a, TB95, TW92a, Wei92, Yaf97, You94a]. Environmentalism [Hag95, Pis94, Wil96]. Environments [NGN96, VAPCO98]. Epidemic [AR97, HG97a, LMS96, LMS97]. epidémicas [LMS96]. Epífitos [NM97]. Epigaeic [IS96]. epigeos [IS96]. Epiphytic [ER98, NM97]. Equal [Win95]. Equalization [BLF^{+}93]. Equalizing [All93b]. Equatorial [TWMM97, FJPC95]. Equus [VW96]. Eradicating [Tem90]. Eradication [Cob97a]. Eretmochelys [HC96]. erizos [MKM^{+}96]. Erosion [Ing90, VW99b, VW99b]. Errata [Ano94c, Ano95b]. Erratum [Ano90h, Ano98c]. Error [Ano99i]. Errors [RHK99]. Erwin [Gas91a]. escala [De 95, HH95a, KS94b, RS93c, RHS94, HM97, SD94]. escalas [RS93c]. escamosos [DS95]. Escape [KEE91]. escarabajo [LCSC95]. Escarabajos [RCM97]. Escarpa [KL97]. Escarpment [KL97]. esfuerzos [BL95a]. Eslogans [Gol99b]. Espacial [BBSK96, HM97, HWG99, EDFM96]. Espaciales [CZN99]. Espacialmente [Sou99, LDP95]. espacio [LNVM94, BPO99]. Espada [RFM97]. especial [CM93]. Especie [BBSK96, CRGGFT97, DHW99, GWH^{+}99, Gel99, GG97, HM97, MFR97, Mor99, PNS99, Wei99, WER^{+}97, DMMG96, GHB95, HCB94, Rab95a, RSR96, SAH^{+}95, Was96, WMW93, YJG^{+}96]. Especies [Aco99, AHMM97, BL99, BLS99, BPO99, BBCR99, CO99, DME95, EE97, FH97, GDK99, GD99a, GJH96, GMF99, GS99, HC99, HRCTS99, Ker97, KSO^{+}99, LCPM97, MLD97, PD97, PDS96, RDE96, SW99a, STA^{+}99, SG99, TSHG95, TCO^{+}96, VW96, WRSP97, WK99, Ang95, BGB96, Bro96, Can96, Cun96, DS95, DP96, Fim94, GWW93b, GB95b, Ham94, HGNK93, KB94, LNVM94, LFOJ96, MNU94, OB96, PTTN^{+}96, QAM96, RH93, SVC96, SDB^{+}96, SL93, Tay95, Tho94a, TTW^{+}94, Ver93, WMW93, WGM^{+}96, WH95, WKS^{+}94, BEG^{+}94, Ber97a, Lam97]. especifica [How94, FESS94]. específicos [Fim94]. Espinosa [Aco99]. Esporádica [WER^{+}97]. esta [KMW95, YJG^{+}96]. estabilidad [BWS94, Cla95]. Establecer [CBAT97, DHW99]. Estableciendo [Lin95]. Establecimiento [BMLW96, Loc99, LRD^{+}96]. Establish [CBAT97]. Establishing [LRD^{+}96]. Establishment [Loc99, SAS95]. Estación [PF99]. Estaciónal [GP97]. estacionales [SB94]. Estadístico [Str99, TG93]. estado [CSBC95, MBW96, PT95, RRM93]. Estados [PS96, RDE96, SB95c, Cur93, Fle94a, FBFO96, GDK99, KL96, Lug97, LM95, Tho93]. Están [TSHG95]. estar [LNA95]. Estatus [HM97]. Este [BRHD97, KL97, KGS93]. Esteparios [YS96]. Estimacion [NE94, SM99]. Estimaciones [LCA99, Har95a]. Estimada [Gar97]. estimar [EDFM96]. Estimate [BBGB92, TD96]. Estimated [Gar97, SS95]. Estimates [Gas91a, Har95a, KF95, LCA99, Pet99, TM96]. Estimating [EDFM96, GF90, JJ99, NE94]. Estimation [VW98]. estocasticidad [Fol94]. estocástico [KTA94]. estocásticos [NGN96, RS95a]. Estrategia [CAH97, RS95a, WKS^{+}94]. Estrategias [RRS96, HBA93]. Estructura [LF97, PKH97, RF99, VPW97, YB96, Bak94, MBH96, FB97, PGW97, YBZ99]. estructurado [PK94]. estructurados [RS95a]. Estructuralmente [CFL96]. estructuras [BC94]. Estuarine [MP99]. Estuarinos [MP99]. estuario [AMGJ95]. Estuary [AMGJ95]. Estudiados [McK99]. Estudio [CC99a, EE97, LOT^{+}99, VT99, BG95b, Can96, Deh96, MNU94, OB96, PB95, PP96, TTW^{+}94]. Estudios [BL99, HD96, PL94, RS93c]. Ethic [OPRCB95]. Ethical [Hab96]. Ethics [Bis94b, Cal90b, Col92, FR93, SFM99, Wei92]. Ethiopia [WRSP97]. Ethnobiomedical [BAR94]. Ethnobotany [CC99b, PGR^{+}94]. Etica [SFM99, FR93]. Eticas [Hab96]. Etiopía [WRSP97]. Etnobotánica [CC99b, PGR^{+}94]. Eucalipto [DHE97]. Eucaliptus [PB94b]. Eucalypt [DHE97]. Eucalyptus [PB94b]. Eudocimus [FBFO96]. Euphagus [GD99b]. Europa [BGB96, BH96, De 95]. Europe [BH96, De 95, Duf94, MWW99, MA98]. European [BGB96, HP94, Sta96, TFHD98]. Europeo [HP94]. Eusattus [BR96]. Evaluación [LOT^{+}99, SM97, BSW^{+}96, HBA93, OB93, PC93, PL94, TB95, WMH94, DBH^{+}99, HP99, MWM^{+}99, MRN99, MSC96, SM99, TB95, Vog95, Wal99b, WMB^{+}97]. evaluaciones [RHS94, TM96]. Evaluando [WA95, WMH94, Rei93, SPP96]. Evaluar [CZN99, RBPB96, SM99, RS95a]. Evaluating [And91, DYS95, KR98, LOT^{+}99, MRN99, MR98b, Pat91, RS95a, Rei93, SM99, SPP96, SM97, Tho94b, Wal99b, Woo99, WMH94]. Evaluation [BCL^{+}98, Göt92, HBA93, HP99, Kei98, LC98b, MWM^{+}99, PCG91, TB95, WMH94]. Evaluations [RHS94, Str94]. Events [HSD93]. Ever [DM92]. Everglades [BHSO92, TTJ^{+}99]. Everything [BNF^{+}98]. Evidence [AR97, ART94, AMW90, Hun98, NY91, Pat94a, QLDV96, RG95]. evidencia [Pat94a, ART94, QLDV96]. Evidencias [McK99]. evolución [Cur93, KO93]. Evolution [Cur93, Ebe95, Win96b]. Evolutionarily [BC94, LRD^{+}96, PSH99, PD97, Wap98]. Evolutionary [KB99, KO93, Nae96, QLDV96, RSMB99, SFLC95, Sun96, VB99, ZDM98].

Evolutivamente [LRD^{+}96, PD97, BC94]. Evolutivas [KB99, SFLC95]. evolutivo [QLDV96, Sun96]. Ex [Ham94]. exactamente [WMW93]. Exactly [WMW93]. exámen [Med93]. Examination [Med93, MS93]. Examining [Ris99]. Example [Ing90, PGFD95, SMGB91, Win96a]. Examples [Bra98, Mil95a]. Exclosures [BBS93]. existentes [PT95]. Existing [PT95]. Exito [Loc99, RMCM97, BLD^{+}95, SPR97, Fim94, KGD96, Lan96, Pat94a, DTFP95, FCM99, MR99]. Exotic [Cob90, Jen96, LV94, Lug90, Lug92, ML94, RFG^{+}98, SW99a, SAS95, TW92a, Wes90]. Exóticas [SW99a, ML94]. exóticos [LV94]. Exotics [Tem90]. Expendable [SB98]. Experiences [VW96]. Experiencias [VW96]. Experiment [LKM90, Wik90]. experimentación [FR93]. Experimental [BBCR99, SGR^{+}98]. Experimentales [BBCR99]. Experimentation [FR93]. Experiments [IU92]. Expert [Fro92]. Explain [ST94]. explicar [ST94]. Explicit [Ano99i, LDP95, Mag96, RHK97, RHK99, Sou99]. explícito [LDP95]. Explícitos [Sou99]. Exploitation [JRP99, Kuc92, SMPN99]. Exploited [Aco99, JB99]. Exploiting [Hab96]. Exploring [HSD93]. Explotación [Hab96, JRP99, SMPN99, Ben93, WSB92]. Explotadas [Aco99]. Exposed [BC99, Woo99]. Expuestos [BC99]. Extensión [BPC97, LFOJ96]. Extensively [FS97]. extensivo [FS97]. externa [GL93b]. Extinción [BER97b, BPC97, BPO99, BBCR99, Cow99, CBAT97, HP99, KB99, MH99a, MFR97, MLD97, McK99, New96, Ree99, RR99b, STA^{+}99, VT99, VPW97, VC99, Weh99, WK99, YBZ99, BBM94b, GHB95, Ang95, BGF95, Bur95, Can96, CB95, CCOO^{+}96, DS92, Fol94, Fra95a, Fri93, KALG94, KC95, LCSC95, MNU94, OO93, PNH93, QLDV96, Rei95, RHCB93, RHSM94, RC94, ST94, SDB^{+}96, SAH^{+}95, Tsc92, TTW^{+}94, WMW93, FC93, GDK99, New95, TSHG95, Tho94a]. Extinciones [GL93a, BWS94, SY94, SFLC95, SS99b, FS97]. Extinct [Rab95a, TSP97, Ver93]. Extinction [Ang95, BER97b, Bot94, BBM94b, BPO99, Bur95, CMW^{+}98, Cow99, CD98, Cut91, Fol94, Fra95a, Fra98, Fri93, GFA90, HS96a, HM98, Kar90a, Kei98, KC95, Kuc92, Lau91, LFOJ96, ML91, MH99a, MG90, MNU94, MLD97, McK99, Mil96, New91, New95, New96, Ree96, RR99b, RC94, Sar95, SY94, ST94, SP98, SJB97, Tho94a, TJ99, Tsc92, Tsu98, TTW^{+}94, VT99, VPW97, VC99, Weh99, WK99]. Extinction-Prone [Ang95]. Extinctions [Ber90, BWS94, Dia95, FS97, GL93a, KALG94, MB92a, Ree99, RBMA98, SFLC95, SS99b]. extinguidas [Ver93]. extinguirse [Rab95a]. Extirpación [LSA99]. Extirpation [FC93, LSA99]. extracción [LLP95]. Extraction [Cob97b, LLP95, SDT93, VJSU98]. extractivas [SDT93]. Extractive [Pel92, SDT93, Sie93]. Extreme [GZWR97]. Extrinsic [LWC95, Mil96]. Exxon [Gar97].

F. [SNP94]. Fabaceae [HVG91]. Faced [GR98, MLD97]. Faces [Yaf99]. facilitar [KGS93]. Facilitate [KGS93, ST96]. Fact [DSM90]. Factor [BD92, KW96, Yu96]. Factores [FM96, LOT^{+}99, MHB^{+}96, KA92]. Factors [FM96, KA92, LWC95, LOT^{+}99, Mil96, MHB^{+}96, MG97, SÅS98]. Faculty [MJ96]. Failed [Bla94]. Failure [Nos97]. Fall [BCP96]. Fallacy [DSM90]. False [Sie96]. familiar [BLF^{+}93]. Family [All93b, BLF^{+}93]. Farming [WCTM98]. Farmland [FGG^{+}95, LV91]. fasciculatus [SAH^{+}95]. Fate [DSM90, MS99c]. Fates [MS97b]. Fauna [DF96, Dud99, Goo95, HSGY98, Lip98, MW90, PRH98, RS91b, Rei93, RR99b, RBMM97, SWK^{+}97, ART94, DPOO94, KMW95, NMGS94]. Faunal [DPOO94, KO93]. Faunas [Bun95, Cut91, LM93, LM93]. faunísticos [KO93]. Favorable [MSHW95]. Faxon [LEMC95]. Features [MC95b, TTJ^{+}99]. fecundación [Sun96]. Fecundidad [PF99]. Fecundity [PF99, SW99b]. Federal [BBS93, Gru90]. Federoff [WRG98]. Feed [PAPA90]. Feeding [PP96]. Feminists [Bis93]. Feminizing [GFP98]. Fencelines [Yos94]. Fencerows [DSB^{+}95a]. fendleri [RMCM97]. fenotípicas [BC94]. Fernandez [CSC^{+}94]. fernandeziana [CSC^{+}94]. Ferret [GW98, RTOB94]. Ferrets [RK93]. fertilidad [Was96]. Fertility [Was96]. ferus [VW96]. Few [Ano99p]. fiber [Har94]. Ficus [AHMM97]. Fidelity [OBY98]. Field [AP99, Bie95, CSC^{+}94, DS92, FR93, Jac93]. Fieldwork [McN95]. Fig [AHMM97]. Fighting [HYL95, Roy95a]. Figs [AHMM97, MCW98]. Filipinas [HC99]. filogenéticos [HD96]. Filogenia [JRP99, Lin95]. Filogeografía [LBR99]. Filográfico [MSO96]. Filters [Nor96]. Finally [Tem93b]. Financial [ABM91]. Finches [GG97]. Finding [Pri96a]. Fine [SÅS98]. Fine-Grained [SÅS98]. fines [Rob95c]. Fingerprint [YAM^{+}98]. Fingerprinting [Gra95, RFDB94]. Finite [SW94]. Finland [RTNK99, VRV^{+}94]. Finlandia [RTNK99, VRV^{+}94]. Fiordland [MJ97, MJ97]. Fir [RFG^{+}98]. Fire [Bak94, BBSK96, Bun95, CRHSS96, DD97, FC99, HS98a, NMV^{+}98, TCK^{+}95]. Fire-Treated [CRHSS96]. Fires [CS98, Hut95c]. firme [CRHSS96, Pat94a]. First [Bis92, BK99, Cua97]. Fish [Aus98, BBGB92, BE94, Cob97b, Cob99, HRV^{+}99, HGNK93, JHHB99, LSA99, LM93, LNPC^{+}95, MKM^{+}96, MW90, OO93, PRH98, QLDV96, Rei93, Rob95b, ROT^{+}99, TW92a, WH95]. Fished [GMSP98]. Fisher [CZN99]. Fisheries [Cam95, HC96, MPC99, SKOO98]. Fishers [Bus92]. Fishery [Bla94, MHL97, MKA96]. Fishes [Ang95, Bir93, BTG93, CCOO^{+}96, CBLV94, Fri93, GW97, JJ99, JRP99, KO93, Moy94, RS91b, SB95a, Wik90]. Fishing [Aus98, Ber95b, Kai98, Rob95a, WN98a, WN98b]. Fitness [BR99, Cla95, LSM99, SHMV97, YAM^{+}98]. fitoplanctívoras [GWW93b]. Five [GAC^{+}95, Sou91]. Flamenco [GB97]. Flamingos [GB97, Sim96]. Flatwoods [MPB96]. flava [GWH97]. Fledging [DH97]. Fleshy [HC99]. Flexibility [Tho94c]. Flies [BMR^{+}92, Tsc92]. floccosa [DS92]. Flooding [MN96]. Floodplain [AMGJ95, BHM94]. Floor [BHM94]. Flora [Sta94, TW92b]. Floral [RS92, SNP94]. Floras [Dud99]. Florida [BSC95, BLD^{+}95, Fra95b, Hed95, JM97, JBB96, KLF98, Mae90, MC95a, MC95b, MHO97, MM92, MPB96, Ogd92, PBKM95, QAM96, RS95b, SB94, TTJ^{+}99]. Floristic [GG90b, RS92]. Floristics [WLDGGZ98]. Flow [Hed95]. Fluctuantes [VW99b]. Fluctuating [Tsu98, VW99b]. Fly [Tsu98, WRSP97]. Flying [CEPR91, FT91, PERC96]. Foca [KGM^{+}97]. Focal [Lam97]. Focales [Lam97]. Focas [KLR97, SRR95]. Focused [SBB^{+}94]. Followed [BBHM95]. Following [Hut95c, Fra97, Mar98]. Food [Ano98d]. Footed [EKL^{+}99, GW98, RK93, RTOB94]. Forage [BD92]. Forager [Ano97c, WL97]. Forager-Resource [Ano97c, WL97]. Foraging [CB98]. Forb [AF98]. Foreign [GL93b]. Forest [AR97, Ama93, AMGJ95, ART94, ARR90, BSC95, Baw92, BS98, BM98, BHM94, BPO99, BS92, Bun95, CHK98, CC99a, CFL96, CS98, Del99, DH97, DM90b, DY93, Duf93c, FESS94, FS95, FL97, FMF95, FEM95, GFR^{+}99, GF90, God93a, God93b, Goe97, Gre92, GS93, Gur92, HVM96, HB90, HCB94, HWB96, HH96, HH98, Has95, HHR99, HE97, Hil95, HEO95, JHHB99, JR99, KALG94, KHS98, KO91, Kin92, KB94, LMNV92, Lau91, LMS96, LFM^{+}98, LLP95, LT92, LF97, LDP95, MVH96, MP97, MW93, Med93, MCO98, MA98, Mil95b, MWCP94, NMV^{+}98, New91, Nie97, OBY98, PVZJCI90, PBG^{+}93, RS93a, RJBB96, Ren99, RDN94, RHSM94, RH93, RR99d, Rob99, RLD99, SGCJ94, SDT93, Sal91, Sto97, SB95b, Thi92, Thi95, Tho93, TCO^{+}96, TWMM97, VTM99, VAPCO98]. Forest [Vog95, WW99, WN98a, WRM^{+}91, Wik90, WCTM98, WER^{+}97, Yah95, Yah96, BAR94]. Forest-Bird [MCO98]. Forest-Dividing [RDN94, Yah95]. Forest-Dwelling [Bun95, FEM95, RS93a]. Forest-Floor [BHM94]. Forest-Interior [Tho93]. Forest-Nesting [BSC95, RDN94]. Forestal [MR99, VTM99, Ben93, KALG94, WSB92]. forestales [LLP95]. forestas [DY93, HCB94, Tho93]. Forested [GHP97, KGD96, MR99, YM96]. Forestry [BH97, Har95b, KK94, Put93, Ric98]. Forests [Ano90g, Ano91m, ADM94, AF98, BM92, BS98, BB97, BEG^{+}94, BEG^{+}95, Bra94, BS92, Bus92, CG95, CC99b, CPN95, DHE97, ER98, Fra94, HH95a, HL99, HY93, HV97, Hut95c, LF97, Lug94, Lug97, Mar92, Mat94, MBBF95, NM97, NABE95, Nos98, Orr93c, Per94, RFG^{+}98, RTS90, SLNP95, SB97, SBL96, SB94, Tho93, VI95, Wel90, Wie99, You94a]. Forewords [Ano96e]. Forms [Sam90, Wal99b]. Forthcoming [Ano96c]. fortis [LEMC95]. Forty [Gil96]. Forty-eight [Gil96]. Forward [Pre94]. Fostering [PC93]. Foundation [CMB96]. Founder [BBHM95, BGK^{+}97]. Founding [MBT90]. Four [GJH96, MS99b]. Fox [PERC96, YJG^{+}96]. Foxes [CEPR91, FT91]. Fraccionados [WK99]. Fractal [WK99]. Fragment [KHS98, LG97, TCO^{+}96]. Fragmentación [HWB96, KSO^{+}99, RJBB96, Ren99, RLD99, VTM99, WSE97, DLK^{+}95, HVM96, IS96, MH93, NHA95, RDN94, Vog95, WLAH92, PB94b, KALG94, Tsc92]. Fragmentadas [HCBH99, YBZ99, FM94]. Fragmentado [FCM99, LCPM97, SS99b, LMNV92, MW93]. Fragmentados [Mor99, PF99, Bur95, WDSA94]. Fragmentation [BH97, BM98, Bra94, BS92, DLK^{+}95, FL97, HVM96, HWB96, Has95, IS96, KALG94, KSO^{+}99, LFM^{+}98, Leb91, LN90, MP97, MH93, MB92b, New91, NHA95, PB94b, RJBB96, Ren99, RDN94, RLD99, SHM91, SB95b, Tsc92, VTM99, Vog95, WLAH92, WSE97, Yah96]. Fragmented [AF98, BWH90, Bur95, Cut91, DHE97, ER98, FM94, FCM99, HCBH99, KR95, LMNV92, LCPM97, MCW98, MW93, Men91, Mor99, PF99, RR99d, SS99b, WLDGGZ98, WDSA94, YM97, YBZ99]. Fragmento [TCO^{+}96]. Fragmentos [BPO99, JR99, TWMM97]. Fragments [BPO99, GF90, JR99, PBG^{+}93, TWMM97]. Frame [VB99]. Framework [And91, HW95, Jen95, LOT^{+}99, MH99b]. France [Deh96]. Francia [Deh96]. franjas [MVH96]. frecuencia [LRT93]. Frecuencias [LCA99, RL96]. Free [CDH95, Jen96, Ros92, Yu94, Yu96]. Frequencies [LCA99, RL96]. Frequency [LRT93, LC98b, WT90]. Freshwater [Ang95, BG95b, HSGY98, MLP^{+}97, RS91b, RR99b, RBMM97, VT99]. Fresno [LF97]. Fringe [Sym96]. Frissell [Sal92]. Fritillary [BBM94b, BBM94c, SO94]. Frog [BHKH96, BTG93, DF96, HM97, KSO^{+}99, LDSH99, PC94, VI95]. Frogs [AR97, BE94, KB98, LMS96, MP97, MGM^{+}98]. Front [Uhl98a]. frugívoras [CUM96]. Frugivores [HC99]. Frugivorous [CUM96]. Fruit [HC99, Tsu98]. Frutas [HC99]. Frutívoros [HC99]. fuego [Bak94, CRHSS96]. Fuelwood [AM99]. fuente [Doa95, DTFP95]. fuentes [BT94, DLK^{+}95]. Fuera [FT96]. fuerte [Was96]. Fuertemente [AHMM97]. Full [Bis93]. Full-Contact [Bis93]. Fully [JFH93]. Función [KS97, HMPK96, SBL96]. Funcional [Gol99b]. funcionamiento [Ris95]. funciones [SL93]. Function [DM90b, HCH^{+}97, HMPK96, KS97, Ris95, Ris99, SBL96]. Functional [Gol99b]. Functions [SL93]. Fund [Tem93a]. Fundamentalism [Orr94c]. Fundamentals [Fle96]. Funding [Pri92]. Fundy [SB99]. Fungus [BEG^{+}94]. Futura [War94a]. Future [Alt90, Ano99c, BWS97, Bes96, Boz93, Hac99, OB97, Orr94a, PT95, RBMA98, War94a, WMB^{+}97]. Futuro [BWS97, Hac99, Boz93, PT95]. Futuros [WMB^{+}97]. Fuzzy [MBW^{+}98]. Fynbos [MBD^{+}99].

G. [RH97a]. Gaddi [Sab96]. Galapagos [Cam95, Mac91, Mau97, WCD97]. galeritus [KO91, KO91]. Gales [PL95]. Gallery [WGO95]. Gallina [PS96]. Game [CLM^{+}95, MS99a, PGFD95, SS98]. ganado [BB93, Fle94a]. Gangas [SFCM92]. gansos [RFDB94]. Gap [CSBC95, DH97, FF95, SB95c, KSC^{+}96]. Gaps [PQL99]. Gardener [WRG98]. Gardens [Ste98]. Garganta [BHM94]. Gargantuan [Rol95]. Gates [Pat94b]. Gazella [RH97a]. Gazelle [RH97a]. Gear [Aus98, WN98a, WN98b]. Geese [RFDB94]. Gene [Hed95, KEE91, LRT93, Ing90]. Generación [MA96]. Generados [WGRT96]. General [CPC96a, Doa95, Gru96, KTA94]. generales [Doa95, SSEM94]. generalidad [FH95]. Generality [FH95, GB96b]. generalización [PL95]. Generalization [PL95]. Generation [MA96, Tro94]. Genero [QLDV96, LBR99, Lin95]. Genes [Rol95, LRT93]. Genetic [Alt94, Bil91, BBCS99, BCL^{+}98, Bre94, BMR^{+}92, BGK^{+}97, BRHD97, Cha95, CRJ^{+}94, Cro92, CBAT97, DDM93, DM90a, DS92, DDM^{+}92, EE97, EKL^{+}99, FBS95, FFL95, FB96, FB97, Fra96, GMMR^{+}96, GWH^{+}99, GHB95, GJH96, GWH97, Gra99, GS99, Gra95, HBD90, HCB94, HWB96, Ham94, HP94, HGK^{+}96, Hed95, HD96, HVG91, HP99, Ing90, KASR99, KR98, KF95, KMR94, KGM^{+}97, LSM99, Leb91, Leb93, LRD^{+}96, MC95a, MWS^{+}97, MBH96, MSO96, MAR96, PWC^{+}98, PEP98, PGW97, PSB98, RVK^{+}94, RF97, RFDB94, RFM97, RTOB94, SCZF96, SMGB91, SLS92, SS95, Sun96, SRV96, SF97, SAH^{+}95, TBL^{+}97, Wap94, Was96, WLAH92, Way92, Wil93a, YB96, YBZ99]. Genetica [SCZF96, BBCS99, BRHD97, EE97, EKL^{+}99, FB97, Fra96, GMMR^{+}96, GWH^{+}99, GJH96, GWH97, GS99, HWB96, HP99, KASR99, KGM^{+}97, LSM99, MSO96, MLP^{+}97, PGW97, PB94b, SF97, TBL^{+}97, YB96, YBZ99, BLF^{+}93, CRJ^{+}94, DS92, FBS95, FB96, GHB95, HCB94, HGK^{+}96, HD96, MBH96, PDC96, RVK^{+}94, Sun96, RGSW96]. Genetically [RL91]. Genéticas [MAR96, Was96, WLAH92]. Genético [BGK^{+}97, CBAT97, RFM97, HP94, RFDB94, SAH^{+}95]. genéticos [DDM93, Ham94, LRD^{+}96]. Genetics [AMW90, Avi95, BLF^{+}93, Ela96, LAF93, Mil96, MAS^{+}97, MLP^{+}97, OG96, Pae99, PDC96, PB94b, RGSW96, SW94, TD96, WT90, WLG^{+}91]. Genic [MBT90]. Gentiana [OLKD98, RVK^{+}94]. Gentry [Hur94]. Genus [Lin95, QLDV96, KB99]. Geocapromys [CLL91]. Geochemical [PCM98]. Geográfica [AFF^{+}97, PL94]. geográfico [MH93, SDC^{+}93]. Geográficos [Vog95]. Geographic [GB96a, Gol97a, MH93, PDSB98, SDC^{+}93, Vog95]. Geographical [AFF^{+}97, PL94, SFM96b]. Geography [Kni99, MA98]. Geometrical [KA92]. geométricos [KA92]. Geomorphological [BABK98, NKA98]. Geopolítica [KB99]. Geopolitical [KB99]. Germinability [Mor99]. Germinación [Mor99]. Germination [Men91, SAS95]. Getting [Tem92a]. Giant [LOT^{+}99]. Gigantes [LOT^{+}99]. Gila [PSH99, DDM93, SHMV97]. Gilia [HP95, PH96]. Gill [RKBP93, SKOO98]. Gillnet [MPC99]. Ginkgos [DLY92]. Ginseng [NGN96]. Gippoliti [SDB^{+}97]. Gir [SGCJ94]. GIS [LDP95, Sil94, SHM97]. glabra [DeM93]. Glacier [TW92b]. Global [BWS94, CB95, Dav96, Dud95, Gas91b, Gin99, HR93, Hei93, KC95, KS97, MB92a, MQ99, OD98, PC98, RT91, SVC96, WR96]. globales [BWS94, CB95]. Globally [Gre96, HH94c]. globalmente [Gre96]. Gnatcatcher [AA97]. Go [God93b, Rab95a]. Goal [Kar90b]. Goals [Ano90i, Ano90j, Ano90k, Ano90l, Ano91i, Pat91]. Goat [Cob97a, Hut95a, Lym94, Sch93a]. Goats [Mac91]. Gobiernos [PDS96]. Goby [LSA99]. Goes [Hun93]. Gold [Pri96a]. Golden [ES94, PC94]. Golden-cheeked [ES94]. Goldstein [BO99, OB97, Ris99, Wal99a]. Gone [Caf95]. Good [GC97, Hig97]. Gophers [SCZF96]. Gorge [BHM94]. gorilas [Har95a]. Gorilla [Har95a]. gorrión [DBCM95]. gorriones [LDP95]. Got [Ewe91]. Götmark [And92]. Government [PDS96, Sto94]. gracilis [CRGGFT97]. Gradient [KASR99, VT99]. Gradiente [KASR99, VT99]. gradientes [Fra94]. Gradients [Fra94]. Grado [DBH^{+}99, KMW95]. Graduate [Eri99, GP98, Jac90, JVM95]. Graham [War94c, War96]. grahamensis [War94c]. Grained [SÅS98]. gran [CLM^{+}95, De 95, RS93c, FS96, FGG^{+}95, GL93a]. Grand [SGR^{+}98]. Grande [FESS94, AL93]. Grandes [RW99, CC96, De 95, FBFO96, New96, You94b, BH93, LM93, MSHW95, RRM93]. grandiflora [PNH93]. granjas [DMMG96]. Grasses [BB93]. Grasshoppers [Sam90]. Grassland [JS93, Mor99, VHM94, War94a, WF99a]. Grassland-Nesting [WF99a]. Grasslands [FS97, Sta98, TW92b, Wei99]. Grassy [PB94b, PSB98]. Gray [GDK99, HMW98, MSHW95, VPW97, WLG^{+}91, WLAH92, WSE97]. Gray-Tailed [WSE97]. Grazing [BB97, BB93, BB99, BD92, BM95, BM97, CFH^{+}95, DRP98, Dud97, Fle94a, FFH96, Hei95b, JB95, MR98a, RF99, Sab96]. Great [BH93, DRP98, Kau93a, KC93, LM93, MS93, MSHW95, AB90, BW91, FS96, GL93a, RRM93, Sta98, SJB97, UCB92]. Greater [BC99, BCL^{+}98, DSM90, DM90a, Nor93, Ber91, Bru91, CAWH91, Job91, MA91, Pat91, Pow91, RT91]. Green [BWBM99, HM97, MBW96, RCM97, Yu94, TBW^{+}98]. Greene [SAH^{+}95]. Greenway [HH94a]. grillo [KA92]. Gris [WSE97, Doa95, MSHW95, WLAH92]. Grises [GDK99, VPW97]. Grizzly [Doa95, MW98, MWM^{+}99, MS92, MR91, MHWP96, MHB^{+}96, Pri96b]. Groenendael [PH96]. Ground [KHS98, NLS93, Pov94, Put97, RS93b, RCM97, YM96]. Ground-Beetle [NLS93]. Ground-Nesting [KHS98]. Groupcut [KDG98]. Grouse [JB99, SÅS98]. Growth [BS92, BRHD97, CC99b, FBS95, Gre92, Jam94, JR99, KB98, Leb93, Lug97, MWCP94, NPM98, Sal99, San99, SLNP95, SB97, Wel90, Zei92, ZG99]. Grulla [Can96, CRHSS96, GSB99, HOF^{+}96]. grupos [NLS93]. Grus [CRHSS96]. guadalupensis [Hel98]. Guam [HBD90]. Guatemala [GBAR97, SDT93]. Guesswork [MSM99]. Guianan [HS98a, Thi92]. Guidelines [LC98a, RRS96, SS99a, SSEM94]. Guides [PCG91]. Guiding [DSM90]. Guinea [FJPC95, MHL97, Sim94]. Gujarat [Jha93, JG91]. Gulf [BHL^{+}92, NPM98, PCM98, RKBP93]. gunii [Rob95c]. gunnii [Rob95c]. Gunung [Sal93]. Gusanos [SB99]. Gymnobelideus [LP96].

H [Hur94]. habitadas [CDH95]. habitados [FEM95]. habitan [Bun95, FGG^{+}95]. Habitantes [RS93a]. Habitat [AA97, ART94, Bar95, Bei93, BN98a, BEG^{+}94, BN97, BLD^{+}95, BFS98, CZN99, CHK98, CD95, CB98, Doa95, DRP98, DLK^{+}95, DTFP95, EDFM96, FH95, FS91, GP97, Gre96, Hun98, KGD96, KTA94, LOT^{+}99, LV91, MWM^{+}99, MP97, MHH98, MM92, ML94, MB92b, Med93, Mil94, MSHW95, OR94, OLKD98, REDH98, RCHP94, RDH97, Sam90, SAM98, SGK95, TFHD98, Tsc92, VHM94, War94a, WPSZ96, WLAH92, WSE97, YJG^{+}96, AA97, BN97, CZN99, GP97, LOT^{+}99, MWM^{+}99, RDH97, War94a, WPSZ96, WSE97, BY96, Bei93, BLD^{+}95, Doa95, DLK^{+}95, FH95, Gre96, KTA94, MM92, ML94, Med93, MSHW95, OR94, Tsc92, VHM94, WLAH92, YJG^{+}96]. Habitat-Based [AA97]. Habitats [Aco99, Aus98, Bur95, CRHSS96, Cut91, Fim94, KR95, MKH98, MP99, PBKM95, Tsc92, Aco99, ART94, BEG^{+}94, CD95, DTFP95, Fim94, RCHP94, MP99]. Hacia [MH99a, Ash97, Fra94, HH95a, RK93, SK93]. Haianan [Son96]. Halfway [Fra92, Mef92]. Handling [Cre97, GAC^{+}95]. Happening [Wym90]. Harbor [RKBP93, SRV96]. Hardwood [Fra94, Tho93]. Harlequin [PC94]. Harperocallis [GWH97]. Harvest [HC96, OK96, SB99, Str98]. Harvested [NGN96]. Harvesting [AB94, BRHD97, GMG94, NLS93, PKH97, PEH93, Pet94, VAPCO98]. Hastings [JDEB98]. Hatcheries [Mef92]. Hatchery [BBGB92, DMMG96]. Hatchling [JBB96, MB99]. Haul [Orr93b]. Hauling [GG90a]. Having [Tem92c]. Hawái [Ama93]. Hawaiana [KGM^{+}97]. Hawaianas [SRR95]. hawaiano [SFLC95]. hawaianos [RFDB94]. Hawaii [Ama93, CLM^{+}95]. Hawaiian [CLM^{+}95, GG90a, KGM^{+}97, RFDB94, SFLC95, SRR95]. Hawk [Wal90]. Hawksbill [BBL93, HC96]. Hazel [SÅS98]. Headed [MG90, TB93, TB93]. Headstart [BCHM94]. Health [CNP^{+}97, Kob99]. Healthy [Jor93, MM96]. Hector [PDSB98]. Helonias [GHB95]. Help [Mac91]. Helping [Rab95a]. Helps [Par95]. Helsinki [RTNK99]. Hemisphere [Sib94]. Hemispheric [SB99]. Hemispherical [WRM^{+}91]. Hemlock [MS93]. Henry [DMD^{+}92]. Herb [PSB98]. herbáceos [PB94b]. Herbaceous [DM92, JFH93]. Herbivores [Ber97a, BV97, De 95, RW99]. Herbivoros [Ber97a, De 95]. Herbs [BHM94, Duf93b]. Here [God93b]. hereditaria [LRT93]. Hereditary [LRT93]. Heritage [Cor96]. Hermana [Gel99]. Herpetofauna [VS95]. Herramienta [BTRM99, DHW99, FDV94]. Herramientas [MPC99]. Herschelia [Lin95]. Herviboros [RW99]. Heschel [OG96]. Heterocigosidad [VW99b, HP94]. Heterogeneidad [HNC97]. Heterogeneity [AB90, BABK98, HNC97, NKA98]. heterostilica [Was96]. Heterozygosity [HP94, VW99b]. Hibridación [LFOJ96]. híbridas [HGK^{+}96]. Híbrido [HNC97]. Hidrológicas [RBPB96]. Hierarchical [MBH96, Nos90b]. hierbas [BHM94]. High [Aug95, ABM91, CLM^{+}95, DM90a, GNH94, HCB94, HSGY98, McK99, Pom93, SG93, TLG^{+}98]. High-Elevation [CLM^{+}95, SG93, TLG^{+}98]. High-Intensity [GNH94]. Higher [KB94, Mat97b]. Highlighting [Mat99]. Highlights [Tem92b]. Highly [FCM99, TSSB95, WS97]. Hills [SB97]. Himachal [Sab96]. Himalayas [YH91]. hipóstesis [ST94]. hipótesis [ST94, Car93]. Hipotético [CFL96]. Historia [HCBH99, RL96, BPS95, CMMLRR95, FBS95, GF97]. Historian [Fra95b]. Historias [JRP99, WP96]. Historic [Lit93]. Histórica [Cow99]. Historical [AS94, BM97, Cow99, JDEB98, PRH98, TSP97]. históricas [AS94]. Histories [BHL^{+}92, JRP99, WP96]. History [BPS95, CMMLRR95, Cla92, FBS95, GF97, Hea94, Man94, Mat94, Mei99, RL96, Riv97, RO99, Tho94c, Web99]. Hluhluwe [Mad99, Mad99]. Hluhluwe-Umfolozi [Mad99, Mad99]. Hobby [Riv97]. Hoc [Pre94]. hojas [OK96]. Holdgate [Rob93a]. Holistic [Fle95]. Holocene [GL93a]. Holoceno [GL93a]. Home [NM95]. Homeless [Cho95, NM95]. homosigocidad [Mer94]. Homozygosity [Mer94]. Honeycreeper [SFLC95]. Honeyguide [DSM90]. Honeysuckle [HV97]. hongkongensis [Sun96]. hongos [BEG^{+}94, HD96]. Hope [Put93]. hormigas [MB96]. Hormones [CCM97, EHBE97]. Horn [BC98, FBS95, RB97a]. horned [DM90a]. Hornero [PF99]. horneros [KGD96]. Hornless [BCGL93]. Horse [Mil95a, Ryd93, VW96]. Horses [Sym96]. Hospederos [MS97a]. Host [Ano95c, MS97a]. Hotspots [DW93, MQ99, MMT^{+}98, NM97, WGM^{+}96, DW93]. House [KW96]. Howling [Sto96]. Huapi [VMMK92]. Hudsonia [GLFM98]. hudsonicus [War94c]. Huevos [KB97, BHKH96]. Human [BT94, CDD96, CNP^{+}97, FM96, FYP96, Geh96, GG99, GS93, Hei96a, Hou95, Hun96a, JD98, KBRB96, KQK93, KC95, Kin92, LOT^{+}99, MH99b, MEE93, Mef94, NMGS94, OPRCB95, Orr94b, Pat91, RS95b, SGCJ94, SK96, San99, TA99, TJ99]. Human-Altered [BT94]. Humana [CDD96, KBRB96, NMGS94, KQK93, KC95]. Humanas [GG99, BT94, FYP96, RS95b, SK96]. Humanism [Sta95]. Humanismo [Sta95]. Humanos [FM96, HD99, LOT^{+}99, MH99b, TA99, SGCJ94]. Humans [Blu93]. húmeda [FS95]. Humedales [FH97, GLFZ96, KSO^{+}99, BG95b, GHB95]. Humid [Uhl98b]. humile [Hol95]. Hummingbirds [RCHP94, SB95b]. Humpback [GD99a]. Hundred [Gat95]. Hunted [PZSG96]. Hunting [ARRK97, BER97b, FJPC95, FMF95, HPB^{+}97, KLF98, KLR97, LA97, MN99, Nos98, TA99, WSB92, WCTM98]. hurón [RK93]. Hurricane [Ogd92]. Hutias [CLL91]. Hybrid [HNC97, HGK^{+}96, Now92]. Hybridization [DC92, GR98, HB94, LFOJ96, RH97a, RG95, VW99a]. Hydrologic [RBPB96]. Hyla [AKG98, AKG98, BE94, MGM^{+}98]. Hyman [Jen95]. Hymenoxys [DeM93]. hyperantus [ST96]. Hypothesis [Car93, CEPR91, LMS97, ST94]. Hypothetical [CFL96]. Hypoxia [CKCM95].

Iberian [ALGPH95, YS96]. Ibericas [YS96]. Ibises [FBFO96]. Ictalurus [TW92a]. Icterus [LBR99]. Idaho [TCK^{+}95, CSBC95, TCK^{+}95, WPSZ96, WMH94]. Identification [BDSD98]. Identified [TFHD98]. Identify [Pae99, Par95]. Identifying [LC98b, PEP98]. Idiot [JD98]. If [Ano99j, Joh94a]. Ignorance [McC95]. igualación [BLF^{+}93]. II [HB96, NKA98, Orr97]. I’iwi [FSCL96, Win96b]. Illegal [HYL95]. Illinois [War94a]. Illustrated [DSB^{+}95a]. Illustration [MB96]. ilustración [MB96]. imbricata [HC96]. Immune [SCZF96]. Impact [Ama93, BWS97, BM98, CBCC93, DC92, FJPC95, GSAZ98, GS93, GW97, HPB^{+}97, MN99, Ogd92, Pet94, PP93, RCHP94, Tem92c, TB93, VS95, VAPCO98]. Impacto [BWS97, CBCC93, TB93, Ama93, GW97, MN99, RCHP94, VS95]. Impactos [GLFZ96, LDP95]. Impacts [AB94, Aus98, KRL98, Nos98, Pat91, PBKM95, SW99b, VMMK92]. Impala [GR98]. Imperative [Blo99a]. Imperfect [Ral92]. Imperiled [RBMM97].

Implementación [GDK99]. Implementation [FWB^{+}99, KM98]. Implementing [GDK99]. implicación [DMMG96]. Implicaciones [Aco99, GMG94, Mer94, OFH94, PSGS94, PGW97, SB94, Ham94, How94, Sto96, Ver93, You94b, GSB97, Hab96, HD96, MS97a, MB99, RB97a, SF97]. Implications [Aco99, Ano97c, BBHM95, BDSD98, BWH90, BMR^{+}92, BR96, Bun95, Bur95, CLM^{+}95, CDV93, DSB^{+}95a, DMMG96, Dud99, GB96a, GMG94, GSB97, Hab96, HB90, Ham94, HD96, HH92, How94, KLF98, MS97a, MB99, Mer94, OFH94, PB95, PSGS94, PGW97, RB97a, Rob95c, RT91, Sch98, ST95, Sto96, SB94, SF97, Tri98, Ver93, WTWS98, WL97, You94b]. Imponderable [Gas91a]. Importance [BHSO92, BC98, DGG^{+}99, DYS95, Ela96, FT91, GB96b, GF97, G96, Got95, HOF^{+}96, Hun98, IU92, Lin95, Loc99, TM96, TB96]. Importancia [HOF^{+}96, Loc99, VTM99, GF97, Lin95, VRV^{+}94, DGG^{+}99]. Important [PH94, SB95a, Sta94]. Importante [STA^{+}99]. Impoundments [VT99]. Improve [OSEC95]. inadecuado [GL93a]. Inadequate [GL93a]. Inappropriate [Cle93]. Inbreeding [BBHM95, EKL^{+}99, Fra95a, Fra98, FNC98, HP95, KHM99, LR91, Mer94, MGB94, PH96, PSGS94, Rym94, RJL99]. Incarnating [Alp95]. incendio [TCK^{+}95]. Incendios [FC99, BBSK96, Bun95, GNH94]. Incentives [Hei96a, TV99]. Incentivos [TV99]. Incertidumbre [SFM99]. Incidence [HMPK96, QAM96]. incidencia [HMPK96]. Incidental [YS96]. Include [Ang94]. Incomplete [SGA98]. Inconsistency [Dal99]. Incorporating [MHH98]. Increase [Ree96]. Increases [Men91, RMCM97]. Incrementa [RMCM97]. Independence [Pat94b]. Independent [MBM^{+}98]. Index [Ano99b, Ano99o, LNPC^{+}95, MB96]. India [DHJG91, JG91, Jha93, SGCJ94, Sab96]. Indian [Mar92]. Indiana [BS92, RHCB93]. indicadoras [BSW^{+}96]. Indicadores [RCM97]. Indicator [Ebe95, PC92, Wea95]. Indicators [NABE95, Nos90b, REDH98, RCM97]. Indice [MB96, LNPC^{+}95]. indígenas [DS95]. Indigenous [Alc93, Ano97c, DS95, MHL97, McL93, Per94, SGW90, WL97]. Individual [VAPCO98]. Indo [DW93, DW93]. Indo-Pacific [DW93]. Indo-pacífica [DW93]. Indonesia [LLP95, Mar98, SDT93, Sal93, Yea97]. Induced [SS94]. Industrial [Bal99, Cam95, CHK98]. Inestabilidad [HCBH99]. Infecciones [KB97]. Infectadas [BC99]. Infection [CD98, KB97]. Inference [NBH^{+}98]. Inferencia [QAM96]. Inferior [BHM94]. Inferring [BGF95, Gol97a, QAM96, Ree96]. Infiriendo [BGF95]. Influence [BABK98, CHK98, GP97, HNC97, Leb91, Leb98, NKA98, PKH97, RDN94, Thi92, VTM99]. Influences [KB97, KLF98, YG92]. Influencia [PKH97, RDN94, GP97, HNC97, KB97]. Influencing [SÅS98]. Información [GMF99, Vog95]. Information [GMF99, Vog95]. Informative [DWM^{+}98]. informativo [HW95]. Informed [HW95]. Infraestructura [Win96a]. Infrastructure [Win96a]. Ingestion [MB99, MB99]. Inglaterra [Vog95]. Ingleborough [MNU94]. ingrahmi [CLL91]. ingreso [FP95]. Inhabited [CDH95]. Initiative [Ehr91, GM91, Sch91, Tho91, Wil91]. Inmune [SCZF96]. Insect [CT99, Gas91b, HE91, PS98, RFG^{+}98, RF99, Sto97]. Insect-Plant [CT99]. Insectívoras [SS99b, BGTB93, SPP96]. Insectivorous [BGTB93, SPP96, SS99b]. Insecto [CT99]. Insectos [RF99, Sto97, Hil95]. Insects [Hil95]. Insight [Nae96]. Insights [CSD98, Fro92, NY91]. Inspiring [Gra98]. Instability [HCBH99]. Instead [Vid94]. Insular [Aco99, BV97, BPC97, DB98, Dud99, VB99]. Insulares [Aco99]. Insularization [New96]. integrado [HH95b]. Integral [MB96]. integrar [KMM94, MWCP94]. Integrate [MWCP94]. Integrated [Blo99a, HH95b, KMM94]. Integrating [Fle90, HJ99, MS92]. integridad [LNPC^{+}95]. Integrity [Aus98, Cam95, FS92, Kar90b, LNPC^{+}95, MB96, MR98b, Nos90a]. Intelligence [Orr92d]. intensidad [GNH94, Sto96]. Intensity [GNH94, Sto96]. Intensive [Bal96]. Inter [Har92]. Inter-Refuge [Har92]. Interacción [CDD96, HC99]. Interacciones [CT99, HGNK93, VC99]. Interaction [CDD96, TLG^{+}98]. Interactions [CT99, HC99, HGNK93, VC99]. Interactive [JS98]. Interactors [CEPR91]. Interdependent [AHMM97]. Interdependientes [AHMM97]. interdisciplinario [HW95]. Interdisciplinary [HW95, KRL98]. Intereses [RS93a]. Interesting [You99]. Interests [RS93a, Sch92]. Interference [HT95, KW96]. Interferencia [HT95]. Interior [BB97, Fra94, Tho93, BB97]. intermedias [RCHP94]. International [Gin97, Med98, MC98, NY91, Tho99]. interpretación [SFM96a]. Interpretation [SFM96a]. Interpretations [Goo94]. Interspecific [HB94]. Intertidal [KQK93]. intertidales [KQK93]. intervalo [CSBC95]. intervalos [SB95c]. Intervention [Ber99]. Intestinal [Sto96]. intestinales [Sto96]. Intra [HNC97]. Intra-plantación [HNC97]. Intracruza [EKL^{+}99, KHM99]. Intraguild [LV96]. Intraspecific [Pae99]. Intratropical [PB95]. Intricate [HGNK93]. intrincadas [HGNK93]. Intrínseca [TA97]. Intrinsic [TA97]. Introducción [LDSH99, GWW93b, BPS95]. Introduced [Ber95b, BTG93, CLL91, CCOO^{+}96, CLM^{+}95, GK96, GMSP98, GWW93b, GK99, GW97, HSGY98, KB98, LDSH99, Loc99, RH97b, SGW90, SSS94, TLG^{+}98, VMMK92]. Introducida [Loc99]. Introducidas [BBCS99, GK99, RH97b, Ber95b, CLM^{+}95]. Introducido [GW97, CCOO^{+}96]. introducion [GK96]. Introduction [BPS95, CPC96b, DWM^{+}98, PD96, TW92a]. Introductions [Jen96, LC98a, MS97b, MS99c]. Introductory [HSS94, Tem93b]. Introgresión [EE97]. Introgression [EE97, MC95a]. inundación [AMGJ95, BHM94]. Inundaciones [MN96]. Invasibilidad [HV97]. Invasibility [HV97, PTTN^{+}96]. Invasion [BS92, BPS95, Gel99, HH92, UCB92, Gel99, BPS95, PTTN^{+}96]. Invasiones [RH97b, HH95b]. Invasions [BHL^{+}92, HH95b, RH97b]. Invasiva [BTRM99]. Invasive [HT95, Ric98]. Invasiveness [GMF99]. Invasividad [GMF99]. invasora [HT95]. Invasoras [HRCTS99]. Inventory [DCL98, JS98, MGM^{+}98]. invernación [CPN95]. invernales [HHH95]. inversiones [DW93, SFCM92]. Invertebrados [Nie97, IS96, OB96]. Invertebrate [HG97b, MFW90, OB96, OBY98, Sam96b]. Invertebrates [HS96a, IS96, Kel93a, Nie97]. Investigación [HP99, MH99b, FDV94, DS92]. Investigate [Way92]. Investigation [MWS^{+}97]. Investments [DW93, SFCM92]. Invierno [BB99]. Invitation [Ano95c]. Iowa [KSO^{+}99]. Ipomopsis [HP95]. Ironwood [Pel92]. irreemplazable [HGNK93]. Irreplaceable [HGNK93]. Isabella [Cob97a]. Isla [Ber95b, Fim94]. Island [Ber95b, CEPR91, EKL^{+}99, Fra98, Jor92, Kar90a, LV91, Rob99, RC94, SG98, SAH^{+}95, AMA^{+}94, Cob97a, Fim94, RG95, Wal90]. Islands [CDH95, MSC96, MS97b, PERC96, RTNK99, SGW90, Cam95, CSC^{+}94, Dud99, JJ99, Mac91, Mau97, RHSM94, Tsu98, WCD97]. Islas [BPC97, RTNK99, CDH95, MSC96, PERC96, RC94, SAH^{+}95, RHSM94]. Isle [WLG^{+}91]. Isleña [EKL^{+}99]. isoenzimas [NCE96]. Isolate [Rob99]. Isolated [Bla91, BD92, CA92, DP96, SB98, Son96, TCO^{+}96, Was96]. Isolation [BTG93, DBCM95, Mil95b, PDSB98]. Isozyme [NCE96]. Isozymes [KR98]. Israel [KASR99]. Issue [GWW93a]. Issues [HSD93, HEO95, HDE95, Wes90]. Italy [CDB99]. Ituri [WCTM98]. Itzá [CW97b]. IUCN [ML91]. Ivory [DP92].

Jacques [Gra98]. Japan [Ano90g]. Japanese [Kel91]. Jay [BLD^{+}95, Dud93]. Jays [ES94]. Jazz [Che99]. jejenes [Tsc92]. Jelmert [BCP96]. Jenkins [Yu96]. jerárquica [MBH96]. Jersey [RDN94]. Joaquin [NTT94]. Jobs [SC97a]. John [DM90b]. Johnson [Duf93c]. Jorobada [GD99a]. Joshua [CK98]. Journal [Ano96d]. Journey [Dra94]. Jóvenes [NM97]. Juan [CSC^{+}94]. jubatus [Mer94]. Judging [BWS94]. judgment [Ste93]. July [Ano99d]. Jumelle [RSR96]. jurisprudencia [DME95]. Juvenil [DH97]. Juvenile [ADFT97, LWC95]. Juveniles [ADFT97, SB94]. Juzgando [BWS94].

Kakerori [RHSM94]. Kalimantan [LLP95, SDT93, Sal93, LLP95]. Kangaroo [GP97, PK94]. Kathmandu [HYL95]. Kea [Ama93, RFM97]. Kearn. [SAH^{+}95]. Keeping [SF96]. Kemp [BCHM94]. Kentucky [GF90]. Kenya [FMF95, Med93]. Kestrel [CJ93, TFHD98]. Key [FT91]. Keys [JS98, BSC95, SB94]. Keystone [Pai95, WH95]. Kibale [NT98]. Kill [Ste93]. Kind [Cua97]. Kinds [SPR97]. Kings [BTG93, RS91a]. Kirtland [MHH98]. Kitchen [Ste98]. Klamath [CZN99, LAF93, CZN99]. Know [Rob93a]. Knowledge [GL93a, HDE95, JM97, Orr96a]. Known [JFH93]. Koopowits [Dia95]. Korean [HOF^{+}96]. Krill [Hol93]. Kruger [MG97].

L. [OO93, OLKD98, RVK^{+}94, SAS95, SFM96b]. Laboratory [BJW95]. Labos [Hab96]. Lacandona [Med94, WGO95]. Lack [AR97, BV97, Gra95, JD98]. Lactondaceae [CSC^{+}94]. Lactoris [CSC^{+}94]. lacustre [FP95]. Lag [BPO99, LS96]. Lago [AM99, SWK^{+}97, DeM93, FC93, ACB^{+}99, CCOO^{+}96, CBCC93, CM93, GWW93b, HGNK93, KO93, OO93, Rei93]. Lagos [BH93, Fra97, LM93, MSHW95]. Lake [FC93, FP95, SWK^{+}97, AM99, ACB^{+}99, CKCM95, CCOO^{+}96, CBCC93, CM93, GWW93b, HGNK93, KO93, OO93, PB94a, Rei93, SKOO98, ZDM98]. Lakes [Fra97, Kau93a, KC93, MSHW95, TLG^{+}98, BH93, LM93, MS93]. Lakeside [DeM93]. Lam. [FCR^{+}98b]. lanceolata [PSB98]. Lanchas [GB97]. Land [All93a, CC99a, Cor96, CD98, DOSP94, DPOO94, FS95, GF97, JM97, Koe98, LS96, Lit93, Mat94, MS99b, OPRCB95, PL95, RCM97, Sad95, SJ98, Sam90, VRV^{+}94, War94a, WRSP97]. Land-Bird [SJ98]. Land-Use [Cor96, DPOO94, FS95, Mat94, WRSP97]. Landbird [Rod92]. Landbirds [DHW99]. Landfill [RH93]. Lands [BB90, CBLV94, Gru90, Kni99, Mae90]. Landscape [ACB^{+}99, AB90, Bak94, BY96, BH97, BSR97, BLD^{+}95, CUM96, CHK98, CFL96, DBH^{+}99, FCM99, GHP97, Haa95, HH95a, HMO98, HNC97, HRCTS99, Hun98, KR95, KSO^{+}99, LMNV92, LB95, LG97, LSV^{+}96, LCPM97, MWM^{+}99, MC95b, MW93, ML94, MN96, MWCP94, MSHW95, MR99, NKA98, Pea97, RR99d, RHK99, SÅS98, SAM98, Ste98, SS99b, VTM99, WMB^{+}97, Wit97a, YM96]. Landscape-Scale [HRCTS99]. Landscapes [DD97, DSB^{+}95a, HMW98, Jai93, KGD96, MCW98, MH99b, PF99, SB97, TA97, WGO95, WK99, YM97]. Langara [Ber95b]. Langosta [Aco99]. Langostino [GW97]. Langtang [FYP96, FYP96]. Lanius [YG92]. Lanka [GG90b, MGB94, Wik90]. Larga [GD99a]. Large [BMR^{+}92, CCR96, CPC96a, CDB99, De 95, EB99, FBFO96, Got95, HPB^{+}97, KL96, KBRB96, LV91, New96, PH94, RH96, RW99, RS93c, WPR96, WR96, WDR^{+}98, You94b]. Large-Scale [CDB99, De 95, Got95, RS93c]. Larger [CC96]. Largo [GD99b, KASR99, PF99, Ren99, FESS94, FC93, Fra94, Har94, JFJ96, Med93]. Lark [YS96]. Larval [TLG^{+}98]. Last [CW97b, Cra94, Cua97, Ste93]. Later [KALG94]. Lates [OO93]. latifoliados [Fra94]. Latimeria [HFS98]. Latin [GL93b, Lan96, Vid93]. Latina [Lan96]. Latino [GL93b]. latirostris [MHO97]. Laundry [Mag92b]. Lava [SSS94]. Law [AM99, DME95, KL96, Sal98]. Lawrence [Sal92]. Laws [LWMG93]. Laying [KB97]. Leadbeater [LP96]. leadbeateri [LP96]. Leading [MHB^{+}96]. Leaf [OK96]. Lecciones [Sym96, TTW^{+}94, NR97]. Lechuza [SGMP99]. Leek [NGN96]. Left [NM95]. Legged [Bes96, Bri95b, KB98, LDSH99, BHKH96]. Legislation [And96, Kar90b, MWW99, RG99]. Legume [SMPN99]. Leguminosae [HE91]. Leguminosas [SMPN99]. Lemur [SHM97, TBL^{+}97]. Lemures [SHM97, GFR^{+}99]. Lemurs [GFR^{+}99]. Leña [AM99]. Leñosas [CC99b, GS99, RH97b]. Lens [VP97]. Leon [EE97]. leones [SGCJ94]. Leopardos [MSO96]. Leopards [MSO96]. Lepilemur [TBL^{+}97]. Leptodactylid [MP97]. leptorrhynchoides [YBZ99]. Lesquerella [RMCM97]. Less [Sab97]. Lesser [TFHD98, LBR99]. Lesson [RO99]. Lessons [CPC96a, Ehr90, Kar90b, Kio91, LKM90, NR97, Sym96, TTW^{+}94, Wie99].

Let [VL91]. Letter [And95, Ano90m, CD90, FVK95, Her90, McL99, Nos95a, Rab95b, Sum95, TN98]. Letters [AWR^{+}98, Ano91j, Ano91k, Ano91l, Ano97d, GWF98, GSB98, MNC^{+}98]. Letting [Wil92]. leucotos [CA92]. Level [Die94, PL94, QAM96]. Levels [BEG^{+}95, DM90a, EKL^{+}99, TB98]. Lewis [MS99a]. Ley [AM99]. leyes [LWMG93, KL96]. liberada [Bus96]. Liberal [Orr91a]. libre [Yu94]. Lichen [ER98, NM97, PBB99]. Lichens [NABE95]. Lie [Rod93a]. Life [BNF^{+}98, CGG^{+}97, CMMLRR95, Hea94, JRP99, Joh94a, Man94, Nae96, Orr92c, RO99, Tho94c, WP96]. Life-History [RO99, Tho94c]. Lifetime [RB95b, SM91]. Like [BCP96, Dun97]. Likelihood [BER97b]. Likely [FSCL96]. Liliaceae [GHB95]. Lily [GWH97]. Limestone [MNU94]. Limit [PM92]. limitación [CC96]. limitaciones [SDB^{+}96, WKS^{+}94]. Limitantes [PS96]. limitar [PM92]. Limitation [CC96]. Limitations [SDB^{+}96, SDB^{+}97, WKS^{+}94]. Limite [Sym96]. Limites [FT96, SMPN99, Rob93b]. Limiting [BD92, PS96]. Limits [Gra99, HM93, Orr98, Rob93b, Tho99]. Limnanthaceae [DS92]. Limnanthes [DS92]. limnológica [BH93]. Limnological [BH93]. linaje [QLDV96]. Linajes [GSB97]. Lineage [QLDV96]. Lineages [GSB97, MBT90, TSSB95]. lineales [Hil95]. lineamientos [SSEM94, RRS96]. Linear [BC97a, Hil95]. Lineares [BC97a]. Linepithema [Hol95]. Lines [HS95]. Link [TD96]. Linkage [BBGB92]. Linkages [HSGY98]. Linked [RS93c]. Linking [CNP^{+}97, LDP95]. Links [You95]. Lion [KM98, MBD95, SGCJ94]. Lion-Human [SGCJ94]. Lions [PPR^{+}91, SJB97]. Líquenes [NM97, PBB99]. List [Kei98, Mag92b, Win98, WMW93]. lista [WMW93]. listadas [Rei95]. Listed [BEG^{+}95, Rei95]. Literacy [Dea95]. Literate [Dun99]. Literatura [GD99b]. Literature [GD99b]. litorales [HGNK93]. Litterfall [FP95]. Little [CLL91]. Littles [SMS97]. Littoral [HGNK93]. Litvaitis [PDFG96]. Lived [CDV93, GD99a]. Livestock [BB97, BB93, BBS93, BM95, DRP98, Dud97, Fle94a, FFH96, HR93, MR98a]. Living [Joh94a, Nae96, NMGS94, Rob93b]. Livistona [OK96]. Lizard [SSS94]. Lizards [DB98, Fit94, VAPCO98]. llanura [AMGJ95]. Llanuras [WER^{+}97, BHM94]. Lluviosa [TWMM97, ART94, SDT93]. Lluvioso [TCO^{+}96]. lluviosos [WDSA94]. Loan [WSB92]. Lobo [FB97, GMMR^{+}96, KHM99, Bro96, MSHW95, WLAH92]. lobophylla [GWH^{+}99]. Lobos [FT96, FB96, VPW97, LRT93, Mec95, RGSW96]. Lobsters [Aco99]. Local [BWS94, Cor96, FS97, KS97, Mun95, New91, NMGS94, New96, PDS96, PM92, RMCM97, SS99b, TJ99, Tsc92]. Locales [FS97, PDS96, SS99b]. localizadas [LNA95]. Localization [TSSB95]. Locally [HH94c]. Locandona [Med94]. Located [LNA95, RS92]. Loci [GMMR^{+}96]. Loftis [Duf93b]. Logged [CC99b]. Loggerhead [BAR^{+}93, CW97a, JBB96, MB99, YG92, Yos94]. Logging [Ano91m, Ben93, Bra94, ELS93, Fri92, Mar98, MGB94, PP93, RR99d, Thi92, WSB92]. Logic [MBW^{+}98]. Long [BB90, CDV93, Del99, FC93, GD99a, GD99b, Har94, JFJ96, KSSM95, Orr93b, PF99, RS91a, Ren99, Rob99]. Long-Lived [CDV93, GD99a]. Long-Range [FC93]. Long-Term [BB90, Del99, GD99b, Har94, JFJ96, KSSM95, Ren99, Rob99]. longirostris [FFL95]. Longleaf [NPM98]. Lonicera [SW99a]. Look [Dai95, McN95]. Looks [Dun97]. Lorantáceos [NR97]. Loranthaceae [RKLS99]. Loranthaceous [NR97]. Loss [Ang95, BMR^{+}92, DB98, FCR^{+}98a, FM96, GSB97, Mil94, MW90, RKLS99, Rob93b, TCO^{+}96, Was96, WSE97]. Losses [KTA94]. Lost [DP96]. Love [Bre94, Orr92c]. Loving [Hol93]. Low [BBHM95, CKCM95, EKL^{+}99, HCB94, KGM^{+}97]. Low-Founder-Number [BBHM95]. Lower [BHM94]. Lowland [FGG^{+}95, GW97, KO98, MBD^{+}99, TCO^{+}96, WER^{+}97]. Loxodonta [AL93]. Luangwa [LWMG93]. Luddites [Bis93]. ludovicianus [YG92]. Lugo [Cob91]. Lupinus [Hel98]. lupus [LR91, LRT93]. Luquillo [Lug94]. Lurking [Dal99]. luteiventris [MGM^{+}98]. Lycaon [GAC^{+}95]. Lychnis [LSM99]. Lyme [Gin94, Gin94]. Lynx [PGFD95].

MacArthur [Dia96]. macdonaldi [CMMLRR95]. macrodactylum [TLG^{+}98]. Macrofungi [RDH97]. Macrohongos [RDH97]. Macroinvertebrados [Fra97]. Macroinvertebrate [Fra97]. maculosa [FCR^{+}98b]. Madagascar [GFR^{+}99, Goo95, KRL98, KRG^{+}99, RSR96, RN96, RS91b, SHM97, TBL^{+}97]. Madera [Fra97]. maderas [Tho93]. maderero [LP96]. Madreselva [HV97]. Maduro [BRHD97]. Maduros [CC99b, Lug97, CPN95, MWCP94]. magister [BY96]. Magnitude [Gas91b]. Maguire [WS97]. maguirei [WS97]. Mahogany [RG95]. Maine [BHL^{+}92, DH98, PCM98, RKBP93, VHM94]. Maintain [Nos90a]. Major [MMT^{+}98]. Making [Blu93, Ehr93, Mag96, Phi94, PC96, RS95a, Saf93, SFM99, Wes91]. mal [Gin94]. Malaclemis [RCMR97]. Malaclemys [RCMR97]. Malacothamnus [SAH^{+}95]. malas [SFCM92]. Malawi [AM99, AM99, Rei93]. Malvaceae [SAH^{+}95]. Mamiferos [BER97b, GSB97, AFF^{+}97, DHE97, TWMM97, WGRT96, CB95, SD94, DY93, Med94, New95, New96, ST94, You94b, GL93a, RW99]. Mammal [FJPC95, FMF95, Med94, MS99b, New95, RBMA98, SD94, Vid93, WDR^{+}98]. Mammalian [Ama93, BW91, CB95, RW99, Sal93, SCAB91, WGRT96]. Mammals [Ano90p, ARR90, AFF^{+}97, BER97b, DY93, DHE97, EB99, GSB97, GL93a, HY93, Lau91, MB92a, New96, ST94, SS94, TWMM97, You94b]. Man [Gur92]. Managed [DBCM95, LLP95, NM97, PTK^{+}98, SÅS98, YM96]. Management [Alp95, AMGJ95, ALF92, ADM94, Aug95, BS98, BHCB99, Blo99a, BMR^{+}92, BR96, BC97b, CSBC95, CUM96, CAWH91, CMB96, CDV93, Cro95, DOSP94, FT96, GLFZ96, Gol99a, GL95, Gru90, Gru94, Gru97, HBD90, HB90, Hei96a, HH95b, HP99, HM96, JM97, KM98, LP96, LF97, LDP95, MS92, MHH98, MHWP96, MH99b, MBM^{+}98, MA96, MWCP94, Nie97, Nor93, NR97, OFH94, Orr90b, Pel95, PGW97, RS95a, RMJ92, RSR96, Rob95c, Saf93, Sal92, SFM96a, Sim96, Sta95, SHS^{+}97, Tho93, TV99, VD94, VL91, War94b, WP97, Wei99, Wes90, WC98, Wil93b, Woo99, Yaf99, You95, Pel92]. Managers [Tem92a]. Managing [AM99, BLS99, How94, Hun96a, Jai93, KK94, LC99, LF97]. Manaos [SS99b]. Manatee [MHO97]. Manatí [MHO97]. manatus [MHO97]. Manaus [SS99b]. Mandate [Bab99]. Manejando [KK94]. Manejo [BHCB99, BLS99, BC97b, GLFZ96, Gru97, HP99, How94, JM97, LC99, LF97, MHWP96, MH99b, MA96, NM97, Nie97, NR97, PGW97, Sta95, TV99, WP97, Wei99, Yaf99, AMGJ95, ADM94, CSBC95, CUM96, DOSP94, DBCM95, GL95, Gru94, HH95b, HM96, LLP95, LP96, LDP95, MWCP94, OFH94, RS95a, RSR96, Rob95c, SFM96a, Sim96, Tho93, VD94, AM99, FT96, LF97]. Mangabey [KO91]. Manglares [Lug97]. Mangrove [Lug97]. Manipulación [GAC^{+}95]. mantelli [BJ95]. Mantled [Sto96]. Manual [Cro95, Qui95]. Many [Erw91, Gol97b, MG90, PAPA90]. Map [Pea92]. Mapeando [FYP96]. Mapimí [Kau93b]. Mapping [FYP96, FS91]. Maps [Ker96]. marco [HW95, LOT^{+}99, MH99b]. Mareas [LSA99]. Margarita [YBZ99, DeM93, SG98]. Marina [Gar97, MB99, RDE96, Rob95b]. Marinas [Aco99, BBCS99, HC96, HWG99, MPC99, Wal99b, EBV95, Ver93]. Marine [Aco99, BL95b, Bra98, EK98, FS92, HSGY98, Hol93, HWG99, IN96, Kau95a, MKA96, MD96, Rob95b, SB95a, Ver93, Wal99b, WN98b]. Marine-Freshwater [HSGY98]. Marino [MKA96]. marinos [BL95b, MKM^{+}96]. marinus [SFM96b]. Mariposa [BGK^{+}97, BBM94b, HMPK96]. Mariposas [Wei99, SSEM94, ST96]. Markers [PEP98, Ran96]. Market [Caf96, FJPC95, Hei96b, WSB92]. Markets [HEO95]. marmorata [Gra95]. Marshall [SAS95]. marsupial [Rob95c]. Marsupialia [SMGB91]. marta [SY94]. Marten [CHK98, SY94]. Martens [Bus92]. Martes [SY94]. Más [GG97, KALG94, DW93]. masa [RHCB93]. Masked [Gel99]. Masoala [KRG^{+}99]. Masola [KRG^{+}99]. Mass [CMW^{+}98, Lip99, RHCB93]. Massachusetts [GF97]. Massachussets [GF97]. Masses [MS99b]. Master [Fra95b]. Maternal [JDEB98]. Mating [HWB96, KEE91, Sun96]. matorral [QAM96]. Matrices [Ano97e, SFM96a, MDW99]. Matrix [MDW99]. Matter [Car94]. Mattered [Joh94a]. Matters [STA^{+}99]. Mature [CPN95]. Maturity [CDV93]. Mauna [Ama93, Ama93, RFM97]. Mauricio [CJ93]. Mauritius [CJ93, SJ98]. mawii [PH94]. máxima [WGM^{+}96]. Maximizing [RS92]. máximos [WGM^{+}96]. May [Fim94, MG90, PB94a]. Maya [CW97b]. Maya-Itzá [CW97b]. Mayor [BC99]. Mayores [CCR96, KBRB96, RH96, WPR96, WR96]. Mbaracayu [HPB^{+}97]. McDonald [Dud97, FFH96]. Meadow [DRP98]. Meadowfoam [DS92]. Measurement [SMGB91]. Measurements [WRM^{+}91]. Measures [KMW95, PS99, WBKH97]. Mecanismos [JR99, SM97]. mecanizada [WSB92]. Mechanisms [JR99, SM97]. Mechanized [WSB92]. Mediated [Sim95]. Medicinal [HEO95]. Medicine [Mef99a]. Medicines [BM92]. Medidas [PS99, KMW95]. medio [FESS94, SK93, RW99, War94a]. Mediterranean [BAR^{+}93, MQ99]. Meeting [Ano90a, Ano95a, Ano95c, GG99, Tem92b, Ano99a]. Meetings [Ano91n]. Meets [Aug95, Rei94]. Megaherbivore [Dud99]. Megalonaias [MLP^{+}97]. megistophylla [MGB94]. Mejillón [Gel99]. Mejillones [VT99]. mejor [BJ95]. mejoramiento [HBA93, MWCP94]. mejorar [OSEC95]. Melanotaenia [ZDM98]. melodus [RRM93]. Memoriam [For94]. Menaces [ALGPH95]. Mesófito [HC99]. Mesopredator [LV96, PDFG96]. Message [Avi94]. Meta [FDV94, HH98]. meta-análisis [FDV94]. Meta-Analysis [FDV94, HH98]. Metabolism [FP95]. metabolismo [FP95]. Metapoblación [AA97, LSA99, HMPK96]. metapoblacional [KA92]. metapoblaciones [QAM96, Tho94a, WKS^{+}94]. Metapopulation [AA97, HMPK96, Hed96, KA92, LSA99, QAM96, SP90, WKS^{+}94, MFW90]. Metapopulations [Tho94a]. Method [Dre94, Mur90, OB93, PP93, RBPB96, SGA98, TM96, WDR^{+}98]. Methods [Ber93, NBH^{+}98, Pae99, RS95a]. método [OB93, TM96]. métodos [RS95a]. Metrioptera [KA92]. Métro [RBPB96]. Metropolitan [DP96]. metropolitana [DP96]. Mexican [AFF^{+}97, DG93, GMMR^{+}96, KHM99, SGMP99]. Mexicana [SGMP99]. Mexicano [GMMR^{+}96, KHM99]. Mexico [TA99, TA99, ADM94, CRGGFT97, CG95, CBLV94, Cue96, DG92, FESS94, FC99, GB97, GLFZ96, LNPC^{+}95, Med94, MT99, REDH98, SP98, VIO98, WGO95, AFF^{+}97, ADM94, CRGGFT97, CBLV94, FESS94, FC99, GB97, GLFZ96, LNPC^{+}95, Med94]. mezcla [IS96]. MHC [GW91, Hug91, MH91, VL91]. Michilía [FC99]. micro [BY96]. Microenvironment [WLDGGZ98]. Microhabitat [BY96, KB98]. Microsatélites [GMMR^{+}96]. Microsatellite [GMMR^{+}96]. Microseris [PSB98]. Microsporidian [CD98]. Midcontinental [SK93]. Middle [MBW96, FESS94, MBW96, RW99]. Midges [Tsc92]. Midwestern [War94a]. Migración [FB97, HOF^{+}96, Gre96, PB95, Sha93, WER^{+}97]. migraciones [SK93]. Migrant [ADFT97, DLK^{+}95, FEM95, FCM99, MA96, MR99, Rod92, RCHP94, Yah95]. Migrante [MA96, MR99]. Migrating [SMHK98]. Migration [BBHM95, FB97, Gre96, HOF^{+}96, PB95, Sha93, WER^{+}97]. Migrations [PBBA90, SK93]. migratoria [ART94]. Migratorias [ADFT97, FCM99, CPN95, DTFP95, FEM95]. migratorios [MH93, RCHP94]. Migratory [CPN95, DTFP95, MH93, PTK^{+}98]. Miguel [Cua97]. Millennium [Mef98a]. mimosoides [YB96]. Minded [Hun94]. mínima [SD94]. Minimal [RF97, SD94]. mínimas [Bei93]. mínimo [Tsc92]. Minimum [Bei93, Tho90, Tsc92]. Minnesota [MCO98]. Minnow [EE97]. Mirlo [GD99b]. Mis [Pel92]. Mishra [Sab98]. Missing [Bal99, Dun97, GL93a, MEE93, Mef94]. Mississippi [LSC^{+}96, MN96]. Missouri [HL99]. Mistletoes [NHA95, NR97, RKLS99]. Mitigación [BN97]. Mitigation [BN97, BN98b, BFS98]. Mitochondrial [AMW90, Avi95, DMMG96, GSB99, Gol97a, LBR99, MAS^{+}97, NTT94, OFH94, PDSB98, Rob95c, TSSB95, WTWS98, WLAH92]. Mitocondrial [LBR99, DMMG96, GSB99, Rob95c, WLAH92]. mitrocondrial [OFH94]. Mixed [Fra94]. Mixedwood [CB98]. mixtos [Fra94]. Mobbing [SRR95]. Mobile [WN98a, WN98b]. mocinno [PB95]. Model [AA97, Ano97c, Aus98, Bar96, Bie95, BJ96, CDB99, FP95, GG99, HMPK96, HC96, Hun98, KEE91, KTA94, LDP95, MB92a, MBW^{+}98, MS93, PKH97, PK94, RWES93, RRM93, SEC98, TB95, WL97]. modelación [SRR95, BLF^{+}93, DLK^{+}95]. Modelando [AHMM97, DOSP94]. Modeling [AHMM97, BLF^{+}93, CSD98, DOSP94, DLK^{+}95, FM96, GLFM98, HMW98, Kin95, SRR95, WCTM98, Woo99]. Modelo [HC96, PKH97, HMPK96, KTA94, LDP95, PK94, RWES93, RRM93, TB95, AA97, FP95, GG99]. Modelos [BHS99, CZN99, LC99, MH99b, MDW99, Sou99, Wit97a, Doa95, EDFM96, FH95, IS96, RS95a, RS93c, BSW^{+}96, SM97]. Models [Ano99i, BHS99, BSW^{+}96, CZN99, CG98, Doa95, EDFM96, FH95, HSD93, IS96, LC99, MH99b, MDW99, PC98, PFSG97, RS95a, RS93c, RHK97, RHK99, SM97, Sou99, TL94, Wit97a, RS92]. modificadas [RHCB93]. modificados [Fim94]. Modification [Kei98]. Modified [Fim94, RHCB93]. Moldeando [CCR96]. Moldelado [FM96]. Molecular [Bro96, CRJ^{+}94, DMD^{+}92, KB99, MWS^{+}97, MSO96, PSH99, QLDV96, Ran96, RFM97, RGSW96, SFM99, SG99, Way92]. Molecular-Genetic [MWS^{+}97]. Moleculares [RFM97]. Mollusc [KQK93]. molothrus [HH95a]. moluscos [KQK93]. Monachus [SRR95, TSP97]. Monarch [WRM^{+}91]. Money [Sab97]. Mongolia [VW96]. Mongooses [PGFD95]. monilifera [SAS95]. Monitor [WW99]. Monitoreo [GDK99, WW99, SSEM94, KMM94]. Monitoring [BG95a, DCL98, GDK99, HFS98, KMM94, KRL98, Nos90b, SSEM94]. Monk [GG90a, KGM^{+}97, SRR95]. Monkey [SS95]. Monkeys [Sto96]. monos [Sto96]. Montana [TW92b, HC99, LF97, RN96, Weh99, MW98, MWM^{+}99, GLFM98]. Montañas [MBBF95, Car93, Lug94, Ash97, FBS95, FB96, FB97, GL93a, HGK^{+}96, KL96, MHWP96, NQH^{+}96, PC96, RJBB96, WPR96]. Montane [Lip98, MP97, MB92a, RN96, Sam90, SLNP95, You94a]. montanos [You94a]. Monte [War94c]. Monterey [EK98]. Morality [And96]. morfoespecies [OB96]. Morphologic [Way92]. Morphology [Win96b]. Morphospecies [OB96]. Mortalidad [Gar97, RCMR97]. Mortality [FL97, Gar97, HE97, LWC95, Lip99, LP98, RCMR97]. Mosaic [ART94, ST94]. mosaico [ART94, ST94]. Mosca [WRSP97]. moscas [Tsc92]. Mosquito [GK99, LDSH99]. Mosquitofish [GK96, GK99, LDSH99]. Most [Ber95a, FSCL96, OD98]. Mostaza [RMCM97]. Moteada [SGMP99]. moteado [LMNV92, LNVM94, TG93]. moth [Ste93]. Moths [Tsc92]. Motorizadas [GB97]. Motorized [GB97]. Mountain [Ber93, BWH90, Goo94, Hut95a, Hut95c, KEL93b, LF97, Lym94, MBD95, PFSG97, RG95, SJB97, Weh99, FB97, MHWP96]. Mountainous [LP98]. Mountains [AB90, Car93, GL93a, Ing90, KL96, Lug94, MW98, New91, RCM97, Ash97, FB96, MBBF95, NQH^{+}96, RJBB96, WPR96]. Mouse [JDEB98]. Movement [DC98, MVH96, SFCM92]. Movements [HMO98, MP99, WPSZ96]. movilidad [WPSZ96]. movimiento [SFCM92]. Movimientos [MP99]. Mt. [War94c, War96]. mtDNA [PSSB95]. Mu [DLY92]. Muérdagos [NR97, RKLS99, NHA95]. muestras [BL95b]. Muestreo [CAH97]. Multi [Ing90, Lam97]. Multi-gene-pool [Ing90]. Multi-Species [Lam97]. Multiespecífica [Lam97]. Multiple [CG98, MR99, OBY98, MR99]. Multiple-Use [MR99, OBY98]. multiples [RS93c]. Multiscale [RS93c]. Mundo [RR99a]. Munro [Rob93a]. munzii [NCE96]. Murciélagos [PP96, RHCB93, SVC96]. muricatus [BR96]. Murphy [Ehr95, Sal92]. Murrelets [Ber95b]. muscosa [BTG93]. Mushrooms [HD96]. Mussel [Gel99, SEC98]. Mussels [MLP^{+}97, VT99]. Must [DP92]. Mustard [RMCM97]. Mustela [RK93]. mustelinus [TBL^{+}97]. Mutación [Lan95]. Mutaciones [GWM^{+}97]. Mutation [GWM^{+}97, Lan95]. Mutualisms [MCW98, NF93]. Mycteria [BHSO92]. Myotis [RHCB93]. Myth [Duf93a]. Myths [Ber91].

N. [Wil93a]. Nabugabo [CCOO^{+}96, CCOO^{+}96, OO93]. Nacional [AM99, BWS97, FYP96, GL95, KRG^{+}99, MM94, Cur93]. Nacionales [AS94, Boz93, WMH94, New95]. Nahuel [VMMK92]. naped [HOF^{+}96]. Narrow [RDN94, SSTL92, WS97]. Nassella [KR98]. Nasty [Tem90]. Nation [Gra99, HHH96]. National [AM99, AS94, AB90, BWS97, BTG93, CK98, Cur93, EK98, FYP96, Gei90, GL95, HS95, Jha93, KRG^{+}99, MBT90, MM94, MG97, MGM^{+}98, NT98, New95, Sal93, Sal91, TW92b, VMMK92, WMH94, Yea97, Boz93]. nativa [HT95]. Nativas [FB97, EE97, HGK^{+}96, ML94]. Native [Ano91m, Ber91, BTG93, DD97, FB97, Fri93, GKA97, GMSP98, Gel99, GB95b, HGK^{+}96, HSGY98, KB98, ML94, MW90, RS93a, SEC98, SAS95, SGK95, UCB92]. natives [EE97]. Nativo [Gel99]. nativos [Fri93, GB95b, RS93a]. Natterjack [DHBG97]. Natural [ALF92, BBGB92, BS98, BC98, BLS99, Bun95, Cor96, CD92, Fra95b, GSB97, GN92, HM96, Hun96a, Mar93, MBM^{+}98, NG92, PS99, Riv97, San99, SCZF96, SB95c, TWMM97, WP97, You94b, CMMLRR95, HCH^{+}97]. Naturales [BLS99, FB96, PT95, PQL99, SCZF96, WP97, Bun95, De 95, GNH94, HM96, Lan96, LNA95, MNU94, You94b]. Naturaleza [Lam97]. Naturalist [PCG91]. Naturalists [Ano99p, BBA^{+}96, Nos96c]. Naturally [BWH90, FB96]. Naturalness [And91, Göt92]. Nature [Ano99k, BD92, Bre94, CPG94, De 95, Edu91, Göt92, Lam97, Lan96, LNA95, MNU94, Nor96, Orr96b, Orr98, Per94, PT95, PQL99, RS92, SB95c]. Near [Gas91a, SJB97, ES94, SS99b]. Near-Extinction [SJB97]. Near-Imponderable [Gas91a]. necesidad [KMM94]. Necesidades [GG99, FYP96]. Necessity [Tem90]. Nectar [PP96]. Nectar-Feeding [PP96]. Nectarívoros [PP96]. Need [Bir93, Bre94, Dun99, KMM94, TD96, VP97]. Needed [Mac91]. Needlegrass [KR98]. Needs [Bri95b, FYP96, GG99, Kau95a]. Negative [ES94]. Neglected [Kni99]. Negligent [Saf93]. Negra [PGW97]. Negras [EKL^{+}99, RK93]. Negro [SF97]. Negros [Ber97a, MAR96, OFH94, PZSG96]. Neighborhood [Cha95]. Neighborhoods [Ano99k]. Nematodes [BL95b]. Neodypsis [RSR96]. neomigratorias [DLK^{+}95]. Neotoma [BY96]. Neotropical [ADFT97, ARR90, CG95, CPN95, DS95, DM90b, DG92, DY93, DLK^{+}95, FEM95, FCM99, MH93, MR99, Rod92, SSEM94, Ste98, WER^{+}97, Yah95]. Neotropicales [ADFT97, CPN95, DY93, FCM99, FEM95, MH93, SSEM94]. Neotrópico [ARRK97]. Neotropics [ARRK97]. Nepal [FYP96, HYL95, HY93, SAM98, YH91]. nesioticus [SAH^{+}95]. Nest [Ama93, Gat94, HH98, Has95, KHS98, KDG98, Pat94a, SW99a, SW99b, Yah96, YS96]. Nested [Bla91, Cut91]. Nesting [BSC95, BBL93, BWBM99, FCM99, JBB96, KHS98, MR99, RDN94, RS95b, SPR97, WF99a]. Nests [BH97, YM96]. Net [SKOO98]. Nets [RKBP93]. Network [RS92]. Neutral [Wit97a]. Neutrality [Ran96]. Neutros [Wit97a]. Neuvo [TA99]. Nevada [DF96, MR98b]. Never [Med98]. Newfoundland [SGR^{+}98]. News [Ano91o, Ano91p, Bru93a, Bru93b, Gin97, Lov90a, Lov90b, Lov90c, Lov90d, Lov91, Tem91]. Newts [GK96, GKA97]. Next [Tro94]. Ngorongoro [MAR96, PPR^{+}91]. Niagara [KL97, KL97]. Nicrophorus [LCSC95]. Nidación [FCM99, MR99, SPR97]. Nidan [WF99a]. nidificación [Pat94a]. nidificadoras [BSC95, RS95b]. nidificantes [RDN94]. Nidos [SW99a, YS96, Ama93]. Niger [Ing90]. nigh [Ste93]. Night [SNP94]. Night-Blooming [SNP94]. Nightmares [Yaf97]. nigripes [RK93]. Nile [GWW93b, OO93, SKOO98]. Nilo [GWW93b, OO93]. niloticus [OO93]. Nitrogen [Wei99]. Nitrógeno [Wei99]. Nivel [HRCTS99, BLD^{+}95, PL94]. Niveles [CT99, HCB94, EKL^{+}99]. No [EE97, Hag92, KHM99, TA99, BTRM99, BC99, CDH95, HT95, PF99, Yu94]. No-nativas [EE97]. Noah [Bab99, Sch93c, SC97a]. noche [SNP94]. Nocturnal [TBL^{+}97]. Nocturno [TBL^{+}97]. nodosum [KMR94]. Nodricismo [SNP94]. nómadas [FBFO96]. Nomadic [FBFO96]. Nomura [RJL99]. Non [Cob97b, EE97, GKA97, God93a, God93b]. Non-Native [GKA97]. Non-natives [EE97]. Non-Sustainable [Cob97b]. Non-Timber [God93a, God93b]. Nonbreeding [GB97, YG92]. Nondeclining [RB95a]. None [OE95]. Nonequilibrium [SB97]. Nongame [TA99]. Nonhuman [Kin92]. Noninfected [BC99]. Noninvasive [BTRM99, WBKH97]. Nonmarine [HS96a]. Nonnative [HT95]. Nontimber [SDT93]. Noreste [CC99b, Sto96, LEMC95]. Normative [CCM99]. Noroeste [Fri93]. Norte [BPS95, RRM93, LMNV92, LNVM94, MSHW95, Wal99b, WPSZ96, BGTB93, CD95, EBV95, Fra94, GD99a, KBRB96, New95, RH97b, RR99b, SPP96, TG93, TCK^{+}95]. North [Fle94a, Fra94, Mef92, New95, RRM93, BGTB93, BPS95, EBV95, GD99a, GF97, Hou95, KBRB96, Meh94, MLP^{+}97, PWC^{+}98, PSSB95, RH97b, RR99b, SPP96, WTWS98, WW98b]. Northeast [PVZJCI90]. Northeastern [CC99b, LEMC95, Sto96]. Northern [Bar95, Gra95, Hol95, HG97b, Ing90, KO93, LMNV92, LNVM94, MCO98, TG93, Wal99b, WBKH97, CD95, Hut95c, MS93, MSHW95, TCK^{+}95]. Northwest [Wie99, Fri93]. Northwestern [BAR^{+}93, DRP98]. Noruega [Sto97]. Norwegian [Sto97]. Norwood [Dud93]. Nosed [SS95]. Noss [Tem93c]. nota [Hes94]. Notables [PD97]. Note [Hes94]. Novel [MPC99]. Nowak [DMD^{+}92, WRG98]. Nuca [HOF^{+}96]. Nueva [MJ97, RDN94, Vog95, PNH93, BJ95, CDH95, MSC96, NR97, PL95, RKLS99]. nuevas [WMH94, BJ95, MPC99]. Nuevo [FESS94]. Null [BSW^{+}96, PFSG97]. nulos [BSW^{+}96]. Number [BBHM95, BPC97, RJL99, VW98]. Numbers [Rym94]. Numero [MLD97, BPC97]. Nurse [SNP94]. Nutria [Gar97, RDE96]. Nutrias [BBCS99]. Nutrient [MBW^{+}98, PTK^{+}98, Wei99]. Nutrient-Poor [Wei99]. nutrientes [FS95]. Nutrients [FS95]. Nutritional [YG92]. Nutt. [SAH^{+}95]. Nuttall [TB93].

o [ADM94, Pre94, RS93a, Rei95, SFCM92]. Oases [SMHK98]. Obituary [Nor91]. Objectives [Ano90i, Ano90j, Ano90k, Ano90l, Ano91i, Lan96]. Objectivity [SF96]. objetivos [Lan96]. Obligate [BR96]. observaciones [JBB96]. Observation [Wea95]. Observations [CSC^{+}94]. Observed [KHM99]. Obstacles [Bra98, FWB^{+}99]. Obtenido [MJ97]. obtenidos [TM96]. Ocasionada [KHM99]. occidental [De 95, LLP95, MB96, NHA95, CC99a, SHM97]. occidentalis [KL97]. Occidente [BB97]. Occupancy [VTM99]. Occupying [PBG^{+}93]. Occurrence [MGM^{+}98, SÅS98]. Ocean [Ang93, Car98, MC98, BAR^{+}93]. Oceanic [MS97b]. Oceans [ROT^{+}99]. O’Connell [Dud92, Roh92]. Ocurrencia [Sto96, QAM96]. Odonata [SS98]. oedipus [CRJ^{+}94]. oeste [Ama93, Fle94a, KS94b, LNPC^{+}95, New95, NLS93, PL95, MWM^{+}99, War94a]. Off [BBA^{+}96, Nos96c, WCD97]. Offs [You94b]. Offset [PP93]. Offshore [MSC96]. Ohio [SB95c, DSB^{+}95a, HV97]. Oil [Gar97]. Oiled [Est98, Jes98]. Okinawa [Tsu98]. Old [BS92, BRHD97, CC99b, JR99, LM92, Lug97, MWCP94, NPM98, SLNP95, SB97, Wel90, ZG99]. Old- [NPM98]. Old-Growth [BS92, BRHD97, CC99b, JR99, Lug97, MWCP94, SLNP95, SB97, Wel90, ZG99]. Older [WGO95]. olive [SAS95]. Oliver [Gol97b]. Olympic [Hut95a, Lym94, Sch93a]. Oncorhynchus [GL95, NTT94, UA94]. One [DM90a, Gat95, MA96, Kau95b]. One-horned [DM90a]. One-Migrant-per-Generation [MA96]. Only [Mef99b]. Onslaught [Sou90]. Ontario [BRHD97, FH97]. Ontogeny [Fur95]. opciones [LP96, Sim96, Tho93]. Open [MC98, ST96]. Operations [UKG^{+}96]. Ophiopogon [GZWR97]. Opler [BBM94c]. oportunidad [Mec95]. Opportunities [LT92]. Opportunity [Mec95]. Options [HBD90, LP96, Mar93, Sim96, Tho93, WW98b, Woo99]. oralis [JDEB98]. Orangutan [Yea97]. Orchid [Lin95, ST95]. Orchids [SP98]. Oregon [BHS95, RLR97, Sch98]. oregona [SSTL92]. organismos [MM92]. Organisms [Cob90, CDV93, Lug90, MM92]. Organización [VC99]. organizadas [JBB96]. Organization [FS92, VC99]. Organized [JBB96]. Oriental [Tsu98, ACB^{+}99]. Oriente [RW99]. origen [MM94, Sun96]. Orígenes [LM93]. Origin [MM94, Sun96, WRG98]. Origins [Gil96, Gol97a, LM93]. Orioles [LBR99]. Ornithological [GD99b]. Ornithology [Gat95]. Ornitológica [GD99b]. orquidea [ST95]. orquídeas [Lin95]. Orthoptera [KA92]. Oso [MWM^{+}99, MHB^{+}96, Pri96b, Doa95]. Osos [MHWP96, PZSG96]. Ostras [MLP^{+}97]. Other [Bal99, Bro96, Sou90]. otras [Bro96]. Otter [EK98, Gar97, RDE96, SGR^{+}98]. Otters [BBCS99]. Ouborg [PH96]. Our [Alt90, Ano99k, Day91, Orr94a]. Outcrossing [RVK^{+}94]. Outplanting [Hel98]. Outreach [Ano99m, GP98, PAG^{+}99]. Outstanding [Baw95a, Los95]. Oveja [Weh99]. Ovenbird [GF90, KGD96]. Ovenbirds [PBG^{+}93, PF99]. Overabundance [GWW93a]. Overcoming [FWB^{+}99]. Overconsumption [Bal99]. Overharvest [SGA98]. Overview [HR93]. Overwinter [CPN95]. Overwintering [WRM^{+}91]. Ovoposición [KB97]. Owl [LMNV92, LNVM94, SGMP99, TG93, WBKH97, WM91]. Owls [Bar95, HSD93]. Own [Uhl98a]. Oxygen [CKCM95]. Oxymoron [Orr90a]. Ozark [HL99, KF95].

Pacifastacus [LEMC95]. Pacific [ABB^{+}97, Bot94, BHS95, CEPR91, CAB^{+}98, DW93, Fri93, GD99a, GK99, Guz91, HSS94, Mef92, MGM^{+}98, PERC96, SGW90, WW98a, WT90, Wap94]. pacífica [DW93]. Pacifico [GK99, ABB^{+}97, Fri93, GD99a, PERC96]. Paddlefishes [Bir93]. Paige [OG96]. Paisaje [BSR97, CFL96, DBH^{+}99, FCM99, HNC97, HRCTS99, KSO^{+}99, LCPM97, MWM^{+}99, MN96, MR99, Pea97, SS99b, WMB^{+}97, Wit97a, Bak94, BY96, CUM96, Haa95, HH95a, KGD96, LMNV92, MW93, ML94, MWCP94, MSHW95]. Paisajes [MH99b, PF99, TA97, WK99]. paisajístico [BLD^{+}95]. paises [BL95a, SK96]. Pájaro [LS96, RWES93, HBA93, RHSM94]. pájaros [Ama93, KALG94, MH93, RDN94, Thi95, Tho93, Tsc92, VHM94]. Palazzo [Vid94]. Paleontological [Dud97, Sar95]. paliata [Sto96]. palliata [Sto96]. Palm [Kin92, OK96, RSR96]. palmera [OK96, RSR96]. palomas [SB94]. Palung [Sal93]. palustris [NPM98]. Panama [Kar90a, Lip99, Rob99]. Pandas [LOT^{+}99, YH91]. pantanos [LRD^{+}96]. Panther [Hed95, KLF98, Mae90, MC95a]. Panthera [MSO96]. Panthers [MC95b]. Papel [GK99, PDS96, Ree99, Ber95b, MW93]. Paper [KM97, OG96]. Papers [Ano95a]. Papua [MHL97, Sim94]. para [Aco99, ABB^{+}97, BMLW96, BO96, Bei93, BSW^{+}96, Ber97a, BTRM99, BHM94, BBCR99, BJ95, Bun95, Bur95, Bus96, CZN99, CAH97, CCOO^{+}96, CCR96, CS97, CBAT97, DHW99, EDFM96, FH95, GG99, GWM^{+}97, GMG94, GB95b, GD99b, HBA93, Har95a, HP94, HC96, Hil95, HW95, KASR99, KGS93, Ker97, KMW95, KTA94, KMM94, Lam97, LNVM94, LWMG93, LRD^{+}96, LA95, LC99, LP96, LOT^{+}99, LSC^{+}96, Loc99, LNPC^{+}95, MB99, MW93, MH99b, MPC99, Mer94, MP99, MHB^{+}96, MBD^{+}99, MWCP94, MSHW95, OSEC95, OFH94, OO93, OB93, PT95, PNS99, PB95, PSGS94, PL95, PK94, PGW97, QLDV96, RS95a, RTNK99, RK93, RBPB96, Rob95c, RS95b, RRS96, RCMR97, SM99, SB99, STA^{+}99, ST94, SFM96a, SFCM92, Sim96, ST95, SM97, SL93, SSEM94, Sto96, Str99, SB95c, SB94, Tay95]. para [TV99, TB95, TA97, TTW^{+}94, VM99, Ver93, VRV^{+}94, War94a, Wei99, WMB^{+}97, WGM^{+}96, WK99, You94b, RDE96]. paradas [RCHP94]. Paradigm [Jop94, Sal90]. Paraguay [HPB^{+}97, KGS93]. Parasite [CD98, Ebe95, GW98]. Parasites [LV94, SGW90, Sto96, Win95]. Parasitic [MS97a]. Parasíticas [MS97a]. Parasitism [HH95a, SW99b, TB93, Woo99]. parasitismo [TB93, HH95a]. parásitos [LV94, Sto96]. parche [LNVM94]. Parches [CFL96, DBCM95, Haa95, QAM96]. Pardo [MHB^{+}96, Pri96b]. Pardos [MHWP96]. pardus [MSO96]. Parents [CC95]. Pariente [GWH97]. Park [Boz93, CK98, Dea95, FYP96, Jha93, Mad99, MKA96, MBT90, NT98, RS91a, Sal93, SMPN99, TW92b, VMMK92, Wes90, Yea97, AM99, AB90, BWS97, FT96, Gei90, GL95, KRG^{+}99, MM94, MG97]. Parker [For94]. Parks [Ano99l, BTG93, CM93, HH96, Hei95a, HS95, HY93, New95, New96, PB94a, Sim95, WMH94, AS94]. Parochialism [Dud95, HH94c]. Parque [MKA96, AM99, BWS97, FT96, FYP96, GL95, KRG^{+}99, Mad99, MM94]. Parques [Boz93, New96, SMPN99, CM93, WMH94, AS94, New95]. Parrots [Gil96, SG98, WKS^{+}94]. Part [Blu93, LF97, McC96, Dun97]. Parte [LF97]. partición [PL95]. particular [YJG^{+}96]. Partir [Gar97, ART94, BGF95, Fol94, QAM96, Rob95c, RC94, TM96]. Partitioning [ML90, PL95]. Partners [Gia93]. Partnerships [KM98, OB97]. Partridge [KASR99]. pasado [Boz93]. Paserina [SPR97]. Paserinas [DH97]. pasos [Pre94]. Passeriform [MS97b, MS99c]. Passerine [KHS98, KR95, Meh94]. Passion [Fle90]. Past [Boz93, Fis94, WCD97]. Pastisales [Wei99]. Pastizal [War94a]. Pastizales [FS97, Mor99, PS99, WF99a, BB93, VHM94]. Pastoral [Sab96]. Pastoreo [BB97, BB99, RF99, Sab96, BB93, Fle94a]. pasturas [CUM96, CD95]. Pastures [CUM96, CD95, PS99]. Patas [EKL^{+}99, RK93]. Patch [BABK98, CFL96, DBCM95, HHR99, LNVM94, QAM96, SLNP95]. Patch-Level [QAM96]. Patch-Scale [BABK98]. Patches [Ano97e, Haa95, MCO98]. Pathogenic [KB97]. Pathology [HM96]. Patógenas [KB97]. patología [HM96]. Patriarch [Wil93a]. Patrón [Cow99]. Patrones [CFL96, RB97a, WF99a, BPS95, Bun95, DPOO94, Gre96, Sha93, AMGJ95, BLD^{+}95, CB95, DS95, Goe97, SWK^{+}97]. Pattern [CHK98, Cow99, Fai96]. Patterns [AMGJ95, BLD^{+}95, BPS95, Bun95, CB95, CFL96, DS95, DPOO94, DC98, Goe97, Gre96, HB95, KR98, LFM^{+}98, NT98, PC92, PC98, RB97a, RT91, SWK^{+}97, Sha93, WF99a]. Pavements [MNU94]. pavimentos [MNU94]. Peces [GW97, HRV^{+}99, JRP99, JHHB99, Ang95, CCOO^{+}96, CBLV94, Fri93, HGNK93, KO93, LM93, LNPC^{+}95, MKM^{+}96, OO93, QLDV96, Rei93, Rob95b, WH95]. Pedagogy [Orr93a, Orr97]. Pedigrees [Mil95a]. Pelagic [Ang93]. Peligro [BBCR99, GDK99, GWM^{+}97, GWH97, MFR97, Mor99, PD97, PS96, SHMV97, STA^{+}99, TSHG95, BEG^{+}94, Bro96, Can96, CCOO^{+}96, CBLV94, DS92, DMMG96, GB95b, HMPK96, HT95, LCSC95, LM95, OO93, PNH93, QLDV96, RSR96, Rei95, RHCB93, RHSM94, SDB^{+}96, SAH^{+}95, Was96, WMW93, YJG^{+}96, DME95, GD99a]. penalidades [LWMG93]. Penalties [LWMG93]. pendiente [BBM94b]. Pending [BBM94b]. Pendleton [CM94]. Peniocereus [SNP94]. Penises [MWS^{+}97]. Pennock [Wap98]. Pennsylvania [PBG^{+}93]. People [Ano99l, Dun97, Fit94, HD99, HY93, NMGS94]. Peoples [Alc93, McL93]. pequeña [HBA93]. Pequeñas [AHMM97, PC93]. Perameles [Rob95c]. Peramelidae [SMGB91]. perca [GWW93b, OO93]. Percentage [Men91]. percepción [HDE95]. Percepciones [Kau93b]. Perceptions [HDE95, Kau93b, Kel91, Kel93a]. Perch [GWW93b, OO93, SKOO98]. percha [MW93]. Perches [MW93]. Pérdida [FM96, GSB97, RKLS99, TCO^{+}96, WSE97, Ang95, DP96, Rob93b, Was96]. pérdidas [KTA94]. perennes [BB93]. Perennial [BB93]. Perfective [Rex93]. Performance [PBG^{+}93, Tho94c]. Perhaps [Mor94]. Periféricas [LSM99, LA95]. Perils [Nos95b]. Period [DH97]. Periodic [WER^{+}97]. Periodo [DH97]. Periods [GS93]. Peripheral [FAP95, LSM99, LA95]. pero [HCB94]. Perro [Mad99, GAC^{+}95, MCR94]. Perros [VC99, CC96]. Persistence [AHMM97, Ber90, Ber93, Ber99, BWS94, Goo94, HMPK96, KEL93b, MCW98, RP99, SP98]. persistencia [BWS94, HMPK96]. Personal [Dra94, Pee96]. Perspectiva [RLD99, BH93]. Perspectivas [AS94, KS94b, CBLV94]. Perspective [BH93, BABK98, Got95, LNA95, Lym94, MBJ^{+}97, NKA98, Pee96, RLD99, Sar95, Sch93a, Str98, WR96]. Perspectives [AS94, BM97, CBLV94, Del99, DG93, Dud97, HHF^{+}97, Hei93, KS94b, LM92, Sal91, Sal92, Uhl98b, Vid93, FNN92]. Perth [RS91a]. Perturbación [ACB^{+}99, LC99, DDM93, Fim94, ML94, NHA95]. perturbaciones [PC94, RS95b]. Perturbation [DDM93, EV98]. Peru [PVZJCI90]. Pesadillas [Yaf97]. pesca [Ber95b]. Pescador [CZN99]. Pesquerias [HC96, MKA96, MPC99]. Pest [NR97]. pesticidas [TB95]. Pesticides [TB95]. Petén [SDT93]. Petroica [AMA^{+}94]. Pez [EE97, GK99, LSA99, LDSH99]. pH [LSS95]. pH/UV [LSS95]. pH/UV-B [LSS95]. Pharmaceutical [Sch92]. Pharomachrus [PB95]. Phenotypic [BC94]. Philippi [KA92]. Philippine [HC99]. Phillips [DMD^{+}92]. Philosopher [Caf95]. Philosophy [Bis93]. Phlox [HCBH99, HCBH99]. Phoca [SRV96]. Phocidae [TSP97]. Phoenix [Kin92]. Photography [WRM^{+}91]. Phragmites [Tsc92, Tsc92]. Phylogenetic [Fai96, HD96, UA94, WOA98]. Phylogenies [JRP99]. Phylogeny [Lin95]. Phylogeographic [DLP^{+}92, MSO96]. Phylogeography [LBR99, WTWS98]. Physiological [WBKH97]. Phytoplanktivorous [GWW93b]. Picture [Sad95]. pictus [GAC^{+}95]. piedra [MNU94]. piernas [BHKH96]. Pigeon [YAM^{+}98]. Pigeons [SB94]. Pigs [Cla92]. Piligro [CBAT97]. pilosa [HCBH99]. Pilot [PCG91]. Pinar [DBCM95]. pinares [MM92]. Pinches [CRJ^{+}94]. pincipios [MWCP94]. Pine [BRHD97, DBCM95, GNH94, MM92, MPB96, NPM98, SB97, TCK^{+}95, Zei92].

Pinniped [MWS^{+}97]. Pino [BRHD97, MPB96]. Pinus [NPM98, TCK^{+}95]. Pinzones [GG97]. Pioneering [Sam96b]. Piping [RRM93]. piquituertos [Ben93]. Pirineos [Deh96]. Pisoteo [MFR97]. Pithecellobium [HWB96]. Plagas [NR97]. Plain [LSC^{+}96, LCPM97, Sha93, YAM^{+}98]. Plains [MBT90, RRM93, SAS95, Sta98]. Plan [LDP95, PAG^{+}99]. Planeación [BN97, SHM97]. Planes [TSHG95]. planicie [LSC^{+}96, Sha93, LCPM97]. Planicies [RRM93]. planificación [HW95, KB94]. Planners [Ano99f]. Planning [BN97, CDB99, Cor96, HW95, KB94, KM98, KCE^{+}93, PC98, SCAB91, Sim94, SHM97]. Plans [TSHG95]. Plant [BEG^{+}94, BL99, BLS99, BS92, CK98, CUM96, CT99, DP96, GJH96, GMF99, GS99, Ham94, HS98b, HE91, HRCTS99, HH95b, HP99, HT95, KB94, KMR94, KS97, LSM99, LFOJ96, MFR97, Mat94, Mau97, MBH96, Men90, MBD^{+}99, Mor99, NHA95, OBY98, Par93, PNH93, PBB99, PM92, RF99, SFM96a, SNP94, TJ99, TTW^{+}94, TCO^{+}96, Wes90, Wil93a, WS97]. Planta [CT99, HP99, LSM99, MFR97, Mor99, WS97, HT95, MBH96, PNH93]. Planta-Insecto [CT99]. plantación [HNC97, Rei95]. Plantaciones [GBAR97, HNC97]. Plantas [BL99, BLS99, CC99b, FS97, GJH96, GS99, HRCTS99, KS97, MS97a, MBD^{+}99, PBB99, RF99, RH97b, TCO^{+}96, BEG^{+}94, DP96, Ham94, HH95b, KB94, LFOJ96, ML94, NHA95, PM92, QAM96, Rei95, SFM96a]. Plantation [HNC97]. Plantations [GBAR97, HNC97]. Planting [Rei95]. Plants [BV97, Dra96, DCL98, FS97, HRCTS99, HEO95, Kei98, LA92, MS97a, ML94, PAPA90, QAM96, RH97b, Rei95, VB99]. Plateada [RFM97]. Plateau [HE97]. Playero [SB99]. Playing [Che99]. Plazo [GD99b, Ren99, Har94, JFJ96]. Plea [CW97b]. Plebejus [BGK^{+}97]. Pledodóntida [HL99]. Pleistocene [Dud99]. Plethodontid [Ash97, HL99]. Pletodóntidas [Ash97]. Plots [Ash97]. Plovers [PC93, RRM93]. Pluralistic [Har93]. pneumonanthe [OLKD98, RVK^{+}94]. Poaching [KSSM95]. pobación [Har94]. Poblacioines [BC99]. Población [ADFT97, EKL^{+}99, HDE95, WA95, BBSK96, HBA93, Har95a, LRT93, MPB96, MAR96, NMGS94, PNH93, PZSG96, PK94, RL96, Tsc92, Was96]. Poblacional [BBCS99, Fra96, GSB99, GJH96, GS99, LSM99, LCA99, MHO97, MFR97, Mor99, PGW97, RLR97, RFM97, Tay95, YB96, FBS95, GMG94, HCB94, Har95a, HHH95, HP95, LP96, MBW96, NGN96, NE94, RVK^{+}94, RWES93, RHS94, RJL95, Sun96, WP96, WLAH92]. Poblacionales [GD99b, PD97, Sou99, Str99, FGG^{+}95, RWES93, SPP96, Sim96, TM96]. Poblacione [VW99b]. Poblaciones [AHMM97, Ber97a, BBCS99, BR99, CS97, FB97, GWM^{+}97, GD99b, GW97, HCBH99, KL97, KLR97, LSM99, LDSH99, MN99, SCZF96, SGMP99, SB99, SF97, Wal99b, WW99, Weh99, YBZ99, AL93, BY96, BWS94, BLF^{+}93, Bur95, Bus96, DDM93, FM94, FBFO96, HGK^{+}96, KQK93, LA95, MKM^{+}96, Mec95, New95, NE94, PERC96, PC93, PB94b, RFDB94, RHS94, SFM96a, SFLC95, Tho93, TB93, GBAR97, WS97]. Pobremente [McK99]. Pobres [Wei99]. Pocket [SCZF96]. poder [TG93, Str99]. podrían [Fim94].

Poecílido [SHMV97]. poecilidos [LV94]. Poeciliopsis [QLDV96]. Point [Cla99, Gre92, Sou91]. Poland [OR94]. Policies [LWMG93, TFHD98]. Policy [Bis94a, Bla94, CAWH91, Mef98b, Mef99b, PC96, Yaf97]. polillas [LRD^{+}96, Tsc92]. polinización [ST95]. polinizador [Was96]. Polinizadores [AHMM97, RKLS99]. Polioptila [AA97]. Political [CKB98, STM95]. Políticas [Yaf97, LWMG93, Sab96]. Político [PC96]. Politics [Orr91b, Sab96]. Pollination [MCW98, ST95]. Pollinator [Ano98d, RKLS99, Was96]. Pollution [CBCC93, OR94]. Pollyannas [Nos95b, SMS97]. Polonia [OR94]. Polymorphism [Avi95, Hug91, MH91, NTT94, Ran96]. Polynesian [FS91]. Polynyas [FS92]. Pond [Gra95, Sem98]. Pond-Breeding [Sem98]. Ponderosa [SB97]. ponerse [MW93]. pool [Ing90]. Poor [SFCM92, Wei99]. Poorly [McK99]. Poplar [HNC97]. Popularity [WMKF98]. Popularization [And94]. Population [ADFT97, Ano90o, Ano97c, Ano99g, Ano99i, AL93, Ber97a, Bil91, BDSD98, BBL93, BBCS99, BAR^{+}93, BBSK96, BBM94c, BR96, Bus96, CA92, Cha95, CG98, Ehr94b, EKL^{+}99, EB99, FBS95, Fra96, FCR^{+}98b, FGG^{+}95, GSAZ98, Geh96, Gil96, GMG94, GSB99, GJH96, Goo94, GS99, GD99b, HBA93, HCB94, Har95a, Har94, HS97, HH94b, HHH95, Hel98, HP95, HFS98, HM98, HDE95, Hun98, JDEB98, JB99, KSSM95, LR91, LRT93, LSM99, Leb93, Leb98, LP96, Lip99, LCA99, MW98, MHO97, MFR97, MCW98, MPB96, MEE93, Mef94, Men90, Men91, MBW96, MAR96, Mor99, MFW90, NGN96, NE94, OLKD98, PPR^{+}91, PNH93, PD97, PZSG96, PGW97, PB94b, PSB98, RVK^{+}94, RLR97, RWES93, RB97b, RFM97, RB95b, RSMB99, RHK99, RHS94, RL91, Rym94, RJL95, Sal99, San99, SPP96]. Population [Sha90, SBB^{+}94, Sim96, SM91, Son96, Sou99, SLS92, SG93, Str99, Sun96, Tay95, Tho90, TM96, Tho93, TJ99, Tri98, Tsc92, VW98, WP96, Wal90, Wap98, Was96, WLAH92, WA95, WL97, Woo99, YB96, RL96]. Populations [Agu94, AHMM97, AMW90, AF98, BBHM95, BBGB92, Ber90, Ber99, BC99, BWS94, BBCS99, BLF^{+}93, BTG93, BMR^{+}92, BR99, Bur95, Bus96, CS97, DDM93, FM94, FMF95, FFL95, Fle94b, FB97, Fra98, FBFO96, FAP95, GGB^{+}96, GF90, GWM^{+}97, Gra95, GBAR97, Gru90, HMW98, HGK^{+}96, HCBH99, KL97, KQK93, KO91, KLR97, Kra95, LSM99, Leb91, LA92, LA95, LC98b, MS92, MKM^{+}96, Mec95, MN99, New95, NY91, Nos98, NE94, PGFD95, PEP98, PDSB98, PERC96, PSSB95, PC93, RFDB94, RB95a, RM97, RR99d, SCZF96, SGMP99, SMGB91, SFM96a, SW94, SFLC95, Son96, Swa98, SF97, TB93, VL91, VW99b, Wal99b, WW99, WDWG94, Weh99, WS97, YBZ99]. Populus [SAS95]. Por [MA96, Bak94, BC97a, BT94, CDH95, DH97, DHE97, DHW99, EKL^{+}99, FC93, FP95, Fra97, FEM95, Gel99, GWW93b, Haa95, HOF^{+}96, HGNK93, KHM99, LS96, MVH96, MLD97, MB99, MBD^{+}99, OO93, PK94, RH93, TA99, TSHG95, TWMM97, WP97, Yaf97]. Porcupines [SJB97]. Porphyrio [BJ95]. Porpoises [RKBP93]. Porque [STA^{+}99]. portorniqueñas [WKS^{+}94]. Posed [TW92a]. posible [OB93]. Positions [Ano90n]. Positive [CPC96a, PGFD95]. Positivism [RP99, SF96]. Possible [OB93]. Possum [LP96]. Post [DDM93, DH97, MB99, TCK^{+}95]. Post-Fire [TCK^{+}95]. Post-Fledging [DH97]. Post-Hatchling [MB99]. Post-Juvenil [DH97]. post-perturbación [DDM93]. Post-Perturbation [DDM93]. posterior [TCK^{+}95]. Postfledging [SB94]. Postmodernism [Sar96]. Pot [Pri96a]. potaninii [GWH^{+}99]. Potato [BKO^{+}95]. Potencial [BWS97, Ama93, DS95, HT95]. potenciales [Hil95, LDP95, WKS^{+}94]. Potential [Ama93, Ano98d, BWS97, CKCM95, DS95, FS91, Hil95, HT95, Joh94b, LDP95, Par93, PCM98, Rod92, Ste98, SFM96b, WKS^{+}94]. Pots [RCMR97]. Poverty [Baw92]. Power [CKB98, HS97, RB95a, RB97b, Sab97, Str99, TG93, TD96, Tho97]. Práctica [BC97b, PQL99, SM99, Har95a]. Practical [Cla99, SM99]. Prácticas [HL99, FS95]. Practice [BC97b, FWB^{+}99, Har95a, Jor92, Orr99b, PQL99, RMJ92, Spe95, WABH95]. Practice-based [BC97b]. Practices [FS95, Fri92, HL99]. Pradera [MN96, PS96]. praderas [How94, MCR94]. Pradesh [Sab96]. Pragmatic [Pri96b]. Pragmático [Pri96b]. Prairie [BCL^{+}98, HCBH99, How94, Men91, MCR94, MN96, PS96, Sta98]. Pre [DD97, KL97, RGSW96]. Pre-1940 [RGSW96]. Pre-colonización [KL97]. Pre-Columbian [DD97]. precaución [Hes94]. Precedente [EKL^{+}99]. Precipitation [PC98, Swa98]. predación [OO93]. predador [CCOO^{+}96]. predadores [Ama93, JFJ96]. Predation [BH97, HH98, Has95, HE91, Jha93, KHS98, KDG98, LV96, OO93, SW99a, SW99b, SPD^{+}98, Woo99, YS96]. Predator [BW91, BJ96, CCOO^{+}96, MP99, PGFD95, Yah96]. Predators [Ama93, CS97, JFJ96, PGFD95]. predecibles [Was96]. Predecir [Loc99]. Predicción [BHS99, SL93, KTA94, MSHW95, Fol94, GMF99, HRCTS99, JRP99, RH97b]. Predicciones [MHB^{+}96, RS95a, Cow99]. Predice [BPC97]. Prediciendo [ML94]. Predict [Loc99, MBW^{+}98]. Predicted [Was96]. Predicting [BHS99, Cow99, Fol94, FCR^{+}98b, GMF99, HRCTS99, JRP99, ML94, MS97b, NT98, RH97b, SPD^{+}98]. Prediction [KTA94, MSHW95, NMV^{+}98, SL93, TJ99]. Predictions [MS99b, MHB^{+}96, RS95a]. Predictive [FP95]. predictivo [FP95]. Predictor [GS99, NMGS94]. Predictors [MGM^{+}98]. Predicts [BPC97]. Prefer [HH95a]. Preferencia [HH95a]. Preocupaciones [DHW99]. Presas [GK99, SB99, VPW97]. Presence [BE94, CZN99, Str99]. Presence-Absence [CZN99, Str99]. Presencia [CZN99, Str99]. Presencia-Ausencia [CZN99, Str99]. Present [Boz93, Fis94, HEO95, RBMA98]. presente [Boz93]. Preservación [SMV96, Lug94]. Preservation [BAR94, Bra98, HEO95, Lug94, Pow91, SMV96]. Preserving [HBD90, LA92]. Presettlement [KL97]. Press [BBM94a, Ehr94a]. préstamos [WSB92]. Presupposition [Hei95b]. Prevalence [Sto96]. Prevention [NMV^{+}98]. Prey [BW91, GK99, LV96, PDFG96, SB99, SPD^{+}98, VPW97]. Primaria [BC97a].

Primario [JR99]. primarios [LLP95, Lug94]. Primary [LLP95, Lug94]. Primate [ABM91, Cow99, KO91, Med93, Str91]. Primates [Cow99, Fim94, Kin92, Med93]. Primavera [RLR97]. primaverales [BHM94]. Primrose [WS97]. Prímula [WS97, Was96, WS97]. Principles [LN93, MWCP94]. Prioridades [BMLW96, CBAT97, DHW99, BSW^{+}96, Lin95, KASR99]. Prioritarias [PNS99]. Priorities [BMLW96, BSW^{+}96, CBAT97, Fai96, Gin99, KASR99, KMR94, LM92, Lin95, Mas91, MQ99, MMT^{+}98, Par95, TB98, WDR^{+}98]. Prioritization [KSC^{+}96]. Prioritize [DW93]. Prioritizing [ABB^{+}97, LSC^{+}96, WW98a]. Priority [DHW99, PNS99]. Priority-Setting [DHW99]. priorización [LSC^{+}96, ABB^{+}97, KSC^{+}96]. priorizar [DW93]. privada [Can96]. Private [Can96, Kni99, Lan96, Mae90, MB95]. Probabilidad [BER97b, MNU94]. Probabilistic [BHS99]. Probabilísticos [BHS99]. Probabilities [Bot94]. Probability [MNU94, Ree96]. Probando [FH95]. Problem [CCR96, Doa95, Geh96, Orr93d, OE95, Pel92, Roj92, TD94]. Problema [CCR96, Doa95]. problemas [BLF^{+}93]. Problems [BLF^{+}93, Orr93d, Par93, Tem94, Vid93, Wes90]. Proboscidean [Ber95a]. Proceso [PC96]. Procesos [Ree99]. Process [Kar90a, Mef98b, PC96, Tsu98]. proclives [Ang95]. Producción [Mor99, JBB96]. Produce [Nos97]. Producing [BDSD98, Dun99]. Product [Pel92, SDT93]. Production [JBB96, Mor99]. Productividad [PS96, TTJ^{+}99]. Productivity [PS96]. productos [LLP95]. Products [God93a, God93b, LLP95]. Professionalism [Orr94b]. Profitable [Phi94]. profundas [BL95b]. Program [DHBG97, GW98, JS93, DSGS93]. Programa [DSGS93]. Programas [KHM99, KMM94]. Programs [BBHM95, CAWH91, Hug91, JVM95, KHM99, KMM94, MJ96, MH91, Mil95a, Tem93a]. Progresos [CJ93]. Progress [CJ93, Sto95]. Project [PCG91, BCHM94]. Projects [GG99, SM99, SS99a, SHS^{+}97]. Prolegomenon [HH94a]. Promote [TA97]. Promover [TA97]. Promoviendo [PC93]. Prone [Ang95]. Proneness [Lau91]. pronosticador [NMGS94]. pronunciadas [You94b]. propagación [DMMG96]. Propagation [Ano99i, DMMG96]. propensión [PTTN^{+}96]. Propensity [HS98a]. Properties [SGR^{+}98]. Property [DME95, HEO95, MB95]. propiedad [DME95]. Proportions [GMSP98]. Proposal [BBS93]. Proposed [IU92, RK93, TW92a, WMH94]. Proposition [GS95]. propuesta [RK93]. propuestas [BJ95, KB94, WMH94]. propuestos [WMH94]. Prospect [Orr94b]. Prospects [Ryd93]. prostrata [HP99]. Protección [DBH^{+}99, STA^{+}99, GF97, PT95, BLS99]. proteccionismo [Yu94]. Protect [CS97, RS95b]. Protected [AM99, Hei95a, Hei96a, HWG99, KMW95, LKM90, NMGS94, NG92, SAQMGQ^{+}99, VRV^{+}94]. Protecting [Bra98, Goo95, Roj92]. Protection [Bar96, BLS99, DBH^{+}99, GF97, GN92, HS96b, KMR94, MFR97, PG91, RS92, Scu99, STA^{+}99, TB95, WF99b]. Protectionism [Yu94]. Proteger [CS97, CCOO^{+}96, RS95b]. Protegerla [MFR97]. Protegidas [SAQMGQ^{+}99, KMW95, NMGS94, VRV^{+}94]. Protein [SS95]. Protejidas [AM99, HWG99, NE94]. Protocol [MBD^{+}99]. Protocolo [MBD^{+}99]. Proveen [LCA99]. Provide [BN98a, LCA99]. Provides [Tom92]. provisto [LS96]. Proximate [Nab94]. Proximity [KHS98]. Proyectos [GG99, SM99]. Prudent [Mar93]. Pruebas [BBCR99]. Przewalski [Mil95a, Ryd93, VW96]. przewalskii [VW96]. Pseudomys [JDEB98]. Pteropodidae [FT91]. Ptilimnium [KMR94]. Ptilochronology [YG92]. Public [Ano99m, BBS93, Can96, GP98, Hun93, JM97, Mef99b, Orr91b, Tem93a]. pública [Can96]. Públicas [JM97]. puede [PM92]. Pueden [RS93c, SDT93]. puerro [NGN96]. Puerto [Lug94, WKS^{+}94]. Pukeko [BJ96]. pulchra [KR98]. pumas [Bei93]. punctatus [CDH95, TW92a]. punto [LNA95]. puntos [WGM^{+}96]. Pupfish [EE97]. Puppet [VM99]. Puppet-Rearing [VM99]. Pupping [GG90a]. Pureza [GMMR^{+}96]. Purging [FNC98]. Purity [GMMR^{+}96]. Purple [KR98]. Purpose [Dal93]. Puting [Yea97]. Putting [Spe95]. PVA [WABH95]. Pyrenees [Deh96].

Quabbin [GF97, GF97]. Quality [BD92, RCHP94, SAM98]. Quantifying [And91]. Quantitative [CRJ^{+}94, GD99a, HMPK96, KR98, PGR^{+}94]. que [BGTB93, Bun95, FM96, FGG^{+}95, KMW95, LN96, MHB^{+}96, NMGS94, WF99a, Yaf97, YJG^{+}96, Gru94, Gru97, Hig97, WMW93]. quebrantahuesos [Bus96]. Queensland [BC97a]. quemada [TCK^{+}95]. Question [Orr90b]. Questions [Bra94]. Quo [Str96]. Quotient [TB95].

Rabbits [PGFD95]. racemosa [GWH97]. radiación [BHKH96]. Radiation [AKG98, BHKH96, LP98]. Radical [Pis94]. Radiocollaring [CCM97, EHBE97]. Radish [KEE91]. Rail [FFL95]. Rails [HBD90]. Rain [AR97, ART94, BM92, BM98, GS93, Hil95, HEO95, Lau91, LMS96, RTS90, SDT93, Thi92, Thi95, TCO^{+}96, TWMM97, VAPCO98, Wik90]. Rainbowfish [ZDM98]. Rainforest [HC99, KO98, MB95, Pel92, Phi93, WLDGGZ98]. Rainforests [HS98a, Tre93, WDSA94]. Raised [SG98]. Rajasthan [JG91]. Rallus [FFL95]. Rana [HM97, LDSH99, BHKH96, PC94, BTG93, KB98, MGM^{+}98]. Ranas [GK99, KSO^{+}99, LMS96]. Ranching [KWR95]. Range [AFF^{+}97, BGTB93, FC93, FGG^{+}95, FAP95, GB96a, LC98a, MH93, PWC^{+}98]. Rangelands [BBS93, BM95, Dud97]. Rango [MH93]. Ranking [LP96]. RAPD [RG95]. Rapid [Ber90, BMR^{+}92, Ing90, OB93, Rob95b, RH93, Weh99, WCD97]. Rápida [RH93, Weh99, OB93]. rápido [Rob95b]. Raptors [REDH98]. Rara [GG97, LSM99, MBH96, ST95, SNP94]. Raras [BL99, BLS99, GJH96, MS97a, PDS96, SG99, LFOJ96, ML94, Tho94a, WGM^{+}96]. Rare [BL99, BLS99, BGK^{+}97, CDH95, GJH96, HE91, HH94c, LSM99, LFOJ96, MS97a, MBH96, ML94, Mil95a, OLKD98, PDS96, Sch98, ST95, SW94, SG99, SP98, SNP94, Tho94a, YB96]. Rarest [Bur90, GG97]. Rareza [BL99, Goe97, DeM93, DY93, KMW95, MM92]. Rarity [ARR90, BL99, DeM93, DY93, DYS95, GB95a, GB96b, Goe97, Kat92, KMW95, MM92, Meh94, WGM^{+}96]. Raro [YB96, CDH95]. Rat [CDH95, GP97, PK94, PGW97]. Rat-Free [CDH95]. Rat-Inhabited [CDH95]. Rata [GP97, PGW97]. ratas [Ber95b, CDH95, PK94]. Rate [RVK^{+}94]. Rates [Bur95, DG92, HS96a, HH98, Has95, Kar90a, KEE91, McK99, RR99b, RC94, SP90, Swa98]. Rating [Orr91d]. Rationale [MC98]. Ratios [MJ96]. raton [Mil95b]. Ratones [WSE97]. Rats [Ber95b]. Rattan [Sie93]. Rattlesnake [Wei92]. Rattus [Goo95]. Rawat [Sab98]. Razorback [DMMG96, MBW96]. Re [LC98a, MS93, WGRT96]. Re-análisis [WGRT96]. Re-examination [MS93]. Re-Introductions [LC98a]. Reader [Gru96]. Reading [Bis94b]. Real [Avi94, KM97, Str95, Tho90]. Realistic [TB98]. Realities [Ber91]. Reanalysis [WGRT96]. Reared [MJ97]. Rearing [PC93, VM99]. Reasons [O’C92, Roh91]. Reassess [Ham94]. Reassessment [Ano99i, Mer94]. Reauthorization [MWN^{+}94]. Rebuttal [GW91]. Received [Ano90f, Ano91g, Ano93b, Ano96b, Ano97b, Ano99d]. Recently [LC98b, Ver93, WGO95]. recientemente [Ver93]. Recientes [Ree99, FS95]. Reciprocal [GG99]. Recíproco [GG99]. reclinata [Kin92]. Reclutamiento [JR99]. Reclutamineto [MJ97]. Recognition [MSO96]. recolección [KQK93]. Recolonización [LSA99]. Recolonization [LSA99]. Recommendations [RS91b]. Reconnaissance [Pri91]. Reconocimiento [MSO96]. Reconstructibility [GFA90]. Recounting [Bla94]. Recover [DM92]. Recovering [GGB^{+}96, Mec95]. Recovery [Can96, DHBG97, DRP98, GW98, HW95, KM98, MKA96, PNH93, Pet99, RHSM94, Roy95b, STM95, SDB^{+}96, TSHG95, Wap94]. Recreacional [HRV^{+}99]. Recreational [HRV^{+}99]. Recruitment [CC95, JR99, LFM^{+}98, MJ97]. Recuperación [TSHG95, HW95, Mec95, PNH93, RHSM94, SDB^{+}96, Can96, MKA96]. Recur [Yaf97]. Recurrentes [Yaf97]. Recursos [WP97, HM96]. Red [BEG^{+}95, BHKH96, Bro96, HBA93, HHR99, KHM99, Kei98, KB98, LDSH99, Mil95b, Now92, NF98, PH92, RWES93, RGSW96, SLS92, Wal90, War94c, WRG98, YH91]. Red-Backed [HHR99, Mil95b]. Red-Cockaded [HBA93, RWES93, SLS92]. Red-Legged [KB98, LDSH99, BHKH96]. Red-Listed [BEG^{+}95]. Red-tailed [Wal90]. Redd [RM97]. Redes [MPC99]. redevelopment [Pat91]. Redford [SGA98]. Redstart [Hun98]. Reducción [LRT93, SM99, MKM^{+}96]. Reduce [MPC99]. Reduced [JR99, PP93, VB99]. Reduced-Impact [PP93]. Reducir [MPC99]. Reduction [GLFM98, LRT93, SM99]. Reductionism [Nab95]. Reductions [MKM^{+}96]. Redundancia [JFJ96]. Redundancy [JFJ96, Nae98, RGH^{+}99, Wal92]. Redwood [ZG99]. Reed [Tem93c]. Reef [Cob97b, Cob99, JJ99, JRP99, MBW^{+}98, MKA96]. Reefs [Guz91, Rob95a]. reevaluación [Mer94]. reevaluar [Ham94]. Reevaluation [Has95, ML91, Sta98]. Reference [CM93, Hun96b, Par93]. referencia [CM93]. Refining [Pai95]. Reflections [Gru97]. Reflects [Nae96].

Reflexiones [Gru97]. reforestación [Deh96]. Reforestation [Deh96]. Refuge [Cur93, Har92, KHP95, PZSG96]. Refugee [Gol97a]. Refuges [SS94]. Refugia [CKCM95, CCOO^{+}96, Sam90, WER^{+}97]. refugio [BHM94, KHP95, PZSG96, CCOO^{+}96]. Refugios [WER^{+}97, Cur93]. Refugium [BHM94]. Regarding [HDE95]. Regeneración [GFR^{+}99, RB97a, TCK^{+}95]. Regeneration [GFR^{+}99, MS93, RB97a, TCK^{+}95]. regilla [AKG98, MGM^{+}98]. Regime [FC99]. Régimen [FC99, FS95]. regímenes [BBSK96, Bun95]. Regimes [BBSK96, Bun95]. Region [BWS97, Bra98, CZN99, DW93, GF97, MS93, NPM98, RS92, BWS97, GBAR97, DW93, MSHW95, GF97, MSHW95, CZN99]. Regional [DF96, KS97, MSHW95, PL94, SHM97, SB95c, Tri98, SB95c]. regionales [PL94]. Regiones [Ash97, BL95b]. Regions [BL95b]. regla [MA96]. Reglas [SM97]. Regular [RC94]. regulares [RC94]. Regulating [Gra99]. Regulatory [Dud92]. Rehabilitation [Est98, Jes98, Yea97]. reina [SNP94]. Reintegrating [BKC^{+}98]. Reintroducción [RFM97, VM99, RK93, VW96]. Reintroduced [FB97, Har94]. reintroducida [Har94]. Reintroducidas [FB97]. Reintroduction [HBD90, Leb93, Ral92, RK93, RFM97, RTOB94, Ryd93, SG98, VM99]. Reintroductions [GDB96, VW96]. Rejoinder [Dud92, FSCL96]. Relación [Fra96, BB93, Cla95, EDFM96, Gre96, LRT93, PL95, RVK^{+}94, DeM93, DSB95b]. Relacionadas [BL99, PKH97]. Relacionando [DPOO94]. Relaciones [BTRM99, CD95, Kau93b, GMMR^{+}96]. Relajada [BR99]. Related [BL99, KDG98]. Relatedness [KF95]. Relating [DPOO94]. Relation [BB93, DSB95b, EDFM96, Gre96, LRT93, PL95, RVK^{+}94]. Relations [Kau93b]. Relationship [DeM93, Fra96, Swa98, WMKF98]. Relationships [BTRM99, Cla95, CD95, GMMR^{+}96, GB96a, Gia93, OLKD98, Ris99, UA94]. Relativa [Loc99, VTM99]. Relative [BJ96, GWH97, Loc99, RS91a, SW99b, SS98, VTM99]. Relativism [SF96]. relativo [PT95]. Relaxed [BR99]. Release [LV96, PDFG96]. Released [Bus96]. Relevance [ALF92, Avi95, Hea94, MH99b, SK99]. Relevancia [MH99b, Sto96]. Relevant [Wes91]. Reliability [MDW99, Nae98, RGH^{+}99, Tay95, TB98]. Relictual [Wel90]. Remaining [BTG93]. Remanentes [BBCS99, FS97, SWK^{+}97, KO93, Mil95b, NHA95]. Remarks [GB95a, HSS94]. Remnant [BBCS99, PSB98, SWK^{+}97, ZG99]. Remnants [BC97a, FS97, KO93, Mil95b, NHA95]. Remoción [CS97, JHHB99]. Remote [Vog95]. remotos [Vog95]. Removal [JHHB99, Lug90]. Removals [SG93]. Removing [CS97]. Renewable [Tom92]. reorganización [MNU94]. Repair [BHKH96]. Reparación [BHKH96]. repentinas [You94b]. Replacement [Hut95c]. Replies [FFH96]. Reply [And92, BO99, Blu93, BBM94c, BM97, BCP96, Cre97, Duf93b, Ehr95, Gas91a, Gol97b, Gol99a, MS99c, OB97, PH96, PDFG96, Pat94b, Ris99, RHK99, RJL99, Sch93b, SDB^{+}97, Wal99a, Wap98, Wil93b, Yu96]. Report [Lyd96]. Reports [Hei95a]. Represamiento [VT99]. Representation [OD98]. Representative [Blo99b, Pre94]. Representativeness [Sto97]. Representatividad [Sto97]. representativos [Pre94]. Represented [NG92]. Reproducción [BMLW96, BBCR99, CBAT97, KHM99, BGTB93, CDH95, DSB95b]. Reproduction [CDH95, OLKD98, SM91]. Reproductiva [BC99, ST95]. Reproductivas [KS97, Sim96]. Reproductive [BC99, DTFP95, GHP97, KGD96, KS97, PS96, PBG^{+}93, PV98, RB95b, RMCM97, ST95, Yos94, YAM^{+}98]. Reproductivo [RMCM97, DTFP95, KGD96, ST95]. Reproductivos [PS96, DeM93]. Reproductoras [BHS99, BSR97, RLD99, TM96]. reptil [CDH95]. Reptile [CDH95, DS95, GNH94, RN96]. Reptiles [RN96, DS95, GNH94]. Republic [Nos98]. requerimientos [PL95]. Required [VW98]. Requirements [PL95, RHK97]. Rescue [DKOW94]. Research [BB90, CP94, Col92, FDV94, Gra99, HB90, HP99, LT92, Mag92b, MH99b, Par93, Pri92, Ste98, UKG^{+}96]. Reserva [Kau93b, SB99, Rob95b, FC99, MN99, SB95c]. Reservaciones [Pre94]. Reservas [LCPM97, PT95, PQL99, Sto97, Wal99b, AL93, DS95, De 95, KB94, LNVM94, Lan96, LNA95, MNU94, PB95, Pre94, PL94, PL95, SDT93, DSGS93]. Reservation [KB94]. Reservations [Pre94]. Reserve [ABM91, BAR94, DSGS93, FC99, Gru90, HPB^{+}97, JS93, LNVM94, LCPM97, MN99, PVZJCI90, PB95, Pre94, PL94, PL95, RS92, Rob95b, Sch98, SS98, SB95c, Wal99b, Kau93b, SB99]. Reserves [AL93, Blo99b, BD92, DS95, De 95, HH96, Lan96, LNA95, MNU94, MC98, Per94, PT95, PQL99, RS92, SDT93, Sie93, Sto97]. Resident [DBCM95]. residente [DBCM95]. Residential [FEM95]. Residents [PCG91]. Residual [MCO98]. Resilience [Wal95, WPR96]. resiliencia [Wal95]. Resistance [BHKH96, Fur95]. resistencia [BHKH96, WPR96]. Resort [And96]. Resource [Ano97c, ALF92, GP98, HM96, MBM^{+}98, Tom92, WP97, WL97]. Resources [Alt94, Bal99, Baw92, Duf94, Gra99, ML90, Mar93, Wil93a]. respecto [DPOO94, HDE95]. Responsabilidades [DHW99]. Response [AA93, Bla93, BB99, Car94, Cob91, Cob99, CM94, CAB^{+}98, DMD^{+}92, Dud97, Duf93c, FNN92, Göt92, HHR99, HL99, HM93, KM97, Lit93, ML94, MEE94, NF98, O’C92, Orr90c, Pet94, RFG^{+}98, Rob93a, Roh92, Sab98, Sal92, SCZF96, SO94, Vid94, WW98a, WRG98, RS93b]. Responses [BJ96, Fra94, Tho93]. Responsibility [DHW99, Kar93]. Respuesta [SCZF96, ML94, Tho93, BB99, HL99]. Respuestas [Fra94]. Restablecimiento [RH93]. Restauración [Hig97, GG99, CJ93, CC99a]. restaurar [LSC^{+}96]. Restitución [Bak94]. Restoration [Ano99n, Bak94, CJ93, CC99a, GG99, Guz91, Hed95, Hig97, LSC^{+}96, Mat95, RH93, SS99a, SGK95, WW98b, WDWG94]. Restos [BC97a]. Restricción [Ber97a]. resultan [LA95]. resultante [RHCB93]. Resulting [RHCB93]. Results [Wik90]. Resuspension [PCM98]. Rethinking [DLP^{+}92]. Retorno [Ash97]. Retrospective [Dud93, Tho97]. Return [Ash97].

Revealed [KO93]. Reveals [Gra95]. Review [MM94, MBM^{+}98, PL94, SHM91]. Reviews [Ano90d, Ano90e, Ano91d, Ano91e, Ano91f, Ano96e]. Revised [RGH^{+}99]. revisión [MM94, PL94, DY93, Weh99]. Revisited [DY93, Dre94, Erw91, Weh99]. rezago [LS96]. Rhamnus [SW99a]. Rhino [BC94, Hut95b, LM94, Rab95a]. Rhinoceros [AMW90, BC98, DM90a, MAR96, MAS^{+}97, OFH94, SF97]. Rhinopithecus [SS95]. Rhinos [BCGL93, BCG94, Ber97a, MBKR92, RB97a]. ribereña [FP95]. ribereñas [PTTN^{+}96]. ribereño [BHM94]. ribereños [FESS94, KS94b]. ribicola [TCK^{+}95]. Rica [BBL93, BWBM99, Boz93, BJW95, CC99b, FS95, Guz91, HDE95, Hun94, PCG91, Sto96]. Rican [WKS^{+}94]. ricas [PTTN^{+}96]. Rich [HS98b, LCPM97, PTTN^{+}96]. Richness [FESS94, FH97, Gas91b, GG90b, Ker97, KSO^{+}99, ML94, NBH^{+}98, PC92, PC98, SVC96, SL93, WGM^{+}96, WRSP97]. Rico [LCPM97, Lug94]. Ridge [Ash97]. Ridley [BCHM94]. Riesgo [KLR97, RBMM97, VPW97, VC99, CB95, Cun96, KC95, Tay95, TB95]. Riesgos [MLD97, WMB^{+}97]. Rifts [Put92]. Right [Bra94]. Rights [Ano91h, Ano91m, DME95, HEO95, Win95]. rigueza [SL93]. Ringlet [ST96]. Rinoceronte [Rab95a, SF97]. Rinocerontes [Ber97a, MAR96, RB97a, BC94, OFH94]. Rio [FESS94, GW97, Med93, MN96, BHM94, FC93, Fra94, GK96, LEMC95, MBW96, RLR97, HW95, FESS94, GLFZ96, LSC^{+}96]. ríos [LNPC^{+}95]. Riparian [BHM94, DRP98, FESS94, FP95, Fra97, JHHB99, KS94b, MVH96, PTTN^{+}96, SMHK98]. riparias [MVH96]. Ripario [JHHB99]. Riqueza [FH97, KSO^{+}99, WRSP97, ML94, SVC96, WGM^{+}96, FESS94, Ker97]. Risk [Kei98, KC95, KHS98, KLR97, Mag91, Neh94, SEC98, Tay95, TB95, TW92a, VPW97, VC99]. Risks [Cun96, GFA90, MLD97, WMB^{+}97]. River [BHM94, DDM93, Fra94, GLFZ96, GW97, Kob99, Med93, MBW96, NTT94, PH94, RLR97, SS98, Bla94, BB98, CBAT97, DDM93, HW95, JDEB98, KO91, LAF93, LSC^{+}96, MN96]. Rivers [GMSP98, LNPC^{+}95]. Road [Pea92, TW92b]. Roads [RJBB96, You94a]. Roanoke [Fra94]. Rob. [SAH^{+}95]. Robin [AMA^{+}94]. Robinson [SGA98]. Rocallosas [FB97, KL96, NQH^{+}96, PC96, RJBB96, WPR96]. Rocallossas [MHWP96]. Rock [CK98, EKL^{+}99, KL97]. Rock-Wallaby [EKL^{+}99]. Rocky [FB96, GSAZ98, Hut95c, KL96, KQK93, FB97, MHWP96, NQH^{+}96, RJBB96, WPR96]. Rocollosas [FB96]. rocosas [KQK93, FBS95]. rodal [DSB95b]. Rodent [Goo95]. Rohlf [Dud92]. roja [HBA93, RWES93, War94c]. rojas [BHKH96]. Rojo [KHM99]. rojos [RGSW96]. rol [GL93b, Thi95]. Role [BAR94, CG95, Gol99a, GK99, GL93b, Mag96, MW93, MSM99, MWW99, PTK^{+}98, PDS96, Ree99, Run95, Sal99, Sta98, Str98, Thi95]. Roles [Ber95b]. Romania [LT92]. Romanian [Pri91]. Rondônia [DOSP94]. Roosting [CB98]. Rorippa [CG98]. rotundifolia [OK96]. Round [Wei92]. Round-Up [Wei92]. Roy [NF98]. Royale [WLG^{+}91]. ruficaudatus [TBL^{+}97]. Rufous [RCHP94]. Rule [MA96, SM97]. Rule-Based [SM97].

Run [BB98, Hea94, HH94b, HHH95]. Running [HVG91]. Runs [NTT94]. Rural [Baw92, PCG91]. Russia [DKOW94]. Russian [SAS95]. Russian-olive [SAS95]. Rust [TCK^{+}95]. Rusty [GD99b]. Rutidosis [YBZ99].

S [Sti97, Hed96]. S. [Sun96]. Sabana [TWMM97]. Sabanas [HD99]. Saberwal [KM97, MR98a]. Sacramento [Bla94, BB98, Hea94, NTT94]. Sage [JB99]. Saguinus [CRJ^{+}94]. Saharan [Lan96]. Salamander [HL99, MPB96]. Salamanders [AB94, Ash97, AP99, PEH93, Pet94, Pet99, Sem98, TLG^{+}98]. Salamandra [HL99, MPB96]. Salamandras [Ash97, PEH93]. Sale [Rol95]. Salmon [ABB^{+}97, Bla94, BB98, CBAT97, CAB^{+}98, Fis94, Fle94b, Hea94, HH94b, HHH95, HW95, Mef92, Neh94, NTT94, RLR97, WW98a, WT90, Wap94, CBAT97, RLR97, HHH95, HW95]. Salmones [ABB^{+}97]. Salmonid [HB94, Man94]. Salmonids [Bot94, HSS94, Tho94c]. salvaje [GAC^{+}95]. salvajes [CC96, Har95a]. salvar [SDT93]. Salvelinus [RM97]. Salwasser [FNN92]. Samoa [PERC96]. Samples [BL95b]. Sampling [AMA^{+}94, CAH97, Gre92]. San [NTT94, KALG94, MBBF95]. Sanctions [Vid94]. Sanctuary [EK98]. Sand [BR96, GNH94, MM92, MSM99]. Sand-Pine [GNH94]. Sandhills [LC99]. Sandpipers [SB99]. sandwicense [RFM97, RFM97]. Sandy [SGR^{+}98]. Sandy-Bottom [SGR^{+}98]. sapo [PC94]. sapos [Car93]. Sarawak [Ano91m]. Sardinian [VS95]. Satélite [HOF^{+}96]. Satellite [HOF^{+}96]. Saudi [RH97a]. saudiya [RH97a]. Savanna [SS98, TWMM97]. Savannas [HD99]. Savants [JD98]. Save [SDT93]. Saving [KM97, Sab97]. Sawyer [Con95]. Saxifragaceae [SSTL92]. Scale [BABK98, CDB99, De 95, Got95, HH95a, HM97, HRCTS99, KS94b, LN90, RS93c, RHS94, SCAB91, Tri98, Wea95]. Scaling [SD94]. Scandinavia [TSSB95]. Scant [IN96]. Scarce [ML90]. Scarlet [HP95, PH96]. SCB [Ano95a, Ano95c, Ano99a, Tem93a]. schauinslani [SRR95]. Scheffer [AA93]. Schindler [Win98]. Science [And94, Ano99n, BO96, BG95a, Blo99a, Con95, Fle90, Hag95, KS94a, Mat95, MHWP96, Mef99b, MM96, Orr99a, Ral95, RR99c, Riv97, TB96, Web99, Wil92]. Scientific [BGF95, Dre94, GL94, MBM^{+}98, Mur90, WP97]. Scientist [LN96]. Scientists [Ano99m, Hau95, Mef98b, Sal99, Sal95]. Sciurus [UCB92]. sclectiva [MGB94]. Scope [PAG^{+}99]. Scrub [BLD^{+}95, GNH94, MM92, QAM96, REDH98]. Scuba [HRV^{+}99]. Sea [BBCS99, BL95b, BCP95, EBV95, Fra92, Gar97, HC96, JOB96, KB99, MB99, MKM^{+}96, RDE96, BAR^{+}93, SRV96]. Seabed [WN98a]. Seabird [MPC99]. Seal [GG90a, KLR97, KGM^{+}97, TSP97]. Seals [SRR95, SRV96]. Search [Tem92d]. Searching [Cob99]. Season [GP97, PF99]. Season-Long [PF99]. Seasonal [SW99b, SB94]. Second [CC99b, NPM98]. Second-Growth [CC99b, NPM98]. Secos [GFR^{+}99]. Section [Ano90o, Ehr91, WN98b]. Secundario [CC99b]. Sedentaria [WER^{+}97]. Sedentary [WER^{+}97]. Sediment [CBCC93, MBW^{+}98, PCM98, SGR^{+}98]. sedimentaria [CBCC93]. Seed [Ham94, HE91, Men91, Mor99]. Seed-Predation [HE91]. Seedling [BM98, CC95]. Seeing [Duf93c, MBKR92]. Seeking [Mat97b]. Segmentos [PD97]. Segments [Wap98, PD97]. Seguimiento [HOF^{+}96]. Seidl [BBM94c]. Seiurus [PF99]. Selección [BR99, Ker97, PQL99, KGS93, KMW95, MNU94, Was96, LCPM97, RTNK99]. Seleccionando [RS95a]. Selected [Ano96e, FL97]. Selecting [PQL99]. Selection [BBHM95, BR99, CB98, KGS93, KMW95, LCPM97, MNU94, RTNK99, San99, Was96]. Selectivamente [CC99b]. Selective [Mar98, Mil95a, MGB94, RR99d, Thi92]. Selectively [CC99b]. Selectivity [RBMA98]. Selva [TWMM97, WGO95, Ama93, ART94, FS95, Hil95, LMS96, RS93a, SDT93, Thi95, Med94]. selvas [Hil95, SBL96]. selvático [Med93]. sembradas [HGK^{+}96]. Semi [HMW98, PS99, BT94]. semi-árida [BT94]. Semi-Natural [PS99]. Semi-Wild [HMW98]. Semiarid [BT94, CRGGFT97, JFJ96, YS96]. Semiárido [CRGGFT97]. Semiaridos [YS96]. semidesérticas [JFJ96]. Semiendemic [G96]. semilias [Ham94]. Semillas [Mor99]. Seminaturales [PS99]. seminuda [DDM93]. Semipalmado [SB99]. Semipalmated [SB99]. Sense [GC97, PC96]. Sensing [Vog95]. Sensitividad [WF99a]. Sensitivity [WF99a]. sensores [Vog95]. September [Ano96b, Ano90f, Ano91g, Ano93b, Ano96b, Ano99d]. Sequel [DM90b]. Sequences [Gol97a, PDSB98, RH97a]. Sequía [BB99]. Sequoia [BTG93]. ser [AL93, Fim94]. Seram [Mar98]. Serengeti [PKH97, PKH97]. serpentina [PDC96]. Serpiente [PGW97, LNA95]. serveros [RL96]. Services [Sal98]. Sesenta [MBBF95]. sesgada [GMG94]. Set [MHB^{+}96, RS95b]. Set-Back [RS95b]. Setting [BMLW96, DHW99, Lin95, Mas91, MQ99, MMT^{+}98]. Settled [WGO95]. Settlement [DOSP94]. Severe [RL96]. Sex [GMG94, Leb98, MJ96]. Sex-Biased [GMG94]. Sexual [CDV93]. sexualmente [GMG94]. Shade [GBAR97]. Shall [BW92]. Shan [DLY92]. Shape [Orr94a]. Sharks [MG90]. Shasta [LEMC95]. Sheep [Ber90, Ber93, Ber99, BWH90, FBS95, Goo94, KEL93b, SG93, Weh99]. Sheepshead [EE97]. Shield [Fra97]. Shifting [ART94]. Shifting-Cultivation [ART94]. Shiitake [HD96]. shivers [Kau95b]. Shock [War94a]. Shorea [MGB94]. Shorebird [SB99, SK93, SB99]. Shorebirds [Sib94]. Shores [KQK93]. Short [BCGL93, BB99]. Short-Duration [BB99]. Short-Term [BCGL93]. Shortcomings [HH94c]. Shortgrass [HS98b]. Should [AL93, LNA95, RS92]. Shouldered [SG98]. Shrike [CW97a]. Shrikes [YG92, Yos94]. Shrinkage [FSCL96, Win96b]. Shrub [YB96]. Shrubland.1 [CLM^{+}95]. Shrubsteppe [KR95, YS96]. sí [BL99]. Sibling [Gel99]. sieboldii [Was96]. Sierra [DF96, DF96, MR98b]. Sieving [MEE94]. SIG [LDP95, SHM97]. Sighting [SDC^{+}93]. Siglo [TCO^{+}96]. significación [HP94]. Significance [HP94, Kai98, VRV^{+}94]. Significant [BC94, KB99, LRD^{+}96, PSH99, PD97, RB97b, Wap98]. Significativa [PD97]. Significativas [KB99, LRD^{+}96, BC94]. Silversword [RFM97]. Silvestre [BHCB99, Fra96, Hac99, LA97, LOT^{+}99, Mad99, MN99, AS94, BT94, Cur93, EDFM96, KHP95, LWMG93, NGN96, NMGS94]. Silvestres [GSB97, MJ97, VC99, Cun96, Ham94, OSEC95]. silvicultura [GNH94, MPB96]. Silvicultural [DH98, GNH94, HL99]. Silviculturales [HL99]. Silviculture [MPB96]. Simpátrico [GP97]. Simple [MH91]. simulación [RS95a, RRM93, SRR95, BBSK96, Tho93]. Simulated [Tho93]. Simulation [BBSK96, Dia95, FNC98, MS93, RS95a, RRM93, SRR95]. simus [RF97]. sin [EKL^{+}99, WP96]. since [HS96a]. Sincrinización [BC99]. sinensis [Sun96]. Singapore [TTW^{+}94]. Singapur [TTW^{+}94]. Sink [Doa95, DLK^{+}95]. Sinks [DTFP95, MS99a, Tri98]. Sistema [HWB96, Cur93, ST95, Sun96, Boz93, LSC^{+}96]. Sistemas [Vog95, DeM93, Pre94, Thi95]. Sistemática [DGG^{+}99, KB99, SFM99, SG99]. Site [Fim94, Lip99]. Site-Specific [Fim94]. Sites [GG90a, KGS93, LSC^{+}96, WRM^{+}91]. sitio [Fim94]. sitios [FESS94, KGS93, LSC^{+}96, NM97]. Situ [Ham94, TNN^{+}99]. Situación [CRGGFT97]. Six [Bes96, BWBM99, Bri95b, Roh91, O’C92]. Six-Legged [Bes96, Bri95b]. Sixty [MBBF95]. Size [All93b, AFF^{+}97, BBGB92, BC98, Bil91, Cha95, Fra96, GB95a, GB96a, GB96b, GJH96, Goo94, Gru90, HH94b, HHH95, HP95, KHS98, LNVM94, LSM99, Leb93, LCA99, Men91, Mor99, NE94, PSB98, RVK^{+}94, RWES93, RB95b, RL91, Rym94, RJL95, SLNP95, SM91, Sun96, Tsc92, VW98, WP96, WMKF98, YG92]. sized [Ber90]. Sizes [BLF^{+}93, Tho90]. Skills [CDD96]. Skink [MSM99]. Skins [Tho99]. Slash [MPB96]. Slow [Orr96a]. Small [AHMM97, Ber99, DC98, HBA93, LA92, LV91, PC93, SB98, SS94, Wal90]. Smaller [PGFD95]. Smallmouth [KB98]. Smoky [AB90]. Snail [CD98]. Snake [HW95, LNA95, PGW97, CBAT97, HW95]. Snapping [PDC96, RSMB99]. Snare [Nos98]. Snow [BCP96]. Snub [SS95]. Snub-Nosed [SS95]. Sobre [BWS97, BRHD97, HNC97, VTM99, VPW97, WSE97, ACB^{+}99, Ama93, BSC95, BB97, BBSK96, CSBC95, Can96, CBCC93, DLK^{+}95, DME95, FS95, FP95, FEM95, GLFZ96, GWW93b, GNH94, Gru97, HVM96, HWB96, HLAS96, JBB96, KC95, KB97, KGD96, KS94b, LWMG93, LDP95, MKM^{+}96, MPB96, MGB94, OK96, OO93, Pat94a, PEH93, PERC96, RDN94, RHCB93, SBL96, TM96, TB93, VS95, Wal99b, MM94]. Sobrevivencia [MJ97]. Social [BC98, CKB98, Job91, Kau93b, VPW97, VC99]. sociales [Kau93b]. Society [Ano90i, Ano90j, Ano90k, Ano90l, Ano91i, Ano91o, Ano91p, Ano98a, Bru93a, Bru93b, Lov90a, Lov90b, Lov90c, Lov90d, Lov91, Sou91, Sto95, Tem91, Tem92c]. Socioeconomic [KRG^{+}99, SDT93]. Socioeconómicos [KRG^{+}99]. Socioeconomics [STM95]. Sociological [MS92]. Sockeye [CBAT97]. Socorro [Wal90]. sodalis [RHCB93]. Softening [Mef98c]. Soil [Deh96, FS95, SS94]. Soils [BB97]. Sokoke [FMF95]. Sol [GBAR97]. Solar [AKG98, BHKH96]. Soluciones [CCR96]. Solution [Blu93, Str96]. Solutions [CCR96, MH91, Par93]. Solving [TD94]. Sombra [GBAR97]. Sombrilla [Lam97, Ber97a]. Some [Ewe91, GB95a, HSD93, IU92, Orr92d]. son [Yaf97, BGTB93]. Songbird [SW99a, SPR97, WGO95, Woo99].

Songbirds [BGTB93, DH97, FEM95, FCM99, SPP96, SW99b]. Sonora [QLDV96, SMPN99]. Sonoran [QLDV96, SMPN99]. soportan [LN96]. sostenibilidad [LN96]. sostenible [Rob93b]. sostenida [RJL95]. sostenido [ART94, GMG94]. Soulé [CMB96]. Sound [Sti97]. Source [Doa95, DLK^{+}95, Hel98]. Sourcebook [Los95]. Sources [BT94, DTFP95]. South [LCPM97, Mad99, MBT90, PL95, RS92, SGW90, WDSA94, BSW^{+}96, CEPR91, IS96, LNA95, Ogd92, PERC96, RLR97]. South-Temperate [WDSA94]. Southeast [BPC97, DG92, LM95, SMHK98]. Southeastern [BC97a, FH97, LP96, Sha93, All93a, BB93, BB99, Zei92]. Southern [RDE96, Ash97, BC99, BSR97, DD97, ELS93, FFL95, GJH96, HGK^{+}96, HHWB90, JHHB99, Lin95, PEH93, Pet99, RFG^{+}98, RDN94, Sam90, SWK^{+}97, Sim96, SF97, VRV^{+}94, Vog95]. Southwest [PBKM95]. Southwestern [BM95, Dud97, DC92, HV97, WRSP97]. Space [HS95, ML90]. Spacing [LNVM94]. Spain [LP98]. Spanish [GMSP98]. Sparrow [DBCM95, TB93]. Sparrows [LDP95]. Spatial [BBSK96, CZN99, Cha95, CG98, EDFM96, HSD93, HM97, HWG99, Koe98, LN90, NBH^{+}98, OBY98, PC98, RS92]. Spatially [Ano99i, LDP95, RHK97, RHK99, Sou99]. Speak [BW92]. Special [Ano90o, CM93, Ehr91, Par93, WN98b]. Species [Aco99, Ang95, AHMM97, ALGPH95, BEG^{+}94, BEG^{+}95, Ber97a, BL99, BLS99, BDSD98, BGB96, BBSK96, BBCR99, Can96, CRGGFT97, CO99, CKCM95, CC99b, Cho95, Cro92, CD98, CKB98, DS95, Dia95, DP96, DHW99, DME95, Erw91, FJPC95, FESS94, Fim94, FH97, FS91, FCR^{+}98b, Gas91b, GZWR97, Gel99, GDK99, GD99a, GHB95, GJH96, GWW93b, Gol99a, G96, GB95b, GMF99, GS99, GDB96, HCB94, Ham94, HS98b, HM97, HE91, HGNK93, HS95, How94, HH94c, Jen96, Jop94, Ker97, KB94, KSO^{+}99, KF95, Kuc92, Lam97, LNVM94, LB95, LFOJ96, LV91, LCPM97, Lug92, ML91, MS92, MNU94, MP97, MHH98, MFR97, Mau97, MLD97, MBKR92, MWW99, Men91, MWN^{+}94, Nab94, Nae98, NBH^{+}98, Nor91, NM95, O’C92, OB96, Pai95, PGFD95, PC92, PC98, PD97, PBBA90, PNS99, PVZJCI90]. Species [PTTN^{+}96, PDS96, Rab95a, RDE96, RA94, RGH^{+}99, RSR96, RG99, Ree96, RH93, Roh91, Roj92, Roy95b, Run95, Saf93, SVC96, STM95, Scu99, STA^{+}99, Sid98, SC97b, SDB^{+}96, SL93, SG99, Sou90, Tay95, TSHG95, Thi92, Tho94a, Tho94b, TTW^{+}94, TCO^{+}96, UCB92, UA94, Ver93, Wap98, Way92, Wea95, Wei99, Wes90, WMW93, WC98, WH95, WKS^{+}94, WRSP97, Win98, WK99, WF99b, YJG^{+}96]. Species-Area [Dia95]. Species-Rich [LCPM97, PTTN^{+}96]. Specific [Fim94]. Specimen [SK99, Win96b]. Sperm [WCD97]. Sphenodon [CDH95]. Spider [RF97]. Spill [Gar97]. Spiny [Aco99]. Spiranthes [ST95, Sun96]. Spotted [Bar95, HSD93, LMNV92, LNVM94, MGM^{+}98, SGMP99, TG93, WBKH97, WM91]. Spread [SEC98]. Spring [RLR97]. Springs [EE97]. springvillensis [MBH96]. Spruce [RFG^{+}98]. Spruce-Fir [RFG^{+}98]. Squamate [DS95]. Squirrel [UCB92, War94c]. Sri [GG90b, MGB94, Wik90]. ssp. [RFM97]. Stability [Ano98d, BWS94, Cla95, LV96, PBBA90]. Stages [PS96]. Stand [BB97, DSB95b, Hut95c, LF97]. Stand-Replacement [Hut95c]. Standards [Cal90a, Leo90]. Standing [TTJ^{+}99]. Stands [CB98, FESS94, KDG98, NPM98, RW99, ZG99]. Stasis [Sto95]. State [Gra99, HHH96, Rei95]. State-Listed [Rei95]. States [All93a, BBS93, BMN92, Cla92, DC92, FBFO96, KMR94, Lug97, LM95, RG99, Sal99, Tho93, Zei92]. Station [Bie95]. Statistical [HS97, Pat94b, Ree96, RB97b, Str99, TG93]. Statistics [Ber93, OG96]. Status [CSBC95, CRGGFT97, Fis94, HH96, HM97, HS96b, JG91, KMW95, Lyd96, MBD95, MBW96, PT95, Pri99, RRM93, Str96, WW98b, Way92, ZDM98]. Stay [DP92]. Steelhead [Bla94]. Steinbeck [Duf93b]. Stephen [GP97, PK94]. Stephens [GP97, PK94]. Steppe [HS98b]. Steps [Pre94]. Sternotherus [WOA98]. Still [Cob99, JFH93, You99]. Stochastic [Dia95, KTA94, NGN96, RS95a]. Stochasticity [Fol94, Leb98]. Stock [Dav96, DLP^{+}92]. Stocked [HGK^{+}96]. Stocks [ABB^{+}97, DMMG96, Neh94, WW98a]. stofoniferum [HVG91]. Stopover [RCHP94]. Stories [Pis94]. Stork [BHSO92]. Storms [PERC96]. Strait [MHL97]. strand [Kau95b]. Strategies [Alt94, BN98b, FNC98, HBA93, Leb93, RO99, RRS96, SJ98, Tri98, VIO98]. Strategy [BMN92, CAH97, DKOW94, RS95a, RS91b]. Streams [JHHB99, LNPC^{+}95, VI95]. Streetfight [Dai95]. Stress [Cre97, CCM97, EHBE97, Fur95, HP95, MBW^{+}98]. striatus [SNP94]. stricta [WDWG94]. Strife [HHF^{+}97]. Strips [Hil95, MVH96]. Strong [CEPR91, Pat94a, Was96]. Strongly [AHMM97]. Structural [NPM98]. Structurally [CFL96]. Structure [Bak94, BAR^{+}93, BR96, Cha95, DM90b, FFL95, FB97, FCR^{+}98b, Fur95, LF97, MBH96, PKH97, PGW97, RF99, Rob95a, RSMB99, SLS92, VPW97, WRM^{+}91, YB96, YBZ99]. Structured [PK94, RS95a]. Structures [BC94]. Structuring [FS92, JDEB98]. Student [All97]. Students [Dun97, Dun99, GP98]. Studied [McK99]. Studies [AP99, HD96, PC92, PL94, RS93c, SC97b, Yah96]. Study [BL99, BG95b, Can96, CC99a, DHJG91, Deh96, EE97, FCR^{+}98b, GSAZ98, Hed95, KEE91, MNU94, MHH98, OB96, PPR^{+}91, PB95, RS91b, TTW^{+}94, VT99, Wei92, WDR^{+}98]. Sturgeon [WW98b]. Sturgeons [Bir93, BDSD98]. suaveolens [YB96]. Sub [Lan96]. Sub-Saharan [Lan96]. subacuáticos [CM93]. Subalpine [HHR99]. Subandean [Ren99]. Subandina [Ren99]. Subdivision [PL94, PL94]. Subdivisions [KWR95, Wue94, Sie96]. Subespecies [MSO96]. Subject [Ano99o]. Subletales [MB99]. Sublethal [MB99]. Sublittoral [GSAZ98, GWW93b]. Submarine [HWG99]. Submarino [HWG99]. Submontane [HC99]. Subsaharica [Lan96]. Subsets [Bla91]. Subsistence [ARRK97, Cob97b, FMF95, JJ99, WCTM98]. Subsistencia [ARRK97]. subsp [SAS95]. subsp. [DS92]. Subspecies [ML90, MSO96]. Subspecific [SS99a]. Substitutions [SS99a]. Substituto [PBB99]. subtidales [GWW93b]. Success [BLD^{+}95, DTFP95, Fim94, FCM99, GHP97, Gat94, Hel98, KGD96, Lan96, Loc99, MR99, Pat94a, PV98, RB95b, RMCM97, SPR97, Yos94]. Successes [Mat99]. Successful [SG98]. Succession [CUM96, FS95, LC99, MW93]. Successional [CPN95, Hun98, Lit93, TA97]. sucesion [FS95, CUM96, CPN95, MW93, LC99]. Sucesionales [TA97]. Sucker [DMMG96, MBW96]. Sudáfrica [LNA95, LCPM97, Mad99]. sudamérica [BSW^{+}96]. Sudeste [BC97a, BB93, RDN94, Sha93]. Sudoeste [HV97]. Suecia [Har94, PS99, RDH97]. Suecos [BEG^{+}94]. Suelo [RCM97, WRSP97, BHM94, FS95, GF97, LS96]. Suelos [BB97]. sugerencias [OO93]. Suggest [RH97a]. Suggestions [OO93]. Suggests [Cob97b]. Suitability [Tsc92]. Suitable [Bar95]. sujeta [PZSG96]. Sumatra [KO98, Rab95a, Thi95]. Sumatran [GS93, Hut95b, MAS^{+}97, Rab95a]. sumidero [Doa95, DTFP95]. sumideros [DLK^{+}95]. Summary [HB90]. Sun [GBAR97]. Sundance [TCK^{+}95]. superficie [VHM94]. Supervivencia [AHMM97, BO96, GWM^{+}97, PF99, GAC^{+}95, ADFT97, CPN95]. Support [BDR^{+}98, LSC^{+}96]. Supported [HGNK93, TSHG95]. Supportive [RL91, Rym94, RJL95, RJL99]. Suppression [Bak94]. supresión [Bak94]. Supuestos [How94, Ham94]. Sureste [BB99, BPC97, FH97, LP96, LM95]. Surficial [SGR^{+}98]. Suroeste [WRSP97]. Surrogate [CO99, PBB99]. Surrogates [OB96]. Surrounding [HNC97]. Survey [Cor96, DS92, PP96, TM96, WGRT96]. Surveys [Ing90]. Survival [ADFT97, AMA^{+}94, BO96, CC95, CPN95, GWM^{+}97, Kar90a, KB98, MJ97, PF99]. Survivorship [BCGL93, EG96, GAC^{+}95]. Sus [AHMM97, BSC95, PL94, SB99]. susceptibilidad [BWS94, LV94]. Susceptibility [BWS94, LV94]. Susquehanna [BHM94, BHM94]. Sustainability [ARRK97, CM97, Duf93a, LN96, MH99a, MHL97, Nos91, Pat91, Sal90, UKG^{+}96, WCTM98]. Sustainable [ART94, Cob97b, CD92, Edu91, Fit94, GMG94, Har95b, LF97, OPRCB95, Pel95, Rob93b, Str98, Tho99, Wil94, Ehr91, GM91, Sch91, Tho91, Wil91]. Sustaining [Sal91]. Sustentabilidad [ARRK97, MH99a, CM97]. Sustentable [LF97]. sustentada [HGNK93]. sustentadas [TSHG95]. Sustituatas [CO99]. sustitutos [OB96]. Swan [MW98]. Sweden [GN92, Har94, NG92, PS99, RDH97]. Swedish [BEG^{+}94, BEG^{+}95]. Sweet [NM95]. Sweetwater [Wei92]. Sympatric [GP97]. Symposium [Ano96d]. Synchrony [BC99]. Synergism [LSS95]. Synonymous [Per93]. Synthliboramphus [Ber95b]. System [BBS93, Boz93, DeM93, DRP98, HWB96, Job91, KEE91, LSC^{+}96, MC98, Rol95, Sal93, ST95, Sun96, Cur93]. Systematic [DGG^{+}99]. Systematics [Con95, KB99, Par95, SFM99, SG99]. Systems [MT99, Pre94, SDT93, Vog95].

T. [Sti97]. Tactics [BMN92]. Tailed [AF98, KLF98, WSE97, Wal90]. Taiwan [HAD^{+}97]. Takahe [BJ95, BJ96, MJ97]. Take [BN97, BFS98, WW99]. Takhi [VW96]. Taking [Dav96, Dra96, Jac93]. tala [KGD96, MGB94, NLS93, PEH93]. Taladas [Ash97]. Talados [CC99b]. Tale [HCBH99]. Talk [Ehr94b]. Talking [BBM94a, Ehr94a]. Tallgrass [How94].

Tamaño [Fra96, GJH96, LSM99, LCA99, Mor99, BLF^{+}93, HP95, LNVM94, NE94, RVK^{+}94, RJL95, Sun96, Tsc92, WP96, AFF^{+}97, HHH95, RWES93]. Tamarins [CRJ^{+}94, KM98]. Tameness [DB98]. Tamiasciuris [War94c]. Tana [KO91, Med93]. Tanager [WER^{+}97]. Tanagers [RLD99]. Tanganyica [ACB^{+}99]. Tanganyika [ACB^{+}99, CBCC93, CM93, HGNK93, PB94a]. Tángaras [RLD99]. Tanjung [Yea97]. Tanzania [New91, NMGS94, New96]. Tanzanian [New96]. tarde [KALG94]. Targeting [TTJ^{+}99]. Taricha [AKG98]. tasa [RVK^{+}94]. Tasas [McK99, Bur95, RR99b, RC94]. Task [Sim95]. Tasmania [KB94]. Taxa [EE97, McK99]. Taxas [EE97]. Taxon [PC92]. Taxones [McK99]. Taxonomía [KB99, Loc99, Bro96]. Taxonomic [GQ99, MBD95, OB97, Par93, Rex93, RBMA98]. Taxonomy [Bro96, KB99, Loc99, PH92]. Taxus [BHS95]. Teaching [Dun97, Qui95, RO99, TD94]. Tears [Tho99, Tur93]. Techniques [SSEM94]. Techno [Mef92]. Techno-Arrogance [Mef92]. Technological [Orr94c]. Technologies [Mef92]. Technology [Aug95, Fra92, Vog95]. Técnicas [CDD96, SSEM94]. tecnologia [Vog95]. Ted [For94]. Tehuacán [CRGGFT97]. Tell [LM94, Tho90]. temas [HDE95]. Temperate [ADM94, BL95b, MW90, WDSA94]. templadas [BL95b]. templados [ADM94, WDSA94]. Temple [Cob91]. Temporal [BBSK96, BPO99, HWG99, LCA99, RL96, LS96, WT90]. Temporales [LCA99, RL96]. Temporary [DC98]. temprana [CPN95]. temprano [HVM96]. Tendencias [BHCB99, GD99b, TM96]. Tension [MM96]. Teoría [PQL99, Har95a]. Terborgh [DM90b]. Term [BCGL93, BB90, Del99, GD99b, Har94, JFJ96, KSSM95, Ren99, Rob99, RS91a]. térmico [RHCB93]. Terrapin [RCMR97, RCMR97]. terreno [FR93]. terreste [DSB95b]. Terrestres [DHW99, RCM97, SS99b, BG95b, Deh96, EDFM96, TB95, VRV^{+}94]. Terrestrial [BG95b, DBH^{+}99, DSB95b, EDFM96, KCE^{+}93, Sem98, SS99b, TB95]. Terrestrial-Breeding [DSB95b]. Territorial [CA92, DBH^{+}99, LNVM94, PBG^{+}93]. territoriales [LNVM94]. Territory [YG92]. Test [CZN99, KMW95, LC98b, ST94, ST94]. Testing [FH95, SC97b]. Tests [BBCR99, PFSG97, Ran96]. Tettigoniidae [KA92]. texanus [DMMG96, MBW96]. Texas [PG91, SVC96]. Text [Baw95a]. Textbook [Die94, Gru96, Tem93b]. Their [BC98, HR93, HRCTS99, KCE^{+}93, LV94, SB99, TNN^{+}99, MCO98, Swa98]. Them [Jor93]. Theory [Har95a, Har92, PQL99, RMJ92, WABH95, Wit97b]. There [Ano99j, Erw91, FNN92, Gol97b, Hag92, RTS90]. Thermal [RHCB93, VAPCO98]. Things [Gol97b, Wil92]. Thinking [Sch96]. Third [Col92]. Thistle [HT95]. Thoughts [Orr92d]. Threat [BEG^{+}95, BGF95, GWM^{+}97, HRCTS99, McK99, SM99, TB98]. Threaten [Cam95]. Threatened [BY96, BSW^{+}96, BEG^{+}94, Bir93, BPC97, DC92, FT91, GHB95, Gre96, HT95, KMW95, LDSH99, ML91, MKH98, MSM99, MWW99, MBD^{+}99, MFW90, NR97, RSR96, Rei95, WOA98, Wei99]. Threats [BEG^{+}95, Gre96, Kuc92, ML91, Mas91, Mau97, MAR96, Nab94, RBMM97]. Three [AMW90, BHL^{+}92, BDSD98, Wal99b, Yaf99]. Threshold [Fra95a]. Thresholds [Tho94c, WK99]. Throw [VL91]. Throwing [SF96]. Thrush [Tri98]. Thuja [KL97]. Tian [DLY92]. Tidewater [LSA99]. Tiempo [Ham94, Fol94]. Tierra [War94a, CRHSS96, DPOO94, FS95]. Tierras [CC99a, GW97, JM97, MBD^{+}99, TCO^{+}96, CBLV94, DOSP94, FGG^{+}95, PL95]. Tiger [KSSM95, PC92, Sab97, SAM98, WDR^{+}98]. Tigers [KM97, ML90, SM91]. Timber [AB94, God93a, God93b, LP96, PEH93, Pet94, VJSU98]. Time [Bis92, BPO99, Car94, Ham94, HS95, Mei99, VB99]. Times [Fol94, LS96]. tin [KGD96]. Tinkering [Sam96a]. Tipo [RCM97, Fim94]. Tipos [SPR97, Wal99b]. Títeres [VM99]. Titers [BSW^{+}96]. Tiwai [Fim94]. Toad [Agu94, DHBG97, KSO^{+}99, LP98, PC94, SFM96b]. Toads [Car93]. Today [LSV^{+}96]. Tofieldia [GWH97]. Together [Bis92, Tem92a]. Tolerance [CKCM95]. Tom [Con95]. Toma [BN97, SFM99, LSC^{+}96, RS95a]. Too [LM94]. Tool [BTRM99, DCL98, DHW99, FDV94, GFP98, Kin95, KRL98, PB94a]. Tools [MPC99, Sta96]. Top [CRJ^{+}94, PGFD95]. Topminnow [PSH99, SHMV97]. Topminnows [LV94]. Topología [Fri93]. Topology [Fri93]. Topos [SCZF96]. tormentas [PERC96]. Toro [LDSH99, Cua97]. torosa [AKG98]. Torres [MHL97]. Tortoise [Bur90]. Tortuga [KB99, MB99, RCMR97, JBB96, PDC96]. Tortugas [HC96, BG95b, EBV95]. Tortuguero [BBL93, BWBM99]. Totoaba [CMMLRR95]. Touch [Kau95b]. Touchstone [UKG^{+}96]. Tourboats [GB97]. Tourists [YH91]. Towhees [PV98]. Toxicology [Gen94, HJ99]. Tracking [HOF^{+}96, Tho94a]. Trade [HYL95, Jen96, MWS^{+}97, Ros92, SB95a, Tho99, Yu94, Yu96]. tradicionales [Thi95]. Traditional [BM92, MT99, Thi95]. Traffic [BH96, GS93]. Tráfico [BH96]. Tragedy [Kay97, San99]. Trail [TW92b]. Training [BHCB99, CDD96, JD98, PCG91, Qui95]. Trait [Fur95]. Traits [KR98, KS97]. Trampas [RCMR97]. Trampling [GLFM98, MFR97]. Transformation [Hou95]. Transient [TA97]. Transitorios [TA97]. Translocación [MBD^{+}99]. translocaciones [Cun96, SMV96, WGRT96]. translocados [WPSZ96]. Translocated [BBCS99, WPSZ96]. Translocation [ALGPH95, FS91, GR98, MBD^{+}99, Wik90]. Translocations [Cun96, SMV96, WGRT96]. transmisión [Cun96]. Transplanting [SG93]. Transport [Loc99]. transportados [FP95]. Transportation [SEC98]. Transporte [Loc99]. Trapdoor [RF97]. traskiae [RG95]. tratados [CRHSS96]. tratamientos [GNH94]. traversi [AMA^{+}94]. través [FP95, ML94, Wal95]. Trawling [EK98, PCM98, SGR^{+}98]. Treated [CRHSS96]. Treatments [GNH94]. Tree [BE94, CK98, DSB95b, FCR^{+}98a, HCB94, HWB96, LFM^{+}98, NABE95, PVZJCI90, VAPCO98]. Tree-Dependent [NABE95]. Tree-Stand [DSB95b]. Treefrogs [GK99, MGM^{+}98]. Trees [Duf93c, FL97, HC99, Orr93c, Ric98, Sch92, SMPN99]. Trend [BWBM99, CG98, MW98, TM96]. Trends [AB90, BHCB99, BBM94c, CPC96a, GD99b, HS97, RM97]. Tres [Wal99b, Yaf99]. Triage [MBKR92]. Tribute [Hur94]. Trichechus [MHO97]. Trifoflium [HVG91]. triguero [NHA95]. Trilepidea [Nor91]. Trillium [JR99, JR99]. tritón [GK96]. Trophy [LA97]. Tropical [ART94, Ano90p, BM92, BS98, BM98, Bie95, BL95b, Bri96, BPO99, CPN95, FS95, HCB94, HWB96, JRP99, Lau91, Lip98, MCW98, MMT^{+}98, New91, PERC96, PFSG97, Put93, RTS90, Rob99, SBL96, TSP97, Tre93, TTW^{+}94, TCO^{+}96, WLDGGZ98, You94a, Hil95, LMS96, SDT93, Thi95]. Tropicales [BPO99, JRP99, BL95b, CPN95, DLK^{+}95, HCB94, Hil95, PERC96, SBL96, You94a]. tropicalis [TSP97]. Trópico [SAQMGQ^{+}99]. trópicos [BL95a, KMM94]. Tropics [BL95a, KMM94, Par93, Uhl98b, SAQMGQ^{+}99]. Trout [DC92, GMSP98, GL95, HGK^{+}96, LAF93, RM97, SGK95, TLG^{+}98]. trucha [GL95, HGK^{+}96]. Truda [Vid94]. True [Med98]. Truly [Hun94]. Truth [BK99]. Tsetse [WRSP97]. tshawytscha [NTT94]. Tsitika [Duf93a]. Tuatara [CDH95]. Tuna [Saf93]. Tupinambis [Fit94]. Turísticas [GB97]. Turkey [Leb91]. Turnover [MNU94]. Turtle [BWBM99, Fra92, GFP98, Gra95, JBB96, KB99, PDC96, PH94, Pri99, RSMB99, WOA98]. Turtles [BBL93, BAR^{+}93, BG95b, CDV93, EBV95, HC96, MB99]. Tuxtlas [DG92]. Twelve [CKCM95]. Twenty [BWBM99]. Twenty-Six [BWBM99]. Twine [Orr94d]. Two [DSB^{+}95a, HHF^{+}97, MP97, NM97, RS95a, SC97a, SGMP99, Son96, SB95a, TSSB95, WPSZ96, Sou91]. Type [RCM97].

U.S. [AS94, Cur93, GDK99, JJ99, KL96, Lyd96, MJ96, MGM^{+}98, PD97, Pri91, RDE96, RA94, Sal91, Tho94b, TB95, WMW93]. U.S.A [Fri93, HE97, Hut95c, MBW96, SB95c, TCK^{+}95, TW92b]. U.S.A. [BHM94, CSBC95, CZN99, KSO^{+}99, NPM98, TA99, WMW93]. Uganda [CCOO^{+}96, Gol97a, NT98]. Ultimate [Kay97, Mil94, Nab94]. umbral [Fra95a]. Umbrales [WK99]. Umbrella [Ber97a, Lam97]. Umfolozi [Mad99, Mad99]. Umpqua [RLR97]. Un-Migrante-Por-Generación [MA96]. Uncertainty [BCG94, MHH98, SFM99, TB98]. Uncompahgre [BBM94b, BBM94c, SO94]. Under-Used [DCL98]. Undergraduate [Tro94]. Underlies [Bra96]. Understand [BTRM99]. Understanding [BJ95]. Understories [DM92, JFH93]. Understory [AF98, New91, SB95b]. Underwater [CM93, PB94a]. Unfished [GMSP98]. Unforeseen [SS94]. Unfragmented [PF99]. Unfulfilled [Joh94b]. Ungulados [BH96, GMG94]. Ungulate [BH96, VW96]. Ungulates [Ber91, GMG94]. único [QLDV96]. Unidad [PD97]. Unidades [DGG^{+}99, LRD^{+}96, PT95, VD94, KB99]. Unidos [Cur93, Fle94a, FBFO96, GDK99, KL96, Lug97, LM95, RDE96, SB95c, Tho93]. Unification [Dia96, Ehr92b]. uniformity [CSC^{+}94]. Unintelligent [Sam96a]. Union [Kei98]. Unique [NTT94, QLDV96]. Unisexual [Kra95].

Unit [PD97]. United [All93a, BBS93, DC92, KMR94, LM95, Zei92, BMN92, Cla92, FBFO96, Lug97, RG99, Sal99, Tho93]. Units [DGG^{+}99, GQ99, KB99, LRD^{+}96, Pae99, PSH99, PT95, RSMB99, VD94, Wal96, Wap98]. Unity [Roy95b]. Universe [BNF^{+}98]. Universidades [BHCB99]. Universities [Nos97]. University [BHCB99, UKG^{+}96, TBW^{+}98]. Unknown [Ano97f]. Unnecessary [Put92]. Unprecedented [EKL^{+}99]. unq [Har94]. Untapped [GP98]. Update [WGRT96]. Upland [BB97, CRHSS96, Lip99]. Urban [Bab99, RTNK99]. Urbanizado [BSR97]. Urbanizing [BSR97, ES94]. Urbano [RTNK99, FEM95]. Urchin [MKM^{+}96]. Ursus [TSSB95]. USA [GF97, DF96]. Usambara [New91]. usando [BLF^{+}93, KMW95, PB95, RFDB94]. Use [ART94, BBGB92, Baw92, BC97a, Ber97a, BN98b, BLD^{+}95, CO99, CHK98, CRHSS96, Cor96, DPOO94, DD97, DMD^{+}92, DDM^{+}92, DHE97, FS95, Fit94, Fro92, GMG94, GP97, GF97, Gre96, GD99b, Haa95, KB98, KGD96, Kin92, KCE^{+}93, LS96, Lit93, MVH96, Mar93, MS97a, Mat94, MWCP94, MR99, OBY98, Pel95, PVZJCI90, RR99a, RR99c, RM97, Sad95, Sal93, Sil94, SMHK98, SL93, TFHD98, Tho99, TWMM97, WGO95, War94a, WPSZ96, WRSP97, ZG99, Way92]. Use-Values [PVZJCI90]. Used [DCL98, FS97, GN92]. USEPA [TB95]. Uses [TG93]. Using [BHS99, BLF^{+}93, CZN99, GQ99, KGS93, KMW95, KSC^{+}96, LB95, LP96, Loc99, MB96, MB92a, OSEC95, Pae99, PRH98, PC98, Phi93, PB95, RB95a, Ree96, Tay95, Vog95, WA95]. Uso [BC97a, Ber97a, FS97, GP97, GD99b, MS97a, MR99, RR99a, WRSP97, ART94, BLD^{+}95, CRHSS96, DPOO94, FS95, GMG94, GF97, Gre96, Haa95, KGS93, SL93, Tay95, CO99, DHE97, LS96, Loc99, War94a]. Usos [TG93]. utilización [KGD96, MWCP94, OSEC95, WPSZ96, CZN99, MVH96, TWMM97]. utilizada [NMGS94]. Utilizando [BHS99, KSC^{+}96, LP96, MB96, Vog95, WA95]. UV [AKG98, BHKH96, LP98]. UV-B [AKG98, BHKH96, LP98, LSS95].

Vacas [Wei99]. Valdez [Ehr90, Gar97]. Validity [NF98]. valiosa [FDV94]. valiosas [LA95]. Valle [CRGGFT97, FS96, MN96, FESS94, LWMG93]. Valley [CRGGFT97, Fis94, LWMG93, MN96, FS96]. Valor [GDK99, HW95, MSC96, SRR95, WA95, PS99]. Valores [BO96, GL95]. Valuable [FDV94, Kin95, LA95, OD98]. Value [BM92, DHJG91, GDK99, HEO95, HW95, MSC96, PS99, SRR95, WA95]. Value-Informed [HW95]. Values [BO96, GL95, Kel93a, Mag96, Mat97a, MM96, Nos96a, Tru94, PVZJCI90]. Valuing [Duf94, Tre93]. Vanishing [YJG^{+}96]. Vaquita [TG93]. var [DeM93, NCE96, SAH^{+}95]. Variabilidad [KGM^{+}97, FBS95, HCB94, DMMG96, WLAH92]. Variability [DMMG96, FBS95, KF95, KGM^{+}97, Leb93, MBT90, MAS^{+}97, SSTL92, WLAH92]. Variable [SCAB91]. Variables [ML94]. Variacion [Fra96, EKL^{+}99, GS99, MSO96, BBSK96, GSB99, HP95, NCE96, RVK^{+}94, RCHP94, RJL95, CRJ^{+}94, FB96, LV94, OFH94, SCZF96]. Variance [Rym94, RJL95, SM91]. Variation [BC98, Bil91, BBSK96, BMR^{+}92, Cha95, CRJ^{+}94, DM90a, Ebe95, EKL^{+}99, FB96, Fra96, GZWR97, GSB99, GS99, Gra95, HCB94, Hea94, HP95, HVG91, KR98, KMR94, LV94, Man94, MSO96, NCE96, NBH^{+}98, OFH94, PWC^{+}98, PSH99, RVK^{+}94, RF97, RCHP94, SCZF96, SMGB91, SLS92, SRV96, YAM^{+}98]. Variegated [ML94]. Variegation [IS96, MB92b]. Variety [NG92]. Vascular [Kei98, PBB99]. Vasculares [PBB99]. Vavilov [Wil93a]. Vectors [PTK^{+}98]. Vegetación [BWS97, BC97a, LC99, BT94, CSBC95, Fra94, Hil95, KGS93, OSEC95, Sha93, ST94]. Vegetados [MP99]. vegetal [CUM96]. vegetales [TTW^{+}94]. Vegetated [MP99]. Vegetation [BWS97, BT94, CSBC95, CLL91, ELS93, FF95, Fra94, HS98b, Hil95, KGS93, LC99, MA91, OSEC95, PS98, Sha93, ST94]. Vegetation-Based [PS98]. Velavadar [Jha93]. Venezuela [RRS96, SG98, TV99]. Venezuelan [TV99]. Veracruz [SP98]. Verbicide [Orr99c]. Verde [HM97, KB99, Yu94]. Verduguillo [CW97a]. Vermont [RCM97]. Vernal [BHM94]. Versus [Gei90, KB99, KWR95, RMJ92, Sie96, Wil93b, Win96b, Wue94]. Vertebrados [GD99a, RF99, Bun95, EDFM96, GB95b, JFJ96, TB95, WH95]. Vertebrate [Bun95, JFJ96, Kra95, OBY98, RF99, WH95]. Vertebrates [EDFM96, GD99a, GB95b, HPB^{+}97, Lit93, TB95]. Very [Bis92]. Via [KEE91, FNC98]. Viabilidad [MHO97, MFR97, MHB^{+}96, PKH97, PF99, RLR97, Tay95, AL93, BBSK96, Bus96, HBA93, Har95a, LMNV92, LNVM94, LP96, NGN96, RHS94]. Viability [Ano90o, AL93, Bar95, BB98, BBSK96, Bus96, CG98, GF90, Gil96, HBA93, Har95a, JB99, KSSM95, LMNV92, LNVM94, Leb98, LP96, MHO97, MFR97, Men90, MHB^{+}96, MFW90, NGN96, PKH97, PF99, RLR97, RHS94, Sha90, Son96, Str94, Tay95]. Viable [Gru90, KO91, Sim96, Tho90, Tsc92, Wal90]. viables [Sim96]. vias [SK96]. Victoria [CKCM95, GWW93b, KO93, SKOO98, SWK^{+}97]. Vida [BHCB99, Fra96, GD99a, Hac99, JRP99, LA97, LOT^{+}99, MN99, AS94, BT94, Cur93, EDFM96, KHP95, LWMG93, Rob93b, WP96]. View [Ano94b, Cla99, Cob99, Fle95, HHH96, JJ99, KWR95, MBH96, UA94]. Viewing [Nor96]. Villafuerte [PDFG96]. Village [LLP95, MN99]. vinculando [LDP95]. Virgin [JJ99, DDM93]. Virginia [Ang95]. Virtual [Orr96b]. Virtue [Orr90d]. Virtues [HH94c]. Virulence [Ebe95]. viscaria [LSM99]. visión [MBH96]. vista [LNA95]. Visual [SB95a]. Vital [KMM94]. vitulina [SRV96]. vive [NMGS94]. Voice [Roy95b]. voladores [PERC96]. Volcano [Mac91]. Voles [HHR99, Mil95b, WSE97]. Volume [Ano99b, Ano99o]. vs [SMS97]. Vulnerabilidad [JRP99, LM93]. Vulnerability [JRP99, Kat92, LM93]. Vulture [Bus96].

Wadden [SRV96]. Wainwright [CAB^{+}98]. Wales [PL95]. Walker [Bla93]. Wallaby [EKL^{+}99]. Want [Sil94]. Wanted [Rol95]. Waples [CAB^{+}98]. War [BK99, MS99a, Wil96]. Warbler [GF90, MHH98]. Warblers [ES94, GHP97]. Wariness [SSS94]. Warming [SVC96]. Wash [SRV96]. Washington [BHS95]. Wasps [AHMM97]. Watches [JBB96]. Water [BHSO92, BT94, CBLV94, GLFZ96]. Waterbird [CD95]. Waterbirds [KHP95]. Waterfowl [PTK^{+}98]. Waterfrogs [ALGPH95]. Waters [EBV95]. Watersheds [MA91, MR98b]. Wax [CLL91]. Wayne [DMD^{+}92]. web [Kau95b, Bra96]. Weed [KEE91]. Were [Ano99j]. Wernstedt [Jen95]. West [Sal93, BB97, LLP95, LNPC^{+}95, MS99a, Saf93, SDT93]. West-Central [LNPC^{+}95]. Western [Gra95, KS94b, RS91a, Sib94, Ama93, BBS93, BHS95, CG95, De 95, Fle94a, HM98, Lip99, MWM^{+}99, MB96, MN99, New95, NLS93, NHA95, PL95, SHM97]. Wet [FS95]. Wetland [BG95b, GHB95, HMO98, KSO^{+}99, MGM^{+}98, PTK^{+}98]. Wetlands [CV95, DC98, FH97, GLFZ96, SB98]. Wetmore [WER^{+}97]. Whale [GD99a]. Whales [BCP96, GDK99, Hol93, WCD97]. Whaling [BCP95, JOB96, WCD97]. Wheatbelt [NHA95]. Where [Ano99p, BCHM94, God93b, LNA95, RS92]. Whimper [Sti97]. White [AF98, BRHD97, CA92, FBFO96, HOF^{+}96, KL97, KLF98, MKH98, PB94b, PSB98, RB97a, SB94, TB93]. White-Backed [CA92, MKH98]. White-Crowned [TB93, SB94]. White-naped [HOF^{+}96]. White-Tailed [AF98, KLF98]. Whitebark [TCK^{+}95]. Whither [Cal90b, Nos93]. Who [BW92, CGG^{+}97]. whole [Kau95b]. Whooping [Can96, CRHSS96, GSB99]. Wide [PC92]. Widespread [FCR^{+}98b, GWH^{+}99]. Width [RDN94]. Wild [Bri95a, Caf95, Cla92, CC96, Cre97, CCM97, EG96, EHBE97, Fle94b, GAC^{+}95, GSB97, HBD90, HMW98, Ham94, Har95a, HS96b, Leb91, Mad99, MJ97, MG97, NGN96, Ryd93, TNN^{+}99, VC99]. Wild-Reared [MJ97]. Wildebeest [PKH97]. Wilderness [Bra98, MMT^{+}98, Nos91]. Wildfire [GNH94, KO98, Uhl98b]. Wildlife [AS94, BHCB99, Bis94a, BT94, Cun96, Cur93, EDFM96, Fra96, GS93, Hac99, HAD^{+}97, HHWB90, Jes98, Kel91, KHP95, LWMG93, LKM90, LA97, LN93, LOT^{+}99, MN99, NT98, NMGS94, Nos98, OSEC95, Per93, RMJ92, Ros92, SS99a, Sim95, Tem92a, You95]. Wildlife* [Est98]. Will [WW99]. Willers [Blu93]. Winker [FSCL96]. Winter [BB99, BB98, Hea94, HH94b, HHH95, Sam90]. Winter-Run [BB98, HH94b, HHH95]. Wiregrass [WDWG94]. Wisconsin [KSO^{+}99]. Wise [Mor95]. Within [HNC97, RBPB96]. Within-Plantation [HNC97]. without [CC95]. Witness [CMW^{+}98]. Wolf [Bro96, CDB99, FB97, GMMR^{+}96, GGB^{+}96, HMW98, JG91, KHM99, LR91, LRT93, Mec95, MSHW95, Now92, NF98, PH92, WLG^{+}91, WLAH92, WRG98]. Wolves [FT96, FB96, Hab96, Jha93, Roy95a, RGSW96, VW99a, VPW97]. Women [HHWB90]. Wood [BHSO92, Tri98, Gei90]. Wooded [Haa95]. Woodland [DBCM95, NHA95, ST96, VTM99, YB96]. Woodlands [AM99, FCR^{+}98a, PB94b, PSB98, WGO95]. Woodlots [Bla91]. Woodpecker [CA92, LS96, MKH98, MA98]. Woodpeckers [HBA93, RWES93, SLS92]. Woodrat [BY96]. Woody [CLL91, CC99b, Fra97, GS99, RH97b]. Work [O’C92, Roh91]. Worked [SC97a]. Working [Sto94]. Workshop [HB90]. World [Bur90, Col92, Gra98, Kei98, Max90, NF93, PC92, PAPA90, RR99a, STM95, Str95, Wil94, Kau93a]. World-Wide [PC92]. Worldwide [NY91]. Worm [GHP97]. Worm-Eating [GHP97]. Worth [LA92]. Wounded [MBKR92]. Wrens [KW96]. Wyoming [PRH98].

xylorrhizus [GZWR97]. Xyrauchen [DMMG96, MBW96].

Yard [Uhl98a]. Years [BWBM99, Gat95, KALG94, MBBF95]. Yellow [SG98]. Yellow-Shouldered [SG98]. Yellowstone [BC99, Bru91, CAWH91, MA91, Pat91, RT91, BWS97, Ber91, BC99, Doa95, GL95, Job91, MR91, MM94, Pow91]. Yes [Ewe91]. Yew [BHS95, Sch92]. Yields [Ano98d]. Yosemite [DF96]. Young [NM97, MS99c]. yroxiphium [RFM97]. Yucatan [GB97, GB97]. Yuma [MBD95]. Yunnan [GZWR97].

Zaire [HHH96, HH96, WCTM98]. Zambia [LWMG93, LKM90, LA97]. Zealand [BJ95, BJ96, CDH95, MJ97, MSC96, NY91, NR97, RKLS99, TW92a]. Zebra [SEC98]. Zelanda [MJ97, NR97]. Zelandia [BJ95, CDH95, MSC96, RKLS99]. Zen [All97]. Zieria [HP99]. zona [DP96, HOF^{+}96]. zonas [BG95b, PTTN^{+}96]. Zone [MW90, Sal93, TSSB95, HOF^{+}96]. Zones [BG95b, MS99a, PTTN^{+}96, Sem98]. Zoning [VJSU98]. Zoo [ML90, WMKF98]. Zoos [GC97]. zorro [YJG^{+}96]. zorros [PERC96]. Zuni [TA99].

References

Anunsen:1993:RS

[AA93] Cathy Sue Anunsen and Roger Anunsen. Response to Scheffer. Conservation Biology, 7(4):954–957, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Akcakaya:1997:HBM

[AA97] H. Reşit Akçakaya and Jonathan L. Atwood. A habitat-based metapopulation model of the California gnatcatcher. Modelo de metapoblación basado en el hábitat de Polioptila c. Californica. Conservation Biology, 11(2):422–434, April 1997. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ambrose:1990:TLH

[AB90] Jonathan P. Ambrose and Susan P. Bratton. Trends in landscape heterogeneity along the borders of Great Smoky Mountains National Park. Conservation Biology, 4(2):135–143, June 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ash:1994:ITH

[AB94] Andrew N. Ash and Richard C. Bruce. Impacts of timber harvesting on salamanders. Conservation Biology, 8(1):300–301, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Allendorf:1997:PPS

[ABB^{+}97] Fred W. Allendorf, David Bayles, Daniel L. Bottom, Kenneth P. Currens, Christopher A. Frissell, David Hankin, James A. Lichatowich, Willa Nehlsen, Patrick C. Trotter, and Thomas H. Williams. Prioritizing Pacific salmon stocks for conservation. Priorización de stocks de salmones del Pacífico para su conservación. Conservation Biology, 11(1):140–152, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See response [WW98a].

Ayres:1991:FCR

[ABM91] José Marcio Ayres, Richard E. Bodmer, and Russell A. Mittermeier. Financial considerations of reserve design in countries with high primate diversity. Conservation Biology, 5(1):109– 114, March 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Alin:1999:ELD

[ACB^{+}99] Simone R. Alin, Andrew S. Cohen, Roger Bills, Masta Mukwaya Gashagaza, Ellinor Michel, Jean-Jacques Tiercelin, Koen Martens, Peter Coveliers, Sima Keita Mboko, Kelly West, Michael Soreghan, Sona Kimbadi, and Gaspard Ntakimazi. Effects of landscape disturbance on animal communities in Lake Tanganyika, East Africa. Efectos de la perturbación del ambiente sobre las comunidades animales en el Lago Tanganyica, Africa Oriental. Conservation Biology, 13(5):1017–1033, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Acosta:1999:BDC

[Aco99] Charles A. Acosta. Benthic dispersal of Caribbean spiny lobsters among insular habitats: Implications for the conservation of exploited marine species. Dispersión béntica de la langosta espinosa del Caribe entre hábitats insulares: Implicaciones para la conservación de especies marinas explotadas. Conservation Biology, 13(3):603–612, June 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anders:1997:JSP

[ADFT97] Angela D. Anders, Donald C. Dearborn, John Faaborg, and Frank R. Thompson III. Juvenile survival in a population of neotropical migrant birds. Supervivencia de juveniles en una población de aves migratorias neotropicales. Conservation Biology, 11(3):698–707, June 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Arriaga:1994:CCM

[ADM94] Laura Arriaga, Sara Díaz, and Carmen Mercado. Conservation or commercial management of temperate forests of Baja California Sur, Mexico. La conservación o el manejo comercial de bosques templados de Baja California Sur, México. Conservation Biology, 8(4):1132–1140, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Augustine:1998:EWT

[AF98] David J. Augustine and Lee E. Frelich. Effects of white-tailed deer on populations of an understory forb in fragmented deciduous forests. Conservation Biology, 12(5):995–1004, October 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Arita:1997:GRS

[AFF^{+}97] Héctor T. Arita, Fernanda Figueroa, Astrid Frisch, Pilar Rodríguez, and Karina Santos-Del-Prado. Geographical range size and the conservation of Mexican mammals. Tamaño del Área de distribución geográfica y la conservación de los mamíferos de México. Conservation Biology, 11(1):92–100, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Aguirre:1994:DTP

[Agu94] Alonso Aguirre. Declining toad populations. Conservation Biology, 8(1):7, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anstett:1997:MPS

[AHMM97] Marie-Charlotte Anstett, Martine Hossaert-McKey, and Doyle McKey. Modeling the persistence of small populations of strongly interdependent species: Figs and fig wasps. Modelando la supervivencia de pequeñas poblaciones de especies fuertemente interdependientes: el caso de los ficus y sus polinizadores. Conservation Biology, 11(1):204–213, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anzalone:1998:ESU

[AKG98] Christopher R. Anzalone, Lee B. Kats, and Malcolm S. Gordon. Effects of solar UV-B radiation on embryonic development in Hyla cadaverina, Hyla regilla, and Taricha torosa. Conservation Biology, 12(3):646–653, June 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Armbruster:1993:PVA

[AL93] Peter Armbruster and Russell Lande. A population viability analysis for African elephant (Loxodonta africana): How big should reserves be? Análisis de la viabilidad de poblaciones del elefante africano (Loxodonta africana): `?Cuán grande deben ser las reservas? Conservation Biology, 7(3):602–610, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Alcorn:1993:IPC

[Alc93] Janis B. Alcorn. Indigenous peoples and conservation. Conservation Biology, 7(2):424–426, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See response [RS93b].

Aplet:1992:RCB

[ALF92] Gregory H. Aplet, Richard D. Laven, and Peggy L. Fiedler. The relevance of conservation biology to natural resource management. Conservation Biology, 6(2):298–300, June 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Arano:1995:STM

[ALGPH95] Begoña Arano, Gustavo Llorente, Mario García-Paris, and Pilar Herrero. Species translocation menaces Iberian waterfrogs. Conservation Biology, 9(1):196–198, February 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Allard:1993:SUS

[All93a] Dorothy J. Allard. The Southeastern United States: Land of biodiversity. Conservation Biology, 7(4):963–965, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Allendorf:1993:DAC

[All93b] Fred Allendorf. Delay of adaptation to captive breeding by equalizing family size. Conservation Biology, 7(2):416–419, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Allendorf:1997:CBZ

[All97] Fred W. Allendorf. The conservation biologist as Zen student. Conservation Biology, 11(5):1045–1046, October 1997. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Alpert:1995:IEM

[Alp95] Peter Alpert. Incarnating ecosystem management. Conservation Biology, 9(4):952–955, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Altleri:1990:HCO

[Alt90] Muguel A. Altleri. How common is our common future? Conservation Biology, 4(1):102–103, March 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Altieri:1994:SCC

[Alt94] Miguel A. Altieri. Strategies for conserving crop genetic resources. Conservation Biology, 8(2):610–611, June 1994. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Abbot:1999:MPW

[AM99] Joanne I. O. Abbot and Ruth Mace. Managing protected woodlands: Fuelwood collection and law enforcement in Lake Malawi National Park. Manejo de areas boscosas protejidas: Colecta de leña y aplicación de la ley en el Parque Nacional Lago Malawi. Conservation Biology, 13(2):418–421, April 1999. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Amarasekare:1993:PIM

[Ama93] Priyanga Amarasekare. Potential impact of mammalian nest predators on endemic forest birds of Western Mauna Kea, Hawaii. Impacto potencial de los predadores de nidos sobre los pájaros de selva endémicos del oeste de Mauna Kea, Hawái. Conservation Biology, 7(2):316–324, June 1993. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ardern:1994:EBS

[AMA^{+}94] Simone L. Ardern, Ian G. McLean, Sandra Anderson, Richard Maloney, and David M. Lambert. The effects of blood sampling on the behavior and survival of the endangered Chatham Island black robin (Petroica traversi). Conservation Biology, 8(3): 857–862, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anderson:1995:FMP

[AMGJ95] Anthony B. Anderson, Pennie Magee, Anne Gély, and Mário Augusto Gonçalves Jardim. Forest management patterns in the floodplain of the Amazon Estuary. Patrones de manejo del bosque en la llanura de inundación del estuario del Amazonas. Conservation Biology, 9(1):47–61, February 1995. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ashley:1990:CGB

[AMW90] Mary V. Ashley, Don J. Melnick, and David Western. Conservation genetics of the black rhinoceros (Diceros bicornis), I: Evidence from the mitochondrial DNA of three populations. Conservation Biology, 4(1):71–77, March 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anderson:1991:CFE

[And91] Jay E. Anderson. A conceptual framework for evaluating and quantifying naturalness. Conservation Biology, 5(3):347–352, September 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anderson:1992:RG

[And92] Jay E. Anderson. Reply to Götmark. Conservation Biology, 6 (3):459, September 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anderson:1994:PS

[And94] E. N. Anderson. Popularization of science. Conservation Biology, 8(3):620–621, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Andau:1995:L

[And95] Mahedi Andau. Letter. Conservation Biology, 9(5):979–981, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anderson:1996:AMR

[And96] E. N. Anderson. Appeal to morality or resort to legislation. Conservation Biology, 10(1):4–5, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Angel:1993:BPO

[Ang93] Martin V. Angel. Biodiversity of the pelagic ocean. Conservation Biology, 7(4):760–772, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Angermeier:1994:DBI

[Ang94] Paul L. Angermeier. Does biodiversity include artificial diversity? Conservation Biology, 8(2):600–602, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Angermeier:1995:EAE

[Ang95] Paul L. Angermeier. Ecological attributes of extinctionprone species: Loss of freshwater fishes of Virginia. Atributos ecológicos de las especies proclives a la extinción: la pérdida de los peces de agua dulce de Virginia. Conservation Biology, 9(1): 143–158, February 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1990:AM

[Ano90a] Anonymous. 1991 annual meeting. Conservation Biology, 4 (4):349, December 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1990:Aa

[Ano90b] Anonymous. Announcements. Conservation Biology, 4(1):11– 14, March 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1990:Ab

[Ano90c] Anonymous. Announcements. Conservation Biology, 4(4): 352–354, December 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1990:BRa

[Ano90d] Anonymous. Book reviews. Conservation Biology, 4(2):205– 211, June 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1990:BRb

[Ano90e] Anonymous. Book reviews. Conservation Biology, 4(3):331– 335, September 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1990:BRS

[Ano90f] Anonymous. Books received (September 1989–August 1990). Conservation Biology, 4(4):462–463, December 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1990:DCJ

[Ano90g] Anonymous. Depletion and conservation of Japan’s forests. Conservation Biology, 4(1):104–111, March 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1990:E

[Ano90h] Anonymous. Erratum. Conservation Biology, 4(4):464, December 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1990:GOSa

[Ano90i] Anonymous. Goals and objectives of the Society for Conservation Biology. Conservation Biology, 4(1):112, March 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1990:GOSb

[Ano90j] Anonymous. Goals and objectives of the Society for Conservation Biology. Conservation Biology, 4(2):212, June 1990. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1990:GOSc

[Ano90k] Anonymous. Goals and objectives of the Society for Conservation Biology. Conservation Biology, 4(3):336, September 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1990:GOSd

[Ano90l] Anonymous. Goals and objectives of the Society for Conservation Biology. Conservation Biology, 4(4):465, December 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1990:L

[Ano90m] Anonymous. Letter. Conservation Biology, 4(3):218, September 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1990:PA

[Ano90n] Anonymous. Positions available. Conservation Biology, 4(3): 337, September 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1990:SSP

[Ano90o] Anonymous. Special section: Population viability analysis. Conservation Biology, 4(1):38, March 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1990:TM

[Ano90p] Anonymous. Tropical mammals. Conservation Biology, 4(4): 454–462, December 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1991:Aa

[Ano91a] Anonymous. Announcements. Conservation Biology, 5(1):17, March 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1991:Ab

[Ano91b] Anonymous. Announcements. Conservation Biology, 5(2):143– 147, June 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1991:Ac

[Ano91c] Anonymous. Announcements. Conservation Biology, 5(4): 445–447, December 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1991:BRa

[Ano91d] Anonymous. Book reviews. Conservation Biology, 5(2):256– 260, June 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1991:BRb

[Ano91e] Anonymous. Book reviews. Conservation Biology, 5(3):426– 430, September 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1991:BRc

[Ano91f] Anonymous. Book reviews. Conservation Biology, 5(4):570– 575, December 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1991:BRS

[Ano91g] Anonymous. Books received (September 1990–August 1991). Conservation Biology, 5(4):575–577, December 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1991:CRA

[Ano91h] Anonymous. Conservation and the rights of animals. Conservation Biology, 5(1):1–3, March 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1991:GOS

[Ano91i] Anonymous. Goals and objectives of the Society for Conservation Biology. Conservation Biology, 5(1):133, March 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1991:La

[Ano91j] Anonymous. Letters. Conservation Biology, 5(2):136–137, June 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1991:Lb

[Ano91k] Anonymous. Letters. Conservation Biology, 5(3):265–267, September 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1991:Lc

[Ano91l] Anonymous. Letters. Conservation Biology, 5(4):435–438, December 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1991:LCN

[Ano91m] Anonymous. Logging, conservation, and native rights in Sarawak forests. Conservation Biology, 5(1):126–130, March 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1991:MC

[Ano91n] Anonymous. Meetings and conferences. Conservation Biology, 5(3):271–272, September 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1991:NSa

[Ano91o] Anonymous. News of the Society. Conservation Biology, 5(2): 135, June 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1991:NSb

[Ano91p] Anonymous. News of the Society. Conservation Biology, 5(3): 263–264, September 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1992:Aa

[Ano92a] Anonymous. Announcements. Conservation Biology, 6(1):16– 17, March 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1992:Ab

[Ano92b] Anonymous. Announcements. Conservation Biology, 6(2):167– 169, June 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1993:A

[Ano93a] Anonymous. Announcements. Conservation Biology, 7(4): 757–759, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1993:BRS

[Ano93b] Anonymous. Books received (September 1992–August 1993). Conservation Biology, 7(4):968–970, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1994:A

[Ano94a] Anonymous. Announcements. Conservation Biology, 8(3):625– 628, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1994:CCB

[Ano94b] Anonymous. Cattle and conservation biology — another view. Conservation Biology, 8(4):919–921, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1994:E

[Ano94c] Anonymous. Errata. Conservation Biology, 8(3):918, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1995:CPS

[Ano95a] Anonymous. Call for papers for the 1996 SCB meeting. Conservation Biology, 9(5):1342, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1995:E

[Ano95b] Anonymous. Errata. Conservation Biology, 9(2):241, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1995:IHS

[Ano95c] Anonymous. An invitation to host the 1998 SCB meeting. Conservation Biology, 9(5):1342, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1996:B

[Ano96a] Anonymous. Books. Conservation Biology, 10(6):1562, December 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1996:BRS

[Ano96b] Anonymous. Books received (September 1995–September 1996). Conservation Biology, 10(6):1568–1571, December 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1996:F

[Ano96c] Anonymous. Forthcoming. Conservation Biology, 10(6):1562, December 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1996:JSA

[Ano96d] Anonymous. Journal and symposium articles. Conservation Biology, 10(6):1560–1561, December 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1996:SEC

[Ano96e] Anonymous. Selected editorials, commentaries, book reviews, forewords. Conservation Biology, 10(6):1561–1562, December 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1997:BBB

[Ano97a] Anonymous. Beating back the brownlash. Conservation Biology, 11(3):821–826, June 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1997:BR

[Ano97b] Anonymous. Books received. Conservation Biology, 11(6): 1465–1467, December 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1997:FRP

[Ano97c] Anonymous. A forager-resource population ecology model and implications for indigenous conservation. Conservation Biology, 11(6):1270–1274, December 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1997:L

[Ano97d] Anonymous. Letters. Conservation Biology, 11(3):593–596, June 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1997:PCM

[Ano97e] Anonymous. Patches, corridors, and matrices. Conservation Biology, 11(1):290–293, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1997:U

[Ano97f] Anonymous. Unknown. Conservation Biology, 11(5):1265, October 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1998:SCB

[Ano98a] Anonymous. 1998 Society for Conservation Biology. Conservation Biology, 12(6):1428–1430, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1998:ECA

[Ano98b] Anonymous. Editor’s comments and acknowledgments. Conservation Biology, 12(6):1173–1174, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1998:E

[Ano98c] Anonymous. Erratum. Conservation Biology, 12(1):257, February 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1998:PCP

[Ano98d] Anonymous. The potential consequences of pollinator declines on the conservation of biodiversity and stability of food crop yields. Conservation Biology, 12(1):8–17, February 1998. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1999:SAM

[Ano99a] Anonymous. 2000 SCB annual meeting. Conservation Biology, 13(5):1234, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1999:AIV

[Ano99b] Anonymous. Author index, volume 13, 1999. Conservation Biology, 13(6):1534–1538, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1999:BFA

[Ano99c] Anonymous. Back to the future with animals and ancestors. Conservation Biology, 13(5):1229–1230, October 1999. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1999:BRS

[Ano99d] Anonymous. Books received September 1998 to July 1999. Conservation Biology, 13(6):1525–1527, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1999:CBA

[Ano99e] Anonymous. Cash and burn in the Amazon. Conservation Biology, 13(5):1231–1232, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1999:CP

[Ano99f] Anonymous. Conservation for planners. Conservation Biology, 13(4):945–946, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1999:CAP

[Ano99g] Anonymous. A conversation about population. Conservation Biology, 13(4):945, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1999:EA

[Ano99h] Anonymous. Editor’s acknowledgments. Conservation Biology, 13(6):1241, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1999:EPS

[Ano99i] Anonymous. Error propagation in spatially explicit population models: a reassessment. Conservation Biology, 13(4):930–933, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1999:ITW

[Ano99j] Anonymous. If there were a cookbook for conservation. Conservation Biology, 13(5):1228–1229, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1999:NON

[Ano99k] Anonymous. Nature in our neighborhoods. Conservation Biology, 13(4):948–950, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1999:PP

[Ano99l] Anonymous. People and parks. Conservation Biology, 13(4): 947–948, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1999:POC

[Ano99m] Anonymous. Public outreach and conservation scientists. Conservation Biology, 13(4):695, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1999:RBS

[Ano99n] Anonymous. Restoration: Becoming a science. Conservation Biology, 13(4):946–947, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1999:SIV

[Ano99o] Anonymous. Subject index, volume 13, 1999. Conservation Biology, 13(6):1528–1533, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:1999:WFN

[Ano99p] Anonymous. Where few naturalists have been. Conservation Biology, 13(5):1230–1231, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ash:1999:CSF

[AP99] Andrew N. Ash and Kenneth H. Pollock. Clearcuts, salamanders, and field studies. Conservation Biology, 13(1):206–208, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Alford:1997:LEE

[AR97] Ross A. Alford and Stephen J. Richards. Lack of evidence for epidemic disease as an agent in the catastrophic decline of Australian rain forest frogs. Conservation Biology, 11(4):1026– 1029, August 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Arita:1990:RNF

[ARR90] Héctor T. Arita, John G. Robinson, and Kent H. Redford. Rarity in neotropical forest mammals and its ecological correlates. Conservation Biology, 4(2):181–192, June 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Alvard:1997:SSH

[ARRK97] Michael S. Alvard, John G. Robinson, Kent H. Redford, and Hillard Kaplan. The sustainability of subsistence hunting in the neotropics. La sustentabilidad de la caza de subsistencia en el neotrópico. Conservation Biology, 11(4):977–982, August 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Andrade:1994:SUT

[ART94] Germán I. Andrade and Heidi Rubio-Torgler. Sustainable use of the tropical rain forest: Evidence from the avifauna in a shifting-cultivation habitat mosaic in the Colombian Amazon. El uso sostenido dc la selva tropical lluviosa. Evidencia a partir de la fauna avícola en un mosaico de hábitats bajo agricultura migratoria. Conservation Biology, 8(2):545–554, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Aguirre:1994:WDU

[AS94] A. Alonso Aguirre and Edward E. Starkey. Wildlife disease in U.S. National Parks: Historical and coevolutionary perspectives. Enfermedades de la vida silvestre en los Parques Nacionales de EEUU: Perspectivas históricas y coevolutivas. Conservation Biology, 8(3):654–661, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ash:1997:DRP

[Ash97] Andrew N. Ash. Disappearance and return of plethodontid salamanders to clearcut plots in the Southern Blue Ridge Mountains. Desaparición y retorno de salamandras pletodóntidas hacia regiones completamente taladas en las Montañas Blue Ridge del Sur. Conservation Biology, 11(4): 983–989, August 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

August:1995:HTM

[Aug95] Peter V. August. High technology meets ecosystem management. Conservation Biology, 9(5):1341–1342, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Auster:1998:CMI

[Aus98] Peter J. Auster. A conceptual model of the impacts of fishing gear on the integrity of fish habitats. Conservation Biology, 12(6):1198–1203, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Avise:1994:RMB

[Avi94] John C. Avise. The real message from Biosphere 2. Conservation Biology, 8(2):327–329, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Avise:1995:MDP

[Avi95] John C. Avise. Mitochondrial DNA polymorphism and a connection between genetics and demography of relevance to conservation. Conservation Biology, 9(3):686–690, June 1995. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anderson:1998:L

[AWR^{+}98] E. N. Anderson, Brian R. Warren, Jesús A. Rivas, William Z. Lidicker, Jr., Gary K. Meffe, Reed F. Noss, and Susan K. Jacobson. Letters. Conservation Biology, 12(6):1168–1172, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Babbitt:1999:NMB

[Bab99] Bruce Babbitt. Noah’s mandate and the birth of urban bioplanning. Conservation Biology, 13(3):677–678, June 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Burnett:1998:IGH

[BABK98] Michael R. Burnett, Peter V. August, James H. Brown, Jr., and Keith T. Killingbeck. The influence of geomorphological heterogeneity on biodiversity. I. A patch-scale perspective. Conservation Biology, 12(2):363–370, April 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Baker:1994:RLS

[Bak94] William L. Baker. Restoration of landscape structure altered by fire suppression. Restitución de la estructura del paisaje alterada por la supresión del fuego. Conservation Biology, 8 (3):763–769, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Baldwin:1995:CAE

[Bal95] Rob Baldwin. Changes in attitude not enough. Conservation Biology, 9(2):240–241, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Baldwin:1996:ICE

[Bal96] Rob Baldwin. Intensive conservation education. Conservation Biology, 10(5):1475–1476, October 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Baltz:1999:ORI

[Bal99] Michael E. Baltz. Overconsumption of resources in industrial countries: The other missing agenda. Conservation Biology, 13 (1):213–215, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bowen:1993:PSL

[BAR^{+}93] Brian Bowen, John C. Avise, James I. Richardson, Anne B. Meylan, Dimitris Margaritoulis, and Sally R. Hopkins-Murphy. Population structure of loggerhead turtles (Caretta caretta) in the Northwestern Atlantic Ocean and Mediterranean Sea. Conservation Biology, 7(4):834–844, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Balick:1994:DEF

[BAR94] Mchael J. Balick, Rosita Arvigo, and Leopoldo Romero. The development of an ethnobiomedical forest reserve in Belize:

Its role in the preservation of biological and cultural diversity. Conservation Biology, 8(1):316–317, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bart:1995:ASH

[Bar95] Jonathan Bart. Amount of suitable habitat and viability of northern spotted owls. Conservation Biology, 9(4):943–946, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Barrows:1996:EMP

[Bar96] Cameron W. Barrows. An ecological model for the protection of a dune ecosystem. Conservation Biology, 10(3):888–891, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bawa:1992:CRP

[Baw92] Kamaljit S. Bawa. Colonialism, rural poverty and the use of forest resources. Conservation Biology, 6(3):477–481, September 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bawa:1995:AOT

[Baw95a] Kamaljit S. Bawa. Another outstanding text. Conservation Biology, 9(1):222–223, February 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bawa:1995:BCD

[Baw95b] Kamaljit S. Bawa. Biodiversity convention and developing countries. Conservation Biology, 9(3):699, June 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bildstein:1990:LLT

[BB90] Keith L. Bildstein and I. Lehr Brisbin, Jr. Lands for long-term research in conservation biology. Conservation Biology, 4(3): 301–308, September 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bock:1993:CPG

[BB93] Carl E. Bock and Jane H. Bock. Cover of perennial grasses in Southeastern Arizona in relation to livestock grazing. Cubierta de pastizales perennes en el sudeste de Arizona en relación con el pastoreo del ganado. Conservation Biology, 7(2):371–377, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Belsky:1997:ELG

[BB97] A. Joy Belsky and Dana M. Blumenthal. Effects of livestock grazing on stand dynamics and soils in upland forests of the Interior West. Efectos del pastoreo sobre la dinámica de Árboles y suelos en bosques en el altiplano del occidente interior. Conservation Biology, 11(2):315–327, April 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Botsford:1998:VSR

[BB98] Louis W. Botsford and John G. Brittnacher. Viability of Sacramento River winter-run Chinook salmon. Conservation Biology, 12(1):65–79, February 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bock:1999:RWB

[BB99] Carl E. Bock and Jane H. Bock. Response of winter birds to drought and short-duration grazing in Southeastern Arizona. Respuesta de aves de invierno a la sequía y pastoreo de corta duración en el sureste de Arizona. Conservation Biology, 13(5): 1117–1123, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bowen:1996:ND

[BBA^{+}96] B. W. Bowen, A. L. Bass, John M. Aguiar, Walt Anderson, David B. Marshall, and Reed F. Noss. Are the naturalists dying off? Conservation Biology, 10(4):923–927, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bryant:1999:ETC

[BBCR99] Edwin H. Bryant, Vickie L. Backus, Michael E. Clark, and David H. Reed. Experimental tests of captive breeding for endangered species. Pruebas experimentales de reproducción en cautiverio para especies en peligro de extinción. Conservation Biology, 13(6):1487–1496, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bodkin:1999:PDG

[BBCS99] J. L. Bodkin, B. E. Ballachey, M. A. Cronin, and K. T. Scribner. Population demographics and genetic diversity in remnant and translocated populations of sea otters. Demografía poblacional y diversidad genética en poblaciones remanentes e introducidas de nutrias marinas. Conservation Biology, 13(6): 1378–1385, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bartley:1992:ULD

[BBGB92] Devin Bartley, Mark Bagley, Graham Gall, and Boyd Bentley. Use of linkage disequilibrium data to estimate effective size of hatchery and natural fish populations. Conservation Biology, 6(3):365–375, September 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Backus:1995:EMI

[BBHM95] Vickie L. Backus, Edwin H. Bryant, Colin R. Hughes, and Lisa M. Meffert. Effect of migration or inbreeding followed by selection on low-founder-number populations: Implications for captive breeding programs. Conservation Biology, 9(5): 1216–1224, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bjorndal:1993:DNP

[BBL93] Karen A. Bjorndal, Alan B. Bolten, and Cynthia J. Lagueux. Decline of the nesting population of hawksbill turtles at Tortuguero, Costa Rica. Conservation Biology, 7(4):925–927, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Britten:1994:MTP

[BBM94a] Hugh B. Britten, Peter F. Brussard, and Dennis D. Murphy. More on talking to the press. Conservation Biology, 8(4): 929–930, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Britten:1994:PEU

[BBM94b] Hugh B. Britten, Peter F. Brussard, and Dennis D. Murphy. The pending extinction of the Uncompahgre fritillary butterfly. La extincion pendiente de la mariposa Boloria acrocnema. Conservation Biology, 8(1):86–94, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See response [SO94].

Britten:1994:PTU

[BBM94c] Hugh B. Britten, Peter F. Brussard, and Dennis D. Murphy. Population trends in the Uncompahgre fritillary butterfly: Reply to Seidl and Opler. Conservation Biology, 8(4):1158–1160, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [SO94] and reply [Ehr95].

Bock:1993:PSF

[BBS93] Carl E. Bock, Jane H. Bock, and Hobart M. Smith. Proposal for a system of federal livestock exclosures on public rangelands in the Western United States. Conservation Biology, 7 (3):731–733, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bradstock:1996:SES

[BBSK96] R. A. Bradstock, M. Bedward, J. Scott, and D. A. Keith. Simulation of the effect of spatial and temporal variation in fire regimes on the population viability of a Banksia species. Simulación del efecto de la variación espacial y temporal en regímenes de incendios, sobre la viabilidad de la población de una especie de Banksia. Conservation Biology, 10(3):776–784, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berger:1994:PAE

[BC94] Joel Berger and Carol Cunningham. Phenotypic alterations, evolutionarily significant structures, and rhino conservation. Alteraciones fenotípicas, estructuras evolutivamente significativas y conservación de rinocerontes. Conservation Biology, 8 (3):833–840, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bentley:1997:UBC

[BC97a] Jocelyn M. Bentley and Carla P. Catterall. The use of bushland, corridors, and linear remnants by birds in southeastern Queensland, Australia. El uso de vegetación primaria, corredores y restos lineares por aves en el sudeste de Queensland, Australia. Conservation Biology, 11(5):1173–1189, October 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brunner:1997:PBA

[BC97b] Ronald D. Brunner and Tim W. Clark. A practice-based approach to ecosystem management. Aproximación al manejo de ecosistemas basada en la práctica. Conservation Biology, 11 (1):48–58, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berger:1998:NVH

[BC98] Joel Berger and Carol Cunningham. Natural variation in Horn size and social dominance and their importance to the conservation of black rhinoceros. Conservation Biology, 12(3):708– 711, June 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berger:1999:RSB

[BC99] Joel Berger and Steven L. Cain. Reproductive synchrony in brucellosis-exposed bison in the Southern Greater Yellowstone ecosystem and in noninfected populations. Sincrinización reproductiva de bisontes expuestos a brucelosis en el ecosistema del Sur del Yellowstone Mayor y en poblacioines no infectadas. Conservation Biology, 13(2):357–366, April 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berger:1994:UDD

[BCG94] Joel Berger, Carol Cunningham, and A. Archie Gawuseb. The uncertainty of data and dehorning black rhinos. Conservation Biology, 8(4):1149–1152, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berger:1993:CST

[BCGL93] Joel Berger, Carol Cunningham, A. Archie Gawuseb, and Malan Lindeque. “Costs” and short-term survivorship of hornless black rhinos. Conservation Biology, 7(4):920–924, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bowen:1994:WTN

[BCHM94] B. W. Bowen, T. A. Conant, and S. R. Hopkins-Murphy. Where are they now? The Kemp’s Ridley Headstart Project. Conservation Biology, 8(3):853–856, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bouzat:1998:GED

[BCL^{+}98] Juan L. Bouzat, Hans H. Cheng, Harris A. Lewin, Ronald L. Westemeier, Jeffrey D. Brawn, and Ken N. Paige. Genetic evaluation of a demographic bottleneck in the greater prairie chicken. Conservation Biology, 12(4):836–843, August 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Butman:1995:WED

[BCP95] Cheryl Ann Butman, James T. Carlton, and Stephen R. Palumbi. Whaling effects on deep-sea biodiversity. Conservation Biology, 9(2):462–464, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Butman:1996:WDF

[BCP96] Cheryl Ann Butman, James T. Carlton, and Stephen R. Palumbi. Whales don’t fall like snow: Reply to Jelmert. Conservation Biology, 10(2):655–656, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bokdam:1992:FQL

[BD92] Jan Bokdam and Michiel F. De Vries Wallis. Forage quality as a limiting factor for cattle grazing in isolated Dutch nature reserves. Conservation Biology, 6(3):399–408, September 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bookbinder:1998:ESB

[BDR^{+}98] Marnie P. Bookbinder, Eric Dinerstein, Arun Rijal, Hank Cauley, and Arup Rajouria. Ecotourism’s support of biodiversity conservation. Conservation Biology, 12(6):1399–1404, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Birstein:1998:PAA

[BDSD98] Vadim J. Birstein, Phaedra Doukakis, Boris Sorkin, and Rob DeSalle. Population aggregation analysis of three caviarproducing species of sturgeons and implications for the species identification of black caviar. Conservation Biology, 12(4):766– 775, August 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bronmark:1994:DPF

[BE94] Christer Brönmark and Per Edenhamn. Does the presence of fish affect the distribution of tree frogs (Hyla arborea)? Conservation Biology, 8(3):841–845, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berg:1994:TPA

[BEG^{+}94] Åke Berg, Beng Ehnström, Lena Gustafsson, Tomas Hallingbäck, Mats Jonsell, and Jan Weslien. Threatened plant, animal, and fungus species in Swedish forests: Distribution and habitat associations. Especies de plantas, animales, y hongos en peligro en bosques suecos: distribución y asociacines de hábitats. Conservation Biology, 8(3):718–731, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berg:1995:TLT

[BEG^{+}95] Åke Berg, Bengt Ehnström, Lena Gustafsson, Tomas Hallingbäck, Mats Jonsell, and Jan Weslien. Threat levels and threats to red-listed species in Swedish forests. Conservation Biology, 9 (6):1629–1633, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beier:1993:DMH

[Bei93] Paul Beier. Determining minimum habitat areas and habitat corridors for cougars. Determinando áreas mínimas de hábitat y hábitat en corredes para pumas. Conservation Biology, 7 (1):94–108, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Benkman:1993:LCC

[Ben93] Craig W. Benkman. Logging, conifers, and the conservation of crossbills. Explotación forestal, coníferas y conservación de piquituertos (“crossbills”). Conservation Biology, 7(3):473–479, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berger:1990:PDS

[Ber90] Joel Berger. Persistence of different-sized populations: an empirical assessment of rapid extinctions in bighorn sheep. Conservation Biology, 4(1):91–98, March 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berger:1991:GYN

[Ber91] Joel Berger. Greater Yellowstone’s native ungulates: Myths and realities. Conservation Biology, 5(3):353–363, September 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berger:1993:PMS

[Ber93] Joel Berger. Persistence of mountain sheep: Methods and statistics. Conservation Biology, 7(2):219–220, June 1993. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berger:1995:MEP

[Ber95a] Joel Berger. The most endangered Proboscidean. Conservation Biology, 9(5):1339–1340, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bertram:1995:RIR

[Ber95b] Douglas F. Bertram. The roles of introduced rats and commercial fishing in the decline of ancient murrelets on Langara Island, British Columbia. El papel de las ratas introducidas y la pesca comercial en la declinación de Synthliboramphus antiquus en la Isla Langara, en la Columbia Británica. Conservation Biology, 9(4):865–872, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berger:1997:PCA

[Ber97a] Joel Berger. Population constraints associated with the use of black rhinos as an umbrella species for desert herbivores. Restricción de poblaciones asociadas con el uso de rinocerontes negros como “Sombrilla de Especies” para herbivoros del desierto. Conservation Biology, 11(1):69–78, February 1997. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bodmer:1997:HLE

[BER97b] Richard E. Bodmer, John F. Eisenberg, and Kent H. Redford. Hunting and the likelihood of extinction of Amazonian mammals. Caza y probabilidad de extinción de mamiferos Amazónicos. Conservation Biology, 11(2):460–466, April 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berger:1999:IPS

[Ber99] Joel Berger. Intervention and persistence in small populations of bighorn sheep. Conservation Biology, 13(2):432–435, April 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bestelmeyer:1996:SLF

[Bes96] Brandon Bestelmeyer. Six-legged future. Conservation Biology, 10(5):1484–1485, October 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Buchanan:1998:MHT

[BFS98] Joseph B. Buchanan, Richard J. Fredrickson, and D. Erran Seaman. Mitigation of habitat “Take” and the core area concept. Conservation Biology, 12(1):238–240, February 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beissinger:1995:MSC

[BG95a] Steven R. Beissinger and James P. Gibbs. Monitoring and science: Comfortable bedfellows. Conservation Biology, 9(2):465– 467, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Burke:1995:TBZ

[BG95b] Vincent J. Burke and J. Whitfield Gibbons. Terrestrial buffer zones and wetland conservation: a case study of freshwater turtles in a Carolina bay. Conservacion de zonas buffer terrestres y humedales: un estudio de caso de las tortugas de agua dulce en la bahía de Carolina. Conservation Biology, 9(6):1365–1369, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bohning-Gaese:1996:CSA

[BGB96] Katrin Böhning-Gaese and Hans-Günther Bauer. Changes in species abundance, distribution, and diversity in a Central European bird community. Cambios en la abundancia de especies, distribución y diversidad en una comunidad de aves de Europa central. Conservation Biology, 10(1):175–187, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Burgman:1995:ITS

[BGF95] Mark A. Burgman, Roger C. Grimson, and Scott Ferson. Inferring threat from scientific collections. Infiriendo la amenaza de extinción a partir de colecciones científicas. Conservation Biology, 9(4):923–928, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brookes:1997:GAF

[BGK^{+}97] Martin I. Brookes, Yvonne A. Graneau, Peter King, Owen C. Rose, Chris D. Thomas, and James L. B. Mallet. Genetic analysis of founder bottlenecks in the rare British butterfly Plebejus argus. Análisis genético de cuellos de botella en la mariposa Británica Plebejus argus. Conservation Biology, 11(3):648–661, June 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bohning-Gaese:1993:DNA

[BGTB93] Katrin Böhning-Gaese, Mark L. Taper, and James H. Brown. Are declines in North American insectivorous songbirds due to causes on the breeding range? `?Son las declinaciones de las aves cantoras insectívoras de Norte América debidas a causas que afectan las áreas de reproducción? Conservation Biology, 7(1):76–86, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bootsma:1993:CAG

[BH93] Harvey A. Bootsma and Robert E. Hecky. Conservation of the African Great Lakes: a limnological perspective. Conservación de los Grandes Lagos Africanos: Una perspectiva limnológica. Conservation Biology, 7(3):644–656, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bruinderink:1996:UTC

[BH96] G. W. T. A. Groot Bruinderink and E. Hazebroek. Ungulate traffic collisions in Europe. Colisiones de tráfico con ungulados en Europa. Conservation Biology, 10(4):1059–1067, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bayne:1997:CEL

[BH97] Erin M. Bayne and Keith A. Hobson. Comparing the effects of landscape fragmentation by forestry and agriculture on predation of artificial nests. Conservation Biology, 11(6):1418–1429, December 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Baxter:1999:TWM

[BHCB99] G. S. Baxter, M. Hockings, R. W. Carter, and R. J. S. Beeton. Trends in wildlife management and the appropriateness of Australian university training tendencias en el manejo de vida silvestre y la aptitud del entrenamiento en universidades Australianas. Conservation Biology, 13(4):842–849, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Blaustein:1996:DRA

[BHKH96] Andrew R. Blaustein, Peter D. Hoffman, Joseph M. Kiesecker, and John B. Hays. DNA repair activity and resistance to solar UV-B radiation in eggs of the red-legged frog. Reparación de ADN y resistencia a radiación solar UV-B en huevos de rana de piernas rojas. Conservation Biology, 10(5):1398–1402, October 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berman:1992:RIG

[BHL^{+}92] Jody Berman, Larry Harris, Walter Lambert, Melanie Buttrick, and Michael Dufresne. Recent invasions of the Gulf of Maine: Three contrasting ecological histories. Conservation Biology, 6 (3):435–441, September 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bratton:1994:LSR

[BHM94] Susan P. Bratton, Jeffrey R. Hapeman, and Austin R. Mast. The Lower Susquehanna River Gorge and floodplain (U.S.A.) as a riparian refugium for vernal, forest-floor herbs. La garganta y las llanuras de inundación del río Susquehanna inferior (E.U.A.) consideradas como un refugio ribereño para hierbas primaverales del suelo de los bosques. Conservation Biology, 8 (4):1069–1077, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Busing:1995:EPY

[BHS95] Richard T. Busing, Charles B. Halpern, and Thomas A. Spies. Ecology of Pacific yew (Taxus brevifolia) in Western Oregon and Washington. Conservation Biology, 9(5):1199–1207, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beard:1999:EPB

[BHS99] Karen H. Beard, Nicolas Hengartner, and David K. Skelly. Effectiveness of predicting breeding bird distributions using probabilistic models. Efectividad en la predicción de distribuciones de aves reproductoras utilizando modelos probabilísticos. Conservation Biology, 13(5):1108–1116, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bancroft:1992:IWC

[BHSO92] G. Thomas Bancroft, Wayne Hoffman, Richard J. Sawicki, and John C. Ogden. The importance of the water conservation areas in the Everglades to the endangered wood stork (Mycteria americana). Conservation Biology, 6(3):392–398, September 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bierregaard:1995:MTF

[Bie95] Richard O. Bierregaard, Jr. A model tropical field station. Conservation Biology, 9(1):221–222, February 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Billington:1991:EPS

[Bil91] H. L. Billington. Effect of population size on genetic variation in a dioecious conifer. Conservation Biology, 5(1):115–119, March 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Birstein:1993:SPT

[Bir93] Vadim J. Birstein. Sturgeons and paddlefishes: Threatened fishes in need of conservation. Conservation Biology, 7(4): 773–787, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bissell:1992:TAV

[Bis92] Steven J. Bissell. Together again for the very first time. Conservation Biology, 6(4):608–610, December 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bissell:1993:FLF

[Bis93] Steven J. Bissell. Feminists, Luddites, and full-contact philosophy. Conservation Biology, 7(3):738–739, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bissell:1994:BWP

[Bis94a] Steven J. Bissell. Broadening wildlife policy. Conservation Biology, 8(1):323–324, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bissell:1994:REE

[Bis94b] Steven J. Bissell. Reading in environmental ethics. Conservation Biology, 8(3):917–918, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bunin:1995:NAT

[BJ95] Judah S. Bunin and Ian G. Jamieson. New approaches toward a better understanding of the decline of takahe (Porphyrio mantelli) in New Zealand. Nuevas propuestas para una mejor comprensión de la declinación del takahe (Porphyrio mantelli) en Nueva Zelandia. Conservation Biology, 9(1):100–106, February 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bunin:1996:RMP

[BJ96] Judah S. Bunin and Ian G. Jamieson. Responses to a model predator of New Zealand’s endangered takahe and its closest relative, the pukeko. Conservation Biology, 10(5):1463–1466, October 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Boza:1995:CRL

[BJW95] Mario A. Boza, Diane Jukofsky, and Chris Wille. Costa Rica is a laboratory, not ecotopia. Conservation Biology, 9(3):684– 685, June 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bowen:1999:WTF

[BK99] Brian W. Bowen and Stephen A. Karl. In war, truth is the first casualty*. Conservation Biology, 13(5):1013–1016, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bengston:1998:REE

[BKC^{+}98] David N. Bengston, Thomas H. Kunz, Philip Cafaro, Jennifer Smith, Lisa L. Manne, and Mary Meagher. Reintegrating ecology and economics. Conservation Biology, 12(4):933–937, August 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brush:1995:PDA

[BKO^{+}95] Stephen Brush, Rick Kesseli, Ramiro Ortega, Pedro Cisneros, Karl Zimmerer, and Carlos Quiros. Potato diversity in the Andean center of crop domestication. Conservation Biology, 9 (5):1189–1198, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Balmford:1995:ACA

[BL95a] Andrew Balmford and Adrian Long. Across-country analyses of biodiversity congruence and current conservation effort in the tropics. Análisis entre paises de la congruencia en la biodiversidad y esfuerzos actuales de conservación en los trópicos. Conservation Biology, 9(6):1539–1547, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Boucher:1995:EBM

[BL95b] Guy Boucher and P. John D. Lambshead. Ecological biodiversity of marine nematodes in samples from temperate, tropical, and deep-sea regions. Biodiversidad ecológica de los nematodes marinos en muestras de regiones templadas, tropicales y de aguas profundas. Conservation Biology, 9(6):1594–1604, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bevill:1999:CRR

[BL99] R. L. Bevill and S. M. Louda. Comparisons of related rare and common species in the study of plant rarity. Comparaciones de especies raras y comunes relacionadas entre sí en estudios de rareza de plantas. Conservation Biology, 13(3):493–498, June 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Blake:1991:NSD

[Bla91] John G. Blake. Nested subsets and the distribution of birds on isolated woodlots. Conservation Biology, 5(1):58–66, March 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Blake:1993:RBW

[Bla93] Greg Blake. A response to Brian Walker. Conservation Biology, 7(1):5–6, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Black:1994:RCF

[Bla94] Michael Black. Recounting a century of failed fishery policy toward California’s Sacramento River salmon and steelhead. Conservation Biology, 8(3):892–894, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Breininger:1995:LPF

[BLD^{+}95] David R. Breininger, Vickie L. Larson, Brean W. Duncan, Rebecca B. Smith, Donna M. Oddy, and Michael F. Goodchild. Landscape patterns of Florida scrub jay habitat use and demographic success. Patrones a nivel paisajístico del uso del hábitat y exito demográfico del arrendajo del chaparral de Florida. Conservation Biology, 9(6):1442–1453, December 1995. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Borlase:1993:MPC

[BLF^{+}93] Suzanne C. Borlase, David A. Loebel, Richard Frankham, Roderick K. Nurthen, David A. Briscoe, and Grant E. Daggard. Modeling problems in conservation genetics using captive Drosophila populations: Consequences of equalization of family sizes. Modelación de problemas en conservación genética usando poblaciones cautivas de drosophila: concequencias de la igualación del tamaño familiar. Conservation Biology, 7 (1):122–131, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Blockstein:1999:ISE

[Blo99a] David E. Blockstein. Integrated science for ecosystem management: an achievable imperative. Conservation Biology, 13 (3):682–685, June 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Blockstein:1999:RBR

[Blo99b] David E. Blockstein. Representative biological reserves. Conservation Biology, 13(6):1239, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bevill:1999:PNE

[BLS99] R. L. Bevill, S. M. Louda, and L. M. Stanforth. Protection from natural enemies in managing rare plant species. Protección contra enemigos naturales en el manejo de especies raras de plantas. Conservation Biology, 13(6):1323–1331, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Blumstein:1993:MHP

[Blu93] Daniel T. Blumstein. Making humans part of the solution: a reply to Willers. Conservation Biology, 7(2):223b–224, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Balick:1992:AEV

[BM92] Michael J. Balick and Robert Mendelsohn. Assessing the economic value of traditional medicines from tropical rain forests. Conservation Biology, 6(1):128–130, March 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brown:1995:LGC

[BM95] James H. Brown and William McDonald. Livestock grazing and conservation on southwestern rangelands. Conservation Biology, 9(6):1644–1647, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brown:1997:HCP

[BM97] James H. Brown and William McDonald. Historical and cultural perspectives on grazing: Reply to Dudley. Conservation Biology, 11(1):270–272, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Benitez-Malvido:1998:IFF

[BM98] Julieta Benitez-Malvido. Impact of forest fragmentation on seedling abundance in a tropical rain forest. Conservation Biology, 12(2):380–389, April 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Balmford:1996:DAS

[BMLW96] Andrew Balmford, Georgina M. Mace, and N. Leader-Williams. Designing the ark: Setting priorities for captive breeding. Diseñando el arca: Establecimiento de prioridades para la reproducción en cautiverio. Conservation Biology, 10(3):719–727, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brussard:1992:STC

[BMN92] Peter F. Brussard, Dennis D. Murphy, and Reed F. Noss. Strategy and tactics for conserving biological diversity in the United States. Conservation Biology, 6(2):157–159, June 1992. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Briscoe:1992:RLG

[BMR^{+}92] D. A. Briscoe, J. M. Malpica, A. Robertson, Gabrielle J. Smith, R. Frankham, R. G. Banks, and J. S. F. Barker. Rapid loss of genetic variation in large captive populations of Drosophila flies: Implications for the genetic management of captive populations. Conservation Biology, 6(3):416–425, September 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bingham:1997:MHT

[BN97] Bruce B. Bingham and Barry R. Noon. Mitigation of habitat “Take”: Application to habitat conservation planning. Aplicación de la mitigación de la “Toma” de hábitat en la planeación de conservación de hábitat. Conservation Biology, 11(1):127–139, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beier:1998:DHC

[BN98a] Paul Beier and Reed F. Noss. Do habitat corridors provide connectivity? Conservation Biology, 12(6):1241–1252, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bingham:1998:UCA

[BN98b] Bruce B. Bingham and Barry R. Noon. The use of core areas in comprehensive mitigation strategies. Conservation Biology, 12(1):241–243, February 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Burkey:1998:LUE

[BNF^{+}98] Tormod V. Burkey, Jari Niemelä, N. N. FitzSimmons, Jeff Langholz, Sharon J. Downes, and T. R. McClanahan. Life, the universe, and everything. Conservation Biology, 12(2):488– 495, April 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Barry:1996:SSV

[BO96] Dwight Barry and Max Oelschlaeger. A science for survival: Values and conservation biology una ciencia para la supervivencia: Valores y biologia de la conservacion. Conservation Biology, 10(3):905–911, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beattie:1999:BBA

[BO99] Andrew J. Beattie and Ian Oliver. Biodiversity buzzwords: Another reply to Goldstein. Conservation Biology, 13(6):1514, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [Gol99b, Gol99a].

Botsford:1994:EPD

[Bot94] Louis W. Botsford. Extinction probabilities and delisting criteria for Pacific salmonids. Conservation Biology, 8(3):873–875, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Boza:1993:CAP

[Boz93] Mario A. Boza. Conservation in action: Past, present, and future of the National Park System of Costa Rica. Conservatión en acción: pasado, presente y futuro del Sistema de Parques Nacionales de Costa Rica. Conservation Biology, 7(2):239–247, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brooks:1997:DPN

[BPC97] Thomas M. Brooks, Stuart L. Pimm, and Nigel J. Collar. Deforestation predicts the number of threatened birds in insular Southeast Asia. La extensión de las deforestaciones predice el número de aves amenazadas de extinción en las islas del sureste de Asia. Conservation Biology, 11(2):382–394, April 1997. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brooks:1999:TLB

[BPO99] Thomas M. Brooks, Stuart L. Pimm, and Joseph O. Oyugi. Time lag between deforestation and bird extinction in tropical forest fragments. Espacio temporal entre la deforestación y la extinción de especies de aves en fragmentos de bosques tropicales. Conservation Biology, 13(5):1140–1150, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Buell:1995:IHI

[BPS95] Ann C. Buell, Andrea J. Pickart, and John D. Stuart. Introduction history and invasion patterns of Ammophila arenaria on the North Coast of California. Historia de la introducción y patrones de invasión de Ammophila arenaria en la Costa Norte de California. Conservation Biology, 9(6):1587–1593, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Britten:1996:PSS

[BR96] Hugh B. Britten and Richard W. Rust. Population structure of a sand dune — obligate beetle, Eusattus muricatus, and its implications for dune management. Conservation Biology, 10 (2):647–652, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bryant:1999:FDU

[BR99] Edwin H. Bryant and David H. Reed. Fitness decline under relaxed selection in captive populations. Disminución de la condición de poblaciones cautivas bajo selección relajada. Conservation Biology, 13(3):665–669, June 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bratton:1994:LFB

[Bra94] Susan P. Bratton. Logging and fragmentation of broadleaved deciduous forests: Are we asking the right ecological questions? Conservation Biology, 8(1):295–297, March 1994. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brand:1996:WCU

[Bra96] Terrel Brand. Web of causality underlies destruction. Conservation Biology, 10(5):1482–1484, October 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brailovskaya:1998:OPM

[Bra98] Tatiana Brailovskaya. Obstacles to protecting marine biodiversity through marine wilderness preservation: Examples from the New England region. Conservation Biology, 12(6):1236– 1240, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brennan:1994:GNL

[Bre94] Andrew Brennan. A genetic need to love nature? Conservation Biology, 8(3):916–917, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Buchert:1997:EHG

[BRHD97] George P. Buchert, Om P. Rajora, James V. Hood, and Bruce P. Dancik. Effects of harvesting on genetic diversity in old-growth eastern white pine in Ontario, Canada. Efecto de la cosecha sobre la diversidad genética de un bosque maduro de pino blanco del este. Conservation Biology, 11(3):747–758, June 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brisbin:1995:CWA

[Bri95a] I. Lehr Brisbin, Jr. Conservation of the wild ancestors of domestic animals. Conservation Biology, 9(5):1327–1328, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Britten:1995:SLN

[Bri95b] Hugh B. Britten. Six-legged needs. Conservation Biology, 9(5): 1338–1339, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Briggs:1996:TDC

[Bri96] John C. Briggs. Tropical diversity and conservation. Diversidad tropical y conservación. Conservation Biology, 10(3):713–718, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brownlow:1996:MTC

[Bro96] C. Alexander Brownlow. Molecular taxonomy and the conservation of the red wolf and other endangered carnivores. Taxonomía molecular y la conservación del lobo colorado y otras especies de carnívoros en peligro. Conservation Biology, 10 (2):390–396, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brussard:1990:A

[Bru90] Peter F. Brussard. Announcement. Conservation Biology, 4 (1):4, March 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brussard:1991:GYE

[Bru91] Peter F. Brussard. The Greater Yellowstone Ecosystem. Conservation Biology, 5(3):336–337, September 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brussard:1993:NSa

[Bru93a] Peter F. Brussard. News of the Society. Conservation Biology, 7(3):450, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brussard:1993:NSb

[Bru93b] Peter F. Brussard. News of the Society. Conservation Biology, 7(4):746, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brothers:1992:FFA

[BS92] Timothy S. Brothers and Arthur Spingarn. Forest fragmentation and alien plant invasion of Central Indiana old-growth forests. Conservation Biology, 6(1):91–100, March 1992. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bawa:1998:NFM

[BS98] Kamaljit S. Bawa and Reinmar Seidler. Natural forest management and conservation of biodiversity in tropical forests. Conservation Biology, 12(1):46–55, February 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bancroft:1995:DEF

[BSC95] G. Thomas Bancroft, Allan M. Strong, and Mary Carrington. Deforestation and its effects on forest-nesting birds in the Florida Keys. Deforestación y sus efectos sobre las aves nidificadoras en bosques de los Cayos de la Florida. Conservation Biology, 9(4):835–844, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bolger:1997:BBA

[BSR97] Douglas T. Bolger, Thomas A. Scott, and John T. Rotenberry. Breeding bird abundance in an urbanizing landscape in coastal Southern California: Abundancia de aves reproductoras en un paisaje urbanizado en la Costa Sur de California. Conservation Biology, 11(2):406–421, April 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Beissinger:1996:NMA

[BSW^{+}96] Steven R. Beissinger, Eric C. Steadman, Timothy Wohlgenant, Geoffrey Blate, and Steve Zack. Null models for assessing ecosystem conservation priorities: Threatened birds as titers of threatened ecosystems in South America. Modelos nulos para la evaluación de prioridades de conservación: aves amenazadas como indicadoras de ecosistemas amenazados en sudamérica. Conservation Biology, 10(5):1343–1352, October 1996. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Burkett:1994:WAH

[BT94] Douglas W. Burkett and Bruce C. Thompson. Wildlife association with human-altered water sources in semiarid vegetation communities. Asociación entre vida silvestre y fuentes de agua alteradas por actividades humanas en comunidades de vegetación semi-árida. Conservation Biology, 8(3):682–690, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bradford:1993:IRP

[BTG93] David F. Bradford, Farinaz Tabatabai, and David M. Graber. Isolation of remaining populations of the native frog, Rana muscosa, by introduced fishes in Sequoia and Kings Canyon National Parks, California. Conservation Biology, 7(4):882– 888, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berger:1999:CEN

[BTRM99] Joel Berger, J. Ward Testa, Tom Roffe, and Steven L. Monfort. Conservation endocrinology: a noninvasive tool to understand relationships between carnivore colonization and ecological carrying capacity. Endocrinología en conservación: una herramienta no invasiva para entender las relaciones entre colonizacíon de carnívoros, y la capacidad de carga ecológica.

Conservation Biology, 13(5):980–989, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bunnell:1995:FDV

[Bun95] Fred L. Bunnell. Forest-dwelling vertebrate faunas and natural fire regimes in British Columbia: Patterns and implications for conservation. Las faunas de vertebrados que habitan los bosques y los régimenes de incendios naturales en la Columbia Británica: patrones y connotaciones para la conservación. Conservation Biology, 9(3):636–644, June 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Burke:1990:CWR

[Bur90] Russell L. Burke. Conservation of the world’s rarest tortoise. Conservation Biology, 4(2):122–126, June 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Burkey:1995:ERA

[Bur95] Tormod Vaaland Burkey. Extinction rates in archipelagoes: Implications for populations in fragmented habitats. Las tasas de extinción en archipelagos: consecuencias para poblaciones en habitats fragmentados. Conservation Biology, 9(3):527–541, June 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Buskirk:1992:CCF

[Bus92] Steven W. Buskirk. Conserving circumboreal forests for martens and fishers. Conservation Biology, 6(3):318–323, September 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bustamante:1996:PVA

[Bus96] Javier Bustamante. Population viability analysis of captive and released bearded vulture populations. Análisis de viabilidad de poblaciones para las poblaciones cautiva y liberada de quebrantahuesos. Conservation Biology, 10(3):822–831, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bowen:1997:IEP

[BV97] Lizabeth Bowen and Dirk Van Vuren. Insular endemic plants lack defenses against herbivores. Conservation Biology, 11(5):

1249–1254, October 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Berger:1991:CMP

[BW91] Joel Berger and John D. Wehausen. Consequences of a mammalian predator–prey disequilibrium in the Great Basin desert. Conservation Biology, 5(2):244–248, June 1991. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bawa:1992:WSS

[BW92] Kamaijit S. Bawa and H. Garrison Wilkes. Who shall speak for biodiversity? Conservation Biology, 6(3):473–474, September 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bjorndal:1999:TSY

[BWBM99] Karen A. Bjorndal, Jerry A. Wetherall, Alan B. Bolten, and Jeanne A. Mortimer. Twenty-six years of green turtle nesting at Tortuguero, Costa Rica: an encouraging trend. Conservation Biology, 13(1):126–134, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bleich:1990:DDM

[BWH90] V. C. Bleich, J. D. Wehausen, and S. A. Holl. Desert-dwelling mountain sheep: Conservation implications of a naturally fragmented distribution. Conservation Biology, 4(4):383–390, December 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Blaustein:1994:ADJ

[BWS94] Andrew R. Blaustein, David B. Wake, and Wayne P. Sousa. Amphibian declines: Judging stability, persistence, and susceptibility of populations to local and global extinctions. Declinación de anfibios: Juzgando estabilidad, persistencia y susceptibilidad de las poblaciones a las extinciones globales. Conservation Biology, 8(1):60–71, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bartlein:1997:FCY

[BWS97] Patrick J. Bartlein, Cathy Whitlock, and Sarah L. Shafer. Future climate in the Yellowstone National Park region and its potential impact on vegetation. Clima futuro en la región del Parque Nacional de Yellowstone y su potencial impacto sobre la vegetación. Conservation Biology, 11(3):782–792, June 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Balcom:1996:MLC

[BY96] Betsie J. Balcom and Richard H. Yahner. Microhabitat and landscape characteristics associated with the threatened allegheny woodrat. Las características del micro hábitat y del paisaje asociadas con las poblaciones amenazada de Neotoma magister. Conservation Biology, 10(2):515–525, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carlson:1992:TDI

[CA92] Allan Carlson and Gustaf Aulén. Territorial dynamics in an isolated white-backed woodpecker (Dendrocopos leucotos) population. Conservation Biology, 6(3):450–454, September 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Currens:1998:CPS

[CAB^{+}98] Kenneth P. Currens, Fred W. Allendorf, David Bayles, Daniel L. Bottom, Christopher A. Frissell, David Hankin, James A. Lichatowich, Patrick C. Trotter, and Thomas A. Williams. Conservation of Pacific salmon: Response to Wainwright and Waples. Conservation Biology, 12(5):1148–1149, October 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cafaro:1995:PGW

[Caf95] Philip Cafaro. A philosopher gone wild. Conservation Biology, 9(4):965–966, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cafaro:1996:MM

[Caf96] Philip Cafaro. To market, to market. Conservation Biology, 10 (2):687–688, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ceska:1997:DSS

[CAH97] Jennifer F. Ceska, James M. Affolter, and J. L. Hamrick. Developing a sampling strategy for Baptisia arachnifera based on allozyme diversity. Desarrollo de una estrategia de muestreo para Baptisia arachnifera basada en diversidad de alozimas. Conservation Biology, 11(5):1133–1139, October 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Callicott:1990:SCT

[Cal90a] J. Baird Callicott. Standards of conservation: Then and now. Conservation Biology, 4(3):229–232, September 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Callicott:1990:WCE

[Cal90b] J. Baird Callicott. Whither conservation ethics? Conservation Biology, 4(1):15–20, March 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Camhi:1995:IFT

[Cam95] Merry Camhi. Industrial fisheries threaten ecological integrity of the Galapagos Islands. Conservation Biology, 9(4):715–719, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cannon:1996:WCR

[Can96] John R. Cannon. Whooping crane recovery: a case study in public and private cooperation in the conservation of endangered species. Recuperación de la grulla americana: un caso de estudio sobre la cooperación pública y privada en la conservación de especies en peligro de extinción. Conservation Biology, 10(3):813–821, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carey:1993:HCC

[Car93] Cynthia Carey. Hypothesis concerning the causes of the disappearance of boreal toads from the mountains of Colorado. Hipótesis acerca de las causas de la desaparición de sapos boreales de las montañas de Colorado. Conservation Biology, 7 (2):355–362, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carey:1994:MTR

[Car94] Cynthia Carey. A matter of time: Response to Aguirre. Conservation Biology, 8(1):7–8, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carlton:1998:AO

[Car98] James T. Carlton. Apostrophe to the ocean. Conservation Biology, 12(6):1165–1167, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Caughley:1993:EE

[Cau93] Graeme Caughley. Elephants and economics. Conservation Biology, 7(4):943–945, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See comments [HLAS96].

Clark:1991:PPE

[CAWH91] Tim W. Clark, Elizabeth Dawn Amato, Donald G. Whittemore, and Ann H. Harvey. Policy and programs for ecosystem management in the Greater Yellowstone Ecosystem: an analysis. Conservation Biology, 5(3):412–422, September 1991. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ceballos:1995:GPM

[CB95] Gerardo Ceballos and James H. Brown. Global patterns of mammalian diversity, endemism, and endangerment. Patrones globales de la diversidad, endemismo y riesgo de extinción de los mamiferos. Conservation Biology, 9(3):559–568, June 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Crampton:1998:SRF

[CB98] Lisa H. Crampton and Robert M. R. Barclay. Selection of roosting and foraging habitat by bats in different-aged aspen mixedwood stands. Conservation Biology, 12(6):1347–1358, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cummings:1997:GAE

[CBAT97] Shawn A. Cummings, Ernest L. Brannon, Karen J. Adams, and Gary H. Thorgaard. Genetic analyses to establish captive breeding priorities for endangered Snake River sockeye salmon. Análisis genético para establecer prioridades de reproducción en cautiverio para el salmón sockeye del Snake River en piligro de extinción. Conservation Biology, 11(3):662–669, June 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cohen:1993:ISP

[CBCC93] Andrew S. Cohen, Roger Bills, Christine Z. Cocquyt, and A. G. Caljon. The impact of sediment pollution on biodiversity in Lake Tanganyika. El impacto de la contaminación sedimentaria sobre la biodiversidad del Lago Tanganyika. Conservation Biology, 7(3):667–677, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Contreras-B:1994:WEF

[CBLV94] Salvador Contreras-B. and M. Lourdes Lozano-V. Water, endangered fishes, and development perspectives in arid lands of Mexico. Aguas, peces en peligro, y perspectivas de desarrollo en las tierras áridas de México. Conservation Biology, 8(2):379– 387, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cafaro:1997:EA

[CBM^{+}97] Philip Cafaro, Nicholas Brokaw, Samuel J. McNaughton, Francis E. Putz, and Richard E. Bodmer. Enough already! Conservation Biology, 11(5):1258–1264, October 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chapman:1995:SDS

[CC95] Colin A. Chapman and Lauren J. Chapman. Survival without dispersers: Seedling recruitment under parents. Conservation Biology, 9(3):675–678, June 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Creel:1996:LAW

[CC96] Scott Creel and Nancy Marusha Creel. Limitation of African wild dogs by competition with larger carnivores. La limitación de los perros salvajes africanos a causa de la competencia con carnívoros grandes. Conservation Biology, 10(2):526–538, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chapman:1999:FRA

[CC99a] Colin A. Chapman and Lauren J. Chapman. Forest restoration in abandoned agricultural land: a case study from East Africa. Restauración de bosques en tierras agriculturales abandonadas: Caso de estudio en Africa Occidental. Conservation Biology, 13(6):1301–1311, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chazdon:1999:EWS

[CC99b] Robin L. Chazdon and Felix G. Coe. Ethnobotany of woody species in second-growth, old-growth, and selectively logged forests of Northeastern Costa Rica. Etnobotánica de plantas leñosas en bosques de crecimiento secundario, bosques maduros, y bosques talados selectivamente en el noreste de Costa Rica. Conservation Biology, 13(6):1312–1322, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Creel:1997:RSH

[CCM97] Scott Creel, Nancy Marusha Creel, and Steven L. Monfort. Radiocollaring and stress hormones in African wild dogs. Conservation Biology, 11(2):544–548, April 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Callicott:1999:CNC

[CCM99] J. Baird Callicott, Larry B. Crowder, and Karen Mumford. Current normative concepts in conservation. Conservation Biology, 13(1):22–35, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chapman:1996:REF

[CCOO^{+}96] Lauren J. Chapman, Colin A. Chapman, Richard Ogutu-Ohwayo, Mark Chandler, Les Kaufman, and Amanda E. Keiter. Refugia for endangered fishes from an introduced predator in Lake Nabugabo, Uganda. Refugio para proteger de un de predador introducido a los peces en peligro de extinción del Lago Nabugabo, en Uganda. Conservation Biology, 10(2):554– 561, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clark:1996:CES

[CCR96] Tim W. Clark, A. Peyton Curlee, and Richard P. Reading. Crafting effective solutions to the large carnivore conservation problem. Moldeando soluciones efectivas para el problema de la conservación de carnívoros mayores. Conservation Biology, 10 (4):940–948, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

CamposDeDuffy:1990:L

[CD90] Maria Jose Campos De Duffy and David Cameron Duffy. Letter. Conservation Biology, 4(2):117–118, June 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Costanza:1992:NCS

[CD92] Robert Costanza and Herman E. Daly. Natural capital and sustainable development. Conservation Biology, 6(1):37–46, March 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Colwell:1995:WCH

[CD95] Mark A. Colwell and Sarah L. Dodd. Waterbird communities and habitat relationships in coastal pastures of Northern California. Comunidades de aves acuáticas y las relaciones entre los hábitats en pasturas costeras del Norte de California. Conservation Biology, 9(4):827–834, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cunningham:1998:ESL

[CD98] Andrew A. Cunningham and Peter Daszak. Extinction of a species of land snail due to infection with a microsporidian parasite. Conservation Biology, 12(5):1139–1141, October 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Corsi:1999:LSM

[CDB99] Fabio Corsi, Eugenio Duprè, and Luigi Boitani. A large-scale model of wolf distribution in Italy for conservation planning. Conservation Biology, 13(1):150–159, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cannon:1996:TCB

[CDD96] John R. Cannon, James M. Dietz, and Lou Ann Dietz. Training conservation biologists in human interaction skills. Capacitación de biólogos de la conservación en técnicas de interacción humana. Conservation Biology, 10(4):1277–1282, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cree:1995:RRN

[CDH95] Alison Cree, Charles H. Daugherty, and Jennifer M. Hay. Reproduction of a rare New Zealand reptile, the tuatara Sphenodon punctatus, on rat-free and rat-inhabited islands. La reproducción en un reptil raro de Nueva Zelandia, la tuatara Sphenodon punctatus en islas habitadas e islas no habitadas por ratas. Conservation Biology, 9(2):373–383, April 1995. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Congdon:1993:DSM

[CDV93] Justin D. Congdon, Arthur E. Dunham, and R. C. Van Loben Sels. Delayed sexual maturity and demographics of Blanding’s turtles (Emydoidea blandingii): Implications for conservation and management of long-lived organisms. Conservation Biology, 7(4):826–833, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cox:1991:FFS

[CEPR91] Paul Alan Cox, Thomas Elmqvist, Elizabeth D. Pierson, and William E. Rainey. Flying foxes as strong interactors in South Pacific island ecosystems: a conservation hypothesis. Conservation Biology, 5(4):448–454, December 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Curtin:1995:GA

[CFH^{+}95] Charles G. Curtin, Thomas L. Fleischner, Carola A. Haas, James D. Fraser, Reed Noss, Thomas L. Fleischner, George Wuerthner, Joy Belsky, Bill Willers, Peter F. Brussard, Dennis D. Murphy, and C. Richard Tracy. Grazing and advocacy. Conservation Biology, 9(2):233–239, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chen:1996:CAS

[CFL96] Jiquan Chen, Jerry F. Franklin, and J. Stephen Lowe. Comparison of abiotic and structurally defined patch patterns in a hypothetical forest landscape. Comparación de patrones de parches definidos Abi’otica y estructuralmente en el paisaje de un bosque hipotético. Conservation Biology, 10(3):854–862, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ceballos:1995:CNB

[CG95] Gerardo Ceballos and Andres García. Conserving neotropical biodiversity: The role of dry forests in Western Mexico. Conservation Biology, 9(6):1349–1356, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Crone:1998:PVR

[CG98] Elizabeth E. Crone and Janet L. Gehring. Population viability of Rorippa columbiae: Multiple models and spatial trend data. Conservation Biology, 12(5):1054–1065, October 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cafaro:1997:WCA

[CGG^{+}97] Philip Cafaro, Jerry A. Griffith, Paul Z. Goldstein, Bill Willers, and John G. Sidle. Who cares about life? Conservation Biology, 11(2):580–586, April 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chambers:1995:SSG

[Cha95] Steven M. Chambers. Spatial structure, genetic variation, and the neighborhood adjustment to effective population size. Conservation Biology, 9(5):1312–1315, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cherfas:1993:BB

[Che93] Jeremy Cherfas. Backgarden biodiversity. Conservation Biology, 7(1):6, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chen:1999:PEJ

[Che99] Linus Chen. Playing Earth jazz. Conservation Biology, 13 (6):1521, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chapin:1998:ILP

[CHK98] Theodore G. Chapin, Daniel J. Harrison, and Donald D. Katnik. Influence of landscape pattern on habitat use by American marten in an industrial forest. Conservation Biology, 12 (6):1327–1337, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chorley:1995:HES

[Cho95] Matthew C. Chorley. Homeless endangered species. Conservation Biology, 9(5):983, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cade:1993:PRM

[CJ93] ToM. J. Cade and Carl G. Jones. Progress in restoration of the Mauritius kestrel. Progresos en la restauración del cernícalo de Mauricio. Conservation Biology, 7(1):169–175, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Camp:1998:ERC

[CK98] Richard J. Camp and Richard L. Knight. Effects of rock climbing on cliff plant communities at Joshua Tree National Park, California. Conservation Biology, 12(6):1302–1306, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Czech:1998:SCP

[CKB98] Brian Czech, Paul R. Krausman, and Rena Borkhataria. Social construction, political power, and the allocation of benefits to endangered species. Conservation Biology, 12(5):1103–1112, October 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chapman:1995:HTT

[CKCM95] Lauren J. Chapman, Leslie S. Kaufman, Colin A. Chapman, and F. Ellis Mckenzie. Hypoxia tolerance in twelve species of East African cichlids: Potential for low oxygen refugia in Lake Victoria. Conservation Biology, 9(5):1274–1288, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clarke:1992:HWP

[Cla92] Colin M. H. Clarke. History of wild pigs in the United States. Conservation Biology, 6(4):610–611, December 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clarke:1995:RBD

[Cla95] Geoffrey M. Clarke. Relationships between developmental stability and fitness: Application for conservation biology. La relación entre la estabilidad del desarrollo y el fitness: Aplicaciones a la biología de conservación. Conservation Biology, 9 (1):18–24, February 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clark:1999:EAP

[Cla99] Jamie Rappaport Clark. The ecosystem approach from a practical point of view. Conservation Biology, 13(3):679–681, June 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cleary:1993:ID

[Cle93] Ruark L. Cleary. Inappropriate debate. Conservation Biology, 7(2):223a, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Campbell:1991:EIH

[CLL91] David G. Campbell, Karen S. Lowell, and Michael E. Lightbourn. The effect of introduced hutias (Geocapromys ingrahmi) on the woody vegetation of Little Wax Cay, Bahamas. Conservation Biology, 5(4):536–541, December 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cole:1995:CII

[CLM^{+}95] F. Russell Cole, Lloyd L. Loope, Arthur C. Medeiros, Jane A. Raikes, and Cynthia S. Wood. Conservation implications of introduced game birds in high-elevation Hawaiian shrubland.1 las concecuencias concervacionistas de aves de caza introducidas en las areas arbustivas de gran altura de Hawaii. Conservation Biology, 9(2):306–313, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Coulter:1993:CLT

[CM93] G. Coulter and R. Mubamba. Conservation in Lake Tanganyika, with special reference to underwater parks. Conservación en el Lago Tanganyika, con referencia especial a parques subacuáticos. Conservation Biology, 7(3):678–685, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Coulter:1994:RPV

[CM94] G. W. Coulter and Raphael Mubamba. Response to Pendleton and van Breda. Conservation Biology, 8(2):331–332, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Callicott:1997:ESC

[CM97] J. Baird Callicott and Karen Mumford. Ecological sustainability as a conservation concept. Sustentabilidad ecológica como concepto de conservación. Conservation Biology, 11(1):32–40, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Coleman:1996:SCB

[CMB96] William G. Coleman, Jack Mattice, and Robert W. Brocksen. Soulé’s conservation biology as the foundation for econometric ecosystem management. Conservation Biology, 10(6):1494– 1499, December 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cisneros-Mata:1995:LHC

[CMMLRR95] Miguel A. Cisneros-Mata, Gabriela Montemayor-López, and Martha J. Román-Rodríguez. Life history and conservation of Totoaba macdonaldi. Historia natural y conservación de Totoaba macdonaldi. Conservation Biology, 9(4):806–814, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chapman:1998:WME

[CMW^{+}98] Lauren J. Chapman, Mark W. Miller, Joan L. Walker, T. R. McClanahan, Diethart Matthies, and Philip Cafaro. Witness to mass extinction. Conservation Biology, 12(1):250–257, February 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cesario:1997:LHH

[CNP^{+}97] Manuel Cesario, Richard B. Norgaard, David R. Peart, Kamaljit S. Bawa, and Thomas V. Dietsch. Linking human health to biological diversity. Conservation Biology, 11(6):1459–1464, December 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Caro:1999:USS

[CO99] T. M. Caro and Gillian O’Doherty. On the use of surrogate species in conservation biology. Uso de especies sustituatas en la conservación biológica. Conservation Biology, 13(4):805–814, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Coblentz:1990:EOD

[Cob90] Bruce E. Coblentz. Exotic organisms: a dilemma for conservation biology. Conservation Biology, 4(3):261–265, September 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Coblentz:1991:RTL

[Cob91] Bruce K. Coblentz. A response to Temple and Lugo. Conservation Biology, 5(1):5–6, March 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Coblentz:1997:GEI

[Cob97a] Bruce E. Coblentz. Goat eradication on Isabella Island. Conservation Biology, 11(5):1047–1048, October 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Coblentz:1997:SCC

[Cob97b] Bruce E. Coblentz. Subsistence consumption of coral reef fish suggests non-sustainable extraction. Conservation Biology, 11 (2):559–561, April 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Coblentz:1999:SSA

[Cob99] Bruce E. Coblentz. Still searching for an alternative view of coral reef fish consumption: a response. Conservation Biology, 13(4):942, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Colvin:1992:CER

[Col92] Jean G. Colvin. A code of ethics for research in the third world. Conservation Biology, 6(3):309–311, September 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Condon:1995:BST

[Con95] Marty Condon. Biodiversity, systematics, and Tom Sawyer science. Conservation Biology, 9(4):711–714, August 1995. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cort:1996:SUN

[Cor96] Cheryl A. Cort. A survey of the use of natural heritage date in local land-use planning. Conservation Biology, 10(2):632– 637, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cowlishaw:1999:PPD

[Cow99] Guy Cowlishaw. Predicting the pattern of decline of African primate diversity: an extinction debt from historical deforestation. Predicciones del patrón de declive en la diversidad de primates Africanos: una deuda de extinción debida a una deforestación histórica. Conservation Biology, 13(5):1183–1193, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

C:1994:RCC

[CP94] Eduardo Santana C. and Enriqne Jardel P. Research for conservation or conservation for research? Conservation Biology, 8(1):6–7, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clark:1996:GLP

[CPC96a] Tim. W. Clark, Paul C. Paquet, and A. Peyton Curlee. General lessons and positive trends in large carnivore conservation. Conservation Biology, 10(4):1055–1058, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clark:1996:I

[CPC96b] Tim W. Clark, Paul C. Paquet, and A. Peyton Curlee. Introduction. Conservation Biology, 10(4):936–939, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Caro:1994:ECB

[CPG94] T. M. Caro, N. Pelkey, and M. Grigione. Effects of conservation biology education on attitudes toward nature. Conservation Biology, 8(3):846–852, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Conway:1995:OSN

[CPN95] Courtney J. Conway, George V. N. Powell, and James D. Nichols. Overwinter survival of neotropical migratory birds in early-successional and mature tropical forests. Supervivencia de invernación de las aves migratorias neotropicales en bosques tropicales de sucesión temprana y maduros. Conservation Biology, 9(4):855–864, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Crade:1994:LDE

[Cra94] Tom J. Crade. Last ditch efforts. Conservation Biology, 8 (1):321–323, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Creel:1997:HAW

[Cre97] Scott Creel. Handling of African wild dogs and chronic stress: Reply to East et al. Conservation Biology, 11(6):1454–1456, December 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [EHBE97].

Cardel:1997:ESB

[CRGGFT97] Yuria Cardel, Victor Rico-Gray, José G. García-Franco, and Leonard B. Thien. Ecological status of Beaucarnea gracilis, an endemic species of the semiarid Tehuacán Valley, México. Situación ecológica de Beaucarnea gracilis, una especie endémica del valle semiárido de Tehuacán, Mexico. Conservation Biology, 11(2):367–374, April 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chavez-Ramirez:1996:ECW

[CRHSS96] Felipe Chavez-Ramirez, Howard E. Hunt, R. Douglas Slack, and Tom V. Stehn. Ecological correlates of whooping crane use of fire-treated upland habitats. Correlaciones ecológicas de la grulla blanca (Grus americana) en el uso de ambientes de tierra firme tratados con fuego. Conservation Biology, 10(1): 217–223, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cheverud:1994:QMG

[CRJ^{+}94] James Cheverud, Eric Routman, Cashell Jaquish, Suzette Tardif, Gloria Peterson, Natasha Belfiore, and Lisa Forman. Quantitative and molecular genetic variation in captive cottontop tamarins (Saguinus oedipus). Variación genética molecular y cuantitativa en los pinches cautivos (Saguinus oedipus). Conservation Biology, 8(1):95–105, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Crother:1992:GCS

[Cro92] Brian I. Crother. Genetic characters, species concepts, and conservation biology. Conservation Biology, 6(3):314, September 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Crow:1995:MBM

[Cro95] Thomas R. Crow. Manual for biodiversity management. Conservation Biology, 9(1):226–227, February 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cote:1997:ERP

[CS97] Isabelle M. Côté and William J. Sutherland. The effectiveness of removing predators to protect bird populations. Efectividad de la remoción de depredadores para proteger poblaciones de aves. Conservation Biology, 11(2):395–405, April 1997. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cochrane:1998:FFB

[CS98] Mark A. Cochrane and Mark D. Schulze. Forest fires in the Brazilian Amazon. Conservation Biology, 12(5):948–950, October 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Caicco:1995:GAM

[CSBC95] Steven L. Caicco, J. M. Scott, B. Butterfield, and B. Csuti. A gap analysis of the management status of the vegetation of Idaho (U.S.A.) un análisis de intervalo sobre el estado del manejo de la vegetación de Idaho. Conservation Biology, 9 (3):498–511, June 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Crawford:1994:LFL

[CSC^{+}94] Daniel J. Crawford, Tod F. Stuessy, Mary Beth Cosner, David W. Haines, Delbert Wiens, and Patricio Peñaillo. Lactoris fernandeziana (Lactondaceae) on the Juan Fernandez Islands: Allozyme uniformity and field observations. Conservation Biology, 8(1):277–280, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Crooks:1998:NIC

[CSD98] Kevin R. Crooks, M. A. Sanjayan, and Daniel F. Doak. New insights on cheetah conservation through demographic modeling. Conservation Biology, 12(4):889–895, August 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Coviella:1999:EEA

[CT99] Carlos E. Coviella and John T. Trumble. Effects of elevated atmospheric carbon dioxide on insect-plant interactions. Efectos de elevados niveles de dióxido de carbono atmosférico en interacciones planta-insecto. Conservation Biology, 13(4):700– 712, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cuaron:1997:MAT

[Cua97] Alfredo D. Cuarón. Miguel Alvarez del Toro: First and last of a kind. Conservation Biology, 11(2):566–568, April 1997. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cueva:1996:BCM

[Cue96] Horacio Cueva. Biodiversity challenges in Mexico. Conservation Biology, 10(4):1302–1303, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

CardosoDaSilva:1996:PSL

[CUM96] José Maria Cardoso Da Silva, Christopher Uhl, and Gregory Murray. Plant succession, landscape management, and the ecology of frugivorous birds in abandoned Amazonian pastures. La sucesión vegetal, el manejo del paisaje y la ecología de las aves frugívoras en pasturas abandonadas de la Amazonía. Conservation Biology, 10(2):491–503, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cunningham:1996:DRW

[Cun96] Andrew A. Cunningham. Disease risks of wildlife translocations. El riesgo de la transmisión de enfermedades en las translocaciones de especies silvestres. Conservation Biology, 10(2):349–353, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Curtin:1993:EUN

[Cur93] Charles G. Curtin. The evolution of the U.S. National Wildlife Refuge System and the doctrine of compatibility. La evolución del sistema nacional de los Estados Unidos de refugios de vida silvestre y la doctrina de compatibilidad. Conservation Biology, 7(1):29–38, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cutler:1991:NFE

[Cut91] Alan Cutler. Nested faunas and extinction in fragmented habitats. Conservation Biology, 5(4):496–504, December 1991. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clark:1995:CAD

[CV95] Heidi Clark and Kristiina Vogt. Conservation atlases of deserts and wetlands. Conservation Biology, 9(1):223–225, February 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cade:1997:CDA

[CW97a] Tom J. Cade and Christopher P. Woods. Changes in distribution and abundance of the loggerhead shrike. Cambios en la distribución y abundancia del verduguillo. Conservation Biology, 11(1):21–31, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chayax:1997:PLM

[CW97b] Reginaldo Chayax and David F. Whitacre. A plea from the last of the Maya-Itzá. Conservation Biology, 11(6):1457–1458, December 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carroll:1999:UPA

[CZN99] Carlos Carroll, William J. Zielinski, and Reed F. Noss. Using presence-absence data to build and test spatial habitat models for the fisher in the Klamath Region, U.S.A. Utilización de datos de presencia-ausencia para construir y evaluar modelos espaciales de hábitat para el pescador de la Región Klamath, U.S.A. Conservation Biology, 13(6):1344–1359, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Daily:1995:LS

[Dai95] Gretchen C. Daily. A look at the streetfight. Conservation Biology, 9(2):465, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Daly:1993:BAP

[Dal93] Herman E. Daly. Beginning again on purpose. Conservation Biology, 7(3):736–738, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Daly:1999:LI

[Dal99] Herman E. Daly. The lurking inconsistency. Conservation Biology, 13(4):693–694, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Davidson:1996:TSG

[Dav96] Ann S. Davidson. Taking stock of global bird conservation. Conservation Biology, 10(5):1478–1479, October 1996. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Day:1991:COD

[Day91] Peter Day. Conserving our daily bread. Conservation Biology, 5(2):133–134, June 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Delibes:1998:TIL

[DB98] Miguel Delibes and M. Carmen Blázquez. Tameness of insular lizards and loss of biological diversity. Conservation Biology, 12 (5):1142–1143, October 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dunning:1995:PIC

[DBCM95] John B. Dunning, Jr., Rene Borgella, Jr., Krista Clements, and Gary K. Meffe. Patch isolation, corridor effects, and colonization by a resident sparrow in a managed pine woodland. Aislamiento de parches, efecto de los corredores y colonización de un gorrión residente en un pinar bajo manejo. Conservation Biology, 9(3):542–550, June 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Duffy:1999:LAD

[DBH^{+}99] David Cameron Duffy, Keith Boggs, Randall H. Hagenstein, Robert Lipkin, and Julie A. Michaelson. Landscape assessment of the degree of protection of Alaska’s terrestrial biodiversity. Evaluación de paisaje en el grado de protección de la biodiversidad territorial de Alaska. Conservation Biology, 13 (6):1332–1343, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dowling:1992:IHT

[DC92] Thomas E. Dowling and Michael R. Childs. Impact of hybridization on a threatened trout of the Southwestern United States. Conservation Biology, 6(3):355–364, September 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dodd:1998:MPC

[DC98] C. Kenneth Dodd, Jr. and Brian S. Cade. Movement patterns and the conservation of amphibians breeding in small, temporary wetlands. Conservation Biology, 12(2):331–339, April 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Droege:1998:CUU

[DCL98] Sam Droege, André Cyr, and Jacques Larivée. Checklists: an under-used tool for the inventory and monitoring of plants and animals. Conservation Biology, 12(5):1134–1138, October 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Delcourt:1997:PCN

[DD97] Hazel R. Delcourt and Paul A. Delcourt. Pre-Columbian native American use of fire on Southern Appalachian landscapes. Conservation Biology, 11(4):1010–1014, August 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dowling:1992:UGC

[DDM^{+}92] Thomas E. Dowling, Bruce D. DeMarais, W. L. Minckley, Michael E. Douglas, and Paul C. Marsh. Use of genetic characters in conservation biology. Conservation Biology, 6(1):7–8, March 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Demarais:1993:PPG

[DDM93] Bruce D. Demarais, Thomas E. Dowling, and W. L. Minckley. Post-perturbation genetic changes in populations of endangered Virgin River chubs. Cambios genéticos post-perturbación en poblaciones amenazadas de “Virgin River chubs” (Gila seminuda). Conservation Biology, 7(2):334–341, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

DeVires:1995:LHD

[De 95] Michiel F. De Vires. Large herbivores and the design of largescale nature reserves in Western Europe. Herbívoros grandes y el diseño de reservas naturales a gran escala en Europa occidental. Conservation Biology, 9(1):25–33, February 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dearden:1995:PLC

[Dea95] Philip Dearden. Park literacy and conservation. Conservation Biology, 9(6):1654–1656, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Deharveng:1996:SCD

[Deh96] Louis Deharveng. Soil collembola diversity, endemism, and reforestation: a case study in the Pyrenees (France). La diversidad de los colémbolos terrestres, el endemismo y la reforestación: un estudio de caso en los Pirineos (Francia). Conservation Biology, 10(1):74–84, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Delissio:1999:LTP

[Del99] Lisa J. Delissio. Long-term perspectives on forest biodiversity. Conservation Biology, 13(6):1522–1523, December 1999. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

DeMauro:1993:RBS

[DeM93] Marcella M. DeMauro. Relationship of breeding system to rarity in the lakeside daisy (Hymenoxys acaulis var. glabra). Relación entre los sistemas reproductivos y la rareza en la margarita del lago (Hymenoxys acaulis var. glabra). Conservation Biology, 7(3):542–550, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Drost:1996:CRF

[DF96] Charles A. Drost and Gary M. Fellers. Collapse of a regional frog fauna in the Yosemite area of the California Sierra Nevada, USA. El colapso de la fauna regional de bactracios en el área de Yosemite en la Sierra Nevada de California (EEUU). Conservation Biology, 10(2):414–425, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dirzo:1992:RDT

[DG92] Rodolfo Dirzo and Maria C. Garcia. Rates of deforestation in Los Tuxtlas, a neotropical area in Southeast Mexico. Conservation Biology, 6(1):84–90, March 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Diego-Gomez:1993:PMB

[DG93] Luis Diego-Goméz. Perspectives on Mexican biodiversity. Conservation Biology, 7(4):967–968, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dimmick:1999:ISB

[DGG^{+}99] Walter Wheaton Dimmick, Michael J. Ghedotti, Michael J. Grose, Anne M. Maglia, Daniel J. Meinhardt, and David S. Pennock. The importance of systematic biology in defining units of conservation. Importancia de la biología sistemática en la definicíon de unidades de conservación. Conservation Biology, 13(3):653–660, June 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Desrochers:1997:GCD

[DH97] André Desrochers and Susan J. Hannon. Gap crossing decisions by forest songbirds during the post-fledging period. Decisíones de cruce de claros por aves paserinas de bosques durante el periodo post-juvenil. Conservation Biology, 11(5):1204–1210, October 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Demaynadier:1998:ESE

[DH98] Phillip G. Demaynadier and Malcolm L. Hunter. Effects of silvicultural edges on the distribution and abundance of amphibians in Maine. Conservation Biology, 12(2):340–352, April 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Denton:1997:RPN

[DHBG97] Jonathan S. Denton, Susan P. Hitchings, Trevor J. C. Beebee, and Antony Gent. A recovery program for the natterjack toad (Bufo calamita) in Britain. Conservation Biology, 11(6): 1329–1338, December 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Downes:1997:UCM

[DHE97] Sharon J. Downes, Kathrine A. Handasyde, and Mark A. Elgar. The use of corridors by mammals in fragmented Australian eucalypt forests. Uso de corredores por mamíferos en bosques de eucalipto en Australia. Conservation Biology, 11(3):718– 726, June 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Daniels:1991:ACV

[DHJG91] R. J. Ranjit Daniels, Malati Hegde, N. V. Joshi, and Madhav Gadgil. Assigning conservation value: a case study from India. Conservation Biology, 5(4):464–475, December 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dunn:1999:PST

[DHW99] Erica H. Dunn, David J. T. Hussell, and Daniel A. Welsh. Priority-setting tool applied to Canada’s landbirds based on concern and responsibility for species herramienta para establecer prioridades basada en preocupaciones y responsabilidades por especie aplicada a aves terrestres de Canada. Conservation Biology, 13(6):1404–1415, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dial:1995:SAC

[Dia95] Roman J. Dial. Species-area curves and Koopowits et al.’s simulation of stochastic extinctions. Conservation Biology, 9 (4):960–961, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dial:1996:BMN

[Dia96] Roman Dial. Beyond MacArthur: New attempts at unification. Conservation Biology, 10(6):1564–1566, December 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dietz:1994:TEL

[Die94] James M. Dietz. A textbook at the ecosystem level. Conservation Biology, 8(4):1170, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dinerstein:1994:ESR

[DKOW94] Eric Dinerstein, Vladimir Krever, David M. Olson, and Laura Williams. An emergency strategy to rescue Russia’s biological diversity. Conservation Biology, 8(4):934–942, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Donovan:1995:MEH

[DLK^{+}95] Therese M. Donovan, Roland H. Lamberson, Allison Kimber, Frank R. Thompson III, and John Faaborg. Modeling the effects of habitat fragmentation on source and sink demography of neotropical migrant birds. Modelación de los efectos de la fragmentación del hábitat sobre la demografia de fuentes y sumideros en aves neomigratorias tropicales. Conservation Biology, 9(6):1396–1407, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dizon:1992:RSC

[DLP^{+}92] Andrew E. Dizon, Christina Lockyer, William F. Perrin, Douglas P. Demaster, and Joyce Sisson. Rethinking the stock concept: a phylogeographic approach. Conservation Biology, 6(1): 24–36, March 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

DelTredici:1992:EGT

[DLY92] Peter Del Tredici, Hsieh Ling, and Guang Yang. The Ginkgos of Tian Mu Shan. Conservation Biology, 6(2):202–209, June 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dinerstein:1990:EGO

[DM90a] Eric Dinerstein and Gary F. Mccracken. Endangered greater one-horned rhinoceros carry high levels of genetic variation. Conservation Biology, 4(4):417–422, December 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dirzo:1990:CND

[DM90b] Rodolfo Dirzo and Alvaro Miranda. Contemporary neotropical defaunation and forest structure, function, and diversity — a sequel to John Terborgh. Conservation Biology, 4(4):444–447, December 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Duffy:1992:DAH

[DM92] David Cameron Duffy and Albert J. Meier. Do Appalachian herbaceous understories ever recover from clearcutting? Conservation Biology, 6(2):196–201, June 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dowling:1992:RWN

[DMD^{+}92] Thomas E. Dowling, W. L. Minckley, Michael E. Douglas, Paul C. Marsh, and Bruce D. Demarais. Response to Wayne, Nowak, and Phillips and Henry: Use of molecular characters in conservation biology. Conservation Biology, 6(4):600–603, December 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dwyer:1995:PRC

[DME95] Lynn E. Dwyer, Dennis D. Murphy, and Paul R. Ehrlich. Property rights case law and the challenge to the Endangered Species Act. La jurisprudencia sobre los derechos a la propiedad y el desafío al Acta de Especies en Peligro. Conservation Biology, 9(4):725–741, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dowling:1996:MDV

[DMMG96] Thomas E. Dowling, W. L. Minckley, Paul C. Marsh, and Elliott S. Goldstein. Mitochondrial DNA variability in the endangered razorback sucker (Xyrauchen texanus): Analysis of hatchery stocks and implications for captive propagation. Variabilidad del ADN mitocondrial en Xyrauchen texanus, una especie en peligro: Análisis de los stocks de las granjas de cría y su implicación en la propagación de animales criados en cautiverio. Conservation Biology, 10(1):120–127, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Doak:1995:SSM

[Doa95] Daniel F. Doak. Source–sink models and the problem of habitat degradation: General models and applications to the Yellowstone grizzly. Los modelos de fuente–sumidero y el problema de la degradación del hábitat: Modelos generales y aplicaciones al oso gris de Yellowstone. Conservation Biology, 9(6):1370–1379, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dale:1994:MEL

[DOSP94] Virginia H. Dale, Robert V. O’Neill, Frank Southworth, and Marcos Pedlowski. Modeling effects of land management in the Brazilian Amazonian settlement of Rondônia. Modelando los efectos del manejo de tierras en los asentamientos de la Amazonia Brasilera de Rondônia. Conservation Biology, 8(1):196– 206, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dobson:1992:IWB

[DP92] Andrew P. Dobson and Joyce H. Poole. Ivory: Why the ban must stay! Conservation Biology, 6(1):149–151, March 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Drayton:1996:PSL

[DP96] Brian Drayton and Richard B. Primack. Plant species lost in an isolated conservation area in metropolitan Boston from 1894 to 1993. La pérdida de especies de plantas desde 1894 a 1993 en un área de conservación aislada dentro de la zona metropolitana de Boston. Conservation Biology, 10(1):30–39, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dale:1994:RPL

[DPOO94] Virginia H. Dale, Scott M. Pearson, Holly L. Offerman, and Robert V. O’Neill. Relating patterns of land-use change to faunal biodiversity in the Central Amazon. Relacionando el cambio de patrones en el uso de la tierra con respecto a la biodiversidad de la fauna en el Amazonas Central. Conservation Biology, 8(4):1027–1036, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Drayton:1994:PJA

[Dra94] Brian Drayton. A personal journey into the Amazon. Conservation Biology, 8(3):913, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Drayton:1996:TCP

[Dra96] Brian Drayton. Taking care of plants. Conservation Biology, 10 (2):690, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Drew:1994:SMR

[Dre94] Gary S. Drew. The scientific method revisited. Conservation Biology, 8(2):596–597, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dobkin:1998:HAR

[DRP98] David S. Dobkin, Adam C. Rich, and William H. Pyle. Habitat and avifaunal recovery from livestock grazing in a riparian meadow system of the Northwestern Great Basin. Conservation Biology, 12(1):209–221, February 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dole:1992:FGS

[DS92] Jefferey A. Dole and Mei Sun. Field and genetic survey of the endangered Butte County meadowfoam — Limnanthes floccosa subsp. californica (Limnanthaceae). Investigación de campo y genética del “Meadowfoam” en peligro de extinción del Condado de Butte — Limnanthes floccosa subsp. californica (Limnanthaceae). Conservation Biology, 6(4):549–558, December 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

DaSilva:1995:PDN

[DS95] Nelson Jorge Da Silva, Jr. and Jack W. Sites, Jr. Patterns of diversity of neotropical squamate reptile species with emphasis on the Brazilian Amazon and the conservation potential of indigenous reserves. Patrones de diversidad de especies de reptiles escamosos, con énfasis en la Amazonía brasilena, y el potencial de conservación de reservas indígenas. Conservation Biology, 9 (4):873–901, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Demers:1995:FEI

[DSB^{+}95a] Michael N. Demers, John W. Simpson, Ralph E. J. Boerner, Alejandra Silva, Leslie Berns, and Francisco Artigas. Fencerows, edges, and implications of changing connectivity illustrated by two contiguous Ohio landscapes. Conservation Biology, 9(5): 1159–1168, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dupuis:1995:RTB

[DSB95b] Linda A. Dupuis, James N. M. Smith, and Fred Bunnell. Relation of terrestrial-breeding amphibian abundance to tree-stand age. Relación entre la abundancia de anfibios con reproducción terreste y la edad del rodal. Conservation Biology, 9(3):645– 653, June 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dunn:1993:EBC

[DSGS93] Christopher P. Dunn, Forest Stearns, Glenn R. Guntenspergen, and David M. Sharpe. Ecological benefits of the Conservation Reserve Program. Beneficios ecológicos del Programa de Conservación de Reservas. Conservation Biology, 7(1):132– 139, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dean:1990:FFF

[DSM90] W. R. J. Dean, W. Roy Siegfried, and I. A. W. Macdonald. The fallacy, fact, and fate of guiding behavior in the greater honeyguide. Conservation Biology, 4(1):99–101, March 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Donovan:1995:RSM

[DTFP95] Therese M. Donovan, Frank R. Thompson III, John Faaborg, and John R. Probst. Reproductive success of migratory birds in habitat sources and sinks. Éxito reproductivo de las aves migratorias en hábitats fuente y sumidero. Conservation Biology, 9(6):1380–1395, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dudley:1992:RRO

[Dud92] Joseph P. Dudley. Rejoinder to Rohlf and O’Connell: Biodiversity as a regulatory criterion. Conservation Biology, 6(4):

587–589, December 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [Roh92, O’C92].

Dudley:1993:JND

[Dud93] Joseph P. Dudley. Jay Norwood “Ding” Darling: a retrospective. Conservation Biology, 7(1):200–203, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dudley:1995:BPG

[Dud95] Joseph P. Dudley. Bioregional parochialism and global activism. Conservation Biology, 9(5):1332–1334, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dudley:1997:PCP

[Dud97] Joseph P. Dudley. Paleontological and cultural perspectives on livestock grazing in southwestern rangelands: Response to Brown and McDonald. Conservation Biology, 11(1):267–269, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dudley:1999:CIE

[Dud99] Joseph P. Dudley. Coevolutionary implications of an endemic Pleistocene megaherbivore fauna for insular floras of the California Channel Islands. Conservation Biology, 13(1):209–210, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Duffus:1993:TBM

[Duf93a] Dacid Duffus. Tsitika to baram: The myth of sustainability. Conservation Biology, 7(2):440–442, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Duffy:1993:HCR

[Duf93b] David Cameron Duffy. Herbs and clearcutting: Reply to Elliot and Loftis and Steinbeck. Conservation Biology, 7(2):221–223, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Duffy:1993:SFT

[Duf93c] David Cameron Duffy. Seeing the forest for the trees: Response to Johnson et al. Conservation Biology, 7(2):436–439, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Duffield:1994:VER

[Duf94] John W. Duffield. Valuing environmental resources in Europe. Conservation Biology, 8(2):605–607, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dunning:1997:MAP

[Dun97] John B. Dunning, Jr. The missing awareness,Part 2: Teaching students what a billion people looks like. Conservation Biology, 11(1):6–10, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dunning:1999:NPE

[Dun99] John B. Dunning, Jr. The need for producing ecologically literate College students. Conservation Biology, 13(6):1239–1240, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dinerstein:1993:BHH

[DW93] Eric Dinerstein and Eric D. Wikramanayake. Beyond “Hotspots”: How to prioritize investments to conserve biodiversity in the Indo-Pacific region. Mas allá de “Hotspots”: como priorizar inversiones en biodiversidad en la región indopacífica. Conservation Biology, 7(1):53–65, March 1993. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Davenport:1998:II

[DWM^{+}98] Leigh A. Davenport, Mark H. Waugh, Karl E. Miller, Campbell O. Webb, Jan Salick, and Les Kaufman. An informative introduction. Conservation Biology, 12(3):733–738, June 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dobson:1993:RNF

[DY93] F. Stephen Dobson and Jinping Yu. Rarity in neotropical forest mammals revisited. Revisión de la rareza en mamíferos de forestas neotropicales. Conservation Biology, 7(3):586–591, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dobson:1995:IER

[DYS95] F. Stephen Dobson, Jinping Yu, and Andrew T. Smith. The importance of evaluating rarity. Conservation Biology, 9(6): 1648–1651, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Erb:1999:DPD

[EB99] John D. Erb and Mark S. Boyce. Distribution of population declines in large mammals. Conservation Biology, 13(1):199–201, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ebert:1995:VPV

[Ebe95] Differ Ebert. Variation in parasite virulence is not an indicator for the evolution of benevolence. Conservation Biology, 9(6): 1652–1653, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Epperly:1995:STN

[EBV95] Sheryan P. Epperly, Joanne Braun, and Allison Veishlow. Sea turtles in North Carolina waters. Las tortugas marinas en las aguas de Carolina del Norte. Conservation Biology, 9(2):384– 394, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Edwards:1996:AWH

[EDFM96] Thomas C. Edwards, Jr., Elena T. Deshler, Dan Foster, and Gretchen G. Moisen. Adequacy of wildlife habitat relation models for estimating spatial distributions of terrestrial vertebrates. La adecuación de los modelos de relación del habitat de la vida silvestre para estimar la distribución espacial de los vertebrados terrestres. Conservation Biology, 10(1):263–270, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Eduardo:1991:NCS

[Edu91] Santana C. Eduardo. Nature conservation and sustainable development in Cuba. Conservation Biology, 5(1):13–17, March 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Echelle:1997:GIE

[EE97] Anthony A. Echelle and Alice F. Echelle. Genetic introgression of endemic taxa by non-natives: a case study with Leon Springs pupfish and sheepshead minnow. Introgresión genética de taxas endémicos debido a especies no-nativas: un caso de estudio del pez cachorro de Leon Springs y el pez bolín. Conservation Biology, 11(1):153–161, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

East:1996:SAW

[EG96] Marion L. East and Joshua Ginsberg. Survivorship in African wild dogs. Conservation Biology, 10(2):313–315, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

East:1997:SHR

[EHBE97] Marion L. East, Heribert Hofer, Roger Burrows, and Marion L. East. Stress hormones and radiocollaring of African wild dogs. Conservation Biology, 11(6):1451–1453, December 1997. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See reply [Cre97].

Ehrenfeld:1990:LVV

[Ehr90] David Ehrenfeld. The lessons of Valdez. Conservation Biology, 4(1):1–3, March 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ehrenfeld:1991:SSS

[Ehr91] David Ehrenfeld. Special section: The Sustainable Biosphere Initiative. Conservation Biology, 5(4):547, December 1991. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ehrenfeld:1992:BC

[Ehr92a] David Ehrenfeld. The business of conservation. Conservation Biology, 6(1):1–3, March 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ehrenfeld:1992:CEU

[Ehr92b] David Ehrenfeld. Conservation and economic unification. Conservation Biology, 6(4):483–484, December 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ehrenfeld:1993:MCB

[Ehr93] David Ehrenfeld. The making of Conservation Biology. Conservation Biology, 7(4):743–745, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ehrenfeld:1994:TP

[Ehr94a] David Ehrenfeld. Talking to the Press. Conservation Biology, 8(3):617, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ehrlich:1994:BTP

[Ehr94b] Anne H. Ehrlich. Blunt talk on population control. Conservation Biology, 8(1):318, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ehrenfeld:1995:RBB

[Ehr95] David Ehrenfeld. Reply to Britten, Brussard, and Murphy. Conservation Biology, 9(1):4, February 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [BBM94c].

Engel:1998:EOT

[EK98] Jonna Engel and Rikk Kvitek. Effects of otter trawling on a benthic community in Monterey Bay National Marine Sanctuary. Conservation Biology, 12(6):1204–1214, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Eldridge:1999:ULL

[EKL^{+}99] Mark D. B. Eldridge, Juliet M. King, Anne K. Loupis, Peter B. S. Spencer, Andrea C. Taylor, Lisa C. Pope, and Graham P. Hall. Unprecedented low levels of genetic variation and inbreeding depression in an island population of the black-footed rockwallaby. Niveles bajos sin precedente de la variación genética y depresión por intracruza del canguro wallaby de patas negras en una población isleña. Conservation Biology, 13(3):531–541, June 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Elam:1996:IGC

[Ela96] Diane R. Elam. Importance of genetics to conservation biology. Conservation Biology, 10(1):308–309, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Elliott:1993:VDA

[ELS93] Katherine J. Elliott, David L. Loftis, and Klaus Steinbeck. Vegetation diversity after logging in the Southern Appalachians. Conservation Biology, 7(2):220–221, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Esseen:1998:EEE

[ER98] Per-Anders Esseen and Karl-Erik Renhorn. Edge effects on an epiphytic lichen in fragmented forests. Conservation Biology, 12(6):1307–1317, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Eriksson:1999:GCE

[Eri99] Lena Eriksson. Graduate conservation education. Conservation Biology, 13(5):955, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Erwin:1991:HMS

[Erw91] Terry L. Erwin. How many species are there?: Revisited. Conservation Biology, 5(3):330–333, September 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Engels:1994:NCB

[ES94] Thomas M. Engels and Charles W. Sexton. Negative correlation of blue jays and golden-cheeked warblers near an urbanizing area. Conservation Biology, 8(1):286–290, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Estes:1998:CAR

[Est98] J. A. Estes. Concerns about rehabilitation of oiled wildlife*. Conservation Biology, 12(5):1156–1157, October 1998. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ehrlen:1998:DPA

[EV98] Johan Ehrlén and Jan Van Groenendael. Direct perturbation analysis for better conservation. Conservation Biology, 12(2):470–474, April 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ewel:1991:YWG

[Ewe91] John J. Ewel. Yes, we got some bananas. Conservation Biology, 5(3):423–425, September 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Faith:1996:CPP

[Fai96] Daniel P. Faith. Conservation priorities and phylogenetic pattern. Conservation Biology, 10(4):1286–1289, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Furlow:1995:PPR

[FAP95] F. Bryant Furlow and Tara Amrijo-Prewitt. Peripheral populations and range collapse. Conservation Biology, 9(6):1345, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Forbes:1996:GVN

[FB96] Stephen H. Forbes and Diane K. Boyd. Genetic variation of naturally colonizing wolves in the Central Rocky Mountains. Variación genética de lobos colonizadores naturales en las Montañas Rocollosas Centrales. Conservation Biology, 10 (4):1082–1090, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Forbes:1997:GSM

[FB97] Stephen H. Forbes and Diane K. Boyd. Genetic structure and migration in native and reintroduced Rocky Mountain wolf populations. Estructura genética y migración de poblaciones nativas y reintroducidas del lobo de las Montañas Rocallosas. Conservation Biology, 11(5):1226–1234, October 1997. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Frederick:1996:CLN

[FBFO96] Peter C. Frederick, Keith L. Bildstein, Bruce Fleury, and John Ogden. Conservation of large, nomadic populations of white ibises (Eudocimus albus) in the United States. La conservación de grandes poblaciones nómadas de ibises blancas (Eudocimus albus) en los Estados Unidos. Conservation Biology, 10(1):

203–216, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

FitzSimmons:1995:PHG

[FBS95] Nancy N. FitzSimmons, Steven W. Buskirk, and Michael H. Smith. Population history, genetic variability, and Horn growth in bighorn sheep. Historia poblacional, variabilidad genética y crecimiento de la cornamenta de los carneros de las Montañas Rocosas. Conservation Biology, 9(2):314–323, April 1995. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

France:1993:ECL

[FC93] R. L. France and N. C. Collins. Extirpation of crayfish in a lake affected by long-range anthropogenic acidification. Extinción del cangrejo de río en un lago afectado por acidificación antropogénica de largo alcance. Conservation Biology, 7(1):184–188, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fule:1999:FRC

[FC99] Peter Z. Fulé and W. Wallace Covington. Fire regime changes in La Michilía Biosphere Reserve, Durango, Mexico. Cambios en el régimen de incendios en la Reserva de la Biosfera de La Michilía, Durango, México. Conservation Biology, 13(3):640– 652, June 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Friesen:1999:NSN

[FCM99] Lyle Friesen, Michael D. Cadman, and R. Jock Mackay. Nesting success of neotropical migrant songbirds in a highly fragmented landscape. Éxito de nidación de aves neotropicales migratorias en un paisaje altamente fragmentado. Conservation Biology, 13(2):338–346, April 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fenton:1998:BLT

[FCR^{+}98a] M. Brock Fenton, David H. M. Cumming, I. L. Rautenbach, Graeme S. Cumming, Meg S. Cumming, Gavin Ford, Russell D. Taylor, Jenna Dunlop, Mark D. Hovorka, Dave S. Johnston, Christine V. Portfors, Matina C. Kalcounis, and Zaccheus Mahlanga. Bats and the loss of tree canopy in African woodlands. Conservation Biology, 12(2):399–407, April 1998. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Freville:1998:PEP

[FCR^{+}98b] Hélène Fréville, Bruno Colas, Joëlle Ronfort, Miquel Riba, and Isabelle Olivieri. Predicting endemism from population structure of a widespread species: Case study in Centaurea maculosa Lam. (Asteraceae). Conservation Biology, 12(6):1269– 1278, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fernandez-Duque:1994:MAV

[FDV94] Eduardo Fernandez-Duque and Claudia Valeggia. Metaanalysis: a valuable tool in conservation research. El metaanálisis: una herramienta valiosa en la investigación conservacionista. Conservation Biology, 8(2):555–561, June 1994. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Friesen:1995:ERD

[FEM95] Lyle E. Friesen, Paul F. J. Eagles, and R. J. Mackay. Effects of residential development on forest-dwelling neotropical migrant songbirds efectos del desarrollo urbano sobre los bosques habitados por aves neotropicales migratorias. Conservation Biology, 9(6):1408–1414, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Farley:1994:ASR

[FESS94] Greg H. Farley, Lisa M. Ellis, James N. Stuart, and Norman J. Scott, Jr. Avian species richness in different-aged stands of riparian forest along the Middle Rio Grande, New Mexico. Riqueza específica de aves en sitios de distinta edad en bosques ribereños a lo largo del valle medio del Río Grande, Nuevo México. Conservation Biology, 8(4):1098–1108, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fearnside:1995:CGA

[FF95] Philip M. Fearnside and João Ferraz. A conservation gap analysis of Brazil’s Amazonian vegetation. Conservation Biology, 9(5):1134–1147, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fleischner:1996:LGR

[FFH96] Thomas L. Fleischner, Denzel Ferguson, and Steven G. Herman. Livestock grazing: Replies to Brown and McDonald.

Conservation Biology, 10(4):927–930, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fleischer:1995:GSE

[FFL95] Robert C. Fleischer, Garth Fuller, and David B. Ledig. Genetic structure of endangered clapper rail (Rallus longirostris) populations in Southern California. Conservation Biology, 9(5): 1234–1243, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fuller:1995:PDR

[FGG^{+}95] R. J. Fuller, R. D. Gregory, D. W. Gibbons, J. H. Marchant, J. D. Wilson, S. R. Baillie, and N. Carter. Population declines and range contractions among lowland farmland birds in Britain. Declinaciones poblacionales y contracciones en el área de distribución de aves que habitan tierras bajas agrícolas en Gran Bretaña. Conservation Biology, 9(6):1425–1441, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fielding:1995:TGB

[FH95] Alan H. Fielding and Paul F. Haworth. Testing the generality of bird-habitat models. Probando la generalidad de los modelos de hábitat para aves. Conservation Biology, 9(6):1466–1481, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Findlay:1997:ACS

[FH97] C. Scott Findlay and Jeff Houlahan. Anthropogenic correlates of species richness in southeastern Ontario wetlands. Correlativos antropogénicos de la riqueza de especies en humedales del sureste de Ontario. Conservation Biology, 11(4):1000–1009, August 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fimbel:1994:ECS

[Fim94] Cheryl Fimbel. Ecological correlates of species success in modified habitats may be disturbance- and site-specific: The primates of Tiwai Island. La correlación ecológica del éxito de las especies en hábitats modificados podrían ser específicos al tipo de perturbación y al sitio: Los primates de la Isla Tiwai. Conservation Biology, 8(1):106–113, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fisher:1994:PPS

[Fis94] Frank W. Fisher. Past and present status of Central Valley Chinook salmon. Conservation Biology, 8(3):870–873, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fitzgerald:1994:TLP

[Fit94] Lee A. Fitzgerald. Tupinambis lizards and people: a sustainable use approach to conservation and development. Conservation Biology, 8(1):12–16, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fa:1995:IMH

[FJPC95] Julia E. Fa, Javier Juste, Jaime Perez Del Val, and Javier Castroviejo. Impact of market hunting on mammal species in Equatorial Guinea. Conservation Biology, 9(5):1107–1115, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ferreira:1997:EFF

[FL97] Leandro V. Ferreira and William F. Laurance. Effects of forest fragmentation on mortality and damage of selected trees in Central Amazonia. Conservation Biology, 11(3):797–801, June 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fleischner:1990:ISP

[Fle90] Thomas L. Fleischner. Integrating science and passion in conservation education. Conservation Biology, 4(4):452–453, December 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fleischner:1994:ECL

[Fle94a] Thomas L. Fleischner. Ecological costs of livestock grazing in Western North America. Costos ecológicos del pastoreo de ganado en el oeste de Estados Unidos. Conservation Biology, 8(3):629–644, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fleming:1994:CBC

[Fle94b] Ian A. Fleming. Captive breeding and the conservation of wild salmon populations. Conservation Biology, 8(3):886–888, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fleishman:1995:HVB

[Fle95] Erica Fleishman. A holistic view of butterflies. Conservation Biology, 9(4):968–969, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fleischner:1996:CBF

[Fle96] Thomas L. Fleischner. Conservation biology fundamentals. Conservation Biology, 10(2):692–693, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fahrig:1994:CFP

[FM94] Lenore Fahrig and Gray Merriam. Conservation of fragmented populations. Conservación de poblaciones fragmentadas. Conservation Biology, 8(1):50–59, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Forester:1996:MHF

[FM96] Deborah J. Forester and Gary E. Machlist. Modeling human factors that affect the loss of biodiversity. Moldelado de factores humanos que afectan la pérdida de biodiversidad. Conservation Biology, 10(4):1253–1263, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fitzgibbon:1995:SHA

[FMF95] Clare D. Fitzgibbon, Hezron Mogaka, and John H. Fanshawe. Subsistence hunting in Arabuko-Sokoke Forest, Kenya, and its effects on mammal populations. Conservation Biology, 9(5): 1116–1126, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fu:1998:CBS

[FNC98] Yong-Bi Fu, Gene Namkoong, and John E. Carlson. Comparison of breeding strategies for purging inbreeding depression via simulation. Conservation Biology, 12(4):856–864, August 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Frissell:1992:TCB

[FNN92] Christopher A. Frissell, Richard K. Nawa, and Reed Noss. Is there any conservation biology in “New Perspectives?”: a response to Salwasser. Conservation Biology, 6(3):461, September 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [Sal92].

Foley:1994:PET

[Fol94] Patrick Foley. Predicting extinction times from environmental stochasticity and carrying capacity. Predicción del tiempo de extinción a partir de estocasticidad ambiental y capacidad de carga. Conservation Biology, 8(1):124–137, March 1994. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Forsyth:1994:TPM

[For94] Adrian Forsyth. Ted Parker: In memoriam. Conservation Biology, 8(1):293–294, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

France:1995:PME

[FP95] R. L. France and R. H. Peters. Predictive model of the effects on lake metabolism of decreased airborne litterfall through riparian deforestation. Modelo predictivo sobre los efectos de la disminución en el ingreso de detritos transportados por el aire a través de la deforestación ribereña sobre el metabolismo lacustre. Conservation Biology, 9(6):1578–1586, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Farnsworth:1993:EEF

[FR93] Elizabeth J. Farnsworth and Judy Rosovsky. The ethics of ecological field experimentation. La ética de la experimentación ecológica en el terreno. Conservation Biology, 7(3):463–472, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Frazer:1992:STC

[Fra92] Nat B. Frazer. Sea turtle conservation and halfway technology. Conservation Biology, 6(2):179–184, June 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fraver:1994:VRA

[Fra94] Shawn Fraver. Vegetation responses along edge-to-interior gradients in the mixed hardwood forests of the Roanoke River Basin, North Carolina. Respuestas de la vegetación a lo largo de gradientes de borde-hacia-el interior de bosques latifoliados mixtos de la cuenca del río Roanoke, en Carolina del Norte. Conservation Biology, 8(3):822–832, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Frankham:1995:IET

[Fra95a] Richard Frankham. Inbreeding and extinction: a threshold effect. Endocría y extinción: un efecto de umbral. Conservation Biology, 9(4):792–799, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Franz:1995:FMN

[Fra95b] Richard Franz. Florida’s master natural historian, Archie Carr. Conservation Biology, 9(4):971–972, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Frankham:1996:RGV

[Fra96] Richard Frankham. Relationship of genetic variation to population size in wildlife. La relación entre la variacion genética y el tamaño poblacional en vida silvestre. Conservation Biology, 10(6):1500–1508, December 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

France:1997:MCW

[Fra97] R. L. France. Macroinvertebrate colonization of woody debris in Canadian Shield lakes following riparian clearcutting. Colonización de detritus de madera por macroinvertebrados en lagos canadienses. Conservation Biology, 11(2):513–521, April 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Frankham:1998:IEI

[Fra98] Richard Frankham. Inbreeding and extinction: Island populations. Conservation Biology, 12(3):665–675, June 1998. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fritz:1992:CLP

[Fri92] Edward C. Fritz. Conservation and logging practices. Conservation Biology, 6(1):6–7, March 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Frissell:1993:TEE

[Fri93] Christopher A. Frissell. Topology of extinction and endangerment of native fishes in the Pacific Northwest and California (U.S.A.). Topología de la extinción y el amenazamiento de peces nativos en el Noroeste Pacífico y California (EEUU). Conservation Biology, 7(2):342–354, June 1993. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Frost:1992:IEU

[Fro92] Thomas M. Frost. Insights from an expert on the use of ecology. Conservation Biology, 6(1):154–155, March 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Franklin:1991:PCP

[FS91] Janet Franklin and David W. Steadman. The potential for conservation of Polynesian birds through habitat mapping and species translocation. Conservation Biology, 5(4):506–521, December 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

France:1992:PCO

[FS92] R. France and M. Sharp. Polynyas as centers of organization for structuring the integrity of Arctic marine communities. Conservation Biology, 6(3):442–446, September 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fernandes:1995:ERL

[FS95] David N. Fernandes and Robert L. Sanford, Jr. Effects of recent land-use practices on soil nutrients and succession under tropical wet forest in Costa Rica. Los efectos de las prácticas recientes de uso de la tierra sobre los nutrientes del suelo y la sucesion bajo un régimen de selva tropical húmeda en Costa Rica. Conservation Biology, 9(4):915–922, August 1995. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fisher:1996:DAC

[FS96] Robert N. Fisher and H. Bradley Shaffer. The decline of amphibians in California’s Great Central Valley. La declinación de anfibios en el Gran Valle Central de California. Conservation Biology, 10(5):1387–1397, October 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fischer:1997:LEP

[FS97] Markus Fischer and Jürg Stöcklin. Local extinctions of plants in remnants of extensively used calcareous grasslands 1950– 1985. Extinciónes locales de plantas en remanentes de pastizales calcáreos de uso extensivo entre 1950 y 1985. Conservation Biology, 11(3):727–737, June 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Freed:1996:SML

[FSCL96] Leonard A. Freed, Thomas B. Smith, John H. Carothers, and Jaan K. Lepson. Shrinkage is not the most likely cause of bill change in i’iwi: a rejoinder to Winker. Conservation Biology, 10(2):659–660, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [Win96b].

Fujita:1991:FFC

[FT91] Marty S. Fujita and Merlin D. Tuttle. Flying foxes (Chiroptera: Pteropodidae): Threatened animals of key ecological and economic importance. Conservation Biology, 5(4):455–463, December 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Forbes:1996:CBM

[FT96] Graham J. Forbes and John B. Theberge. Cross-boundary management of Algonquin Park wolves. Manejo de lobos fuera de los limites del Parque Algonquin. Conservation Biology, 10 (4):1091–1097, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Furlow:1995:ETO

[Fur95] F. Bryant Furlow. Effects of trait ontogeny and structure on stress resistance. Conservation Biology, 9(3):477–478, June 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Foose:1995:L

[FVK95] Thomas J. Foose, Nico Van Strien, and Mohd Khan bin Momin Khan. Letter. Conservation Biology, 9(5):977–978, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fleishman:1999:CPO

[FWB^{+}99] Erica Fleishman, Gary H. Wolff, Carol L. Boggs, Paul R. Ehrlich, Alan E. Launer, John O. Niles, and Taylor H. Ricketts. Conservation in practice: Overcoming obstacles to implementation. Conservation Biology, 13(2):450–452, April 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fox:1996:MCB

[FYP96] Jefferson Fox, Pralad Yonzon, and Nancy Podger. Mapping conflicts between biodiversity and human needs in Langtang National Park, Nepal. Mapeando conflictos entre la biodiversidad y las necesidades humanas en el Parque Nacional de Langtang, Nepal. Conservation Biology, 10(2):562–569, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

GomezDeSilvaGarza:1996:CIS

[G96] Héctor Gómez De Silva Garza. The conservation importance of semiendemic species. Conservation Biology, 10(2):674–675, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ginsberg:1995:HSA

[GAC^{+}95] J. R. Ginsberg, K. A. Alexander, S. Creel, P. W. Kat, J. W. Mcnutt, and M. G. L. Mills. Handling and survivorship of African wild dog (Lycaon pictus) in five ecosystems. Manipulación y supervivencia del perro salvaje africano (Lycaon pictus) en cinco ecosistemas. Conservation Biology, 9(3):665–674, June 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Garshelis:1997:SOM

[Gar97] David L. Garshelis. Sea otter mortality estimated from carcasses collected after the Exxon Valdez oil spill. Mortalidad de nutria marina estimada a partir de cuerpos colectados después del derrame del Exxon Valdez. Conservation Biology, 11(4): 905–916, August 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gaston:1991:ENI

[Gas91a] Kevin J. Gaston. Estimates of the near-imponderable: a reply to Erwin. Conservation Biology, 5(4):564–566, December 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gaston:1991:MGI

[Gas91b] Kevin J. Gaston. The magnitude of global insect species richness. Conservation Biology, 5(3):283–296, September 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gates:1994:EEA

[Gat94] J. Edward Gates. Edge effects on avian nest success. Conservation Biology, 8(3):618–620, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gates:1995:OHY

[Gat95] Simon Gates. One hundred years of ornithology. Conservation Biology, 9(3):698–699, June 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gaston:1995:RBS

[GB95a] Kevin J. Gaston and Tim M. Blackburn. Rarity and body size: Some cautionary remarks. Conservation Biology, 9(1): 210–213, February 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Goodrich:1995:CAN

[GB95b] J. M. Goodrich and S. W. Buskirk. Control of abundant native vertebrates for conservation of endangered species. El control de vertebrados nativos abundavies para la conservacion de especies en peligro. Conservation Biology, 9(6):1357–1364, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gaston:1996:CIG

[GB96a] Kevin J. Gaston and Tim M. Blackburn. Conservation implications of geographic range size–body size relationships. Conservation Biology, 10(2):638–646, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gaston:1996:RBS

[GB96b] Kevin J. Gaston and Tim M. Blackburn. Rarity and body size: Importance of generality. Conservation Biology, 10(4): 1295–1298, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Galicia:1997:EMT

[GB97] Eduardo Galicia and Guy A. Baldassarre. Effects of motorized tourboats on the behavior of nonbreeding American flamingos in Yucatan, Mexico. Efecto de lanchas turísticas motorizadas en el comportamiento del flamenco americano en Yucátan, México. Conservation Biology, 11(5):1159–1165, October 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Greenberg:1997:BPS

[GBAR97] Russell Greenberg, Peter Bichier, Andrea Cruz Angon, and Robert Reitsma. Bird populations in shade and sun coffee plantations in Central Guatemala. Poblaciones de aves en plantaciones cafetaleras en sombra y sol en la región central de Guatemala. Conservation Biology, 11(2):448–459, April 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gippoliti:1997:CBZ

[GC97] Spartaco Gippoliti and Giuseppe Maria Carpaneto. Captive breeding, zoos, and good sense. Conservation Biology, 11 (3):806–807, June 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gerber:1999:QAE

[GD99a] Leah R. Gerber and Douglas P. Demaster. A quantitative approach to Endangered Species Act classification of long-lived vertebrates: Application to the North Pacific humpback whale. Aproximación cuantitativa a la clasificación de vertebrados de vida larga del Acta de Especies en Peligro: Aplicación a la ballena jorobada del Pacífico Norte. Conservation Biology, 13(5): 1203–1214, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Greenberg:1999:DRB

[GD99b] Russell Greenberg and Sam Droege. On the decline of the rusty blackbird and the use of ornithological literature to document long-term population trends. Declive de poblaciones del mirlo Euphagus carolensis y uso de la literatura ornitológica para documentar tendencias poblacionales de largo plazo. Conservation Biology, 13(3):553–559, June 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Griffiths:1996:SR

[GDB96] Huw I. Griffiths, Angus Davison, and Johnny Birks. Species reintroductions. Conservation Biology, 10(4):923, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gerber:1999:GWV

[GDK99] Leah R. Gerber, Douglas P. Demaster, and Peter M. Kareiva. Gray whales and the value of monitoring data in implementing the U.S. Endangered Species Act. Ballenas grises y el valor de datos de monitoreo en la implementación de el Acta de Especies en Peligro de Extinción de los Estados Unidos. Conservation Biology, 13(5):1215–1219, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gehrt:1996:HPP

[Geh96] Stanley D. Gehrt. The human population problem: Educating and changing behavior. Conservation Biology, 10(3):900–903, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Geist:1990:AVB

[Gei90] V. Geist. Agriculture versus bison in Canada’s Wood Buffalo National Park. Conservation Biology, 4(4):345–346, December 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Geller:1999:DNM

[Gel99] Jonathan B. Geller. Decline of a native mussel masked by sibling species invasion. Disminuciones de un mejillón nativo encubierta por la invasión de una especie hermana. Conservation Biology, 13(3):661–664, June 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gensemer:1994:ETA

[Gen94] Robert W. Gensemer. Environmental toxicology: Approaches and definitions. Conservation Biology, 8(4):1171–1172, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gibbs:1990:EVO

[GF90] James P. Gibbs and John Faaborg. Estimating the viability of ovenbird and Kentucky warbler populations in forest fragments. Conservation Biology, 4(2):193–196, June 1990. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Golodetz:1997:HIL

[GF97] Alisa D. Golodetz and David R. Foster. History and importance of land use and protection in the North Quabbin region of Massachusetts (USA). Historia e importancia del uso y protección de suelo en la regíon North Quabbin de Massachussets (E.U.A.). Conservation Biology, 11(1):227–235, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ginzburg:1990:RDD

[GFA90] Lev R. Ginzburg, Scott Ferson, and H. Reşit Akçakaya. Reconstructibility of density dependence and the conservative assessment of extinction risks. Conservation Biology, 4(1):63–70, March 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Girondot:1998:FTE

[GFP98] Marc Girondot, Hélène Fouillet, and Claude Pieau. Feminizing turtle embryos as a conservation tool. Conservation Biology, 12(2):353–362, April 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ganzhorn:1999:LRD

[GFR^{+}99] Jörg U. Ganzhorn, Joanna Fietz, Edmond Rakotovao, Dorothea Schwab, and Dietmar Zinner. Lemurs and the regeneration of dry deciduous forest in Madagascar. Lemures y la regeneración de bosques deciduos secos en Madagascar. Conservation Biology, 13(4):794–804, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gerrodette:1990:DCC

[GG90a] Tim Gerrodette and William G. Gilmartin. Demographic consequences of changed pupping and hauling sites of the Hawaiian monk seal. Conservation Biology, 4(4):423–430, December 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gunatilleke:1990:DFR

[GG90b] I. A. U. N. Gunatilleke and C. V. S. Gunatilleke. Distribution of floristic richness and its conservation in Sri Lanka. Conservation Biology, 4(1):21–31, March 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grant:1997:RDF

[GG97] Peter R. Grant and B. Rosemary Grant. The rarest of Darwin’s finches. La especie más rara de los pinzones de Darwin. Conservation Biology, 11(1):119–126, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Geist:1999:RMM

[GG99] Cathy Geist and Susan M. Galatowitsch. Reciprocal model for meeting ecological and human needs in restoration projects. Modelo recíproco para alcanzar necesidades ecológicas y humanas en proyectos de restauracíon. Conservation Biology, 13 (5):970–979, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gehring:1996:RWP

[GGB^{+}96] Thomas M. Gehring, Joelle L. Gehring, Mark A. Beckel, Margaret A. Callahan, and L. David Mech. Recovering wolf populations. Conservation Biology, 10(1):5–7, February 1996. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Godt:1995:GDT

[GHB95] Mary Jo W. Godt, J. L. Hamrick, and Susan Bratton. Genetic diversity in a threatened wetland species, Helonias bullata (Liliaceae). Diversidad genética en una especie de humedales, Helonias bullata (Liliaceae) amenazada de extincion. Conservation Biology, 9(3):596–604, June 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gale:1997:RSW

[GHP97] George A. Gale, Lise A. Hanners, and Stephen R. Patton. Reproductive success of worm-eating warblers in a forested landscape. Conservation Biology, 11(1):246–250, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Giannecchini:1993:ENP

[Gia93] Joan Giannecchini. Ecotourism: New partners, new relationships. Conservation Biology, 7(2):429–432, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gilpin:1996:FEP

[Gil96] Michael Gilpin. Forty-eight parrots and the origins of population viability analysis. Conservation Biology, 10(6):1491–1493, December 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ginsberg:1994:LDC

[Gin94] Howard S. Ginsberg. Lyme disease and conservation. El mal de “Lyme” y la conservación. Conservation Biology, 8(2):343– 353, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ginsberg:1997:NAI

[Gin97] Joshua Ginsberg. A new approach to international conservation news. Conservation Biology, 11(6):1275, December 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ginsberg:1999:GCP

[Gin99] Joshua Ginsberg. Global conservation priorities. Conservation Biology, 13(1):5, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Godt:1996:GDP

[GJH96] Mary Jo W. Godt, Bart R. Johnson, and J. L. Hamrick. Genetic diversity and population size in four rare Southern Appalachian plant species. Diversidad genética y tamaño poblacional de cuatro especies de plantas raras en el sur de los apalaches. Conservation Biology, 10(3):796–805, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gamradt:1996:EIC

[GK96] Seth C. Gamradt and Lee B. Kats. Effect of introduced crayfish and mosquitofish on California newts. Efecto de la introducion del cangrejo de río y del “mosquitofish” al tritón de California. Conservation Biology, 10(4):1155–1162, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Goodsell:1999:EIM

[GK99] Jeff A. Goodsell and Lee B. Kats. Effect of introduced mosquitofish on Pacific treefrogs and the role of alternative prey. Efectos del pez mosquito y ranas arborícolas del Pacifico introducidas y su papel como presas alternativas. Conservation Biology, 13(4):921–924, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gamradt:1997:ANN

[GKA97] Seth C. Gamradt, Lee B. Kats, and Christopher B. Anzalone. Aggression by non-native crayfish deters breeding in California newts. Conservation Biology, 11(3):793–796, June 1997. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grayson:1993:MMG

[GL93a] Donald K. Grayson and Stephanie D. Livingston. Missing mammals on Great Basin mountains: Holocene extinctions and inadequate knowledge: Mamíferos desaparecidos en las Montañas de la Gran Cuenca: Extinciones durante el Holoceno e inadecuado conocimiento. Conservation Biology, 7(3): 527–532, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gullison:1993:RFD

[GL93b] Raymond E. Gullison and Elizabeth C. Losos. The role of foreign debt in deforestation in Latin America. El rol de la deuda externa en la deforestación de Latino América. Conservation Biology, 7(1):140–147, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Goodman:1994:SC

[GL94] Steve M. Goodman and Scott M. Lanyon. Scientific collecting. Conservation Biology, 8(1):314–315, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gresswell:1995:VAM

[GL95] Robert E. Gresswell and William J. Liss. Values associated with management of Yellowstone cutthroat trout in Yellowstone National Park. Valores asociados con el manejo de la trucha (Oncorhynchus clarki bouvieri) del Parque Nacional Yellowstone. Conservation Biology, 9(1):159–165, February 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gross:1998:MCB

[GLFM98] Kevin Gross, John R. Lockwood III, Cecil C. Frost, and William F. Morris. Modeling controlled burning and trampling reduction for conservation of Hudsonia montana. Conservation Biology, 12(6):1291–1301, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Glenn:1996:EWM

[GLFZ96] Edward P. Glenn, Christopher Lee, Richard Felger, and Scott Zengel. Effects of water management on the wetlands of the Colorado River Delta, Mexico. Impactos del manejo de agua sobre los humedales del Río Colorado, México. Conservation Biology, 10(4):1175–1186, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grubb:1991:CSB

[GM91] Peter J. Grubb and Robert M. May. Comments on the Sustainable Biosphere Initiative. Conservation Biology, 5(4):548–549, December 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See comments [Sch91, Tho91, Wil91].

Goodwin:1999:PIP

[GMF99] Brett J. Goodwin, Andrew J. McAllister, and Lenore Fahrig. Predicting invasiveness of plant species based on biological information. Predicción de la invasividad de especies en base a información biológica. Conservation Biology, 13(2):422–426, April 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ginsberg:1994:SBH

[GMG94] J. R. Ginsberg and E. J. Milner-Gulland. Sex-biased harvesting and population dynamics in ungulates: Implications for conservation and sustainable use. Cosecha sexualmente sesgada y dinámica poblacional en ungulados: Implicaciones para la conservación y el uso sostenido. Conservation Biology, 8(1):157– 166, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Garcia-Moreno:1996:RGP

[GMMR^{+}96] Jaime García-Moreno, Majorie D. Matocq, Michael S. Roy, Eli Geffen, and Robert K. Wayne. Relationships and genetic purity of the endangered Mexican wolf based on analysis of microsatellite loci. Relaciones y pureza genética del lobo mexicano basadas en el análisis de loci microsatélites. Conservation Biology, 10(2):376–389, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Garcia-Marin:1998:PNI

[GMSP98] J. L. García-Marín, N. Sanz, and C. Pla. Proportions of native and introduced brown trout in adjacent fished and unfished Spanish rivers. Conservation Biology, 12(2):313–319, April 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gotmark:1992:CUP

[GN92] Frank Götmark and Christer Nilsson. Criteria used for protection of natural areas in Sweden 1909–1986. Conservation Biology, 6(2):220–231, June 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Greenberg:1994:EHI

[GNH94] Cathryn H. Greenberg, Daniel G. Neary, and Larry D. Harris. Effect of high-intensity wildfire and silvicultural treatments on reptile communities in sand-pine scrub efecto de los incendios naturales de alta intensidad y de los tratamientos de silvicultura sobre las comunidades de reptiles en un brozal de ambiente arenoso. Conservation Biology, 8(4):1047–1057, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Godoy:1993:NTFa

[God93a] Ricardo Godoy. Non-timber forest products. Conservation Biology, 7(1):210–211, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Godoy:1993:NTFb

[God93b] Ricardo Godoy. Non-timber forest products: Where to go from here? Conservation Biology, 7(4):965–966, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Goerck:1997:PRB

[Goe97] Jaqueline M. Goerck. Patterns of rarity in the birds of the Atlantic forest of Brazil. Patrones de rareza en las aves del bosque Atlántico de Brasil. Conservation Biology, 11(1):112– 118, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Goldberg:1997:IGO

[Gol97a] Tony L. Goldberg. Inferring the geographic origins of “Refugee” chimpanzees in Uganda from mitochondrial DNA sequences. Conservation Biology, 11(6):1441–1446, December 1997. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Goldstein:1997:HMT

[Gol97b] Paul Z. Goldstein. How many things are there? A reply to Oliver and Beattie, Beattie and Oliver, Oliver and Beattie, and Oliver and Beattie. Conservation Biology, 11(2):571–574, April 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [OB93, OB96, OB97].

Goldstein:1999:CRS

[Gol99a] Paul Z. Goldstein. Clarifying the role of species in ecosystem management: a reply. Conservation Biology, 13(6):1515–1517, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [BO99].

Goldstein:1999:FEB

[Gol99b] Paul Z. Goldstein. Functional ecosystems and biodiversity buzzwords ecosistema funcional y biodiversidad como eslogans. Conservation Biology, 13(2):247–255, April 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See comments [Wal99a, Ris99, BO99] and reply [Gol99a].

Goodson:1994:PPS

[Goo94] Nike J. Goodson. Persistence and population size in mountain sheep: Why different interpretations? Conservation Biology, 8 (3):617–618, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Goodman:1995:RMD

[Goo95] Steven M. Goodman. Rattus on Madagascar and the dilemma of protecting the endemic rodent fauna. Conservation Biology, 9(2):450–453, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gotmark:1992:NEC

[Göt92] Frank Götmark. Naturalness as an evaluation criterion in nature conservation: a response to Anderson. Conservation Biology, 6(3):455, September 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gotelli:1995:ILS

[Got95] Nicholas J. Gotelli. Importance of a large-scale perspective. Conservation Biology, 9(2):469–470, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Goldingay:1997:ISS

[GP97] Ross L. Goldingay and Mary V. Price. Influence of season and a sympatric congener on habitat use by Stephens’ kangaroo rat. Influencia estaciónal y de un congenere simpátrico en el uso de hábitat de la rata canguro de Stephen. Conservation Biology, 11(3):708–717, June 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Giblin:1998:GSU

[GP98] David Giblin and Rich Pagen. Graduate students as an untapped resource for public outreach. Conservation Biology, 12 (6):1421–1422, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grady:1999:UCC

[GQ99] James M. Grady and Joseph M. Quattro. Using character concordance to define taxonomic and conservation units. Conservation Biology, 13(5):1004–1007, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Green:1998:THE

[GR98] Wendy C. H. Green and Aron Rothstein. Translocation, hybridization, and the endangered black-faced impala. Conservation Biology, 12(2):475–480, April 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gray:1995:DFR

[Gra95] Elizabeth M. Gray. DNA fingerprinting reveals a lack of genetic variation in northern populations of the western pond turtle (Clemmys marmorata). Conservation Biology, 9(5):1244–1255, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grajal:1998:JCH

[Gra98] Alejandro Grajal. Jacques Cousteau and his world’s inspiring beauty. Conservation Biology, 12(2):487, April 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grajal:1999:BNS

[Gra99] Alejandro Grajal. Biodiversity and the nation state: Regulating access to genetic resources limits biodiversity research in developing countries. Conservation Biology, 13(1):6–10, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Green:1992:AFG

[Gre92] Edwin J. Green. Assessing forest growth with point sampling data. Conservation Biology, 6(2):296–297, June 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Green:1996:AGT

[Gre96] Andy J. Green. Analyses of globally threatened anatidae in relation to threats, distribution, migration patterns, and habitat use. Análisis de la globalmente amenazada anatidae en relación con amenazas, distribución, patrones de migración y uso del hábitat. Conservation Biology, 10(5):1435–1445, October 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grumbine:1990:VPR

[Gru90] R. Edward Grumbine. Viable populations, reserve size, and federal lands management: a critique. Conservation Biology, 4(2):127–134, June 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grumbine:1994:WEM

[Gru94] R. Edward Grumbine. What is ecosystem management? Qué es manejo del ecosistema? Conservation Biology, 8(1):27–38, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grumbine:1996:CBT

[Gru96] Ed Grumbine. A conservation biology textbook for the general reader. Conservation Biology, 10(1):309–310, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grumbine:1997:RWE

[Gru97] R. Edward Grumbine. Reflections on “What is ecosystem management?” Reflexiones sobre “`?Que es el manejo de ecosistemas?”. Conservation Biology, 11(1):41–47, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Griffiths:1993:IHT

[GS93] Mike Griffiths and Carel P. Schaik. The impact of human traffic on the abundance and activity periods of Sumatran rain forest wildlife. Conservation Biology, 7(3):623–626, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gavin:1995:P

[GS95] Thomas A. Gavin and Paul W. Sherman. Proposition 80. Conservation Biology, 9(6):1343–1344, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gram:1999:PDP

[GS99] Wendy K. Gram and Victoria L. Sork. Population density as a predictor of genetic variation for woody plant species. Densidad poblacional como predictor de la variación genética de especies de plantas leñosas. Conservation Biology, 13(5):1079–1087, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Garrabou:1998:IDR

[GSAZ98] Joaquim Garrabou, Enric Sala, Antoni Arcas, and Mikel Zabala. The impact of diving on rocky sublittoral communities: a case study of a bryozoan population. Conservation Biology, 12(2):302–312, April 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gompper:1997:CIN

[GSB97] Matthew E. Gompper, Peter B. Stacey, and Joel Berger. Conservation implications of the natural loss of lineages in wild mammals and birds. Implicaciones de la conservación en la pérdida natural de linajes en mamiferos y aves silvestres. Conservation Biology, 11(4):857–867, August 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Guariguata:1998:L

[GSB98] Manuel R. Guariguata, Reinmar Seidler, and Kamal S. Bawa. Letters. Conservation Biology, 12(5):941, October 1998. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Glenn:1999:EPB

[GSB99] Travis C. Glenn, Wolfgang Stephan, and Michael J. Braun. Effects of a population bottleneck on whooping crane mitochondrial DNA variation. Efectos de un cuello de botella poblacional en la variación ADN mitocondrial de la grulla americana. Conservation Biology, 13(5):1097–1107, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Guries:1992:BBF

[Gur92] Raymond Guries. Of bulldozers and bunnies: Forest biodiversity for the common man. Conservation Biology, 6(1):153–154, March 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Guzman:1991:RCR

[Guz91] Héctor M. Guzmán. Restoration of coral reefs in Pacific Costa Rica. Conservation Biology, 5(2):189–194, June 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gilpin:1991:MCB

[GW91] Michael Gilpin and Christopher Wills. MHC and captive breeding: a rebuttal. Conservation Biology, 5(4):554–555, December 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Guan:1997:EII

[GW97] Rui-Zhang Guan and Peter Roy Wiles. Ecological impact of introduced crayfish on benthic fishes in a British lowland river. Impacto ecológico de un langostino de río introducido en poblaciones de peces bentónicos de un río de las tierras bajas Británicas. Conservation Biology, 11(3):641–647, June 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gompper:1998:PCB

[GW98] Matthew E. Gompper and Elizabeth S. Williams. Parasite conservation and the black-footed Ferret recovery program. Conservation Biology, 12(3):730–732, June 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gardner:1998:L

[GWF98] Alfred L. Gardner, Jerry T. Warren, and Robert France. Letters. Conservation Biology, 12(3):499–501, June 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Godt:1997:GDE

[GWH97] Mary Jo W. Godt, Joan Walker, and J. L. Hamrick. Genetic diversity in the endangered lily Harperocallis flava and a close relative, Tofieldia racemosa. Diversidad genética de la azucena en peligro Harperocallis flava, y un pariente cercano Tofieldia racemosa. Conservation Biology, 11(2):361–366, April 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ge:1999:CGD

[GWH^{+}99] Song Ge, Ke-Qing Wang, De-Yuan Hong, Wen-Hui Zhang, and Yuan-Gang Zu. Comparisons of genetic diversity in the endangered Adenophora lobophylla and its widespread congener, A. potaninii. Comparaciones de la diversidad genética de la especie amenazada Adenophora lobophylla y su congénere de distribución amplia A. Potaninii. Conservation Biology, 13 (3):509–513, June 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gilligan:1997:MAT

[GWM^{+}97] Dean M. Gilligan, Lynn M. Woodworth, Margaret E. Montgomery, David A. Briscoe, and Richard Frankham. Is mutation accumulation a threat to the survival of endangered populations? `?Es la Acumulación de mutaciones una amenaza para la supervivencia de poblaciones en peligro? Conservation Biology, 11(5):1235–1241, October 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Garrott:1993:OIC

[GWW93a] Robert A. Garrott, P. J. White, and Callie A. Vanderbilt White. Overabundance: an issue for conservation biologists? Conservation Biology, 7(4):946–949, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Goldschmidt:1993:CEI

[GWW93b] Tijs Goldschmidt, Frans Witte, and Jan Wanink. Cascading effects of the introduced Nile perch on the detritivorous/ phytoplanktivorous species in the sublittoral areas of Lake Victoria. Efectos en cascada por la introduccion de la perca del Nilo sobre las especies detritívoras/fitoplanctívoras en las áreas subtidales del Lago Victoria. Conservation Biology, 7(3):686–700, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ge:1997:AVO

[GZWR97] Song Ge, Da-Ming Zhang, Hai-Qun Wang, and Guang-Yuan Rao. Allozyme variation in Ophiopogon xylorrhizus, an extreme endemic species of Yunnan, China. Conservation Biology, 11 (2):562–565, April 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Haas:1995:DUC

[Haa95] Carola A. Haas. Dispersal and use of corridors by birds in wooded patches on an agricultural landscape. La dispersión y el uso de corredores por aves en parches boscosos de un paisaje agrícola. Conservation Biology, 9(4):845–854, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Haber:1996:BCE

[Hab96] Gordon C. Haber. Biological, conservation, and ethical implications of exploiting and controlling wolves. Implicaciones biológicas, conservacionistas y eticas de la explotación y control de labos. Conservation Biology, 10(4):1068–1081, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hackel:1999:CCF

[Hac99] Jeffrey D. Hackel. Community conservation and the future of Africa’s wildlife. Conservación con comunidades y el futuro de la conservación de la vida silvestre africana. Conservation Biology, 13(4):726–734, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hsu:1997:WCT

[HAD^{+}97] Minna J. Hsu, Govindasamy Agoramoorthy, Konjev Desender, Leon Baert, and Héctor Reyes Bonilla. Wildlife conservation in Taiwan. Conservation Biology, 11(4):834–838, August 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hagan:1992:CBW

[Hag92] John M. Hagan. Conservation biology when there is no crisis — yet. Conservation Biology, 6(3):475–476, September 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hagan:1995:ESC

[Hag95] John M. Hagan. Environmentalism and the science of conservation biology. Conservation Biology, 9(5):975–976, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hamilton:1994:ESC

[Ham94] Matthew B. Hamilton. Ex situ conservation of wild plant species: Time to reassess the genetic assumptions and implications of seed banks. Conservacion ex situ de especies de plantas silvestres: Tiempo de reevaluar los supuestos genéticos y las implicaciones de los bancos de semilias. Conservation Biology, 8 (1):39–49, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harrison:1992:TTI

[Har92] Robert L. Harrison. Toward a theory of inter-refuge corridor design. Conservation Biology, 6(2):293–295, June 1992. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hartshorn:1993:PC

[Har93] Gary Hartshorn. Pluralistic conservation. Conservation Biology, 7(1):211–212, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hartman:1994:LTP

[Har94] Göran Hartman. Long-term population development of a reintroduced beaver (Castor fiber) population in Sweden. Desarrollo a largo plazo de unq pobación reintroducida de castor (Castor fiber) en Suecia. Conservation Biology, 8(3):713–717, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harcourt:1995:PVE

[Har95a] A. H. Harcourt. Population viability estimates: Theory and practice for a wild gorilla population. Estimaciones de la viabilidad poblacional: teoría y práctica para una población de gorilas salvajes. Conservation Biology, 9(1):134–142, February 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hartshorn:1995:SF

[Har95b] Gary S. Hartshorn. Sustainable forestry? Conservation Biology, 9(1):225–226, February 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Haskell:1995:REF

[Has95] David G. Haskell. A reevaluation of the effects of forest fragmentation on rates of bird-nest predation. Conservation Biology, 9(5):1316–1318, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Haupt:1995:SCA

[Hau95] Lyanda L. Haupt. Scientists in conservation activism. Conservation Biology, 9(3):691–693, June 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hall:1990:EEI

[HB90] Charles A. S. Hall and Dennis P. Bradley. Ecological economics: Its implications for forest management and research (a workshop summary). Conservation Biology, 4(3):221–226, September 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hindar:1994:SCI

[HB94] Kjetil Hindar and Torveig Balstad. Salmonid culture and interspecific hybridization. Conservation Biology, 8(3):881–882, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hernandez:1995:ECC

[HB95] Héctor M. Hernández and Rolando T. Bárcenas. Endangered cacti in the Chihuahuan desert: I. Distribution patterns. Conservation Biology, 9(5):1176–1188, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hernandez:1996:ECC

[HB96] Héctor M. Hernández and Rolando T. Bárcenas. Endangered cacti in the Chihuahuan desert: II. Biogeography and conservation. Cactácuas amenazadas en el desierto chihuahuense: II. Biogeografía y conservación. Conservation Biology, 10(4):1200– 1209, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Haig:1993:PVA

[HBA93] Susan M. Haig, James R. Belthoff, and David H. Allen. Population viability analysis for a small population of red-cockaded woodpeckers and an evaluation of enhancement strategies analisis de viabilidad de una pequeña población del pájaro carpintero cabeza roja y evaluación de estrategias para el mejoramiento. Conservation Biology, 7(2):289–301, June 1993. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Haig:1990:MOP

[HBD90] Susan M. Haig, Jonathan D. Ballou, and Scott R. Derrickson. Management options for preserving genetic diversity: Reintroduction of Guam rails to the wild. Conservation Biology, 4(3): 290–300, September 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Heppell:1996:AFM

[HC96] Selina S. Heppell and Larry B. Crowder. Analysis of a fisheries model for harvest of hawksbill sea turtles (Eretmochelys imbricata). Analisis de un modelo de pesquerias para las tortugas marinas de Cary (Eretmochelys imbricata). Conservation Biology, 10(3):874–880, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hamann:1999:IAF

[HC99] Andreas Hamann and Eberhard Curio. Interactions among frugivores and fleshy fruit trees in a Philippine submontane rainforest. Interacción entre especies de frutívoros y Árboles de frutas carnosas en un bosque mesófito de montaña Filipinas. Conservation Biology, 13(4):766–773, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hall:1994:LGV

[HCB94] Pamela Hall, Michael R. Chase, and Kamaljit S. Bawa. Low genetic variation but high population differentiation in a common tropical forest tree species. Bajos niveles de variabilidad genética pero altos niveles de diferenciación poblacional en una especie común de arboles de forestas tropicales. Conservation Biology, 8(2):471–482, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Heard:1999:DIF

[HCBH99] Stephen B. Heard, Michael A. Campbell, Mary L. Bonine, and Stephen D. Hendrix. Developmental instability in fragmented populations of prairie phlox: a cautionary tale. Inestabilidad en el desarrollo de poblaciones fragmentadas de Phlox pilosa: una historia de advertencia. Conservation Biology, 13(2):274– 281, April 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Haila:1997:NBE

[HCH^{+}97] Yrjö Haila, Patrick J. Comer, Malcolm Hunter, Michael J. Samways, Clive Hambler, Martin R. Speight, Paul Hendricks, Stephen Herrero, F. Stephen Dobson, Andrew T. Smith, and Jinping Yu. A natural “Benchmark” for ecosystem function. Conservation Biology, 11(2):300–307, April 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hibbett:1996:IPS

[HD96] David S. Hibbett and Michael J. Donoghue. Implications of phylogenetic studies for conservation of genetic diversity in shiitake mushrooms. Implicaciones de estudios filogenéticos en la conservación de la diversidad genética de hongos shiitake. Conservation Biology, 10(5):1321–1327, October 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hoare:1999:CBP

[HD99] Richard E. Hoare and Johan T. Du Toit. Coexistence between people and elephants in African savannas. coexistencia entre humanos y elefantes en sabanas africanas. Conservation Biology, 13(3):633–639, June 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Holl:1995:KPC

[HDE95] Karen D. Holl, Gretchen C. Daily, and Paul R. Ehrlich. Knowledge and perceptions in Costa Rica regarding environment, population, and biodiversity issues. El conocimiento y la percepción en Costa Rica con respecto a temas del ambiente, la poblacion y la biodiversidad. Conservation Biology, 9(6):1548– 1558, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hegazy:1991:EIS

[HE91] Ahmad K. Hegazy and Naeem Mohamed Eesa. On the ecology, insect seed-predation, and conservation of a rare and endemic plant species: Ebenus amnitagei (Leguminosae). Conservation Biology, 5(3):317–324, September 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hiers:1997:EAD

[HE97] John Kevin Hiers and Jonathan P. Evans. Effects of anthracnose on dogwood mortality and forest composition of the Cumberland Plateau (U.S.A.). Conservation Biology, 11(6):1430– 1435, December 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Healey:1994:VLH

[Hea94] M. C. Healey. Variation in the life history characteristics of Chinook salmon and its relevance to conservation of the Sacramento winter run of Chinook salmon. Conservation Biology, 8 (3):876–877, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hedrick:1995:GFG

[Hed95] Philip W. Hedrick. Gene flow and genetic restoration: The Florida panther as a case study. Conservation Biology, 9(5): 996–1007, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hedrick:1996:BMC

[Hed96] Philip W. Hedrick. Bottleneck(s) or metapopulation in cheetahs. Conservation Biology, 10(3):897–899, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Heinen:1993:GPB

[Hei93] Joel T. Heinen. Global perspectives on biodiversity. Conservation Biology, 7(3):740–741, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Heinen:1995:RPP

[Hei95a] Joel T. Heinen. Reports on parks and protected areas. Conservation Biology, 9(2):471–472, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Heinz:1995:CGP

[Hei95b] Dan Heinz. Continued grazing a presupposition. Conservation Biology, 9(4):708–709, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Heinen:1996:HBI

[Hei96a] Joel T. Heinen. Human behavior, incentives, and protected area management. Conservation Biology, 10(2):681–684, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Heinen:1996:MB

[Hei96b] Joel T. Heinen. The market for biodiversity. Conservation Biology, 10(4):1303–1305, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Helenurm:1998:ODS

[Hel98] Kaius Helenurm. Outplanting and differential source population success in Lupinus guadalupensis. Conservation Biology, 12(1):118–127, February 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hodson:1995:RFP

[HEO95] Thomas J. Hodson, Fred Englander, and Hsu O’Keefe. Rain forest preservation, markets, and medicinal plants: Issues of property rights and present value. Conservation Biology, 9(5): 1319–1321, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Herrero:1990:L

[Her90] Stephen Herrero. Letter. Conservation Biology, 4(1):6–7, March 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hess:1994:CCC

[Hes94] George R. Hess. Conservation corridors and contagious disease: a cautionary note. Corredores de conservación y enfermedades contagiosas: una nota de precaución. Conservation Biology, 8 (1):256–262, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hissmann:1998:PMC

[HFS98] Karen Hissmann, Hans Fricke, and Jürgen Schauer. Population monitoring of the coelacanth (Latimeria chalumnae). Conservation Biology, 12(4):759–765, August 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hero:1997:EDA

[HG97a] Jean-Marc Hero and Graeme R. Gillespie. Epidemic disease and amphibian declines in Australia. Conservation Biology, 11 (4):1023–1025, August 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Human:1997:EAA

[HG97b] Kathleen G. Human and Deborah M. Gordon. Effects of Argentine ants on invertebrate biodiversity in northern California. Conservation Biology, 11(5):1242–1248, October 1997. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hayes:1996:GDN

[HGK^{+}96] John P. Hayes, Stanley Z. Guffey, Frank J. Kriegler, Gary F. McCracken, and Charles R. Parker. The genetic diversity of native, stocked, and hybrid populations of brook trout in the Southern Appalachians. La diversidad genética de poblaciones de trucha de arroyo nativas, sembradas e híbridas en el sur de las Montañas Apalaches. Conservation Biology, 10(5): 1403–1412, October 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hori:1993:LFC

[HGNK93] Michio Hori, Masta Mukwaya Gashagaza, Muderhwa Nshombo, and Hiroya Kawanabe. Littoral fish communities in Lake Tanganyika: Irreplaceable diversity supported by intricate interactions among species comunidades de peces litorales en el Lago Tanganyika: Diversidad irreemplazable sustentada por intrincadas interacciones entre especies. Conservation Biology, 7(3): 657–666, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hobbs:1992:DDI

[HH92] Richard J. Hobbs and Laura F. Huenneke. Disturbance, diversity, and invasion: Implications for conservation. Conservation Biology, 6(3):324–337, September 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harris:1994:GEP

[HH94a] Larry D. Harris and Thomas Hector. Greenway ecology, the prolegomenon. Conservation Biology, 8(2):603–605, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hedrick:1994:EPS

[HH94b] Philip W. Hedrick and Dennis Hedgecock. Effective population size in winter-run Chinook salmon. Conservation Biology, 8(3): 890–892, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hunter:1994:VSP

[HH94c] Malcolm L. Hunter, Jr. and Alan Hutchinson. The virtues and shortcomings of parochialism: Conserving species that are locally rare, but globally common. Conservation Biology, 8(4): 1163–1165, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hahn:1995:PLS

[HH95a] D. Caldwell Hahn and Jeff S. Hatfield. Parasitism at the landscape scale: Cowbirds prefer forests. Parasitismo a escala de paisaje: Preferencia de molothrus ater hacia los bosques. Conservation Biology, 9(6):1415–1424, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hobbs:1995:IAE

[HH95b] Richard J. Hobbs and Stella E. Humphries. An integrated approach to the ecology and management of plant invasions. Un enfoque integrado a la ecología y manejo de las invasiones de plantas. Conservation Biology, 9(4):761–770, August 1995. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hart:1996:SEZ

[HH96] John A. Hart and Jefferson S. Hall. Status of Eastern Zaire’s forest parks and reserves. Conservation Biology, 10(2):316– 327, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hartley:1998:MAF

[HH98] Mitschka J. Hartley and Malcolm L. Hunter. A meta-analysis of forest cover, edge effects, and artificial nest predation rates. Conservation Biology, 12(2):465–469, April 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hart:1997:CCS

[HHF^{+}97] Terese Hart, John Hart, Cheryl Fimbel, Robert Fimbel, William F. Laurance, Carrie Oren, Thomas T. Struhsaker, Howard C. Rosenbaum, Peter D. Walsh, Yvette Razafindrakoto, Michel Vely, and Rob DeSalle. Conservation and civil strife: Two perspectives from Central Africa. Conservation Biology, 11(2):308–314, April 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hedrick:1995:EPS

[HHH95] Philip W. Hedrick, Dennis Hedgecock, and Scott Hamelberg. Effective population size in winter-run Chinook salmon. Tamaño poblacional efectivo en las corridas invernales del salmón Chinook. Conservation Biology, 9(3):615–624, June 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hart:1996:CDN

[HHH96] Terese B. Hart, John A. Hart, and Jefferson S. Hall. Conservation in the declining nation state: a view from Eastern Zaire. Conservation Biology, 10(2):685–686, April 1996. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hayward:1999:RRB

[HHR99] Gregory D. Hayward, Stephen H. Henry, and Leonard F. Ruggiero. Response of red-backed voles to recent patch cutting in subalpine forest. Conservation Biology, 13(1):168–176, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hunter:1990:WWS

[HHWB90] Malcolm L. Hunter, Jr., Robert K. Hitchcock, and Barbara Wyckoff-Baird. Women and wildlife in Southern Africa. Conservation Biology, 4(4):448–451, December 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Higgs:1997:WGE

[Hig97] Eric S. Higgs. What is good ecological restoration? `?Que es una buena restauración ecológica? Conservation Biology, 11 (2):338–348, April 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hill:1995:LSR

[Hil95] C. J. Hill. Linear strips of rain forest vegetation as potential dispersal corridors for rain forest insects. Bandas de lineales vegetación de selvas tropicales como potenciales corredores dispersores para insectos de la selva tropical. Conservation Biology, 9(6):1559–1566, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hansen:1999:CTI

[HJ99] Lara J. Hansen and Michael L. Johnson. Conservation and toxicology: Integrating the disciplines*. Conservation Biology, 13(5):1225–1227, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hazan:1994:ACC

[HK94] Bruce W. Hazan and Thomas H. Kunz. Addressing the chiropteran crisis. Conservation Biology, 8(3):914, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Herbeck:1999:PSR

[HL99] Laura A. Herbeck and David R. Larsen. Plethodontid salamander response to silvicultural practices in Missouri Ozark forests. Respuesta de una salamandra pledodóntida a las prácticas silviculturales en los bosques de Missouri Ozark. Conservation Biology, 13(3):623–632, June 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hedrick:1996:DCB

[HLAS96] Philip W. Hedrick, Robert C. Lacy, Fred W. Allendorf, and Michael E. Soulé. Directions in conservation biology: Comments on Caughley. Direcciones en la biología de la conservación: Comentarios sobre Caughley. Conservation Biology, 10 (5):1312–1320, October 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [Cau93].

Holdgate:1993:LCR

[HM93] Martin Holdgate and David A. Munro. Limits to caring: a response. Conservation Biology, 7(4):938–940, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See response [Rob93a].

Holling:1996:CCP

[HM96] C. S. Holling and Gary K. Meffe. Command and control and the pathology of natural resource management. Comando-y-control y la patología del manejo de los recursos naturales. Conservation Biology, 10(2):328–337, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hecnar:1997:SSD

[HM97] Stephen J. Hecnar and Robert T. M’Closkey. Spatial scale and determination of species status of the green frog. Escala espacial y determinación del estatus de la especie de rana verde. Conservation Biology, 11(3):670–682, June 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hobbs:1998:BED

[HM98] Richard J. Hobbs and Harold A. Mooney. Broadening the extinction debate: Population deletions and additions in California and Western Australia. Conservation Biology, 12(2):271– 283, April 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Haig:1998:AMW

[HMO98] Susan M. Haig, David W. Mehlman, and Lewis W. Oring. Avian movements and wetland connectivity in landscape conservation. Conservation Biology, 12(4):749–758, August 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hanski:1996:QIF

[HMPK96] Ilkka Hanski, Atte Moilanen, Timo Pakkala, and Mikko Kuussaari. The quantitative incidence function model and persistence of an endangered butterfly metapopulation. El modelo cuantitativo de la función de incidencia y la persistencia de una metapoblación de una mariposa en peligro. Conservation Biology, 10(2):578–590, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Haight:1998:MDG

[HMW98] Robert G. Haight, David J. Mladenoff, and Adrian P. Wydeven. Modeling disjunct gray wolf populations in semi-wild landscapes. Conservation Biology, 12(4):879–888, August 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hanowski:1997:IWP

[HNC97] Joann M. Hanowski, Gerald J. Niemi, and Donald C. Christian. Influence of within-plantation heterogeneity and surrounding landscape composition on avian communities in hybrid poplar plantations. Influencia de la heterogeneidad intra-plantación y de la composición del paisaje circundante sobre comunidades de aves en plantaciones de alamo híbrido. Conservation Biology, 11(4):936–944, August 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hobbs:1996:BA

[Hob96] Richard J. Hobbs. Biodiversity in Australia. Conservation Biology, 10(5):1485–1486, October 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Higuchi:1996:STW

[HOF^{+}96] Hiroyoshi Higuchi, Kiyoaki Ozaki, Go Fujita, Jason Minton, Mutsuyuki Ueta, Masaki Soma, and Nagahisa Mita. Satellite tracking of white-naped crane migration and the importance of the Korean Demilitarized Zone. Seguimiento por satélite de la migración de la grulla nuca blanca y la importancia de la Zona Desmilitarizada Coreana. Conservation Biology, 10 (3):806–812, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Holt:1993:LWK

[Hol93] Sidney Holt. On loving whales, krill, and marine ecosystems. Conservation Biology, 7(3):451–452, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Holway:1995:DAA

[Hol95] David A. Holway. Distribution of the Argentine ant (Linepithema humile) in northern California. Conservation Biology, 9 (6):1634–1637, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Houghton:1995:HTN

[Hou95] R. A. Houghton. Human transformation of North America. Conservation Biology, 9(6):1661–1663, December 1995. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Howe:1994:MSD

[How94] Henry F. Howe. Managing species diversity in tallgrass prairie: Assumptions and implications manejo de diversidad especifica en praderas de “Tallgrass”: Supuestos e implicaciones. Conservation Biology, 8(3):691–704, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hartl:1994:GDE

[HP94] Günther B. Hartl and Zdzislaw Pucek. Genetic depletion in the European bison (Bison bonasus) and the significance of electrophoretic heterozygosity for conservation. Agotamiento genético en el bisón europeo (Bison bonasus) y la significación de la heterocigosidad electroforética para la conservación. Conservation Biology, 8(1):167–174, March 1994. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Heschel:1995:IDE

[HP95] M. Shane Heschel and Ken N. Paige. Inbreeding depression, environmental stress, and population size variation in scarlet gilia (Ipomopsis aggregata). Depresión de endocria, stress ambiental y variación en el tamaño poblacional de Ipomopsis aggregata. Conservation Biology, 9(1):126–133, February 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See comments [OG96].

Hogbin:1999:ECG

[HP99] Patricia M. Hogbin and Rod Peakall. Evaluation of the contribution of genetic research to the management of the endangered plant Zieria prostrata. Evaluación de la contribución de la investigación genética en el manejo de la planta Zieria prostrata amenazada de extinción. Conservation Biology, 13(3):514–522, June 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hill:1997:IHL

[HPB^{+}97] Kim Hill, Jonathan Padwe, Carlos Bejyvagi, Ambrosio Bepurangi, Felipe Jakugi, Roberto Tykuarangi, and Tito Tykuarangi. Impact of hunting on large vertebrates in the mbaracayu reserve, Paraguay. Conservation Biology, 11(6): 1339–1353, December 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hall:1993:LBT

[HR93] Stephen J. G. Hall and John Ruane. Livestock breeds and their conservation: a global overview. Conservation Biology, 7(4): 815–825, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Higgins:1999:PLS

[HRCTS99] Steven I. Higgins, David M. Richardson, Richard M. Cowling, and Terry H. Trinder-Smith. Predicting the landscape-scale distribution of alien plants and their threat to plant diversity. Predicción de la distribución de plantas invasoras y su amenaza contra la diversidad de especies a nivel de paisaje. Conservation Biology, 13(2):303–313, April 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hawkins:1999:ERS

[HRV^{+}99] Julie P. Hawkins, Callum M. Roberts, Tom Van’t Hof, Kalli De Meyer, Jamie Tratalos, and Chloe Aldam. Effects of recreational scuba diving on Caribbean coral and fish communities. Efectos del buceo recreacional en comunidades de coral y peces del Caribe. Conservation Biology, 13(4):888–897, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Humphrey:1990:BAC

[HS90] Stephen R. Humphrey and Bradley M. Stith. A balanced approach to conservation. Conservation Biology, 4(4):341–343, December 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Houston:1995:ASN

[HS95] Douglas B. Houston and Edward G. Schreiner. Alien species in national parks: Drawing lines in space and time. Conservation Biology, 9(1):204–209, February 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hambler:1996:ERB

[HS96a] Clive Hambler and Martin R. Speight. Extinction rates in British nonmarine invertebrates since 1900. Conservation Biology, 10(3):892–896, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Heinen:1996:SPA

[HS96b] Joel T. Heinen and Sompoad Srikosamatara. Status and protection of Asian wild cattle and buffalo. Conservation Biology, 10(4):931–935, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hayes:1997:SPA

[HS97] John P. Hayes and Robert J. Steidl. Statistical power analysis and amphibian population trends. Conservation Biology, 11(1): 273–275, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hammond:1998:PFG

[HS98a] David S. Hammond and Hans Ter Steege. Propensity for fire in guianan rainforests. Conservation Biology, 12(5):944–947, October 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hazlett:1998:DAR

[HS98b] Donald L. Hazlett and Neil W. Sawyer. Distribution of alkaloidrich plant species in shortgrass steppe vegetation. Conservation Biology, 12(6):1260–1268, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harrison:1993:SMS

[HSD93] Susan Harrison, Andy Stahl, and Daniel Doak. Spatial models and spotted owls: Exploring some biological issues behind recent events. Conservation Biology, 7(4):950–953, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Holmquist:1998:HDM

[HSGY98] Jeff G. Holmquist, Jutta M. Schmidt-Gengenbach, and Beverly Buchanan Yoshioka. High dams and marine-freshwater linkages: Effects on native and introduced fauna in the Caribbean. Conservation Biology, 12(3):621–630, June 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hedgecock:1994:CBE

[HSS94] Dennis Hedgecock, Paul Siri, and Donald R. Strong. Conservation biology of endangered Pacific salmonids: Introductory remarks. Conservation Biology, 8(3):863–864, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Huenneke:1995:PIB

[HT95] Laura Foster Huenneke and James K. Thomson. Potential interference between a threatened endemic thistle and an invasive nonnative plant. Interferencia potencial entre un cardo endémico en peligro y una planta invasora no nativa. Conservation Biology, 9(2):416–425, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hughes:1991:MPD

[Hug91] Austin L. Hughes. MHC polymorphism and the design of captive breeding programs. Conservation Biology, 5(2):249–251, June 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See comments [VL91].

Hunter:1993:BGP

[Hun93] Malcolm I. Hunter, Jr. Biodiversity goes public. Conservation Biology, 7(1):208–209, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hunter:1994:CRT

[Hun94] J. Robert Hunter. Is Costa Rica truly conservation-minded? Conservation Biology, 8(2):592–595, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hunter:1996:BME

[Hun96a] Malcolm Hunter, Jr. Benchmarks for managing ecosystems: Are human activities natural? Conservation Biology, 10(3): 695–697, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hunter:1996:DRC

[Hun96b] Malcolm L. Hunter, Jr. A desktop reference for conservationists. Conservation Biology, 10(5):1477–1478, October 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hunt:1998:ELP

[Hun98] P. D. Hunt. Evidence from a landscape population model of the importance of early successional habitat to the American redstart. Conservation Biology, 12(6):1377–1389, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hurlbert:1994:TAH

[Hur94] Kathryn J. Hurlbert. A tribute to alwyn h. Gentry. Conservation Biology, 8(1):291–292, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hutchins:1995:OMG

[Hut95a] Michael Hutchins. Olympic mountain goat controversy continues. Conservation Biology, 9(5):1324–1326, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hutchins:1995:SRC

[Hut95b] Michael Hutchins. Sumatran rhino conservation. Conservation Biology, 9(5):977, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hutto:1995:CBC

[Hut95c] Richard L. Hutto. Composition of bird communities following stand-replacement fires in Northern Rocky Mountain (U.S.A.) conifer forests. Conservation Biology, 9(5):1041–1058, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hutchinson:1997:IEA

[HV97] Todd F. Hutchinson and John L. Vankat. Invasibility and effects of Amur honeysuckle in Southwestern Ohio forests. Invasibilidad y efectos de la madreselva de Amur en bosques del sudoeste de Ohio. Conservation Biology, 11(5):1117–1124, October 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hickey:1991:GVR

[HVG91] R. James Hickey, Michael A. Vincent, and Sheldon I. Guttman. Genetic variation in running buffalo clover (Trifoflium stofoniferum, Fabaceae). Conservation Biology, 5(3):309–316, September 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hagan:1996:EDF

[HVM96] John M. Hagan, W. Matthew Vander Haegen, and Peter S. McKinley. The early development of forest fragmentation effects on birds. Los efectos del desarrollo temprano de la fragmentación del bosque sobre las aves. Conservation Biology, 10 (1):188–202, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hyman:1995:VIF

[HW95] Jeffrey B. Hyman and Kris Wernstedt. A value-informed framework for interdisciplinary analysis: Application to recovery planning for Snake River salmon. Un marco para el análisis interdisciplinario basado en el valor informativo: Su aplicación a la planificación para la recuperación del salmón del Rio “Snake”. Conservation Biology, 9(3):625–635, June 1995. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hall:1996:EFF

[HWB96] Pamela Hall, Sarah Walker, and Kamaljit Bawa. Effect of forest fragmentation on genetic diversity and mating system in a tropical tree, Pithecellobium elegans. Efectos de la fragmentación del bosque sobre la diversidad genética y el sistema de cruza en un arbol tropical. Conservation Biology, 10(3):757–768, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hooker:1999:MPA

[HWG99] Sascha K. Hooker, Hal Whitehead, and Shannon Gowans. Marine protected area design and the spatial and temporal distribution of cetaceans in a submarine canyon. Diseño de areas marinas protejidas y la distribución espacial y temporal de cetaceos en un cañón submarino. Conservation Biology, 13 (3):592–602, June 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hunter:1993:ADB

[HY93] Malcolm L. Hunter, Jr. and Pralad Yonzon. Altitudinal distributions of birds, mammals, people, forests, and parks in Nepal. Conservation Biology, 7(2):420–423, June 1993. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Heinen:1995:FIT

[HYL95] Joel T. Heinen, Pralad B. Yonzon, and Blair Leisure. Fighting the illegal fur trade in Kathmandu, Nepal. Conservation Biology, 9(2):246–254, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Irish:1996:SEM

[IN96] Kerry E. Irish and Elliott A. Norse. Scant emphasis on marine biodiversity. Conservation Biology, 10(2):680–681, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ingram:1990:MGP

[Ing90] Gordon Brent Ingram. Multi-gene-pool surveys in areas with rapid genetic erosion: an example from the Aïr Mountains, Northern Niger. Conservation Biology, 4(1):78–90, March 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ingham:1996:AFV

[IS96] Deidre S. Ingham and Michael J. Samways. Application of fragmentation and variegation models to epigaeic invertebrates in South Africa. La aplicación de los modelos de fragmentación y mezcla a los invertebrados epigeos de Africa del Sur. Conservation Biology, 10(5):1353–1358, October 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Inglis:1992:CSD

[IU92] G. Inglis and A. J. Underwood. Comments on some designs proposed for experiments on the biological importance of corridors. Conservation Biology, 6(4):581–586, December 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jacobson:1990:GEC

[Jac90] Susan K. Jacobson. Graduate education in conservation biology. Conservation Biology, 4(4):431–440, December 1990. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jacobson:1993:TCB

[Jac93] Mathew Jacobson. Taking conservation biology to the field. Conservation Biology, 7(1):8–9, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jain:1993:MLB

[Jai93] Subodh Jain. Managing landscapes for biodiversity. Conservation Biology, 7(4):966–967, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

James:1994:EGE

[Jam94] Paul C. James. On economic growth and ecological decay. Conservation Biology, 8(4):1161–1162, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Joslyn:1995:GA

[JB95] Leif Joslyn and Rich Baker. Grazing and advocacy. Conservation Biology, 9(3):475–476, June 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Johnson:1999:VCE

[JB99] Kris Harold Johnson and Clait E. Braun. Viability and conservation of an exploited sage grouse population. Conservation Biology, 13(1):77–84, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Johnson:1996:EOT

[JBB96] Steve A. Johnson, Karen A. Bjorndal, and Alan B. Bolten. Effects of organized turtle watches on loggerhead (Caretta caretta) nesting behavior and hatchling production in Florida. El efecto de observaciones organizadas sobre el comportamiento de anidación y la producción de crías de la tortuga cahuama en Florida. Conservation Biology, 10(2):570–577, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jacobson:1998:TIS

[JD98] Susan K. Jacobson and Mallory D. Duff. Training Idiot Savants: The lack of human dimensions in conservation biology. Conservation Biology, 12(2):263–267, April 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jerry:1998:HCM

[JDEB98] Dean R. Jerry, Tanya A. Dow, Martin S. Elphinstone, and Peter R. Baverstock. Historical and contemporary maternal population structuring in the endangered Hastings River mouse (Pseudomys oralis). Conservation Biology, 12(5):1017–1022, October 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jenkins:1995:HWA

[Jen95] Peter Jenkins. Hyman and Wernstedt’s analytical framework. Conservation Biology, 9(5):984, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jenkins:1996:FTE

[Jen96] Peter T. Jenkins. Free trade and exotic species introductions. Conservation Biology, 10(1):300–302, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jessup:1998:ROW

[Jes98] David A. Jessup. Rehabilitation of oiled wildlife. Conservation Biology, 12(5):1153–1155, October 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Johnson:1993:ECH

[JFH93] A. Sydney Johnson, William M. Ford, and Philip E. Hale. The effects of clearcutting on herbaceous understories are still not fully known. Conservation Biology, 7(2):433–435, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jaksic:1996:ERL

[JFJ96] Fabian M. Jaksic, Peter Feinsinger, and Jaime E. Jimenez. Ecological redundancy and long-term dynamics of vertebrate predators in semiarid Chile. Redundancia ecológica y dinámica al largo plazo de los vertebrados de predadores en las areas semidesérticas de Chile. Conservation Biology, 10(1):252–262, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jhala:1991:SCW

[JG91] Yadvendradev V. Jhala and Robert H. Giles, Jr. The status and conservation of the wolf in Gujarat and Rajasthan, India. Conservation Biology, 5(4):476–483, December 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jhala:1993:PBW

[Jha93] Yadvendradev V. Jhala. Predation on blackbuck by wolves in Velavadar National Park, Gujarat, India. Conservation Biology, 7(4):874–881, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jones:1999:ERF

[JHHB99] E. B. Dale Jones III, Gene S. Helfman, Joshua O. Harper, and Paul V. Bolstad. Effects of riparian forest removal on fish assemblages in Southern Appalachian streams. Efectos de la remoción del bosque ripario en ensamblajes de peces en arroyos de los Apalaches del Sur. Conservation Biology, 13(6):1454–1465, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jeffrey:1999:AVE

[JJ99] Christopher F. G. Jeffrey and Cecil A. Jennings. An alternative view on estimating subsistence consumption of coral reef fishes in the U.S. Virgin Islands. Conservation Biology, 13(4):939– 941, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jacobson:1997:PAK

[JM97] Susan K. Jacobson and Susan B. Marynowski. Public attitudes and knowledge about ecosystem management on department of defense land in Florida. Actidudes públicas y conocimiento del manejo de ecosistemas de las tierras del departamento de defensa en Florida. Conservation Biology, 11(3):770–781, June 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jobes:1991:GYS

[Job91] Patrick C. Jobes. The Greater Yellowstone social system. Conservation Biology, 5(3):387–394, September 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jelmert:1996:WDS

[JOB96] Anders Jelmert and Dag Oscar Oppen-Berntsen. Whaling and deep-sea biodiversity. Conservation Biology, 10(2):653–654, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Johns:1994:LIL

[Joh94a] David M. Johns. Living as if life mattered. Conservation Biology, 8(1):320–321, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Johnson:1994:UP

[Joh94b] Nels C. Johnson. Unfulfilled potential. Conservation Biology, 8 (4):1168–1169, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jope:1994:PSC

[Jop94] Katherine L. Jope. Paradigm of species conservation. Conservation Biology, 8(4):924–925, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jorge:1992:IBC

[Jor92] Soberón M. Jorge. Island biogeography and conservation practice. Conservation Biology, 6(2):161, June 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jordan:1993:HEH

[Jor93] Carl F. Jordan. Healthy ecosystems: How do we define them? Conservation Biology, 7(2):444–445, June 1993. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jules:1999:MRT

[JR99] Erik S. Jules and Beverly J. Rathcke. Mechanisms of reduced trillium recruitment along edges of old-growth forest fragments mecanismos de reclutamiento de Trillium en bordes de fragmentos de bosque primario. Conservation Biology, 13(4):784– 793, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jennings:1999:PVT

[JRP99] Simon Jennings, John D. Reynolds, and Nicholas V. C. Polunin. Predicting the vulnerability of tropical reef fishes to exploitation with phylogenies and life histories. Predicción de la vulnerabilidad a la explotación en peces de arrecifes tropicales: un enfoque basado en la filogenia e historias de vida. Conservation Biology, 13(6):1466–1475, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Johnson:1993:CRP

[JS93] Douglas H. Johnson and Michael D. Schwartz. The conservation reserve program and grassland birds. Conservation Biology, 7(4):934–937, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jarvie:1998:IKI

[JS98] James K. Jarvie and P. F. Stevens. Interactive keys, inventory, and conservation. Conservation Biology, 12(1):222–224, February 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jacobson:1995:NDC

[JVM95] Susan K. Jacobson, Eleta Vaughan, and Sarah Webb Miller. New directions in conservation biology: Graduate programs.

Conservation Biology, 9(1):5–17, February 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kindvall:1992:GFM

[KA92] Oskar Kindvall and Ingemar Ahlén. Geometrical factors and metapopulation dynamics of the bush cricket, Metrioptera bicolor Philippi (Orthoptera: Tettigoniidae). Factores geométricos y dinámica metapoblacional del grillo Metrioptera bicolor Philippi (Orthoptera: Tettigoniidae). Conservation Biology, 6(4):520–529, December 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kaiser:1998:SBF

[Kai98] Michel J. Kaiser. Significance of bottom-fishing disturbance. Conservation Biology, 12(6):1230–1235, December 1998. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kattan:1994:FFB

[KALG94] Gustavo H. Kattan, Humberto Alvarez-López, and Manuel Giraldo. Forest fragmentation and bird extinctions: San Antonio eighty years later. Fragmentación forestal y extinción de pájaros: San Antonio 80 años más tarde. Conservation Biology, 8(1):138–146, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Karr:1990:ASR

[Kar90a] James R. Karr. Avian survival rates and the extinction process on Barro Colorado Island, Panama. Conservation Biology, 4(4): 391–397, December 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Karr:1990:BIG

[Kar90b] James R. Karr. Biological integrity and the goal of environmental legislation: Lessons for conservation biology. Conservation Biology, 4(3):244–250, September 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Karr:1993:AR

[Kar93] James R. Karr. Advocacy and responsibility. Conservation Biology, 7(1):8, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kark:1999:CPC

[KASR99] Salit Kark, Philip U. Alkon, Uriel N. Safriel, and Ettore Randi. Conservation priorities for chukar partridge in Israel based on genetic diversity across an ecological gradient. Prioridades de consevación para Alectoris chukar en Israel, basadas en la diversidad genética a lo largo de un gradiente ecológico. Conservation Biology, 13(3):542–552, June 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kattan:1992:RVB

[Kat92] Gustavo H. Kattan. Rarity and vulnerability: The birds of the Cordillera Central of Colombia. Conservation Biology, 6(1):64– 70, March 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kaufman:1993:CWG

[Kau93a] Les Kaufman. The challenge of the World’s great lakes. Conservation Biology, 7(3):447–449, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kaus:1993:EPS

[Kau93b] Andrea Kaus. Environmental perceptions and social relations in the Mapimí Biosphere Reserve. Percepciones ambientales y relaciones sociales en la reserva de la Biósfera de Mapimí. Conservation Biology, 7(2):398–406, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kaufman:1995:BAM

[Kau95a] Les Kaufman. Building an agreement on marine conservation needs. Conservation Biology, 9(3):696–698, June 1995. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kaufman:1995:TOS

[Kau95b] Les Kaufman. “Touch one strand, and the whole web shivers.”. Conservation Biology, 9(2):470–471, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kay:1997:UTC

[Kay97] Charles E. Kay. The ultimate tragedy of commons. Conservation Biology, 11(6):1447–1448, December 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kirkpatrick:1994:CDE

[KB94] J. B. Kirkpatrick and M. J. Brown. A comparison of direct and environmental domain approaches to planning reservation of forest higher plant communities and species in Tasmania. Una comparación de propuestas directas y de dominio ambiental en la planificación de reservas de bosques de las comunidades y especies de plantas altas en Tasmania. Conservation Biology, 8 (1):217–224, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kiesecker:1997:IEL

[KB97] Joseph M. Kiesecker and Andrew R. Blaustein. Influences of egg laying behavior on pathogenic infection of amphibian eggs. Influencia de la conducta de ovoposición sobre infecciones patógenas en huevos de anfibio. Conservation Biology, 11(1): 214–220, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kiesecker:1998:EIB

[KB98] Joseph M. Kiesecker and Andrew R. Blaustein. Effects of introduced bullfrogs and smallmouth bass on microhabitat use, growth, and survival of native red-legged frogs (Rana aurora). Conservation Biology, 12(4):776–787, August 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Karl:1999:ESU

[KB99] Stephen A. Karl and Brian W. Bowen. Evolutionary significant units versus geopolitical taxonomy: Molecular systematics of an endangered sea turtle (genus Chelonia). Unidades evolutivas significativas contra taxonomía geopolítica: Sistemática molecular de la tortuga verde amenazada de extinción. Conservation Biology, 13(5):990–999, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kaufman:1998:CB

[KBM98] J. H. Kaufman, D. Brodbeck, and O. R. Melroy. Critical biodiversity. Conservation Biology, 12(3):521–532, June 1998. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kellert:1996:HCL

[KBRB96] Stephen R. Kellert, Matthew Black, Colleen Reid Rush, and Alistair J. Bath. Human culture and large carnivore conservation in North America. Cultura humana y conservación de carnívoros mayores en Norte América. Conservation Biology, 10(4):977–990, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kaufman:1993:GLA

[KC93] Les Kaufman and Andrew S. Cohen. The great lakes of Africa. Conservation Biology, 7(3):632–633, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kerr:1995:EHA

[KC95] Jeremy T. Kerr and David J. Currie. Effects of human activity on global extinction risk. Efectos de la actividad humana sobre el riesgo de extinción global. Conservation Biology, 9(6): 1528–1538, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kremen:1993:TAA

[KCE^{+}93] C. Kremen, R. K. Colwell, T. L. Erwin, D. D. Murphy, R. F. Noss, and M. A. Sanjayan. Terrestrial arthropod assemblages: Their use in conservation planning. Conservation Biology, 7(4): 796–808, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

King:1998:ERN

[KDG98] David I. King, Richard M. DeGraaf, and Curtice R. Griffin. Edge-related nest predation in clearcut and groupcut stands. Conservation Biology, 12(6):1412–1415, December 1998. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Klinger:1991:RMS

[KEE91] Terrie Klinger, Diane R. Elam, and Norman C. Ellstrand. Radish as a model system for the study of engineered gene escape rates via crop-weed mating. Conservation Biology, 5(4): 531–535, December 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Keith:1998:EMW

[Kei98] David A. Keith. An evaluation and modification of world conservation union red list criteria for classification of extinction risk in vascular plants. Conservation Biology, 12(5):1076–1090, October 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kellert:1991:JPW

[Kel91] Stephen R. Kellert. Japanese perceptions of wildlife. Conservation Biology, 5(3):297–308, September 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kellert:1993:VPI

[Kel93a] Stephen R. Kellert. Values and perceptions of invertebrates. Conservation Biology, 7(4):845–855, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Krausman:1993:PMS

[KEL93b] Paul R. Krausman, Richard C. Etchberger, and Raymond M. Lee. Persistence of mountain sheep. Conservation Biology, 7 (2):219, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kershaw:1996:MC

[Ker96] M. Kershaw. Maps for conservation. Conservation Biology, 10 (4):1301–1302, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kerr:1997:SRE

[Ker97] Jeremy T. Kerr. Species richness, endemism, and the choice of areas for conservation. Riqueza de especies, endemismo y selección de areas para conservación. Conservation Biology, 11 (5):1094–1100, October 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Koppelman:1995:GEV

[KF95] Jeffrey B. Koppelman and Dennis E. Figg. Genetic estimates of variability and relatedness for conservation of an Ozark cave crayfish species complex. Conservation Biology, 9(5): 1288–1294, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

King:1996:ECH

[KGD96] David I. King, Curtice R. Griffin, and Richard M. Degraff. Effects of clearcutting on habitat use and reproductive success of the ovenbird in forested landscapes. Efectos de la tala sobre la utilización de habitat y el éxito reproductivo de horneros en tin paisaje boscoso. Conservation Biology, 10(5):1380–1386, October 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kretzmann:1997:LGV

[KGM^{+}97] Maria B. Kretzmann, William G. Gilmartin, Axel Meyer, Gerard P. Zegers, Steven R. Fain, Bruce F. Taylor, and Daniel P. Costa. Low genetic variability in the Hawaiian monk seal. Baja variabilidad genética en la foca Hawaiana. Conservation Biology, 11(2):482–490, April 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Keel:1993:UVA

[KGS93] Shirley Keel, Alwyn H. Gentry, and Lucio Spinzi. Using vegetation analysis to facilitate the selection of conservation sites in Eastern Paraguay. El uso del análisis de vegetación para facilitar la selección de sitios a conservar en el este de Paraguay. Conservation Biology, 7(1):66–75, March 1993. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kalinowski:1999:NID

[KHM99] Steven T. Kalinowski, Philip W. Hedrick, and Philip S. Miller. No inbreeding depression observed in Mexican and red wolf captive breeding programs. Depresión en el lobo rojo y el lobo mexicano no ocasionada por intracruza en los programas de reproducción en cautiverio. Conservation Biology, 13(6):1371–1377, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Klein:1995:EED

[KHP95] Mary L. Klein, Stephen R. Humphrey, and H. Franklin Percival. Effects of ecotourism on distribution of waterbirds in a wildlife refuge. El efecto del ecotourismo en la distribución de las aves acuáticas de un refugio de vida silvestre. Conservation Biology, 9(6):1454–1465, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Keyser:1998:EFF

[KHS98] Amber J. Keyser, Geoffrey E. Hill, and Eric C. Soehren. Effects of forest fragment size, nest density, and proximity to edge on the risk of predation to ground-nesting passerine birds. Conservation Biology, 12(5):986–994, October 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kinnaird:1992:CFP

[Kin92] Margaret F. Kinnaird. Competition for a forest palm: Use of Phoenix reclinata by human and nonhuman primates. Conservation Biology, 6(1):101–107, March 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kingston:1995:VMT

[Kin95] Tigga Kingston. Valuable modeling tool. Conservation Biology, 9(4):966–968, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kiosman:1991:EEC

[Kio91] Sharon Kiosman. Education and empowerment: Conservation lessons from children. Conservation Biology, 5(1):9–10, March 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kuusipalo:1994:MBF

[KK94] Jussi Kuusipalo and Jyrki Kangas. Managing biodiversity in a forestry environment manejando la biodiversidad en un ambiente de bosque. Conservation Biology, 8(2):450–460, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Keiter:1996:LLC

[KL96] Robert B. Keiter and Harvey Locke. Law and large carnivore conservation in the rocky mountains of the U.S. and Canada. Leyes y conservación de carnívoros en las Montañas Rocallosas de los Estados Unidos y Canadá. Conservation Biology, 10(4): 1003–1012, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kelly:1997:ERC

[KL97] P. E. Kelly and D. W. Larson. Effects of rock climbing on populations of presettlement eastern white cedar (Thuja occidentalis) on cliffs of the Niagara Escarpment, Canada. Efectos del alpinismo en poblaciones pre-colonización de cedro blanco del este (Thuja occidentalis) en los acantilados de la Escarpa del Niagara, Canada. Conservation Biology, 11(5):1125–1132, October 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kilgo:1998:IHB

[KLF98] John C. Kilgo, Ronald F. Labisky, and Duane E. Fritzen. Influences of hunting on the behavior of white-tailed deer: Implications for conservation of the Florida panther. Conservation Biology, 12(6):1359–1364, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kokko:1997:RAH

[KLR97] Hanna Kokko, Jan Lindström, and Esa Ranta. Risk analysis of hunting of seal populations in the Baltic. Análisis de riesgo de la caza de poblaciones de focas en el Báltico. Conservation Biology, 11(4):917–927, August 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Karanth:1997:APT

[KM97] K. Ullas Karanth and M. D. Madhusudan. Avoiding paper tigers and saving real tigers: Response to Saberwal. Conservation Biology, 11(3):818–820, June 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [Sab97].

Kleiman:1998:RMC

[KM98] Devra G. Kleiman and Jeremy J. C. Mallinson. Recovery and management committees for lion tamarins: Partnerships in conservation planning and implementation. Conservation Biology, 12(1):27–38, February 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kremen:1994:EMV

[KMM94] Claire Kremen, Adina M. Merenlender, and Dennis D. Murphy. Ecological monitoring: a vital need for integrated conservation and development programs in the tropics. Monitoreo ecológico: una necesidad vital para integrar programas de conservación y desarrollo en los trópicos. Conservation Biology, 8(2):388– 397, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kress:1994:GVP

[KMR94] W. John Kress, G. David Maddox, and Cheryl S. Roesel. Genetic variation and protection priorities in Ptilimnium nodosum (Apiaceae), an endangered plant of the Eastern United States. Conservation Biology, 8(1):271–276, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kershaw:1995:TSR

[KMW95] Melanie Kershaw, Georgina M. Mace, and P. H. Williams. Threatened status, rarity, and diversity as alternative selection measures for protected areas: a test using Afrotropical antelopes. El grado en que la fauna esta amenazada, la rareza y la diversidad como medidas de selección alternativas para áreas protegidas; un test usando antílopes Afrotropicales. Conservation Biology, 9(2):324–334, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Knight:1999:PLN

[Kni99] Richard L. Knight. Private lands: The neglected geography. Conservation Biology, 13(2):223–224, April 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kinnaird:1991:VPE

[KO91] Margaret F. Kinnaird and Timothy G. O’Brien. Viable populations for an endangered forest primate, the Tana River crested mangabey (Cercocebus galeritus galeritus). Conservation Biology, 5(2):203–213, June 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kaufman:1993:ECB

[KO93] Les Kaufman and Peter Ochumba. Evolutionary and conservation biology of cichlid fishes as revealed by faunal remnants in northern Lake Victoria. Evolución y conservación biológica de los peces cíclidos de acuerdo a remanentes faunísticos en el Lago Victoria. Conservation Biology, 7(3):719–730, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kinnaird:1998:EEW

[KO98] Margaret F. Kinnaird and Timothy G. O’Brien. Ecological effects of wildfire on lowland rainforest in Sumatra. Conservation Biology, 12(5):954–956, October 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kobori:1999:DRH

[Kob99] Hiromi Kobori. Diagnosing river health. Conservation Biology, 13(6):1523–1524, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Koenig:1998:SAC

[Koe98] Walter D. Koenig. Spatial autocorrelation in California land birds. Conservation Biology, 12(3):612–620, June 1998. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Keough:1993:CBH

[KQK93] Michael J. Keough, Gerald P. Quinn, and Alice King. Correlations between human collecting and intertidal mollusc populations on rocky shores. Correlaciones entre la recolección humana y las poblaciones de moluscos intertidales de costas rocosas. Conservation Biology, 7(2):378–390, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Knick:1995:LCF

[KR95] Steven T. Knick and John T. Rotenberry. Landscape characteristics of fragmented shrubsteppe habitats and breeding passerine birds. Conservation Biology, 9(5):1059–1071, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Knapp:1998:CIQ

[KR98] Eric E. Knapp and Kevin J. Rice. Comparison of isozymes and quantitative traits for evaluating patterns of genetic variation in purple needlegrass (Nassella pulchra). Conservation Biology, 12(5):1031–1041, October 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kraus:1995:CUV

[Kra95] Fred Kraus. The conservation of unisexual vertebrate populations. Conservation Biology, 9(4):956–959, August 1995. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kremen:1999:DMN

[KRG^{+}99] Claire Kremen, Vincent Razafimahatratra, R. Philip Guillery, Jocelyn Rakotomalala, Andrew Weiss, and Jean-Solo Ratsisompatrarivo. Designing the Masoala National Park in Madagascar based on biological and socioeconomic data. Diseño del Parque Nacional Masola en Madagascar en base datos biológicos y socioeconómicos. Conservation Biology, 13(5): 1055–1068, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kremen:1998:ITM

[KRL98] Claire Kremen, Isaia Raymond, and Kate Lance. An interdisciplinary tool for monitoring conservation impacts in Madagascar. Conservation Biology, 12(3):549–563, June 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Knopf:1994:BDS

[KS94a] Fritz L. Knopf and Fred B. Samson. Biological diversity-science and action. Conservation Biology, 8(3):909–911, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Knopf:1994:SPA

[KS94b] Fritz L. Knopf and Fred B. Samson. Scale perspectives on avian diversity in western riparian ecosystems perspectivas de escala sobre la diversidad de aves en ecosistemas ribereños del oeste. Conservation Biology, 8(3):669–676, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kunin:1997:PRT

[KS97] William E. Kunin and Avi Shmida. Plant reproductive traits as a function of local, regional, and global abundance. Características reproductivas de plantas en función de la abundancia local, regional y global. Conservation Biology, 11(1):183–192, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kiester:1996:CPU

[KSC^{+}96] A. Ross Kiester, J. Michael Scott, Blair Csuti, Reed F. Noss, Bart Butterfield, Kevin Sahr, and Denis White. Conservation prioritization using GAP data. Priorización de la conservación utilizando datos GAP. Conservation Biology, 10(5):1332–1342, October 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Knutson:1999:ELC

[KSO^{+}99] Melinda G. Knutson, John R. Sauer, Douglas A. Olsen, Michael J. Mossman, Lisa M. Hemesath, and Michael J. Lannoo. Effects of landscape composition and wetland fragmentation on frog and toad abundance and species richness in Iowa and Wisconsin, U.S.A. Efectos de la composición del paisaje y la fragmentación de humedales en la abundancia de ranas y la riqueza de especies en Iowa y Wisconsin, U.S.A. Conservation Biology, 13(6):1437–1446, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kenney:1995:LTE

[KSSM95] John S. Kenney, James L. D. Smith, Anthony M. Starfield, and Charles W. Mcdougal. The long-term effects of tiger poaching on population viability. Conservation Biology, 9(5):1127–1133, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Koopowitz:1994:GSM

[KTA94] Harold Koopowitz, Alan D. Thornhill, and Mark Andersen. A general stochastic model for the prediction of biodiversity losses based on habitat conversion. Un modelo estocástico general para la predicción de las pérdidas de la biodiversidad basadas en la conversión del hábitat. Conservation Biology, 8(2):425– 438, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kuchling:1992:AET

[Kuc92] Gerald Kuchling. Assessing extinction threats and exploitation of species. Conservation Biology, 6(1):7, March 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kennedy:1996:ICH

[KW96] E. Dale Kennedy and Douglas W. White. Interference competition from house wrens as a factor in the decline of Bewick’s wrens. Conservation Biology, 10(1):281–284, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Knight:1995:RVS

[KWR95] Richard L. Knight, George N. Wallace, and William E. Riebsame. Ranching the view: Subdivisions versus agriculture. Conservation Biology, 9(2):459–461, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lesica:1992:SPP

[LA92] Peter Lesica and Fred W. Allendorf. Are small populations of plants worth preserving? Conservation Biology, 6(1):135– 139, March 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lesica:1995:WPP

[LA95] Peter Lesica and Fred W. Allendorf. When are peripheral populations valuable for conservation? `?Cuándo resultan las poblaciones periféricas valiosas para la conservación? Conservation Biology, 9(4):753–760, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lewis:1997:THW

[LA97] Dale M. Lewis and Peter Alpert. Trophy hunting and wildlife conservation in Zambia. Caza deportiva y conservación de la vida silvestre en Zambia. Conservation Biology, 11(1):59–68, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Leary:1993:CGB

[LAF93] Robb F. Leary, Fred W. Allendorf, and Stephen H. Forbes. Conservation genetics of bull trout in the Columbia and Klamath River drainages. Conservation Biology, 7(4):856–865, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lambeck:1997:FSM

[Lam97] Robert J. Lambeck. Focal species: a multi-species umbrella for nature conservation. Especies focales: una sombrilla multiespecífica para conservar la naturaleza. Conservation Biology, 11(4):849–856, August 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lande:1995:MCM

[Lan95] Russell Lande. Mutation and conservation. Mutación y conservación. Conservation Biology, 9(4):782–791, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Langholz:1996:EOS

[Lan96] Jeff Langholz. Economics, objectives, and success of private nature reserves in Sub-Saharan Africa and Latin America. La economía, los objetivos y el éxito de las reservas naturales en Africa subsaharica y America Latina. Conservation Biology, 10(1):271–280, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Laurance:1991:ECE

[Lau91] William F. Laurance. Ecological correlates of extinction proneness in Australian tropical rain forest mammals. Conservation Biology, 5(1):79–89, March 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lapin:1995:ULE

[LB95] Marc Lapin and Burton V. Barnes. Using the landscape ecosystem approach to assess species and ecosystem diversity. Conservation Biology, 9(5):1148–1158, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lovette:1999:MDP

[LBR99] Irby J. Lovette, Eldredge Bermingham, and Robert E. Ricklefs. Mitochondrial DNA phylogeography and the conservation of endangered Lesser Antillean Icterus orioles. Filogeografía basada en ADN mitocondrial y la conservación de orioles antillanos del género Icterus. Conservation Biology, 13(5):1088– 1096, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lomolino:1998:RCR

[LC98a] Mark V. Lomolino and Rob Channell. Range collapse, reintroductions, and biogeographic guidelines for conservation. Conservation Biology, 12(2):481–484, April 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Luikart:1998:EET

[LC98b] Gordon Luikart and Jean-Marie Cornuet. Empirical evaluation of a test for identifying recently bottlenecked populations from allele frequency data. Conservation Biology, 12(1):228–237, February 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lesica:1999:SDS

[LC99] Peter Lesica and Stephen V. Cooper. Succession and disturbance in sandhills vegetation: Constructing models for managing biological diversity. Sucesión y perturbación en la vegetación de colinas de arena: Construcción de modelos para el manejo de la diversidad biológica. Conservation Biology, 13 (2):293–302, April 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Luikart:1999:TCA

[LCA99] Gordon Luikart, Jean-Marie Cornuet, and Fred W. Allendorf. Temporal changes in allele frequencies provide estimates of population bottleneck size. Cambios temporales en frecuencias de alelos proveen estimaciones del tamaño de un cuello de botella poblacional. Conservation Biology, 13(3):523–530, June 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lombard:1997:RSS

[LCPM97] A. T. Lombard, R. M. Cowling, R. L. Pressey, and P. J. Mustart. Reserve selection in a species-rich and fragmented landscape on the Agulhas Plain, South Africa. Selección de reservas en un paisaje fragmentado rico en especies de la Planicie Agulhas, Sudáfrica. Conservation Biology, 11(5):1101–1116, October 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lomolino:1995:ECE

[LCSC95] Mark V. Lomolino, J. Curtis Creighton, Gary D. Schnell, and David L. Certain. Ecology and conservation of the endangered American burying beetle (Nicrophorus americanus). Ecología y conservación del escarabajo Nicrophorus americanus en peligro de extinción. Conservation Biology, 9(3):605–614, June 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Liu:1995:PEF

[LDP95] Jianguo Liu, John B. Dunning, Jr., and H. Ronald Pulliam. Potential effects of a forest management plan on Bachman’s sparrows (Aimophila aestivalis): Linking a spatially explicit model with GIS. Impactos potenciales de un plan de manejo del bosque sobre los gorriones de bachman (Aimophila aestivalis): vinculando un modelo espacialmente explícito con SIG. Conservation Biology, 9(1):62–75, February 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lawler:1999:EIM

[LDSH99] Sharon P. Lawler, Deborah Dritz, Terry Strange, and Marcel Holyoak. Effects of introduced mosquitofish and bullfrogs on the threatened California red-legged frog. Effectos de la introducción del pez mosquito y rana toro en poblaciones de la rana. Conservation Biology, 13(3):613–622, June 1999. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Leberg:1991:IFB

[Leb91] Paul L. Leberg. Influence of fragmentation and bottlenecks on genetic divergence of wild Turkey populations. Conservation Biology, 5(4):522–530, December 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Leberg:1993:SPR

[Leb93] Paul L. Leberg. Strategies for population reintroduction: Effects of genetic variability on population growth and size. Conservation Biology, 7(1):194–199, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Leberg:1998:ICS

[Leb98] Paul Leberg. Influence of complex sex determination on demographic stochasticity and population viability. Conservation Biology, 12(2):456–459, April 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Light:1995:DSC

[LEMC95] Theo Light, Don C. Erman, Chris Myrick, and Jay Clarke. Decline of the Shasta crayfish (Pacifastacus fortis Faxon) of Northeastern California. Declinaciones del cangrejo de río de Shasta (Pacifastacus fortis Faxon) del Noreste de California. Conservation Biology, 9(6):1567–1577, December 1995. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Leopold:1990:SC

[Leo90] Aldo Leopold. Standards of conservation. Conservation Biology, 4(3):227–228, September 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lindenmayer:1997:MSS

[LF97] David B. Lindenmayer and Jerry F. Franklin. Managing stand structure as part of ecologically sustainable forest management in Australian mountain ash forests. Manejo de la estructura como parte del manejo ecológicamente sustentable de los bosques de fresno de montaña en Australia. Conservation Biology, 11(5):1053–1068, October 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Laurance:1998:EFF

[LFM^{+}98] William F. Laurance, Leandro V. Ferreira, Judy M. Rankin-De Merona, Susan G. Laurance, Roger W. Hutchings, and Thomas E. Lovejoy. Effects of forest fragmentation on recruitment patterns in Amazonian tree communities. Conservation Biology, 12(2):460–464, April 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Levin:1996:HER

[LFOJ96] Donald A. Levin, Javier Francisco-Ortega, and Robert K. Jansen. Hybridization and the extinction of rare plant species. Hibridación y extensión de especies raras the plantas. Conservation Biology, 10(1):10–16, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Laurance:1997:HCF

[LG97] William F. Laurance and Claude Gascon. How to creatively fragment a landscape. Conservation Biology, 11(2):577–579, April 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Linder:1995:SCP

[Lin95] H. P. Linder. Setting conservation priorities: The importance of endemism and phylogeny in the Southern African orchid genus Herschelia. Estableciendo las prioridades de conservación: la importancia del endemismo y la filogenia en el género de orquídeas del sur de Africa. Conservation Biology, 9(3):585–595, June 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lips:1998:DTM

[Lip98] Karen R. Lips. Decline of a tropical montane amphibian fauna. Conservation Biology, 12(1):106–117, February 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lips:1999:MMP

[Lip99] Karen R. Lips. Mass mortality and population declines of anurans at an upland site in Western Panama. Conservation Biology, 13(1):117–125, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Litvaitis:1993:RES

[Lit93] John A. Litvaitis. Response of early successional vertebrates to historic changes in land use. Conservation Biology, 7(4): 866–873, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lewis:1990:WCO

[LKM90] Dale Lewis, Gilson B. Kaweche, and Ackim Mwenya. Wildlife conservation outside protected areas — lessons from an experiment in Zambia. Conservation Biology, 4(2):171–180, June 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lawrence:1995:AEF

[LLP95] Deborah C. Lawrence, Mark Leighton, and David R. Peart. Availability and extraction of forest products in managed and primary forest around a Dayak village in West Kalimantan, Indonesia. Disponibilidad y extracción de productos forestales en bosques bajo manejo y bosques primarios en los alrededores de una aldea de Dayak, en el Kalimantan occidental, en Indonesia. Conservation Biology, 9(1):76–88, February 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lawrence:1992:NPO

[LM92] Nathaniel Lawrence and Dennis Murphy. New perspectives or old priorities? Conservation Biology, 6(3):465, September 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lowe-McConnell:1993:FFA

[LM93] Rosemary H. Lowe-McConnell. Fish faunas of the African Great Lakes: Origins, diversity, and vulnerability. Faunas de peces de los Grandes Lagos Africanos: Orígenes, diversidad y vulnerabilidad. Conservation Biology, 7(3):634–643, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Loutit:1994:ERD

[LM94] Blythe Loutit and Sharon Montgomery. The efficacy of rhino dehorning: Too early to tell! Conservation Biology, 8(4): 923–924, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lydeard:1995:DEA

[LM95] Charles Lydeard and Richard L. Mayden. A diverse and endangered aquatic ecosystem of the Southeast United States un ecosistema diverso y aquatico del sureste de los Estados Unidos en peligro. Conservation Biology, 9(4):800–805, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lamberson:1992:DAN

[LMNV92] Roland H. Lamberson, Robert McKelvey, Barry R. Noon, and Curtis Voss. A dynamic analysis of northern spotted owl viability in a fragmented forest landscape. Análisis dinámico de la viabilidad del búo moteado del norte en un paisaje de bosque fragmentado. Conservation Biology, 6(4):505–512, December 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Laurance:1996:EDC

[LMS96] William F. Laurance, Keith R. McDonald, and Richard Speare. Epidemic disease and the catastrophic decline of Australian rain forest frogs. Las enfermedades epidémicas y la declinación catastrófica de las ranas de la selva tropical Australiana. Conservation Biology, 10(2):406–413, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Laurance:1997:DED

[LMS97] William F. Laurance, Keith R. Mcdonald, and Richard Speare. In defense of the epidemic disease hypothesis. Conservation Biology, 11(4):1030–1034, August 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lord:1990:SSC

[LN90] Janice M. Lord and David A. Norton. Scale and the spatial concept of fragmentation. Conservation Biology, 4(2):197–202, June 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lindenmayer:1993:EPD

[LN93] David B. Lindenmayer and Henry A. Nix. Ecological principles for the design of wildlife corridors. Conservation Biology, 7(3): 627–631, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lele:1996:SSB

[LN96] Sharachchandra Lélé and Richard B. Norgaard. Sustainability and the Scientist’s burden. La sostenibilidad y la carga que soportan los científicos. Conservation Biology, 10(2):354–365, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lombard:1995:WSN

[LNA95] A. T. Lombard, A. O. Nicholls, and P. V. August. Where should nature reserves be located in South Africa? A snake’s perspective. `?Donde deberían estar localizadas las reservas naturales en Sudáfrica? El punto de vista de una serpiente. Conservation Biology, 9(2):363–372, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lyons:1995:IBI

[LNPC^{+}95] John Lyons, Sonia Navarro-Pérez, Philip A. Cochran, Eduardo C. Santana, and Manuel Guzmán-Arroyo. Index of biotic integrity based on fish assemblages for the conservation of streams and rivers in West-Central Mexico. Indice de integridad biótica basado en las características de las comunidades de peces para la conservación de arroyos y ríos en el centro-oeste de México. Conservation Biology, 9(3):569–584, June 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lamberson:1994:RDT

[LNVM94] Roland H. Lamberson, Barry R. Noon, Curtis Voss, and Kevin S. McKelvey. Reserve design for territorial species: The effects of patch size and spacing on the viability of the northern spotted owl. Diseño de reservas para especies territoriales: El efecto del tamaño de parche y del espacio en la viabilidad del búho moteado del norte. Conservation Biology, 8(1):185– 195, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lockwood:1999:UTP

[Loc99] Julie L. Lockwood. Using taxonomy to predict success among introduced avifauna: Relative importance of transport and establishment. Uso de la taxonomía para predecir el exito de avifauna introducida: Importancia relativa del transporte y establecimiento. Conservation Biology, 13(3):560–567, June 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Losos:1995:OSD

[Los95] Elizabeth C. Losos. Outstanding sourcebook for data and analyses. Conservation Biology, 9(3):695–696, June 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Liu:1999:FEE

[LOT^{+}99] Jianguo Liu, Zhiyun Ouyang, William W. Taylor, Richard Groop, Yingchun Tan, and Heming Zhang. A framework for evaluating the effects of human factors on wildlife habitat: the case of giant pandas. Marco para la evaluación de los efectos de factores humanos en el hábitat de vida silvestre: Caso de estudio de pandas gigantes. Conservation Biology, 13(6):1360–1370, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lovejoy:1990:NSa

[Lov90a] Thomas E. Lovejoy. News of the Society. Conservation Biology, 4(1):4, March 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lovejoy:1990:NSb

[Lov90b] Thomas E. Lovejoy. News of the Society. Conservation Biology, 4(2):116, June 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lovejoy:1990:NSc

[Lov90c] Thomas E. Lovejoy. News of the Society. Conservation Biology, 4(3):217, September 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lovejoy:1990:NSd

[Lov90d] Thomas E. Lovejoy. News of the Society. Conservation Biology, 4(4):344, December 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lovejoy:1991:NS

[Lov91] Thomas E. Lovejoy. News of the Society. Conservation Biology, 5(1):4, March 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lindenmayer:1996:RCT

[LP96] David B. Lindenmayer and Hugh P. Possingham. Ranking conservation and timber management options for Leadbeater’s possum in southeastern Australia using population viability analysis. Categorizacíon de las opciones para la conservación y manejo maderero de Gymnobelideus leadbeateri en el sureste de Australia, utilizando análisis de viabilidad poblacional. Conservation Biology, 10(1):235–251, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lizana:1998:EUB

[LP98] Miguel Lizana and Emilio M. Pedraza. The effects of UV-B radiation on toad mortality in mountainous areas of Central Spain. Conservation Biology, 12(3):703–707, June 1998. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Laikre:1991:IDC

[LR91] Linda Laikre and Nils Ryman. Inbreeding depression in a captive wolf (Canis lupus) population. Conservation Biology, 5 (1):33–40, March 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Legge:1996:GCE

[LRD^{+}96] John T. Legge, Richard Roush, Rob Desalle, Alfred P. Vogler, and Bernie May. Genetic criteria for establishing evolutionarily significant units in Cryan’s buckmoth. Criterios genéticos para el establecimiento de unidades evolutivamente significativas para las polillas de los pantanos. Conservation Biology, 10 (1):85–98, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Laikre:1993:HBC

[LRT93] Linda Laikre, Nils Ryman, and Elizabeth A. Thompson. Hereditary blindness in a captive wolf (Canis lupus) population: Frequency reduction of a deleterious allele in relation to gene conservation. Ceguera hereditaria en una población cautiva de lobos (Canis lupus): Reducción en la frecuencia de un alelo deletereo en relación con la conservación de genes. Conservation Biology, 7(3):592–601, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ligon:1996:LUL

[LS96] J. David Ligon and Peter B. Stacey. Land use, lag times and the detection of demographic change: The case of the acorn woodpecker. Uso de suelo, rezago temporal y la detección de cambios demográficos: un ejemplo provisto por el pájaro carpintero bellotero. Conservation Biology, 10(3):840–846, June 1996. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lafferty:1999:ERM

[LSA99] Kevin D. Lafferty, Cam C. Swift, and Richard F. Ambrose. Extirpation and recolonization in a metapopulation of an endangered fish, the tidewater goby. Extirpación y recolonización de una metapoblación de un pez amenazado, el goby de mareas. Conservation Biology, 13(6):1447–1453, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Llewellyn:1996:DSS

[LSC^{+}96] Daniel W. Llewellyn, Gary P. Shaffer, Nancy Jo Craig, Lisa Creasman, David Pashley, Mark Swan, and Cindy Brown. A decision-support system for prioritizing restoration sites on the Mississippi River alluvial plain. Sistema de apoyo a la toma de decisiones para la priorización de sitios a restaurar en la planicie aluvial del Río Mississippi. Conservation Biology, 10 (5):1446–1455, October 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lammi:1999:GDP

[LSM99] Antti Lammi, Pirkko Siikamäki, and Kaisa Mustajärvi. Genetic diversity, population size, and fitness in central and peripheral populations of a rare plant Lychnis viscaria. Diversidad genética, tamaño poblacional y condición de poblaciones periféricas de la planta rara Lychnis viscaria. Conservation Biology, 13(5):1069–1078, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Long:1995:PUB

[LSS95] Leslie E. Long, Loralei S. Saylor, and Michael E. Soulé. A pH/ UV-B synergism in amphibians. Conservation Biology, 9(5): 1301–1303, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Litvaitis:1996:LET

[LSV^{+}96] J. A. Litvaitis, D. F. Smith, R. Villafuerte, J. Oehler, and J. Carlson. Landscape ecology today. Conservation Biology, 10(1):306–308, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Leslie:1992:BDD

[LT92] David M. Leslie, Jr. and Alexandru Tissescu. Beyond the Danube Delta: Forest conservation and research opportunities in Romania. Conservation Biology, 6(2):165–167, June 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lugo:1990:REO

[Lug90] Ariel E. Lugo. Removal of exotic organisms. Conservation Biology, 4(4):345, December 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lugo:1992:MES

[Lug92] Ariel E. Lugo. More on exotic species. Conservation Biology, 6 (1):6, March 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lugo:1994:PPF

[Lug94] Ariel E. Lugo. Preservation of primary forests in the Luquillo Mountains, Puerto Rico. Preservación de bosques primarios en las montañas de Luquillo, Puerto Rico. Conservation Biology, 8 (4):1122–1131, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lugo:1997:OGM

[Lug97] Ariel E. Lugo. Old-growth mangrove forests in the United States: Bosques maduros de manglares en los Estados Unidos. Conservation Biology, 11(1):11–20, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Loman:1991:BFM

[LV91] Jon Loman and Torbjrn Von Schantz. Birds in a farmland — more species in small than in large habitat island. Conservation Biology, 5(2):176–188, June 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Leberg:1994:VAD

[LV94] Paul L. Leberg and Robert C. Vrijenhoek. Variation among desert topminnows in their susceptibility to attack by exotic parasites. Variación entre poecilidos del desierto en su susceptibilidad a ataques de parásitos exóticos. Conservation Biology, 8(2):419–424, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Litvaitis:1996:IPM

[LV96] John A. Litvaitis and Rafael Villafuerte. Intraguild predation, mesopredator release, and prey stability. Conservation Biology, 10(2):676–677, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Laurenson:1995:EFJ

[LWC95] M. Karen Laurenson, Nadja Wielebnowski, and T. M. Caro. Extrinsic factors and juvenile mortality in cheetahs. Conservation Biology, 9(5):1329–1331, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Leader-Williams:1993:PEW

[LWMG93] N. Leader-Williams and E. J. Milner-Gulland. Policies for the enforcement of wildlife laws: The balance between detection and penalties in Luangwa Valley, Zambia. Políticas para el cumplimiento de leyes sobre vida silvestre: Balance entre detección y penalidades en el valle de Luangwa, Zambia. Conservation Biology, 7(3):611–617, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lydeard:1996:UBS

[Lyd96] Charles Lydeard. U.S. biodiversity status report. Conservation Biology, 10(5):1480–1482, October 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lyman:1994:OMG

[Lym94] R. Lee Lyman. The Olympic mountain goat controversy: a different perspective. Conservation Biology, 8(3):898–901, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marston:1991:WVG

[MA91] Richard A. Marston and Jay E. Anderson. Watersheds and vegetation of the Greater Yellowstone Ecosystem. Conservation Biology, 5(3):338–346, September 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mills:1996:OMP

[MA96] L. Scott Mills and Fred W. Allendorf. The one-migrant-per-generation rule in conservation and management. La regla de un-migrante-por-generación en conservación y manejo. Conservation Biology, 10(6):1509–1518, December 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mikusinski:1998:EGF

[MA98] Grzegorz Mikusiński and Per Angelstam. Economic geography, forest distribution, and woodpecker diversity in Central Europe. Conservation Biology, 12(1):200–208, February 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

MacFarland:1991:GAV

[Mac91] Craig MacFarland. Goats on Acedo Volcano in the Galapagos Islands: Help needed. Conservation Biology, 5(1):6–7, March 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Maddock:1999:WDD

[Mad99] Ant Maddock. Wild dog demography in Hluhluwe-Umfolozi Park, South Africa. Demografía del perro silvestre en el Parque Hluhluwe-Umfolozi, Sudáfrica. Conservation Biology, 13 (2):412–417, April 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Maehr:1990:FPP

[Mae90] David S. Maehr. The Florida panther and private lands. Conservation Biology, 4(2):167–170, June 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Maguire:1991:RAC

[Mag91] Lynn A. Maguire. Risk analysis for conservation biologists. Conservation Biology, 5(1):123–125, March 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Maguire:1992:CCB

[Mag92a] Lynn A. Maguire. Conserving cultural and biological diversity. Conservation Biology, 6(2):303–304, June 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Maguire:1992:LLB

[Mag92b] Lynn A. Maguire. A laundry list for biodiversity research. Conservation Biology, 6(2):304–305, June 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Maguire:1996:MRV

[Mag96] Lynn A. Maguire. Making the role of values in conservation explicit: Values and conservation biology. Conservation Biology, 10(3):914–916, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mangel:1994:LHV

[Man94] Marc Mangel. Life history variation and salmonid conservation. Conservation Biology, 8(3):879–880, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marcot:1992:CIF

[Mar92] Bruce G. Marcot. Conservation of Indian forests. Conservation Biology, 6(1):12–16, March 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Markandya:1993:OED

[Mar93] A. Markandya. Options for economic development and prudent use of natural resources. Conservation Biology, 7(1):212–213, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moehlman:1996:GDT

[MAR96] Patricia D. Moehlman, George Amato, and Victor Runyoro. Genetic and demographic threats to the black rhinoceros population in the Ngorongoro Crater. Amenazas genéticas y demográificas de la población de rinocerontes negros en el Cráter Ngorongoro. Conservation Biology, 10(4):1107–1114, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marsden:1998:CBA

[Mar98] Stuart J. Marsden. Changes in bird abundance following selective logging on seram, Indonesia. Conservation Biology, 12 (3):605–611, June 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Master:1991:ATS

[Mas91] Lawrence L. Master. Assessing threats and setting priorities for conservation. Conservation Biology, 5(4):559–563, December 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Morales:1997:MDV

[MAS^{+}97] Juan Carlos Morales, Patrick Mahedi Andau, Jatna Supriatna, Zainal-Zahari Zainuddin, and Don J. Melnick. Mitochondrial DNA variability and conservation genetics of the Sumatran rhinoceros. Conservation Biology, 11(2):539–543, April 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Matlack:1994:PDL

[Mat94] Glenn Matlack. Plant demography, land-use history, and the commercial use of forests. Conservation Biology, 8(1):298–299, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mattei:1995:RES

[Mat95] Jennifer H. Mattei. Restoration ecology is a science. Conservation Biology, 9(5):1340–1341, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Matsuda:1997:CBV

[Mat97a] Brent M. Matsuda. Conservation biology, values, and advocacy. Conservation Biology, 11(6):1449–1450, December 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mattingly:1997:SBH

[Mat97b] Hayden T. Mattingly. Seeking balance in higher education. Conservation Biology, 11(5):1049–1052, October 1997. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Matthies:1999:HCS

[Mat99] Diethart Matthies. Highlighting conservation successes. Conservation Biology, 13(4):945–946, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mauchamp:1997:TAP

[Mau97] A. Mauchamp. Threats from alien plant species in the Galápagos Islands. Conservation Biology, 11(1):260–263, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Max:1990:BW

[Max90] Peter Max. The better world. Conservation Biology, 4(1):3, March 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McDonald:1992:UMM

[MB92a] Kelly A. McDonald and James H. Brown. Using montane mammals to model extinctions due to global change. Conservation Biology, 6(3):409–415, September 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mcintyre:1992:HVA

[MB92b] S. Mcintyre and G. W. Barrett. Habitat variegation, an alternative to fragmentation. Conservation Biology, 6(1):146–147, March 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mendelsohn:1995:PPR

[MB95] Robert Mendelsohn and Michael Balick. Private property and rainforest conservation. Conservation Biology, 9(5):1322–1323, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Majer:1996:BII

[MB96] J. D. Majer and G. Beeston. The biodiversity integrity index: an illustration using ants in Western Australia. El indice de la biodiversidad integral: una ilustración utilizando hormigas en Australia occidental. Conservation Biology, 10(1):65–73, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McCauley:1999:CID

[MB99] Shannon J. McCauley and Karen A. Bjorndal. Conservation implications of dietary dilution from debris ingestion: Sublethal effects in post-hatchling loggerhead sea turtles. Implicaciones para la conservación, dilución de dietas por ingestión de basura: Efectos subletales en crías de la tortuga marina Caretta caretta. Conservation Biology, 13(4):925–929, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Minnich:1995:SYC

[MBBF95] Richard A. Minnich, Michael G. Barbour, Jack H. Burk, and Robert F. Fernau. Sixty years of change in Californian conifer forests of the San Bernardino Mountains. Sesenta años de cambios en los bosques de conífera californianos de las montañas de San Bernardino. Conservation Biology, 9(4):902–914, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mcivor:1995:TCS

[MBD95] Donald E. Mcivor, John A. Bissonette, and Gary S. Drew. Taxonomic and conservation status of the Yuma mountain lion. Conservation Biology, 9(5):1033–1040, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Milton:1999:PPC

[MBD^{+}99] S. J. Milton, W. J. Bond, M. A. Du Plessis, D. Gibbs, C. Hilton-Taylor, H. P. Linder, L. Raitt, J. Wood, and J. S. Donaldson. A protocol for plant conservation by translocation in threatened lowland fynbos. Protocolo para la conservación de plantas por translocación: Fynbos amenazados de tierras bajas. Conservation Biology, 13(4):735–743, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McCue:1996:HVG

[MBH96] Kimberlie A. McCue, Edward S. Buckler, and Timothy P. Holtsford. A hierarchical view of genetic structure in the rare annual plant Clarkia springvillensis. Una visión jerárquica de la estructura genética de la planta anual rara Clarkia springvillensis. Conservation Biology, 10(5):1425–1434, October 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Maddockret:1997:ACA

[MBJ^{+}97] Ant Maddockret, Grant A. Benn, Dave N. Johnson, C. Rob Scott-Shaw, Truman Young, Alexander H. Harcourt, Michael Clinchy, and Charles J. Krebs. An African conservation Agency’s perspective on advocacy. Conservation Biology, 11 (4):831–833, August 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McIntyre:1992:STS

[MBKR92] S. McIntyre, G. W. Barrett, R. L. Kitching, and H. F. Recher. Species triage — seeing beyond wounded rhinos. Conservation Biology, 6(4):604–606, December 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meffe:1998:ISR

[MBM^{+}98] Gary K. Meffe, P. Dee Boersma, Dennis D. Murphy, Barry R. Noon, H. Ronald Pulliam, Michael E. Soulé, and Donald M. Waller. Independent scientific review in natural resource management. Conservation Biology, 12(2):268–270, April 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mcclenaghan:1990:FLG

[MBT90] Leroy R. Mcclenaghan, Jr., Joel Berger, and Harold D. Truesdale. Founding lineages and genic variability in plains bison (Bison bison) from Badlands National Park, South Dakota. Conservation Biology, 4(3):285–289, September 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Modde:1996:PSR

[MBW96] Timothy Modde, Kenneth P. Burnham, and Edmund J. Wick. Population status of the razorback sucker in the Middle Green River (U.S.A.). El estado poblacional de Xyrauchen texanus en el río “Middle Green” (EEUU). Conservation Biology, 10(1): 110–119, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meesters:1998:FLM

[MBW^{+}98] Erik H. Meesters, Rolf P. M. Bak, Susie Westmacott, Mark Ridgley, and Steve Dollar. A fuzzy logic model to predict coral reef development under nutrient and sediment stress. Conservation Biology, 12(5):957–965, October 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Maehr:1995:DGI

[MC95a] David S. Maehr and Gerard B. Caddick. Demographics and genetic introgression in the Florida panther. Conservation Biology, 9(5):1295–1298, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Maehr:1995:LFP

[MC95b] David S. Maehr and James A. Cox. Landscape features and panthers in Florida. Conservation Biology, 9(5):1008–1019, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mills:1998:RSI

[MC98] Claudia E. Mills and James T. Carlton. Rationale for a system of international reserves for the open ocean. Conservation Biology, 12(1):244–247, February 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McCoy:1995:CI

[McC95] Earl D. McCoy. The costs of ignorance. Conservation Biology, 9(3):473–474, June 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McCoy:1996:APC

[McC96] Earl D. McCoy. Advocacy as part of conservation biology. Conservation Biology, 10(3):919–920, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McKinney:1999:HRE

[McK99] Michael L. McKinney. High rates of extinction and threat in poorly studied taxa. Evidencias de altas tasas de extinción y amenaza en taxones pobremente estudiados. Conservation Biology, 13(6):1273–1281, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McLarney:1993:IPC

[McL93] William O. McLarney. Indigenous peoples and conservation. Conservation Biology, 7(4):748–749, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McLarney:1999:L

[McL99] William O. McLarney. Letter. Conservation Biology, 13(1): 4, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McNaughton:1995:ILA

[McN95] S. J. McNaughton. Inside look at African fieldwork. Conservation Biology, 9(6):1658–1659, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Merrill:1998:RPT

[MCO98] Samuel B. Merrill, Francesca J. Cuthbert, and Gary Oehlert. Residual patches and their contribution to forest-bird diversity on northern Minnesota aspen clearcuts. Conservation Biology, 12(1):190–199, February 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Miller:1994:PDB

[MCR94] Brian Miller, Gerardo Ceballos, and Richard Reading. The prairie dog and biotic diversity. El perro de las praderas y la diversidad biótica. Conservation Biology, 8(3):677–681, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mawdsley:1998:PPP

[MCW98] Nick A. Mawdsley, Stephen G. Compton, and Robert J. Whittaker. Population persistence, pollination mutualisms, and figs in fragmented tropical landscapes. Conservation Biology, 12 (6):1416–1420, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Murphy:1996:CBM

[MD96] Dennis D. Murphy and David A. Duffus. Conservation biology and marine biodiversity. Conservation Biology, 10(2):311–312, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mills:1999:RCA

[MDW99] L. Scott Mills, Daniel F. Doak, and Michael J. Wisdom. Reliability of conservation actions based on elasticity analysis of matrix models. Confiabilidad de las acciones de conservación basadas en análisis de elasticidad de modelos de matríces.

Conservation Biology, 13(4):815–829, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mech:1995:COR

[Mec95] L. David Mech. The challenge and opportunity of recovering wolf populations. El desafio y la oportunidad de las poblaciones de lobos en recuperación. Conservation Biology, 9(2):270–278, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Medley:1993:PCA

[Med93] Kimberly E. Medley. Primate conservation along the Tana River, Kenya: an examination of the forest habitat. Conservacion de primates a lo largo del rio Tana, Kenya: exámen del hábitat selvático. Conservation Biology, 7(1):109–121, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Medellin:1994:MDC

[Med94] Rodrigo A. Medellín. Mammal diversity and conservation in the Selva Lacandona, Chiapas, Mexico. Diversidad de mamíferos y conservación en la Selva Locandona, Chiapas, México. Conservation Biology, 8(3):780–799, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Medellin:1998:TIC

[Med98] Rodrigo A. Medellín. True international collaboration: Now or never. Conservation Biology, 12(5):939–940, October 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meffe:1993:HPC

[MEE93] Gary K. Meffe, Anne H. Ehrlich, and David Ehrenfeld. Human population control: The missing agenda. Conservation Biology, 7(1):1–3, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meffe:1994:RSA

[MEE94] Gary K. Meffe, Anne H. Ehrlich, and David Ehrenfeld. Response to Sieving et al. Conservation Biology, 8(1):308–309, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meffe:1992:TAH

[Mef92] Gary K. Meffe. Techno-arrogance and halfway technologies: Salmon hatcheries on the Pacific Coast of North America. Conservation Biology, 6(3):350–354, September 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meffe:1994:HPC

[Mef94] Gary K. Meffe. Human population control: The missing awareness. Conservation Biology, 8(1):310–313, March 1994. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meffe:1998:CBM

[Mef98a] Gary K. Meffe. Conservation Biology: Into the Millennium. Conservation Biology, 12(1):1–3, February 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meffe:1998:CSP

[Mef98b] Gary K. Meffe. Conservation scientists and the policy process. Conservation Biology, 12(4):741–742, August 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meffe:1998:SB

[Mef98c] Gary K. Meffe. Softening the boundaries. Conservation Biology, 12(2):259–260, April 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meffe:1999:CM

[Mef99a] Gary K. Meffe. Conservation medicine. Conservation Biology, 13(5):953–954, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meffe:1999:CSP

[Mef99b] Gary K. Meffe. Conservation science and public policy: Only the beginning. Conservation Biology, 13(3):463–464, June 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mehlman:1994:RNA

[Meh94] David W. Mehlman. Rarity in North American passerine birds. Conservation Biology, 8(4):1141–1145, December 1994. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meine:1999:ATC

[Mei99] Curt Meine. It’s about time: Conservation biology and history. Conservation Biology, 13(1):1–3, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Menges:1990:PVA

[Men90] Eric S. Menges. Population viability analysis for an endangered plant. Conservation Biology, 4(1):52–62, March 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Menges:1991:SGP

[Men91] Eric S. Menges. Seed germination percentage increases with population size in a fragmented prairie species. Conservation Biology, 5(2):158–164, June 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Merola:1994:RHC

[Mer94] Michele Merola. A reassessment of homozygosity and the case for inbreeding depression in the cheetah, Acinonyx jubatus: Implications for conservation. Una reevaluación de la homosigocidad y un argumento para la depresión de endocría en el cheetah Acinonyx jubatus: Implicaciones para la conservación. Conservation Biology, 8(4):961–971, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Maschinski:1997:DPV

[MFR97] Joyce Maschinski, Robert Frye, and Sue Rutman. Demography and population viability of an endangered plant species before and after protection from trampling. Demografiá y viabilidad poblacional de un especie de planta en peligro de extinción antes y después de protegerla contra pisoteo. Conservation Biology, 11(4):990–999, August 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Murphy:1990:EMA

[MFW90] Dennis D. Murphy, Kathy E. Freas, and Stuart B. Weiss. An environment-metapopulation approach to population viability analysis for a threatened invertebrate. Conservation Biology, 4(1):41–51, March 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Manire:1990:MSM

[MG90] Charles A. Manire and Samuel H. Gruber. Many sharks may be headed toward extinction. Conservation Biology, 4(1):10– 11, March 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mills:1997:FAD

[MG97] Michael G. L. Mills and Martyn L. Gorman. Factors affecting the density and distribution of wild dogs in the Kruger National Park. Conservation Biology, 11(6):1397–1406, December 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Murawski:1994:ESL

[MGB94] Darlyne A. Murawski, I. A. U. Nimal Gunatilleke, and Kamaljit S. Bawa. The effects of selective logging on inbreeding in Shorea megistophylla (Dipterocarpaceae) from Sri Lanka. El efecto de la tala sclectiva sobre la depresión de endogamia de Shorea megistophylla (Dipterocarpacea) de Sri Lanka. Conservation Biology, 8(4):997–1002, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Munger:1998:UNW

[MGM^{+}98] James C. Munger, Mark Gerber, Katy Madrid, Martha-Ann Carroll, Wade Petersen, and Lisa Heberger. U.S. National Wetland Inventory classifications as predictors of the occurrence of Columbia spotted frogs (Rana luteiventris) and Pacific treefrogs (Hyla regilla). Conservation Biology, 12(2):320–330, April 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Miller:1991:MPD

[MH91] Philip S. Miller and Philip W. Hedrick. MHC polymorphism and the design of captive breeding programs: Simple solutions are not the answer. Conservation Biology, 5(4):556–558, December 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Maurer:1993:GRF

[MH93] Brian A. Maurer and S. Greg Heywood. Geographic range fragmentation and abundance in neotropical migratory birds.

Rango geográfico, fragmentación y abundancia en pájaros migratorios neotropicales. Conservation Biology, 7(3):501–509, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Morris:1997:BBB

[MH97] Douglas W. Morris and Lawrence Heidinga. Balancing the books on biodiversity. Conservation Biology, 11(1):287–289, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mace:1999:ATS

[MH99a] Georgina M. Mace and Elodie J. Hudson. Attitudes toward sustainability and extinction. Actitudes hacia la sustentabilidad y la extinción. Conservation Biology, 13(2):242–246, April 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McIntyre:1999:FCH

[MH99b] S. McIntyre and Richard Hobbs. A framework for conceptualizing human effects on landscapes and its relevance to management and research models. Marco para la conceptualización de los efectos humanos en paisajes y su relevancia en modelos de manejo y investigación. Conservation Biology, 13(6): 1282–1292, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mills:1996:FLD

[MHB^{+}96] L. Scott Mills, Stephen G. Hayes, Calib Baldwin, Michael J. Wisdom, John Citta, David J. Mattson, and Kerry Murphy. Factors leading to different viability predictions for a grizzly bear data set factores que conducen a diferentes predicciones de viabilidad para un conjunto de datos del oso pardo. Conservation Biology, 10(3):863–873, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marshall:1998:IEU

[MHH98] Elizabeth Marshall, Robert Haight, and Frances R. Homans. Incorporating environmental uncertainty into species management decisions: Kirtland’s warbler habitat management as a case study. Conservation Biology, 12(5):975–985, October 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marsh:1997:SID

[MHL97] H. Marsh, A. N. M. Harris, and I. R. Lawler. The sustainability of the indigenous dugong fishery in Torres Strait, Australia/ Papua New Guinea. Conservation Biology, 11(6):1375–1386, December 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marmontel:1997:PVA

[MHO97] Miriam Marmontel, Stephen R. Humphrey, and Thomas J. O’Shea. Population viability analysis of the Florida manatee (Trichechus manatus latirostris), 1976–1991. Análisis de viabilidad poblacional del manatí de Florida (Trichechus manatus latirostris), 1976–1991. Conservation Biology, 11(2):467–481, April 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mattson:1996:SMR

[MHWP96] David J. Mattson, Stephen Herrero, R. Gerald Wright, and Craig M. Pease. Science and management of Rocky Mountain grizzly bears ciencia y manejo de los osos pardos de las Montañas rocallossas. Conservation Biology, 10(4):1013–1025, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Miller:1994:UCH

[Mil94] Monique A. Miller. The ultimate cause of habitat loss. Conservation Biology, 8(3):620, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Miller:1995:SBP

[Mil95a] Philip S. Miller. Selective breeding programs for rare alleles: Examples from the Przewalski’s horse and California condor pedigrees. Conservation Biology, 9(5):1262–1273, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mills:1995:EEI

[Mil95b] L. Scott Mills. Edge effects and isolation: Red-backed voles on forest remnants. El efecto de los bordes y el aislamiento: el raton campestre (Clethrionomys californicus) en remanentes de bosque. Conservation Biology, 9(2):395–403, April 1995. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mills:1996:CEG

[Mil96] L. Scott Mills. Cheetah extinction: Genetics or extrinsic factors? Conservation Biology, 10(2):315, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McLean:1996:FSR

[MJ96] Ian G. McLean and Susan K. Jacobson. Faculty sex ratios in U.S. conservation education programs. Conservation Biology, 10(1):4, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Maxwell:1997:SRC

[MJ97] Jane M. Maxwell and Ian G. Jamieson. Survival and recruitment of captive-reared and wild-reared takahe in Fiordland, New Zealand: Sobrevivencia y reclutamineto del takahe obtenido en cautiverio y en condiciones silvestres en fiordland, nueva zelanda. Conservation Biology, 11(3):683–691, June 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McClanahan:1996:FRC

[MKA96] T. R. McClanahan and B. Kaunda-Arara. Fishery recovery in a coral-reef marine park and its effect on the adjacent fishery. Recuperación de pesquerías en un parque marino de arrecife coralino. Conservation Biology, 10(4):1187–1199, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Martikainen:1998:TBW

[MKH98] Petri Martikainen, Lauri Kaila, and Yrjö Haila. Threatened beetles in white-backed woodpecker habitats. Conservation Biology, 12(2):293–301, April 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McClanahan:1996:ESU

[MKM^{+}96] T. R. McClanahan, A. T. Kamukuru, N. A. Muthiga, M. Gilagabher Yebio, and D. Obura. Effect of sea urchin reductions on algae, coral, and fish populations. Efectos de la reducción en la abundancia de erizos marinos sobre las poblaciones de algas, corales y peces. Conservation Biology, 10(1):136–154, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Maguire:1990:ASR

[ML90] Lynn A. Maguire and Robert C. Lacy. Allocating scarce resources for conservation of endangered subspecies: Partitioning zoo space for tigers. Conservation Biology, 4(2):157–166, June 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mace:1991:AET

[ML91] Georgina M. Mace and Russell Lande. Assessing extinction threats: Toward a reevaluation of IUCN threatened species categories. Conservation Biology, 5(2):148–157, June 1991. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McIntyre:1994:PRN

[ML94] S. McIntyre and S. Lavorel. Predicting richness of native, rare, and exotic plants in response to habitat and disturbance variables across a variegated landscape. Prediciendo riqueza de plantas nativas, raras y exóticas en respuesta al hábitat y a variables de perturbación a través de un paisaje diversificado. Conservation Biology, 8(2):521–531, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McCarthy:1997:EDR

[MLD97] Michael A. McCarthy, David B. Lindenmayer, and Martin Drechsler. Extinction debts and risks faced by abundant species. Deudas de extinción y riesgos enfrentados por un numero abundante de especies. Conservation Biology, 11(1): 221–226, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mulvey:1997:CGN

[MLP^{+}97] Margaret Mulvey, Charles Lydeard, Debra L. Pyer, Kimberly M. Hicks, Jayne Brim-Box, James D. Williams, and Robert S. Butler. Conservation genetics of North American freshwater mussels Amblema and Megalonaias. Conservación genética de las ostras americanas de agua dulce Amblema y Megalonaias. Conservation Biology, 11(4):868–878, August 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McCoy:1992:ROS

[MM92] Earl D. McCoy and Henry R. Mushinsky. Rarity of organisms in the sand pine scrub habitat of Florida. Rareza de los organismos en un hábitat arenoso de pinares enanos en la Florida. Conservation Biology, 6(4):537–548, December 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meagher:1994:OBB

[MM94] Mary Meagher and Margaret E. Meyer. On the origin of brucellosis in bison of Yellowstone National Park: a review. Sobre el origen de la brucelosis en el bisonte del Parque Nacional Yellowstone: una revisión. Conservation Biology, 8(3):645–653, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meine:1996:CVC

[MM96] Curt Meine and Gary K. Meffe. Conservation values, conservation science: a healthy tension. Conservation Biology, 10 (3):916–917, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mittermeier:1998:BHM

[MMT^{+}98] Russell A. Mittermeier, Norman Myers, Jorgen B. Thomsen, Gustavo A. B. Da Fonseca, and Silvio Olivieri. Biodiversity hotspots and major tropical wilderness areas: Approaches to setting conservation priorities. Conservation Biology, 12 (3):516–520, June 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Miller:1996:PLC

[MN96] Mark W. Miller and Thomas D. Nudds. Prairie landscape change and flooding in the Mississippi River valley. Cambio del paisaje de pradera e inundaciones en el valle del rio Mississippi. Conservation Biology, 10(3):847–853, June 1996. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Muchaal:1999:IVH

[MN99] Pia K. Muchaal and Germain Ngandjui. Impact of village hunting on wildlife populations in the Western Dja Reserve, Cameroon. Impacto de la cacería en poblaciones de vida silvestre de la Reserva Dja, Camerún. Conservation Biology, 13 (2):385–396, April 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Male:1998:L

[MNC^{+}98] Timothy D. Male, Reed F. Noss, Timothy M. Crowe, Jon Paul Rodríguez, and Gary K. Meffe. Letters. Conservation Biology, 12(4):743–745, August 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Margules:1994:AST

[MNU94] C. R. Margules, A. O. Nicholls, and M. B. Usher. Apparent species turnover, probability of extinction and the selection of nature reserves: a case study of the Ingleborough Limestone Pavements. Aparente reorganización de las especies, probabilidad de extinción y de selección de las reservas naturales: un estudio de caso de los pavimentos de piedra caliza de Ingleborough. Conservation Biology, 8(2):398–409, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Morimoto:1994:CP

[Mor94] David C. Morimoto. A chance, perhaps. Conservation Biology, 8(2):334, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Morimoto:1995:BEW

[Mor95] David Morimoto. On becoming Earth wise. Conservation Biology, 9(5):1336–1338, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Morimoto:1996:BA

[Mor96] David C. Morimoto. Boarding the ark. Conservation Biology, 10(4):1299–1300, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Morgan:1999:EPS

[Mor99] John W. Morgan. Effects of population size on seed production and germinability in an endangered, fragmented grassland plant. Efectos del tamaño poblacional en la producción y germinación de semillas en una especie de planta en peligro dentro de pastizales fragmentados. Conservation Biology, 13(2):266– 273, April 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moyle:1994:DAF

[Moy94] Peter B. Moyle. The decline of anadromous fishes in California. Conservation Biology, 8(3):869–870, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marsh:1997:EHF

[MP97] David M. Marsh and Peter B. Pearman. Effects of habitat fragmentation on the abundance of two species of leptodactylid frogs in an Andean montane forest. Conservation Biology, 11 (6):1323–1328, December 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Micheli:1999:EVH

[MP99] Fiorenza Micheli and Charles H. Peterson. Estuarine vegetated habitats as corridors for predator movements. Hábitats estuarinos vegetados como corredores para movimientos de depredadores. Conservation Biology, 13(4):869–881, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Means:1996:ESP

[MPB96] D. Bruce Means, John G. Palis, and Mary Baggett. Effects of slash pine silviculture on a Florida population of flatwoods salamander. El efecto de la silvicultura de clareo de pino sobre una población de salamandra Ambystoma cingulatum de los bosques bajos de Florida. Conservation Biology, 10(2):426– 437, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Melvin:1999:NTR

[MPC99] Edward F. Melvin, Julia K. Parrish, and Loveday L. Conquest. Novel tools to reduce seabird bycatch in coastal gillnet fisheries. Nuevas herramientas para reducir la captura accidental de aves marinas con redes agalleras de pesquerías costeras. Conservation Biology, 13(6):1386–1397, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Medail:1999:BHM

[MQ99] Frédéric Médail and Pierre Quézel. Biodiversity hotspots in the Mediterranean basin: Setting global conservation priorities. Conservation Biology, 13(6):1510–1513, December 1999.

CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mattson:1991:CYG

[MR91] David J. Mattson and Matthew M. Reid. Conservation of the Yellowstone grizzly bear. Conservation Biology, 5(3):364–372, September 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mishra:1998:LGB

[MR98a] Charudutt Mishra and G. S. Rawat. Livestock grazing and biodiversity conservation: Comments on Saberwal. Conservation Biology, 12(3):712–714, June 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [Sab97] and response [Sab98].

Moyle:1998:EBI

[MR98b] Peter B. Moyle and Paul J. Randall. Evaluating the biotic integrity of watersheds in the Sierra Nevada, California. Conservation Biology, 12(6):1318–1326, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Morse:1999:NSN

[MR99] Solon F. Morse and Scott K. Robinson. Nesting success of a neotropical migrant in a multiple-use, forested landscape. Éxito de nidación de un migrante neotropical en un paisaje forestal de uso múltiple. Conservation Biology, 13(2):327–337, April 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Main:1999:ECC

[MRN99] Martin B. Main, Fritz M. Roka, and Reed F. Noss. Evaluating costs of conservation. Evaluación de costos de conservación. Conservation Biology, 13(6):1262–1272, December 1999. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Maguire:1992:IBS

[MS92] Lynn A. Maguire and Christopher Servheen. Integrating biological and sociological concerns in endangered species management: Augmentation of grizzly bear populations. Conservation Biology, 6(3):426–434, September 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mladenoff:1993:EHR

[MS93] David J. Mladenoff and Forest Stearns. Eastern hemlock regeneration and deer browsing in the Northern Great Lakes region: a re-examination and model simulation. Conservation Biology, 7(4):889–900, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marvier:1997:CIH

[MS97a] Michelle A. Marvier and David L. Smith. Conservation implications of host use for rare parasitic plants. Implicaciones del uso de hospederos en la conservación de plantas parasíticas raras. Conservation Biology, 11(4):839–848, August 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moulton:1997:PFP

[MS97b] Michael P. Moulton and James G. Sanderson. Predicting the fates of passeriform introductions on oceanic islands. Conservation Biology, 11(2):552–558, April 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Martin:1999:WZG

[MS99a] Paul S. Martin and Christine R. Szuter. War zones and game sinks in Lewis and Clark’s West. Conservation Biology, 13(1): 36–45, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Medellin:1999:PMD

[MS99b] Rodrigo A. Medellín and Jorge Soberón. Predictions of mammal diversity on four land masses. Conservation Biology, 13(1): 143–149, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moulton:1999:FPI

[MS99c] Michael P. Moulton and James G. Sanderson. Fate of passeriform introductions: Reply to Duncan and Young. Conservation Biology, 13(4):937–938, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mortimer:1996:ACV

[MSC96] Russell Mortimer, Basil Sharp, and John Craig. Assessing the conservation value of New Zealand’s offshore islands. Evaluación del valor de la conservación de las islas costeras de Nueva Zelandia. Conservation Biology, 10(1):25–29, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mladenoff:1995:RLA

[MSHW95] David J. Mladenoff, Theodore A. Sickley, Robert G. Haight, and Adrian P. Wydeven. A regional landscape analysis and prediction of favorable gray wolf habitat in the Northern Great Lakes Region. Un análisis regional del paisaje y una prediccíon del hábitat favorable para el lobo gris en la región norte de los Grandes Lagos. Conservation Biology, 9(2):279–294, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McCoy:1999:RGC

[MSM99] Earl D. McCoy, Phillip E. Sutton, and Henry R. Mushinsky. The role of guesswork in conserving the threatened sand skink. Conservation Biology, 13(1):190–194, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Miththapala:1996:PSR

[MSO96] Sriyanie Miththapala, John Seidensticker, and Stephen J. O’Brien. Phylogeographic subspecies recognition in leopards (Panthera pardus): Molecular genetic variation. Reconocimiento filográfico de subespecies en leopardos (Panthera pardus): Variación genética molecular. Conservation Biology, 10(4):1115–1132, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moguel:1999:BCT

[MT99] Patricia Moguel and Victor M. Toledo. Biodiversity conservation in traditional coffee systems of Mexico. Conservation Biology, 13(1):11–21, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Munn:1995:CLC

[Mun95] Charles A. Munn. Conservation by, for, and with local communities. Conservation Biology, 9(5):1335–1336, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Murphy:1990:CBS

[Mur90] Dennis D. Murphy. Conservation biology and scientific method. Conservation Biology, 4(2):203–204, June 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Machtans:1996:URB

[MVH96] Craig S. Machtans, Marc-André Villard, and Susan J. Hannon. Use of riparian buffer strips as movement corridors by forest birds. Utilización de franjas riparias amortiguadoras como corredores por aves de bosque. Conservation Biology, 10(5): 1366–1379, October 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moyle:1990:BLT

[MW90] Peter B. Moyle and Jack E. Williams. Biodiversity loss in the temperate zone: Decline of the native fish fauna of California. Conservation Biology, 4(3):275–284, September 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McClanahan:1993:AFS

[MW93] T. R. McClanahan and R. W. Wolfe. Accelerating forest succession in a fragmented landscape: The role of birds and perches. Aceleración de la sucesión del bosque en un paisaje fragmentado: el papel de las aves y los cetros para ponerse en percha. Conservation Biology, 7(2):279–288, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mace:1998:DPT

[MW98] Richard D. Mace and John S. Waller. Demography and population trend of grizzly bears in the Swan Mountains, Montana. Conservation Biology, 12(5):1005–1016, October 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mladenoff:1994:APL

[MWCP94] David J. Mladenoff, Mark A. White, Thomas R. Crow, and John Pastor. Applying principles of landscape design and management to integrate old-growth forest enhancement and commodity use. Aplicando pincipios de diseño y manejo del paisaje para integrar el mejoramiento de bosques maduros y su utilización para consumo. Conservation Biology, 8(3):752–762, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mace:1999:LEG

[MWM^{+}99] Richard D. Mace, John S. Waller, Timothy L. Manley, Katherine Ake, and William T. Wittinger. Landscape evaluation of grizzly bear habitat in Western Montana. Evaluación de paisaje del hábitat del oso grizzly en el Oeste de Montana. Conservation Biology, 13(2):367–377, April 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Murphy:1994:RES

[MWN^{+}94] Dennis Murphy, David Wilcove, Reed Noss, John Harte, Carl Safina, Jane Lubchenco, Terry Root, Victor Sher, Les Kaufman, Michael Bean, and Stuart Pimm. On reauthorization of the Endangered Species Act. Conservation Biology, 8(1):1–3, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Malik:1997:PPT

[MWS^{+}97] S. Malik, P. J. Wilson, R. J. Smith, D. M. Lavigne, and B. N. White. Pinniped penises in trade: a molecular-genetic investigation. Conservation Biology, 11(6):1365–1374, December 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McLean:1999:RLC

[MWW99] Ian F. G. McLean, Andrew D. Wight, and Gwyn Williams. The role of legislation in conserving Europe’s threatened species. Conservation Biology, 13(5):966–969, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nabhan:1994:PUT

[Nab94] Gary Paul Nabhan. Proximate and ultimate threats to endangered species. Conservation Biology, 8(4):928–929, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nabhan:1995:DRB

[Nab95] Gary Paul Nabhan. The dangers of reductionism in biodiversity conservation. Conservation Biology, 9(3):479–481, June 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nilsson:1995:TDL

[NABE95] Sven G. Nilsson, Ulf Arup, Rickard Baranowski, and Stefan Ekman. Tree-dependent lichens and beetles as indicators in conservation forests. Conservation Biology, 9(5):1208–1215, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Naess:1996:LLR

[Nae96] Arne Naess. Living a life that reflects evolutionary insight. Conservation Biology, 10(6):1557–1559, December 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Naeem:1998:SRE

[Nae98] Shahid Naeem. Species redundancy and ecosystem reliability. Conservation Biology, 12(1):39–45, February 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nichols:1998:IMS

[NBH^{+}98] James D. Nichols, Thierry Boulinier, James E. Hines, Kenneth H. Pollock, and John R. Sauer. Inference methods for spatial variation in species richness and community composition when not all species are detected. Conservation Biology, 12(6):1390–1398, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Neel:1996:IVE

[NCE96] Maile C. Neel, Janet Clegg, and Norman C. Ellstrand. Isozyme variation in Echinocereus engelmannii var. munzii (Cacteceae). La variación de isoenzimas en Echinocereus engelmannii var. munzii (Cactaceae). Conservation Biology, 10(2):622–631, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nunney:1994:EEP

[NE94] Leonard Nunney and Diane R. Elam. Estimating the effective population size of conserved populations estimacion del tamaño poblacional efectivo de poblaciones protejidas. Conservation Biology, 8(1):175–184, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nehlsen:1994:SSR

[Neh94] Willa Nehlsen. Salmon stocks at risk: Beyond 214. Conservation Biology, 8(3):867–869, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Newmark:1991:TFF

[New91] William D. Newmark. Tropical forest fragmentation and the local extinction of understory birds in the Eastern Usambara Mountains, Tanzania. Conservation Biology, 5(1):67–78, March 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Newmark:1995:EMP

[New95] William D. Newmark. Extinction of mammal populations in Western North American national parks. Extinción de poblaciones de mamíferos en el los Parques Nacionales del oeste de Norte América. Conservation Biology, 9(3):512–526, June 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Newmark:1996:ITP

[New96] William D. Newmark. Insularization of Tanzanian parks and the local extinction of large mammals. Aislamiento de parques de Tanzania y la extinción local de mamíferos grandes. Conservation Biology, 10(6):1549–1556, December 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nabhan:1993:CNW

[NF93] Gary Paul Nabhan and Ted Fleming. The conservation of new world mutualisms. Conservation Biology, 7(3):457–462, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nowak:1998:VRW

[NF98] R. M. Nowak and N. E. Federoff. Validity of the red wolf: Response to Roy et al. Conservation Biology, 12(3):722–725, June 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [RGSW96] and response [WRG98].

Nilsson:1992:PAS

[NG92] Christer Nilsson and Frank Götmark. Protected areas in Sweden: Is natural variety adequately represented? Conservation Biology, 6(2):232–242, June 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nantel:1996:PVA

[NGN96] Patrick Nantel, Daniel Gagnon, and Andrée Nault. Population viability analysis of American ginseng and wild leek harvested in stochastic environments. Análisis de viabilidad poblacional del ginseng americano y el puerro silvestre cosechados en ambientes estocásticos. Conservation Biology, 10(2):608–621, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Norton:1995:FDP

[NHA95] David A. Norton, Richard J. Hobbs, and Lyn Atkins. Fragmentation, disturbance, and plant distribution: Mistletoes in woodland remnants in the Western Australian wheatbelt. La fragmentación, perturbación y distribución de las plantas: los muérdagos en remanentes del bosque en el cinturón triguero de Australia occidental. Conservation Biology, 9(2):426–438, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Niemela:1997:IBF

[Nie97] Jari Niemelä. Invertebrates and boreal forest management: Invertebrados y manejo de bosques boreales. Conservation Biology, 11(3):601–610, June 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nichols:1998:IGH

[NKA98] William F. Nichols, Keith T. Killingbeck, and Peter V. August. The influence of geomorphological heterogeneity on biodiversity. II. A landscape perspective. Conservation Biology, 12(2):371–379, April 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Niemela:1993:ECC

[NLS93] Jari Niemelä, David Langor, and John R. Spence. Effects of clear-cut harvesting on boreal ground-beetle assemblages (Coleoptera: Carabidae) in Western Canada. Efecto de la tala en grupos de carábidos boreales (Coleoptera: Carabidae) en el oeste de Canadá. Conservation Biology, 7(3):551–561, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noss:1995:ESL

[NM95] Reed F. Noss and Dennis D. Murphy. Endangered species left homeless in sweet home. Conservation Biology, 9(2):229–331, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Neitlich:1997:HEL

[NM97] Peter N. Neitlich and Bruce McCune. Hotspots of epiphytic lichen diversity in two young managed forests. Sitios críticos de diversidad de líquenes epífitos en dos bosques jóvenes bajo manejo. Conservation Biology, 11(1):172–182, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Newmark:1994:CBW

[NMGS94] William D. Newmark, David N. Manyanza, Deo-Gratias M. Gamassa, and Hashan I. Sariko. The conflict between wildlife and local people living adjacent to protected areas in Tanzania: Human density as a predictor. El conflicto entre la fauna silvestre y la población que vive adyacente a áreas protegidas en Tanzania: Densidad humana utilizada como pronosticador. Conservation Biology, 8(1):249–255, March 1994. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nepstad:1998:FFP

[NMV^{+}98] Daniel Nepstad, Adriana Moreira, Adalberto Veríssimo, Paul Lefebvre, Peter Schlesinger, Christopher Potter, Carlos Nobre, Alberto Setzer, Thelma Krug, Ana Cristina Barros, Ane Alencar, and João Raposo Pereira. Forest fire prediction and prevention in the Brazilian Amazon. Conservation Biology, 12(5): 951–953, October 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Norton:1991:TAO

[Nor91] David A. Norton. Trilepidea adamsii: an obituary for a species. Conservation Biology, 5(1):52–57, March 1991. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Norse:1993:GBE

[Nor93] Elliott A. Norse. Greater (and better) ecosystem management. Conservation Biology, 7(2):443–444, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

North:1996:VNT

[Nor96] Malcom North. Viewing nature through filters. Conservation Biology, 10(2):688–690, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noss:1990:CWM

[Nos90a] Reed F. Noss. Can we maintain biological and ecological integrity? Conservation Biology, 4(3):241–243, September 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noss:1990:IMB

[Nos90b] Reed F. Noss. Indicators for monitoring biodiversity: a hierarchical approach. Conservation Biology, 4(4):355–364, December 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noss:1991:SW

[Nos91] Reed F. Noss. Sustainability and wilderness. Conservation Biology, 5(1):120–122, March 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noss:1993:WCB

[Nos93] Reed F. Noss. Whither Conservation Biology? Conservation Biology, 7(2):215–217, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noss:1994:CCB

[Nos94] Reed F. Noss. Cows and conservation biology. Conservation Biology, 8(3):613–616, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noss:1995:L

[Nos95a] Reed F. Noss. Letter. Conservation Biology, 9(5):983, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noss:1995:PP

[Nos95b] Reed F. Noss. The perils of pollyannas. Conservation Biology, 9(4):701–703, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noss:1996:CBV

[Nos96a] Reed F. Noss. Conservation biology, values, and advocacy. Conservation Biology, 10(3):904, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noss:1996:CC

[Nos96b] Reed F. Noss. Conservation or convenience? Conservation Biology, 10(4):921–922, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noss:1996:ND

[Nos96c] Reed F. Noss. The naturalists are dying off. Conservation Biology, 10(1):1–3, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noss:1997:FUP

[Nos97] Reed F. Noss. The failure of universities to produce conservation biologists. Conservation Biology, 11(6):1267–1269, December 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noss:1998:ICS

[Nos98] Andrew J. Noss. The impacts of cable snare hunting on wildlife populations in the forests of the Central African Republic. Conservation Biology, 12(2):390–398, April 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nowak:1992:RWH

[Now92] Ronald M. Nowak. The red wolf is not a hybrid. Conservation Biology, 6(4):593–595, December 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See comments [PH92].

Noel:1998:SCO

[NPM98] Jill M. Noel, William J. Platt, and E. Barry Moser. Structural characteristics of old- and second-growth stands of longleaf pine (Pinus palustris) in the Gulf coastal region of the U.S.A. Conservation Biology, 12(3):533–548, June 1998. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noss:1996:CBC

[NQH^{+}96] Reed F. Noss, Howard B. Quigley, Maurice G. Hornocker, Troy Merrill, and Paul C. Paquet. Conservation biology and carnivore conservation in the Rocky Mountains. Biología de la conservación y conservacíon de carnivoros en las Montañas Rocallosas. Conservation Biology, 10(4):949–963, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Norton:1997:LEM

[NR97] David A. Norton and Nick Reid. Lessons in ecosystem management from management of threatened and pest loranthaceous mistletoes in New Zealand and Australia. Lecciones de manejo de ecosistemas manejo de muérdagos lorantáceos amenazados y plagas en Nueva Zelanda y Australia. Conservation Biology, 11(3):759–769, June 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Naughton-Treves:1998:PPC

[NT98] Lisa Naughton-Treves. Predicting patterns of crop damage by wildlife around Kibale National Park, Uganda. Conservation Biology, 12(1):156–168, February 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nielsen:1994:MDP

[NTT94] Jennifer L. Nielsen, Douglas Tupper, and W. Kelley Thomas. Mitochondrial DNA polymorphism in unique runs of Chinook salmon (Oncorhynchus tshawytscha) from the Sacramento–San Joaquin River Basin. Conservation Biology, 8(3):882–884, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noon:1991:ECW

[NY91] Barry R. Noon and Kimberley Young. Evidence of continuing worldwide declines in bird populations: Insights from an international conference in New Zealand. Conservation Biology, 5 (2):141–143, June 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Oliver:1993:PMR

[OB93] Ian Oliver and Andrew J. Beattie. A possible method for the rapid assessment of biodiversity. Un método posible para la evaluación rápida de la biodiversidad. Conservation Biology, 7 (3):562–568, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [Gol97b].

OBrien:1994:CCC

[O’B94] Stephen J. O’Brien. The Cheetah’s conservation controversy. Conservation Biology, 8(4):1153–1155, December 1994. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Oliver:1996:IMS

[OB96] Ian Oliver and Andrew J. Beattie. Invertebrate morphospecies as surrogates for species: a case study las morfoespecies de invertebrados como sustitutos de las especies: un caso de estudio. Conservation Biology, 10(1):99–109, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [Gol97b].

Oliver:1997:FTP

[OB97] Ian Oliver and Andrew J. Beattie. Future taxonomic partnerships: Reply to Goldstein. Conservation Biology, 11(2):575– 576, April 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [Gol97b].

Oliver:1998:SFP

[OBY98] Ian Oliver, Andrew J. Beattie, and Alan York. Spatial fidelity of plant, vertebrate, and invertebrate assemblages in multipleuse forest in Eastern Australia. Conservation Biology, 12(4): 822–835, August 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

OConnell:1992:RSB

[O’C92] Michael O’Connell. Response to: “Six biological reasons why the Endangered Species Act doesn’t work and what to do about it”. Conservation Biology, 6(1):140–143, March 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [Roh91] and rejoinder [Dud92].

Olson:1998:GRA

[OD98] David M. Olson and Eric Dinerstein. The global 200: a representation approach to conserving the Earth’s most biologically valuable ecoregions. Conservation Biology, 12(3):502–515, June 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1995:NBP

[OE95] David W. Orr and David Ehrenfeld. None so blind: The problem of ecological denial. Conservation Biology, 9(5):985–987, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

ORyan:1994:MDV

[OFH94] C. O’Ryan, J. R. B. Flamand, and E. H. Harley. Mitochondrial DNA variation in black rhinoceros (Diceros bicornis): Conservation management implications. Variación de ADN mitrocondrial en rinocerontes negros (diceros bicornis): Implicaciones para el manejo de conservación. Conservation Biology, 8(2):495–500, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ouborg:1996:DGS

[OG96] N. Joop Ouborg and Jan M. Groenendael. Demography, genetics, or statistics: Comments on a paper by Heschel and Paige. Conservation Biology, 10(4):1290–1291, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [HP95].

Ogden:1992:IHA

[Ogd92] John C. Ogden. The impact of Hurricane Andrew on the ecosystems of South Florida. Conservation Biology, 6(4):488–492, December 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

OBrien:1996:EHL

[OK96] Timothy G. O’Brien and Margaret F. Kinnaird. Effect of harvest on leaf development of the Asian palm Livistona rotundifolia. Efecto de la cosecha sobre el desarrollo de las hojas de la palmera Asiática Livistona rotundifolia. Conservation Biology, 10(1):53–58, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Oostermeijer:1998:RBP

[OLKD98] J. Gerard B. Oostermeijer, Sheila H. Luijten, Zdenka V. Křenová, and Hans C. M. Den Nijs. Relationships between population and habitat characteristics and reproduction of the rare Gentiana pneumonanthe L. Conservation Biology, 12(5):

1042–1053, October 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ogutu-Ohwayo:1993:EPN

[OO93] Richard Ogutu-Ohwayo. The effects of predation by Nile perch, Lates niloticus L., on the fish of Lake Nabugabo, with suggestions for conservation of endangered endemic cichlids. Efectos de la predación por la perca delNilo, Lates niloticus L. sobre los peces del Lago Nabugabo, con sugerencias para la conservación de cíclidos endémicos en peligro de extinción. Conservation Biology, 7(3):701–711, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

ONeal:1995:HEL

[OPRCB95] Amy E. O’Neal, Anand S. Pandian, Satya V. Rhodes-Conway, and Alan H. Bornbusch. Human economies, the land ethic, and sustainable conservation. Conservation Biology, 9(1):217–220, February 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Oleksyn:1994:PHD

[OR94] Jacek Oleksyn and Peter B. Reich. Pollution, habitat destruction, and biodiversity in Poland. Contaminación, destrucción del hábitat y biodiversidad en Polonia. Conservation Biology, 8(4):943–960, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1990:CEO

[Orr90a] David W. Orr. Is conservation education an oxymoron? Conservation Biology, 4(2):119–121, June 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1990:QM

[Orr90b] David W. Orr. The question of management. Conservation Biology, 4(1):8–9, March 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1990:R

[Orr90c] David W. Orr. Response. Conservation Biology, 4(4):348–349, December 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1990:VCE

[Orr90d] David W. Orr. The virtue of conservation education. Conservation Biology, 4(3):219–220, September 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1991:BDA

[Orr91a] David Orr. Biological diversity, agriculture, and the liberal arts. Conservation Biology, 5(3):268–270, September 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1991:PCP

[Orr91b] David Orr. Politics, conservation, and public education. Conservation Biology, 5(1):10–12, March 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1991:EC

[Orr91c] David W. Orr. The economics of conservation. Conservation Biology, 5(4):439–441, December 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1991:RC

[Orr91d] David W. Orr. Rating colleges. Conservation Biology, 5(2):138– 140, June 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1992:BE

[Orr92a] David W. Orr. The business of education. Conservation Biology, 6(3):315–317, September 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1992:EED

[Orr92b] David W. Orr. Education and the ecological design arts. Conservation Biology, 6(2):162–164, June 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1992:LL

[Orr92c] David W. Orr. For the love of life. Conservation Biology, 6(4): 486–487, December 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1992:STI

[Orr92d] David W. Orr. Some thoughts on intelligence. Conservation Biology, 6(1):9–11, March 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1993:AP

[Orr93a] David W. Orr. Architecture as pedagogy. Conservation Biology, 7(2):226–228, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1993:ECL

[Orr93b] David W. Orr. Educating a constituency for the long haul. Conservation Biology, 7(4):752–754, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1993:FT

[Orr93c] David W. Orr. Forests and trees. Conservation Biology, 7(3): 454–456, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1993:PDD

[Orr93d] David W. Orr. The problem of disciplines/the discipline of problems. Conservation Biology, 7(1):10–12, March 1993. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1994:ESO

[Orr94a] David W. Orr. The effective shape of our future. Conservation Biology, 8(3):622–624, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1994:PHP

[Orr94b] David W. Orr. Professionalism and the human prospect. Conservation Biology, 8(1):9–11, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1994:TF

[Orr94c] David W. Orr. Technological fundamentalism. Conservation Biology, 8(2):335–337, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1994:TB

[Orr94d] David W. Orr. Twine in the baler. Conservation Biology, 8(4): 931–933, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1995:CC

[Orr95] David W. Orr. Conservation and conservatism. Conservation Biology, 9(2):242–245, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1996:SK

[Orr96a] David W. Orr. Slow knowledge. Conservation Biology, 10 (3):699–702, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1996:VN

[Orr96b] David W. Orr. Virtual nature. Conservation Biology, 10(1):8– 9, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1997:API

[Orr97] David W. Orr. Architecture as pedagogy II. Conservation Biology, 11(3):597–600, June 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1998:LNE

[Orr98] David W. Orr. The limits of nature and the educational nature of limits. Conservation Biology, 12(4):746–748, August 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1999:AS

[Orr99a] David Orr. The architecture of science. Conservation Biology, 13(2):228–231, April 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1999:ECC

[Orr99b] David W. Orr. Education, careers, and callings: the practice of conservation biology. Conservation Biology, 13(6):1242–1245, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Orr:1999:V

[Orr99c] David W. Orr. Verbicide. Conservation Biology, 13(4):696– 699, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

ONeil:1995:UWC

[OSEC95] Thomas A. O’Neil, Robert J. Steidl, W. Daniel Edge, and Blair Csuti. Using wildlife communities to improve vegetation classification for conserving biodiversity. La utilización de comunidades silvestres para mejorar la clasificación de la vegetación para la conservación de la biodiversidad. Conservation Biology, 9(6):1482–1491, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Paetkau:1999:UGI

[Pae99] David Paetkau. Using genetics to identify intraspecific conservation units: a critique of current methods. Conservation Biology, 13(6):1507–1509, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pyke:1999:POD

[PAG^{+}99] Christopher R. Pyke, Peter S. Alagona, Noah Goldstein, Britta Bierwagen, Jennifer Merrick, Heather Rosenberg, and John Gallo. A plan for outreach: Defining the scope of conservation education. Conservation Biology, 13(6):1238–1239, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Paine:1995:CRC

[Pai95] R. T. Paine. A conversation on refining the concept of Keystone species. Conservation Biology, 9(4):962–964, August 1995. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Prescott-Allen:1990:HMP

[PAPA90] Robert Prescott-Allen and Christine Prescott-Allen. How many plants feed the world? Conservation Biology, 4(4):365–374, December 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Parnell:1993:PTR

[Par93] J. Parnell. Plant taxonomic research, with special reference to the tropics: Problems and potential solutions. Conservation Biology, 7(4):809–814, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Parnell:1995:SHI

[Par95] John Parnell. Systematics helps identify conservation priorities. Conservation Biology, 9(4):972–973, August 1995. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Patten:1991:HIG

[Pat91] Duncan T. Patten. Human impacts in the Greater Yellowstone Ecosystem: Evaluating sustainability goals and ecoredevelopment. Conservation Biology, 5(3):405–411, September 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Paton:1994:EEA

[Pat94a] Peter W. C. Paton. The effect of edge on avian nest success: How strong is the evidence? El efecto de borde sobre el éxito en la nidificación: `?Cuan firme es la evidencia? Conservation Biology, 8(1):17–26, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Paton:1994:SIR

[Pat94b] Peter W. C. Paton. Statistical independence: Reply to Gates. Conservation Biology, 8(3):620, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pendleton:1994:UPM

[PB94a] Linwood H. Pendleton and Anita Breda. Underwater parks may not be the best conservation tool for Lake Tanganyika. Conservation Biology, 8(2):330–331, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Prober:1994:CGW

[PB94b] Suzanne M. Prober and A. H. D. Brown. Conservation of the grassy white box woodlands: Population genetics and fragmentation of Eucalyptus albens. Conservación de los bosques herbáceos de Eucaliptus albens: Genética de poblaciones y fragmentacíon en Eucaliptus albens. Conservation Biology, 8(4): 1003–1013, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Powell:1995:IIM

[PB95] George V. N. Powell and Robin Bjork. Implications of intratropical migration on reserve design: a case study using Pharomachrus mocinno. Las concecuencias de la migración intratropical para el diseño de reservas: un caso de estudio usando Pharomachrus mocinno. Conservation Biology, 9(2):354–362, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pharo:1999:VPD

[PBB99] Emma J. Pharo, Andrew J. Beattie, and Doug Binns. Vascular plant diversity as a surrogate for bryophyte and lichen diversity. Diversidad de plantas vasculares como substituto de la diversidad de briofitas y líquenes. Conservation Biology, 13 (2):282–292, April 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Perry:1990:SME

[PBBA90] D. A. Perry, J. G. Borchers, S. L. Borchers, and M. P. Amaranthus. Species migrations and ecosystem stability during climate change: The belowground connection. Conservation Biology, 4 (3):266–274, September 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Porneluzi:1993:RPT

[PBG^{+}93] Paul Porneluzi, James C. Bednarz, Laurie J. Goodrich, Nancy Zawada, and Jeff Hoover. Reproductive performance of territorial ovenbirds occupying forest fragments and a contiguous forest in Pennsylvania. Conservation Biology, 7(3):618–622, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pearlstine:1995:ICD

[PBKM95] Leonard G. Pearlstine, Laura A. Brandt, Wiley M. Kitchens, and Frank J. Mazzotti. Impacts of citrus development on habitats of southwest Florida. Conservation Biology, 9(5): 1020–1032, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pearson:1992:WWS

[PC92] David L. Pearson and Fabio Cassola. World-wide species richness patterns of tiger beetles (Coleoptera: Cicindelidae): Indicator taxon for biodiversity and conservation studies. Conservation Biology, 6(3):376–391, September 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Powell:1993:ASP

[PC93] Abby N. Powell and Francesca J. Cuthbert. Augmenting small populations of plovers: an assessment of cross-fostering and captive-rearing. Promoviendo el crecimiento de pequeñas poblaciones de chorlitejos: una evaluación de la adopción cruzada y la crianza en cautiverio. Conservation Biology, 7 (1):160–168, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pounds:1994:ADC

[PC94] J. Alan Pounds and Martha L. Crump. Amphibian declines and climate disturbance: The case of the golden toad and the harlequin frog. Declinación de anfibios y perturbaciones climáticas: El caso del sapo dorado y la rana harlequin. Conservation Biology, 8(1):72–85, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Primm:1996:MSP

[PC96] Steven A. Primm and Tim W. Clark. Making sense of the policy process for carnivore conservation. Conservación de carnívoros en las Montañas Rocallosas: Comprensión del proceso político. Conservation Biology, 10(4):1036–1045, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pearson:1998:GPS

[PC98] David L. Pearson and Steven S. Carroll. Global patterns of species richness: Spatial models for conservation planning using bioindicator and precipitation data. Conservation Biology, 12 (4):809–821, August 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Paaby:1991:TRR

[PCG91] Pia Paaby, David B. Clark, and Héctor González. Training rural residents as naturalist guides: Evaluation of a pilot project in Costa Rica. Conservation Biology, 5(4):542–546, December 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pilskaln:1998:RSB

[PCM98] Cynthia H. Pilskaln, James H. Churchill, and Lawrence M. Mayer. Resuspension of sediment by bottom trawling in the Gulf of Maine and potential geochemical consequences. Conservation Biology, 12(6):1223–1229, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Press:1996:I

[PD96] Daniel Press and Daniel F. Doak. Introduction. Conservation Biology, 10(6):1489–1490, December 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pennock:1997:CES

[PD97] David S. Pennock and Walter W. Dimmick. Critique of the evolutionarily significant unit as a definition for “Distinct Population Segments” under the U.S. Endangered Species Act. Crítica de la unidad evolutivamente significativa como una definición de los “Segmentos poblacionales notables” en el Acta de Especies en Peligro de E.U.A. Conservation Biology, 11(3):611– 619, June 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See reply [Wap98].

Phillips:1996:CGC

[PDC96] Christopher A. Phillips, Walter W. Dimmick, and John L. Carr. Conservation genetics of the common snapping turtle (Chelydra serpentina). La conservación genética de la tortuga Chelydra serpentina. Conservation Biology, 10(2):397–405, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Palomares:1996:MRP

[PDFG96] Francisço Palomares, Miguel Delibes, Pablo Ferreras, and Pilar Gaona. Mesopredator release and prey abundance: Reply to Litvaitis and Villafuerte. Conservation Biology, 10(2):678–679, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Press:1996:RLG

[PDS96] Daniel Press, Daniel F. Doak, and Paul Steinberg. The role of local government in the conservation of rare species. El papel de los gobiernos locales en la conservación de especies raras. Conservation Biology, 10(6):1538–1548, December 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pichler:1998:GIH

[PDSB98] F. B. Pichler, S. M. Dawson, E. Slooten, and C. S. Baker. Geographic isolation of Hector’s Dolphin populations described by mitochondrial DNA sequences. Conservation Biology, 12 (3):676–682, June 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pearl:1992:RMC

[Pea92] Mary Pearl. A road map for conservationists. Conservation Biology, 6(2):305–306, June 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pearman:1997:CAD

[Pea97] Peter B. Pearman. Correlates of amphibian diversity in an altered landscape of Amazonian Ecuador. Correlaciones de la diversidad de anfibios en un paisaje alterado de la Amazonía Ecuatoriana. Conservation Biology, 11(5):1211–1225, October 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peet:1996:PPB

[Pee96] Robert K. Peet. A personal perspective on biodiversity. Conservation Biology, 10(2):691–692, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Petranka:1993:ETH

[PEH93] James W. Petranka, Matthew E. Eldridge, and Katherine E. Haley. Effects of timber harvesting on Southern Appalachian salamanders efectos de la tala del bosque sobre las salamandras en el sur de los apalaches. Conservation Biology, 7(2):363–370, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peluso:1992:IPM

[Pel92] Nancy Lee Peluso. The ironwood problem: (mis)Management and development of an extractive rainforest product. Conservation Biology, 6(2):210–219, June 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pell:1995:SUE

[Pel95] Bill Pell. Sustainable use and ecosystem management. Conservation Biology, 9(2):239–240, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Petit:1998:IPC

[PEP98] Rémy J. Petit, Abdelhamid El Mousadik, and Odile Pons. Identifying populations for conservation on the basis of genetic markers. Conservation Biology, 12(4):844–855, August 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Perry:1993:BWS

[Per93] David A. Perry. Biodiversity and wildlife are not synonymous. Conservation Biology, 7(1):204–205, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peres:1994:IRN

[Per94] Carlos A. Peres. Indigenous reserves and nature conservation in Amazonian forests. Conservation Biology, 8(2):586–588, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pierson:1996:ETC

[PERC96] E. D. Pierson, T. Elmqvist, W. E. Rainey, and P. A. Cox. Effects of tropical cyclonic storms on flying fox populations on the South Pacific islands of Samoa. El efecto de las tormentas ciclónicas tropicales sobre las poblaciones de zorros voladores en las islas del Pacífico Sur de Samoa. Conservation Biology, 10(2):438–451, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Petranka:1994:RIT

[Pet94] James W. Petranka. Response to impact of timber harvesting on salamanders. Conservation Biology, 8(1):302–303, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Petranka:1999:RSA

[Pet99] James W. Petranka. Recovery of salamanders after clearcutting in the Southern Appalachians: a critique of Ash’s estimates.

Conservation Biology, 13(1):203–205, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Porneluzi:1999:SLF

[PF99] Paul A. Porneluzi and John Faaborg. Season-long fecundity, survival, and viability of ovenbirds in fragmented and unfragmented landscapes. Fecundidad a lo largo de la estación, supervivencia y viabilidad del hornero (Seiurus aurocapillus) en paisajes fragmentados y no fragmentados. Conservation Biology, 13(5):1151–1161, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pounds:1997:TNM

[PFSG97] J. Alan Pounds, Michael P. L. Fogden, Jay M. Savage, and George C. Gorman. Tests of null models for amphibian declines on a tropical mountain. Conservation Biology, 11(6): 1307–1322, December 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pease:1991:PBT

[PG91] Craig M. Pease and Joseph A. Grzybowaki. Protection of biodiversity in Texas. Conservation Biology, 5(1):7–8, March 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Palomares:1995:PEG

[PGFD95] F. Palomares, P. Gaona, P. Ferreras, and M. Delibes. Positive effects on game species of top predators by controlling smaller predator populations: an example with Lynx, mongooses, and rabbits. Conservation Biology, 9(2):295–305, April 1995. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Phillips:1994:QEA

[PGR^{+}94] O. Phillips, A. H. Gentry, C. Reynel, P. Wilkin, and C. Galvez-Durand B. Quantitative ethnobotany and Amazonian conservation. Etnobotánica cuantitativa y la conservación de la Amazonia. Conservation Biology, 8(1):225–248, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Prior:1997:PGS

[PGW97] K. A. Prior, H. L. Gibbs, and P. J. Weatherhead. Population genetic structure in the black rat snake: Implications for management. Estructura genética poblacional en la serpiente rata negra: Implicaciones para su manejo. Conservation Biology, 11(5):1147–1158, October 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Phillips:1992:CRW

[PH92] Michael K. Phillips and V. Gary Henry. Comments on red wolf taxonomy. Conservation Biology, 6(4):596–599, December 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [Now92].

Polisar:1994:CLE

[PH94] John Polisar and Robert H. Horwich. Conservation of the large, economically important river turtle Dermatemys mawii in Belize. Conservation Biology, 8(2):338–342, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Paige:1996:IDS

[PH96] Ken N. Paige and M. Shane Heschel. Inbreeding depression in scarlet gilia: a reply to Ouborg and Van Groenendael. Conservation Biology, 10(4):1292–1294, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Phillips:1993:UCR

[Phi93] Oliver Phillips. Using and conserving the rainforest. Conservation Biology, 7(1):6–7, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Phillips:1994:MBP

[Phi94] Oliver Phillips. Making biodiversity profitable. Conservation Biology, 8(1):319–320, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pister:1994:REA

[Pis94] E. P. Pister. Radical environmentalism: Anthologies and stories. Conservation Biology, 8(2):611–612, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Price:1994:ASD

[PK94] Mary V. Price and Patrick A. Kelly. An age-structured demographic model for the endangered Stephens’ kangaroo rat. Un modelo demográfico estructurado por edades para la población de ratas canguros de Stephen. Conservation Biology, 8(3): 810–821, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pascual:1997:IMS

[PKH97] Miguel A. Pascual, Peter Kareiva, and Ray Hilborn. The influence of model structure on conclusions about the viability and harvesting of Serengeti wildebeest. La influencia de la estructura de un modelo en la conclusiones relacionadas con la viabilidad y cosecha del Ña del Serengetí. Conservation Biology, 11(4):966–976, August 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pressey:1994:LGS

[PL94] R. L. Pressey and V. S. Logan. Level of geographical subdivision and its effects on assessments of reserve coverage: a review of regional studies. Nivel de subdivisión geográfica y sus efectos en la evaluación de la cobertura de reservas: una revisión de estudios regionales. Conservation Biology, 8(4):1037–1046, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pressey:1995:RCR

[PL95] R. L. Pressey and V. S. Logan. Reserve coverage and requirements in relation to partitioning and generalization of land classes: Analyses for Western New South Wales. Cobertura y requerimientos de reservas en relación con la partición y generalización de las clases de tierras para el oeste de Nueva Gales del Sur. Conservation Biology, 9(6):1506–1517, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Primack:1992:DCL

[PM92] Richard B. Primack and S. L. Miao. Dispersal can limit local plant distribution. La dispersión puede limitar la distribución local de las plantas. Conservation Biology, 6(4):513–519, December 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pavlik:1993:REP

[PNH93] Bruce M. Pavlik, Daniel L. Nickrent, and Ann M. Howald. The recovery of an endangered plant. I. Creating a new population of Amsinckia grandiflora. La recuperación de una planta en peligro de extinción. I. Creando una nueva población de Amsinckia grandiflora. Conservation Biology, 7(3):510–526, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peterson:1999:ASC

[PNS99] A. Townsend Peterson and Adolfo G. Navarro-Sigüenza. Alternate species concepts as bases for determining priority conservation areas. Conceptos alternos de especie como base para la determinación de areas de conservación prioritarias. Conservation Biology, 13(2):427–431, April 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pomeroy:1993:CHB

[Pom93] Derek Pomeroy. Centers of high biodiversity in Africa. Conservation Biology, 7(4):901–907, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Povilitis:1994:CGC

[Pov94] Tony Povilitis. Common ground for conservation biologists? Conservation Biology, 8(2):598–599, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Power:1991:EPE

[Pow91] Thomas Michael Power. Ecosystem preservation and the economy in the Greater Yellowstone area. Conservation Biology, 5 (3):395–404, September 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Putz:1993:RIL

[PP93] Francis E. Putz and Michelle A. Pinard. Reduced-impact logging as a carbon-offset method. Conservation Biology, 7(4): 755–757, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Petit:1996:SCC

[PP96] Sophie Petit and Leon Pors. Survey of columnar cacti and carrying capacity for nectar-feeding bats on Curaçao. Estudio de los cactus columnares y la capacidad de carga de los murciélagos nectarívoros de Curaçao. Conservation Biology, 10(3):769–775, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Packer:1991:CSP

[PPR^{+}91] C. Packer, A. E. Pusey, H. Rowley, D. A. Gilbert, J. Martenson, and S. J. O’Brien. Case study of a population bottleneck: Lions of the Ngorongoro Crater. Conservation Biology, 5(2):219–230, June 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Prendergast:1999:GBT

[PQL99] John R. Prendergast, Rachel M. Quinn, and John H. Lawton. The gaps between theory and practice in selecting nature reserves. Distanciamiento entre teoría y práctica en la selección de reservas naturales. Conservation Biology, 13(3):484–492, June 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pressey:1994:AHR

[Pre94] R. L. Pressey. Ad Hoc reservations: Forward or backward steps in developing representative reserve systems? Reservaciones ad hoc: pasos adelante o atrás en el desarrollo de sistemas de reservas representativos? Conservation Biology, 8(3):662–668, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Patton:1998:UHD

[PRH98] Timothy M. Patton, Frank J. Rahel, and Wayne A. Hubert. Using historical data to assess changes in Wyoming’s fish fauna. Conservation Biology, 12(5):1120–1128, October 1998. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pringle:1991:URE

[Pri91] Catherine M. Pringle. U.S.–Romanian environmental reconnaissance of the Danube Delta. Conservation Biology, 5(4): 442–445, December 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Primack:1992:FBR

[Pri92] Richard B. Primack. Funding for biodiversity research. Conservation Biology, 6(2):307–308, June 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Primack:1996:FPG

[Pri96a] Richard B. Primack. Finding the pot of gold. Conservation Biology, 10(2):690–691, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Primm:1996:PAG

[Pri96b] Steven A. Primm. A pragmatic approach to grizzly bear conservation. Un acercamiento pragmático a la conservación del oso pardo. Conservation Biology, 10(4):1026–1035, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pritchard:1999:SBT

[Pri99] Peter C. H. Pritchard. Status of the black turtle. Conservation Biology, 13(5):1000–1003, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peterson:1996:RSL

[PS96] Markus J. Peterson and Nova J. Silvy. Reproductive stages limiting productivity of the endangered Attwater’s prairie chicken estados reproductivos limitantes de la productividad de la gallina de pradera attwater en peligro. Conservation Biology, 10 (4):1264–1276, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Panzer:1998:EVB

[PS98] Ron Panzer and Mark W. Schwartz. Effectiveness of a vegetation-based approach to insect conservation. Conservation Biology, 12(3):693–702, June 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Part:1999:CVS

[PS99] Tomas Pärt and Bo Söderström. Conservation value of seminatural pastures in Sweden: Contrasting botanical and avian measures. Valor de la conservación de pastizales seminaturales en Suecia: Contraste de medidas botánicas y aviares. Conservation Biology, 13(4):755–765, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Prober:1998:CGW

[PSB98] Suzanne M. Prober, Linden H. Spindler, and A. H. D. Brown. Conservation of the grassy white box woodlands: Effects of remnant population size on genetic diversity in the allotetraploid herb Microseris lanceolata. Conservation Biology, 12 (6):1279–1290, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pray:1994:EDI

[PSGS94] Leslie A. Pray, James M. Schwartz, Charles J. Goodnight, and Lori Stevens. Environmental dependency of inbreeding depression: Implications for conservation biology. Dependencia ambiental de la depresión de endocría: Implicaciones para la conservación biológica. Conservation Biology, 8(2):562–568, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Parker:1999:MVE

[PSH99] Karen M. Parker, Ruby J. Sheffer, and Philip W. Hedrick. Molecular variation and evolutionarily significant units in the endangered gila topminnow. Conservation Biology, 13(1):108– 116, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Polziehn:1995:BMD

[PSSB95] Renee O. Polziehn, Curtis Strobeck, Jane Sheraton, and Robin Beech. Bovine mtDNA discovered in North American bison populations. Conservation Biology, 9(6):1638, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peres:1995:ANR

[PT95] Carlos A. Peres and John W. Terborgh. Amazonian nature reserves: an analysis of the defensibility status of existing conservation units and design criteria for the future. Las reservas naturales Amazónicas: un análisis del estado relativo de protección de las unidades de conservación existentes y del criterio de diseño para el futuro. Conservation Biology, 9(1):34–46, February 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Post:1998:RMW

[PTK^{+}98] D. M. Post, J. P. Taylor, J. F. Kitchell, M. H. Olson, D. E. Schindler, and B. R. Herwig. The role of migratory waterfowl as nutrient vectors in a managed wetland. Conservation Biology, 12(4):910–920, August 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Planty-Tabacchi:1996:ISR

[PTTN^{+}96] Anne-Marie Planty-Tabacchi, Eric Tabacchi, Robert J. Naiman, Collette Deferrari, and Henri Décamps. Invasibility of species-rich communities in riparian zones. La propensión a la invasión de las comunidades de zonas ribereñas ricas en especies. Conservation Biology, 10(2):598–607, April 1996. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Putz:1992:UR

[Put92] Francis E. Putz. Unnecessary rifts. Conservation Biology, 6 (2):301–302, June 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Putz:1993:HTF

[Put93] Francis E. Putz. Hope for tropical forestry and conservation. Conservation Biology, 7(3):734–736, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Putz:1997:BGC

[Put97] Francis E. Putz. A breeding ground for conservation biologists. Conservation Biology, 11(3):813–814, June 1997. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Purcell:1998:DRS

[PV98] Kathryn L. Purcell and Jared Verner. Density and reproductive success of California towhees. Conservation Biology, 12(2):442– 450, April 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pinedo-Vasquez:1990:UVT

[PVZJCI90] Miguel Pinedo-Vasquez, Daniel Zarin, Peter Jipp, and Jomber Chota-Inuma. Use-values of tree species in a communal forest reserve in northeast Peru. Conservation Biology, 4(4):405–416, December 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Paetkau:1998:VGD

[PWC^{+}98] David Paetkau, Lisette P. Waits, Peter L. Clarkson, Lance Craighead, Ernie Vyse, Ryk Ward, and Curtis Strobeck. Variation in genetic diversity across the range of North American brown bears. Conservation Biology, 12(2):418–429, April 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Powell:1996:DAH

[PZSG96] Roger A. Powell, John W. Zimmerman, D. Erran Seaman, and James F. Gilliam. Demographic analyses of a hunted black bear population with access to a refuge. Análisis demográfico de una población de osos negros sujeta a la caza y con acceso a un refugio. Conservation Biology, 10(1):224–234, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Quintana-Ascencio:1996:IMD

[QAM96] Pedro F. Quintana-Ascencio and Eric S. Menges. Inferring metapopulation dynamics from patch-level incidence of Florida scrub plants. Inferencia de dinámica de metapoblaciones a partir de la ocurrencia de especies de plantas en parches de matorral en Florida. Conservation Biology, 10(4):1210–1219, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Quattro:1996:MEU

[QLDV96] Joseph M. Quattro, Paul L. Leberg, Michael E. Douglas, and Robert C. Vrijenhoek. Molecular evidence for a unique evolutionary lineage of endangered Sonoran desert fish (genus Poeciliopsis). Evidencia molecular de un linaje evolutivo único para peces en peligro de extinción (genero Poeciliopsis) del desierto de Sonora. Conservation Biology, 10(1):128–135, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Quigley:1995:TBC

[Qui95] Howard B. Quigley. Teaching the basics: a conservation training manual. Conservation Biology, 9(1):223, February 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ranker:1994:ASU

[RA94] Tom A. Ranker and Anna M. Arft. Allopolyploid species and the U.S. Endangered Species Act. Conservation Biology, 8(3): 895–897, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rabinowitz:1995:HSG

[Rab95a] Alan Rabinowitz. Helping a species go extinct: The Sumatran rhino in Borneo. Ayudando a una especie a extinguirse: El rinoceronte de Sumatra en Borneo. Conservation Biology, 9 (3):482–488, June 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rabinowitz:1995:L

[Rab95b] Alan Rabinowitz. Letter. Conservation Biology, 9(5):981–982, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ralls:1992:IAR

[Ral92] Katherine Ralls. The imperfect art of reintroduction. Conservation Biology, 6(3):481–482, September 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ralls:1995:IS

[Ral95] Katherine Ralls. But is it science? Conservation Biology, 9(5): 983–984, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rand:1996:NTM

[Ran96] David M. Rand. Neutrality tests of molecular markers and the connection between DNA polymorphism, demography, and conservation biology. Conservation Biology, 10(2):665–671, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reed:1995:ANA

[RB95a] J. Michael Reed and Andrew R. Blaustein. Assessment of “Nondeclining” amphibian populations using power analysis. Conservation Biology, 9(5):1299–1300, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rockwell:1995:EPS

[RB95b] R. F. Rockwell and G. F. Barrowclough. Effective population size and lifetime reproductive success. Conservation Biology, 9 (5):1225–1233, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rachlow:1997:CIP

[RB97a] Janet L. Rachlow and Joel Berger. Conservation implications of patterns of Horn regeneration in dehorned white rhinos. Implicaciones conservacionsitas de los patrones de regeneración de cuernos en rinocerontes descornados. Conservation Biology, 11 (1):84–91, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reed:1997:BSP

[RB97b] J. Michael Reed and Andrew R. Blaustein. Biologically significant population declines and statistical power. Conservation Biology, 11(1):281–282, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Russell:1998:PFT

[RBMA98] Gareth J. Russell, Thomas M. Brooks, Michael M. McKinney, and C. Gregory Anderson. Present and future taxonomic selectivity in bird and mammal extinctions. Conservation Biology, 12(6):1365–1376, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Richter:1997:TIF

[RBMM97] Brian D. Richter, David P. Braun, Michael A. Mendelson, and Lawrence L. Master. Threats to imperiled freshwater fauna. Amenazas a la fauna dulceacuícola en riesgo. Conservation Biology, 11(5):1081–1093, October 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Richter:1996:MAH

[RBPB96] Brian D. Richter, Jeffrey V. Baumgartner, Jennifer Powell, and David P. Braun. A method for assessing hydrologic alteration within ecosystems. Un métro para evaluar alteraciones hidrológicas dentro de ecosistemas. Conservation Biology, 10 (4):1163–1174, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rosenzweig:1994:IER

[RC94] Michael L. Rosenzweig and Colin W. Clark. Island extinction rates from regular censuses. Tasas de extinción de islas a partir de censos regulares. Conservation Biology, 8(2):491–494, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Russell:1994:IVS

[RCHP94] Robert W. Russell, F. Lynn Carpenter, Mark A. Hixon, and David C. Paton. The impact of variation in stopover habitat quality on migrant rufous hummingbirds. Impacto de la variación en la calidad de los hábitats de las paradas intermedias en colibríes (“rufous hummingbirds”) migratorios. Conservation Biology, 8(2):483–490, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rykken:1997:GBI

[RCM97] Jessica J. Rykken, David E. Capen, and Sean P. Mahabir. Ground beetles as indicators of land type diversity in the green mountains of Vermont: Escarabajos terrestres como indicadores del tipo de diversidad del suelo en green mountains, Vermont. Conservation Biology, 11(2):522–530, April 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roosenburg:1997:DTM

[RCMR97] Willem M. Roosenburg, William Cresko, Michael Modesitte, and Matthew B. Robbins. Diamondback terrapin (Malaclemys terrapin) mortality in crab pots. Mortalidad de la tortuga Malaclemis terrapin en trampas para cangrejos. Conservation Biology, 11(5):1166–1172, October 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ralls:1996:DCD

[RDE96] Katherine Ralls, Douglas P. Demaster, and James A. Estes. Developing a criterion for delisting the southern sea otter under the U.S. Endangered Species Act. Desarrollo de un criterio Para desenlistar la nutria marina del sur bajo el Acta de Especies Amenazadas de los Estados Unidos. Conservation Biology, 10 (6):1528–1537, December 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rydin:1997:BCH

[RDH97] Håkan Rydin, Martin Diekmann, and Tomas Hallingbäck. Biological characteristics, habitat associations, and distribution of macrofungi in Sweden. Características biológicas, asociaciones de hábitat y distribución de macrohongos en Suecia. Conservation Biology, 11(3):628–640, June 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rich:1994:DFF

[RDN94] Adam C. Rich, David S. Dobkin, and Lawrence J. Niles. Defining forest fragmentation by corridor width: The influence of narrow forest-dividing corridors on forest-nesting birds in Southern New Jersey. Definiendo la fragmentación del bosque en base al ancho de los corredores: Influencia de los corredores divisores del bosque angostos sobre los pájaros nidificantes del bosque en el sudeste de nueva Jersey. Conservation Biology, 8 (4):1109–1121, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rodriguez-Estrella:1998:RIE

[REDH98] Ricardo Rodríguez-Estrella, José Antonio Donázar, and Fernando Hiraldo. Raptors as indicators of environmental change in the scrub habitat of baja California sur, Mexico. Conservation Biology, 12(4):921–925, August 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reed:1996:USP

[Ree96] J. Michael Reed. Using statistical probability to increase confidence of inferring species extinction. Conservation Biology, 10 (4):1283–1285, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reed:1999:RBR

[Ree99] J. Michael Reed. The role of behavior in recent avian extinctions and endangerments. El papel de la conducta en procesos recientes de extinción y amenaza de aves. Conservation Biology, 13(2):232–241, April 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reinthal:1993:EBC

[Rei93] Peter Reinthal. Evaluating biodiversity and conserving Lake Malawi’s cichlid fish fauna. Evaluando la biodiversidad y conservando la fauna de peces cíclidos del Lago Malawi. Conservation Biology, 7(3):712–718, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reid:1994:CMC

[Rei94] Walter V. Reid. Conservation meets California. Conservation Biology, 8(2):607–608, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reinartz:1995:PSL

[Rei95] James A. Reinartz. Planting state-listed endangered and threatened plants. La plantación de plantas listadas como amenazadas o en peligro de extinción. Conservation Biology, 9(4): 771–781, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Renjifo:1999:CCS

[Ren99] Luis Miguel Renjifo. Composition changes in a subandean avifauna after long-term forest fragmentation. Cambios en la composición de la avifauna subandina después de una fragmentación de largo plazo del bosque. Conservation Biology, 13 (5):1124–1139, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rex:1993:TPB

[Rex93] Michael A. Rex. A taxonomic perfective on biodiversity. Conservation Biology, 7(1):209–210, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ramirez:1997:MGV

[RF97] Martin G. Ramirez and Julie L. Froehlig. Minimal genetic variation in a coastal dune arthropod: The trapdoor spider Aptostichus simus (Cyrtaucheniidae). Conservation Biology, 11(1):256–259, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rambo:1999:EVG

[RF99] Jennie L. Rambo and Stanley H. Faeth. Effect of vertebrate grazing on plant and insect community structure efecto del pastoreo de vertebrados en la estructura comunitaria de plantas e insectos. Conservation Biology, 13(5):1047–1054, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rave:1994:GAT

[RFDB94] Elizabeth H. Rave, Robert C. Fleischer, Fern Duvall, and Jeffrey M. Black. Genetic analyses through DNA fingerprinting of captive populations of Hawaiian geese. Análisis genético de poblaciones cautivas de gansos hawaianos usando “fingerprinting” de ADN. Conservation Biology, 8(3):744–751, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rabenold:1998:RAC

[RFG^{+}98] Kerry N. Rabenold, Peter T. Fauth, Bradley W. Goodner, Jennifer A. Sadowski, and Patricia G. Parker. Response of avian communities to disturbance by an exotic insect in spruce-fir forests of the Southern Appalachians. Conservation Biology, 12(1):177–189, February 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robichaux:1997:MGC

[RFM97] Robert H. Robichaux, Elizabeth A. Friar, and David W. Mount. Molecular genetic consequences of a population bottleneck associated with reintroduction of the Mauna Kea silversword (Arg yroxiphium sandwicense ssp. sandwicense [Asteraceae]). Consecuencias genético–moleculares de un cuello de botella poblacional asociado con la reintroducción de la espada plateada de Mauna Kea (Argyroxiphium sandwicense ssp. sandwicense [Asteraceae]). Conservation Biology, 11(5):1140–1146, October 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rieseberg:1995:HCI

[RG95] Loren H. Rieseberg and Daniel Gerber. Hybridization in the Catalina Island mountain mahogany (Cercocarpus traskiae): RAPD evidence. Conservation Biology, 9(1):199–203, February 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ray:1999:ESL

[RG99] Justina C. Ray and Joshua R. Ginsberg. Endangered species legislation beyond the borders of the United States. Conservation Biology, 13(5):956–958, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rastetter:1999:RAS

[RGH^{+}99] Edward B. Rastetter, Laura Gough, Anne E. Hartley, Darrell A. Herbert, Knute J. Nadelhoffer, and Mathew Williams. A revised assessment of species redundancy and ecosystem reliability. Conservation Biology, 13(2):440–443, April 1999. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roy:1996:MGP

[RGSW96] Michael S. Roy, Eli Geffen, Deborah Smith, and Robert K. Wayne. Molecular genetics of pre-1940 red wolves. Genética molecular de lobos rojos antes de 1940. Conservation Biology, 10(5):1413–1424, October 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See response [NF98].

Robinson:1993:FRC

[RH93] George R. Robinson and Steven N. Handel. Forest restoration on a closed landfill: Rapid addition of new species by bird dispersal. Restablecimiento del bosque en una clausura: Rápida adición de especies por aves dispersoras. Conservation Biology, 7(2):271–278, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rasker:1996:EDC

[RH96] Raymond Rasker and Arlin Hackman. Economic development and the conservation of large carnivores. Desarrollo económico y la conservación de carnivoros mayores. Conservation Biology, 10(4):991–1002, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rebholz:1997:CBS

[RH97a] W. E. R. Rebholz and E. H. Harley. Cytochrome b sequences from the endangered Saudi gazelle (Gazella saudiya) suggest hybridization with chinkara (G. bennetti). Conservation Biology, 11(1):251–255, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reichard:1997:PIW

[RH97b] Sarah Hayden Reichard and Clement W. Hamilton. Predicting invasions of woody plants introduced into North America. Predicción de invasiones de plantas leñosas introducidas a Norte américa. Conservation Biology, 11(1):193–203, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Richter:1993:MCE

[RHCB93] Andreas R. Richter, Stephen R. Humphrey, James B. Cope, and Virgil Brack, Jr. Modified cave entrances: Thermal effect on body mass and resulting decline of endangered Indiana bats (Myotis sodalis). Entradas modificadas de cavernas, efecto térmico resultante sobre la masa corporal y declinación de los murciélagos en peligro de extinción de Indiana (Myotis sodalis). Conservation Biology, 7(2):407–415, June 1993. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ruckelshaus:1997:ADR

[RHK97] Mary Ruckelshaus, Cynthia Hartway, and Peter Kareiva. Assessing the data requirements of spatially explicit dispersal models. Conservation Biology, 11(6):1298–1306, December 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ruckelshaus:1999:DLE

[RHK99] Mary Ruckelshaus, Cynthia Hartway, and Peter Kareiva. Dispersal and landscape errors in spatially explicit population models: a reply. Conservation Biology, 13(5):1223–1224, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ruggiero:1994:VAB

[RHS94] Leonard F. Ruggiero, Gregory D. Hayward, and John R. Squires. Viability analysis in biological evaluations: Concepts of population viability analysis, biological population, and ecological scale. Análisis de viabilidad en evaluaciones biológicas: Conceptos del análisis de viabilidad poblacional, poblaciones biológicas y escala ecológica. Conservation Biology, 8(2):364– 372, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robertson:1994:RKE

[RHSM94] Hugh A. Robertson, J. Rod Hay, Edward K. Saul, and Gerald V. Mccormack. Recovery of the kakerori: an endangered forest bird of the Cook Islands. La recuperación del kakerori: un pájaro del bosque de las Islas Cook, en peligro de extinción. Conservation Biology, 8(4):1078–1086, December 1994. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Richardson:1998:FTI

[Ric98] David M. Richardson. Forestry trees as invasive aliens. Conservation Biology, 12(1):18–26, February 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Risser:1995:BEF

[Ris95] Paul G. Risser. Biodiversity and ecosystem function biodiversidad y el funcionamiento del ecosistema. Conservation Biology, 9(4):742–746, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Risser:1999:ERB

[Ris99] Paul G. Risser. Examining relationships between ecosystem function and biodiversity: Reply to Goldstein. Conservation Biology, 13(2):438–439, April 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [Gol99b, Gol99a].

Rivas:1997:NHH

[Riv97] Jesús Antonio Rivas. Natural history: Hobby or science? Conservation Biology, 11(3):811–812, June 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reed:1996:CRF

[RJBB96] Rebecca A. Reed, Julia Johnson-Barnard, and William L. Baker. Contribution of roads to forest fragmentation in the Rocky Mountains. La contribución de los caminos a la fragmentación de bosques en las Montañas Rocallosas. Conservation Biology, 10(4):1098–1106, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ryman:1995:SBV

[RJL95] Nils Ryman, Per Erik Jorde, and Linda Laikre. Supportive breeding and variance effective population size. La crianza sostenida y la variacíon en el tamaño efectivo poblacional. Conservation Biology, 9(6):1619–1628, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ryman:1999:SBI

[RJL99] Nils Ryman, Per Erik Jorde, and Linda Laikre. Supportive breeding and inbreeding effective number: Reply to Nomura. Conservation Biology, 13(3):673–676, June 1999. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reading:1993:ATP

[RK93] Richard P. Reading and Stephen R. Kellert. Attitudes toward a proposed reintroduction of black-footed ferrets (Mustela nigripes) Actitudes hacia una propuesta para la reintroducción del hurón de patas negras (Mustela nigripes). Conservation Biology, 7(3):569–580, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Read:1993:HPG

[RKBP93] Andrew J. Read, Scott D. Kraus, Kathryn D. Bisack, and Debra Palka. Harbor porpoises and gill nets in the Gulf of Maine. Conservation Biology, 7(1):189–193, March 1993. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robertson:1999:EPL

[RKLS99] Alastair W. Robertson, Dave Kelly, Jenny J. Ladley, and Ashley D. Sparrow. Effects of pollinator loss on endemic New Zealand mistletoes (Loranthaceae). Efectos de la pérdida de polinizadores en muérdagos (Loranthaceae) de Nueva Zelandia. Conservation Biology, 13(3):499–508, June 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ryman:1991:ESB

[RL91] Nils Ryman and Linda Laikre. Effects of supportive breeding on the genetically effective population size. Conservation Biology, 5(3):325–329, September 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Richards:1996:TCA

[RL96] Christine Richards and Paul L. Leberg. Temporal changes in allele frequencies and a Population’s history of severe bottlenecks. Cambios temporales en frecuencias de alelos y la historia de una población con serveros cuellos de botella. Conservation Biology, 10(3):832–839, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rosenberg:1999:EFF

[RLD99] Kenneth V. Rosenberg, James D. Lowe, and André A. Dhondt. Effects of forest fragmentation on breeding tanagers: a continental perspective. Efectos de la fragmentación de bosque en tángaras reproductoras: una perspectiva continental. Conservation Biology, 13(3):568–583, June 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ratner:1997:PVA

[RLR97] Susan Ratner, Russell Lande, and Brett B. Roper. Population viability analysis of spring Chinook salmon in the South Umpqua River, Oregon. Análisis de viabilidad poblacional del salmón Chinook de primavera en el río South Umpqua, Oregon. Conservation Biology, 11(4):879–889, August 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rieman:1997:URC

[RM97] B. E. Rieman and D. L. Myers. Use of redd counts to detect trends in bull trout (Salvelinus confluentus) populations. Conservation Biology, 11(4):1015–1018, August 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roll:1997:RSI

[RMCM97] Julia Roll, Randall J. Mitchell, Robert J. Cabin, and Diane L. Marshall. Reproductive success increases with local density of conspecifics in a desert mustard (Lesquerella fendleri). El exito reproductivo incrementa con la densidad local de coespecificos en la mostaza del desierto (Lesquerella fendleri). Conservation Biology, 11(3):738–746, June 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rasker:1992:ETV

[RMJ92] Raymond Rasker, Michael V. Martin, and Rebecca L. Johnson. Economics: Theory versus practice in wildlife management. Conservation Biology, 6(3):338–349, September 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Raxworthy:1996:MAR

[RN96] Christopher J. Raxworthy and Ronald A. Nussbaum. Montane amphibian and reptile communities in Madagascar. Comunidades de anfibios y reptiles de montaña en Madagascar. Conservation Biology, 10(3):750–756, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rivas:1999:TCE

[RO99] Jesús Antonio Rivas and Renée Yvonne Owens. Teaching conservation effectively: a lesson from life-history strategies. Conservation Biology, 13(2):453–454, April 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robinson:1993:BWY

[Rob93a] John G. Robinson. “Believing what you know ain’t so”: Response to Holdgate and Munro. Conservation Biology, 7(4):

941–942, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [HM93].

Robinson:1993:LCS

[Rob93b] John G. Robinson. The limits to caring: Sustainable living and the loss of biodiversity. Los limites de cuidari: vida sostenible y la pérdida de biodiversidad. Conservation Biology, 7(1):20– 28, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roberts:1995:EFE

[Rob95a] Callum M. Roberts. Effects of fishing on the ecosystem structure of coral reefs. Conservation Biology, 9(5):988–995, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roberts:1995:RBF

[Rob95b] Callum M. Roberts. Rapid build-up of fish biomass in a Caribbean marine reserve. Crecimiento rápido en la biomasa de peces en una reserva marina del Caribe. Conservation Biology, 9(4):815–826, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robinson:1995:IMD

[Rob95c] Nicholas Andrew Robinson. Implications from mitochondrial DNA for management to conserve the eastern barred bandicoot (Perameles gunnii). Deducciones a partir del ADN mitocondrial para el manejo con fines de conservar el marsupial Perameles gunii. Conservation Biology, 9(1):114–125, February 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robinson:1999:LTC

[Rob99] W. Douglas Robinson. Long-term changes in the avifauna of Barro Colorado Island, Panama, a tropical forest isolate. Conservation Biology, 13(1):85–97, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rodenhouse:1992:PEC

[Rod92] Nicholas L. Rodenhouse. Potential effects of climatic change on a neotropical migrant landbird. Conservation Biology, 6 (2):263–272, June 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rodda:1993:HLB

[Rod93a] Gordon H. Rodda. How to Lie with biodiversity. Conservation Biology, 7(4):959–960, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rodriguez:1993:WA

[Rod93b] Jon Paul Rodríguez. What is America? Conservation Biology, 7(2):223, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rohlf:1991:SBR

[Roh91] Daniel J. Rohlf. Six biological reasons why the Endangered Species Act doesn’t work — and what to do about it. Conservation Biology, 5(3):273–282, September 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See response [O’C92].

Rohlf:1992:RO

[Roh92] Daniel J. Rohlf. Response to O’Connell. Conservation Biology, 6(1):144–145, March 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See rejoinder [Dud92].

Rojas:1992:SPC

[Roj92] Martha Rojas. The species problem and conservation: What are we protecting? Conservation Biology, 6(2):170–178, June 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rolston:1995:GSG

[Rol95] Holmes Rolston III. Genes for sale: Gargantuan computer system wanted. Conservation Biology, 9(6):1659–1660, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rose:1992:FTW

[Ros92] Debra A. Rose. Free trade and wildlife trade. Conservation Biology, 6(1):148–150, March 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roberts:1999:EOF

[ROT^{+}99] Callum M. Roberts, Alison Ormsby, John Tschirhart, Joel Berger, Jennifer L. O’Hara, David Johns, and Walter V. Reid.

Emptying the oceans of fish. Conservation Biology, 13(1): 216–222, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roy:1995:FW

[Roy95a] Michael Roy. Fighting for wolves. Conservation Biology, 9(6): 1660–1661, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roy:1995:UVE

[Roy95b] Michael Roy. Unity of voice in endangered species recovery. Conservation Biology, 9(2):457–458, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roebuck:1999:PPC

[RP99] Paul Roebuck and Paul Phifer. The persistence of positivism in conservation biology. Conservation Biology, 13(2):444–446, April 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Redford:1999:CBW

[RR99a] Kent H. Redford and Brian D. Richter. Conservation of biodiversity in a world of use. Conservación de la biodiversidad en un mundo de uso. Conservation Biology, 13(6):1246–1256, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ricciardi:1999:ERN

[RR99b] Anthony Ricciardi and Joseph B. Rasmussen. Extinction rates of North American freshwater fauna. Tasas de extinción de fauna de agua dulce en Norte América. Conservation Biology, 13(5):1220–1222, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Richter:1999:ASB

[RR99c] Brian D. Richter and Kent H. Redford. The art (and science) of brokering deals between conservation and use. Conservation Biology, 13(6):1235–1237, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robinson:1999:ESL

[RR99d] W. Douglas Robinson and Scott K. Robinson. Effects of selective logging on forest bird populations in a fragmented landscape. Conservation Biology, 13(1):58–66, February 1999. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ryan:1993:SPP

[RRM93] Mark R. Ryan, Brian G. Root, and Paul M. Mayer. Status of piping plovers in the Great Plains of North America: a demographic simulation model. Estado de los chorlitos (Charadrius melodus) en las Grandes Planicies de Norte América: un modelo demográfico de simulación. Conservation Biology, 7(3): 581–585, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rodriguez:1996:GDC

[RRS96] Jon Paul Rodríguez and Franklin Rojas-Suárez. Guidelines for the design of conservation strategies for the animals of Venezuela. Lineamientos para el diseño de estrategias de conservación de animales para los Venezuela. Conservation Biology, 10(4):1245–1252, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Recher:1991:LTC

[RS91a] H. F. Recher and D. L. Serventy. Long term changes in the relative abundances of birds in Kings Park, Perth, Western Australia. Conservation Biology, 5(1):90–102, March 1991. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reinthal:1991:FFM

[RS91b] P. N. Reinthal and M. L. J. Stiassny. The freshwater fishes of Madagascar: a study of an endangered fauna with recommendations for a conservation strategy. Conservation Biology, 5 (2):231–243, June 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rebelo:1992:WSN

[RS92] A. G. Rebelo and W. R. Siegfried. Where should nature reserves be located in the Cape Floristic Region, South Africa? Models for the spatial configuration of a reserve network aimed at maximizing the protection of floral diversity. Conservation Biology, 6(2):243–252, June 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Redford:1993:FDN

[RS93a] Kent H. Redford and Allyn Maclean Stearman. Forest-dwelling native Amazonians and the conservation of biodiversity: Interests in common or in collision? Habitantes nativos de la selva Amazónica y la conservación de la biodiversidad: Intereses en común o en conflicto? Conservation Biology, 7(2):248–255, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Redford:1993:CGR

[RS93b] Kent H. Redford and Allyn Maclean Stearman. On common ground? Response to Alcorn. Conservation Biology, 7(2):427– 428, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [Alc93].

Root:1993:CLS

[RS93c] Terry L. Root and Stephen H. Schneider. Can large-scale climatic models be linked with multiscale ecological studies? `?Pueden los modelos climáticos a gran escala conectarse con estudios ecológicos a multiples escalas? Conservation Biology, 7(2):256–270, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ralls:1995:CMS

[RS95a] Katherine Ralls and Anthony M. Starfield. Choosing a management strategy: Two structured decision-making methods for evaluating the predictions of stochastic simulation models. Seleccionando una estrategia de manejo: dos métodos estructurados de toma de decisiones para evaluar las predicciónes de modelos de simulación estocásticos. Conservation Biology, 9(1): 175–181, February 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rodgers:1995:SBD

[RS95b] James A. Rodgers, Jr. and Henry T. Smith. Set-back distances to protect nesting bird colonies from human disturbance in Florida distancia de alejamiento para proteger de las perturbaciones humanas a las colonias de aves nidificadoras en Florida. Conservation Biology, 9(1):89–99, February 1995. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rehm-Switky:1990:CE

[RSM90] Kathy M. Rehm-Switky and Dennis Murphy. Conservation education. Conservation Biology, 4(4):347–348, December 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roman:1999:PSC

[RSMB99] Joseph Roman, Steven D. Santhuff, Paul E. Moler, and Brian W. Bowen. Population structure and cryptic evolutionary units in the alligator snapping turtle. Conservation Biology, 13(1):135–142, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ratsirarson:1996:CMT

[RSR96] Joelisoa Ratsirarson, John A. Silander, Jr., and Alison F. Richard. Conservation and management of a threatened Madagascar palm species, Neodypsis decaryi, Jumelle. Conservación y manejo de la especie de palmera de Madagascar en peligro Neodypsis decaryi, Jumelle. Conservation Biology, 10(1):40– 52, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Romme:1991:IGC

[RT91] William H. Romme and Monica G. Turner. Implications of global climate change for biogeographic patterns in the Greater Yellowstone Ecosystem. Conservation Biology, 5(3):373–386, September 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ranta:1999:SIC

[RTNK99] Pertti Ranta, Antti Tanskanen, Jari Niemelä, and Arto Kurtto. Selection of islands for conservation in the urban archipelago of Helsinki, Finland. Selección de islas para conservación en el archipiélago urbano de Helsinki, Finlandia. Conservation Biology, 13(6):1293–1300, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Russell:1994:GBB

[RTOB94] William C. Russell, E. Tom Thorne, Robert Oakleaf, and Jonathan D. Ballou. The genetic basis of black-footed Ferret reintroduction. Conservation Biology, 8(1):263–266, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Redford:1990:TMB

[RTS90] Kent H. Redford, Andrew Taber, and Javier A. Simonetti. There is more to biodiversity than the tropical rain forests. Conservation Biology, 4(3):328–330, September 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rundel:1995:RSE

[Run95] Philip W. Rundel. The role of species in ecosystems. Conservation Biology, 9(2):467–469, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Raijmann:1994:GVO

[RVK^{+}94] Léon E. L. Raijmann, Nico C. Van Leeuwen, Ralph Kersten, J. Gerard B. Oostermeijer, Hans C. M. Den Nijs, and Steph B. J. Menken. Genetic variation and outcrossing rate in relation to population size in Gentiana pneumonanthe L.. La variación genética y la tasa de endocría en relación con el tamaño poblacional en Gentiana pneumonanthe L. Conservation Biology, 8 (4):1014–1026, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rohner:1999:LMH

[RW99] Christoph Rohner and David Ward. Large mammalian herbivores and the conservation of arid Acacia stands in the Middle East. Mamíferos herviboros grandes y la conservación de Acacia en el Medio Oriente. Conservation Biology, 13(5):1162–1171, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reed:1993:EPS

[RWES93] J. Michael Reed, Jeffrey R. Walters, Ted E. Emigh, and D. Erran Seaman. Effective population size in red-cockaded woodpeckers: Population and model differences. Tamaño poblacional efectivo en el pajaro carpintero cabeza roja: Diferencias poblacionales y de modelo. Conservation Biology, 7(2):302– 308, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ryder:1993:PHP

[Ryd93] Oliver A. Ryder. Przewalski’s horse: Prospects for reintroduction into the wild. Conservation Biology, 7(1):13–19, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ryman:1994:SBE

[Rym94] Nils Ryman. Supportive breeding and effective population size: Differences between inbreeding and variance effective numbers. Conservation Biology, 8(3):888–890, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Saberwal:1996:PPG

[Sab96] Vasant K. Saberwal. Pastoral politics: Gaddi grazing, degradation, and biodiversity conservation in Himachal Pradesh, India. Políticas agrarias: Pastoreo de las comunidades de Gaddi, degradación y conservación de la biodiversidad en Himachal Pradesh, India. Conservation Biology, 10(3):741–749, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Saberwal:1997:STM

[Sab97] Vasant K. Saberwal. Saving the tiger: More money or less power? Conservation Biology, 11(3):815–817, June 1997. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See response [KM97] and reply [Sab98].

Saberwal:1998:DEC

[Sab98] Vasant K. Saberwal. Degradation and environmental conservation: Response to Mishra and Rawat. Conservation Biology, 12(3):715–717, June 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [Sab97, MR98a].

Sader:1995:BPL

[Sad95] Steven Sader. Big picture of land use change. Conservation Biology, 9(6):1657–1658, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Safina:1993:BTW

[Saf93] Carl Safina. Bluefin tuna in the West Atlantic: Negligent management and the making of an endangered species. Conservation Biology, 7(2):229–238, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Swensen:1995:GAE

[SAH^{+}95] Susan M. Swensen, Gerard J. Allan, Melanie Howe, Wayne J. Elisens, Steven A. Junak, and Loren H. Rieseberg. Genetic analysis of the endangered island endemic Malacothamnus fasciculatus (Nutt.) Greene var. nesioticus (Rob.) Kearn. (Malvaceae). Análisis genético de Malacothamnus fasciculatus (Nutt.) Greene var. nesioticus (Rob.) Kearn. (Malvaceae), una especie endémica a islas y en peligro de extinción. Conservation Biology, 9(2):404–415, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Salwasser:1990:SCP

[Sal90] Hal Salwasser. Sustainability as a conservation paradigm. Conservation Biology, 4(3):213–216, September 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Salwasser:1991:NPS

[Sal91] Hal Salwasser. New perspectives for sustaining diversity in U.S. national forest ecosystems. Conservation Biology, 5(4): 567–569, December 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See response [FNN92].

Salwasser:1992:NPE

[Sal92] Hal Salwasser. From new perspectives to ecosystem management: Response to Frissell et al. and Lawrence and Murphy. Conservation Biology, 6(3):469, September 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Salafsky:1993:MUB

[Sal93] Nick Salafsky. Mammalian use of a buffer zone agroforestry system bordering Gunung Palung National Park, West Kalimantan, Indonesia. Conservation Biology, 7(4):928–933, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Salzman:1995:SA

[Sal95] Lorna Salzman. Scientists and advocacy. Conservation Biology, 9(4):709–710, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Salzman:1998:ESL

[Sal98] James Salzman. Ecosystem services and the law. Conservation Biology, 12(3):497–498, June 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Salonius:1999:PGU

[Sal99] Peter Salonius. Population growth in the United States and Canada: a role for scientists. Conservation Biology, 13(6): 1518–1519, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Samways:1990:LFW

[Sam90] Michael J. Samways. Land forms and winter habitat refugia in the conservation of montane grasshoppers in Southern Africa. Conservation Biology, 4(4):375–382, December 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Samways:1996:AUT

[Sam96a] Michael J. Samways. The art of unintelligent tinkering. Conservation Biology, 10(5):1307, October 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Samways:1996:PBI

[Sam96b] Michael J. Samways. Pioneering book on invertebrate conservation. Conservation Biology, 10(5):1476–1477, October 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Smith:1998:LAT

[SAM98] James L. David Smith, Sean C. Ahearn, and Charles McDougal. Landscape analysis of tiger distribution and habitat quality in Nepal. Conservation Biology, 12(6):1338–1346, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sandvik:1999:HPG

[San99] Hanno Sandvik. On human population growth, natural selection, and the tragedy of the commons. Conservation Biology, 13(2):447–449, April 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sanchez-Azofeifa:1999:PAC

[SAQMGQ^{+}99] G. Arturo Sánchez-Azofeifa, Carlos Quesada-Mateo, Pablo Gonzalez-Quesada, S. Dayanandan, and Kamaljit S. Bawa. Protected areas and conservation of biodiversity in the Tropics. Areas protegidas y conservación de biodiversidad en el trópico. Conservation Biology, 13(2):407–411, April 1999. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sarkar:1995:PPE

[Sar95] Sahotra Sarkar. A paleontological perspective on extinction. Conservation Biology, 9(3):694–695, June 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sarkar:1996:PBC

[Sar96] Sahotra Sarkar. Postmodernism and biodiversity conservation. Conservation Biology, 10(1):305–306, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shafroth:1995:GEN

[SAS95] Patrick B. Shafroth, Gregor T. Auble, and Michael L. Scott. Germination and establishment of the native plains cottonwood (Populus deltoides Marshall subsp. monilifera) and the exotic Russian-olive (Elaeagnus angustifolia L.). Conservation Biology, 9(5):1169–1175, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Saari:1998:FID

[SÅS98] Lennart Saari, Johan Åberg, and Jon E. Swenson. Factors influencing the dynamics of occurrence of the hazel grouse in a fine-grained managed landscape. Conservation Biology, 12 (3):586–592, June 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Strong:1994:PDW

[SB94] Allan M. Strong and G. Thomas Bancroft. Postfledging dispersal of white-crowned pigeons: Implications for conservation of deciduous seasonal forests in the Florida Keys. Dispersión de palomas cabeziblancas juveniles; implicaciones para la conservación de bosques deciduos estacionales en los callos de la Florida. Conservation Biology, 8(3):770–779, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Spotte:1995:VCT

[SB95a] Stephen Spotte and Patricia M. Bubucis. Visual censusing of two coral fishes important to the marine aquarium trade. Conservation Biology, 9(5):1304–1306, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stouffer:1995:EFF

[SB95b] Philip C. Stouffer and Richard O. Bierregaard, Jr. Effects of forest fragmentation on understory hummingbirds in Amazonian Brazil. Conservation Biology, 9(5):1085–1094, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Strittholt:1995:ABG

[SB95c] James R. Strittholt and Ralph E. J. Boerner. Applying biodiversity gap analysis in a regional nature reserve design for the edge of Appalachia, Ohio (U.S.A.). Aplicacíon de análisis de intervalos de biodiversidad en el diseño de una Reserva Natural Regional para en el borde de Appalachia, Ohio, Estados Unidos. Conservation Biology, 9(6):1492–1505, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shinneman:1997:NDB

[SB97] Douglas J. Shinneman and William L. Baker. Nonequilibrium dynamics between catastrophic disturbances and old-growth forests in ponderosa pine landscapes of the black hills. Conservation Biology, 11(6):1276–1288, December 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Semlitsch:1998:SIW

[SB98] Raymond D. Semlitsch and J. Russell Bodie. Are small, isolated wetlands expendable? Conservation Biology, 12(5):1129–1133, October 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shepherd:1999:ECB

[SB99] Philippa C. F. Shepherd and J. Sherman Boates. Effects of a commercial baitworm harvest on semipalmated sandpipers and their prey in the Bay of Fundy Hemispheric Shorebird Reserve. Efectos de la colecta comercial de gusanos para carnada en poblaciones del playero semipalmado y sus presas en la Bahía de la Reserva Fundy Hemispheric Shorebird. Conservation Biology, 13(2):347–356, April 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sieving:1994:PCD

[SBB^{+}94] Katie Sieving, BIO SCI 296A, Scan Bechta, Daniel Burmester, Michael Casazza, Michael Finan, Joyce Hunting, Shawna Martinez, Joseph Merz, Kathleen Norton, Barbara Rocco, Tamara Sasaki, Robert Stafford, Bret Stedman, Michael Wenzel, and Brian Williams. The population crisis demands a focused agenda. Conservation Biology, 8(1):305–307, March 1994. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Silver:1996:ECB

[SBL96] Whendee L. Silver, Sandra Brown, and Ariel E. Lugo. Effects of changes in biodiversity on ecosystem function in tropical forests. Efectos de los cambios en la biodiversidad sobre la función de los ecosistemas en las selvas tropicales. Conservation Biology, 10(1):17–24, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Scott:1997:NWT

[SC97a] J. Michael Scott and Blair Csuti. Noah worked two jobs. Conservation Biology, 11(5):1255–1257, October 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sites:1997:TSB

[SC97b] Jack W. Sites, Jr. and Keith A. Crandall. Testing species boundaries in biodiversity studies. Conservation Biology, 11 (6):1289–1297, December 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schonewald-Cox:1991:SVD

[SCAB91] Christine Schonewald-Cox, Rahman Azari, and Stephanie Blume. Scale, variable density, and conservation planning for mammalian carnivores. Conservation Biology, 5(4):491–495, December 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schindler:1991:CSB

[Sch91] D. W. Schindler. Comments on the Sustainable Biosphere Initiative. Conservation Biology, 5(4):550–551, December 1991.

CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See comments [GM91, Tho91, Wil91].

Schwartz:1992:CPI

[Sch92] Mark W. Schwartz. Conservation and pharmaceutical interests: The case of yew trees. Conservation Biology, 6(1):152–153, March 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Scheffer:1993:OGC

[Sch93a] Victor B. Scheffer. The Olympic goat controversy: a perspective. Conservation Biology, 7(4):916–919, December 1993. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Scheffer:1993:RA

[Sch93b] Victor B. Scheffer. Reply to the anunsens. Conservation Biology, 7(4):958, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Scheuer:1993:BBN

[Sch93c] James H. Scheuer. Biodiversity: Beyond Noah’s ark. Conservation Biology, 7(1):206–207, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schmutz:1996:DTC

[Sch96] Josef K. Schmutz. Disciplinary thinking in conservation biology. Conservation Biology, 10(3):698, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schultz:1998:DBI

[Sch98] Cheryl B. Schultz. Dispersal behavior and its implications for reserve design in a rare Oregon butterfly. Conservation Biology, 12(2):284–292, April 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Scudder:1999:ESP

[Scu99] G. G. E. Scudder. Endangered species protection in Canada. Conservation Biology, 13(5):963–965, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sanjayan:1996:GVI

[SCZF96] M. A. Sanjayan, Kevin Crooks, Gerard Zegers, and David Foran. Genetic variation and the immune response in natural populations of pocket gophers. Variación genetica y respuesta inmune en poblaciones naturales de topos. Conservation Biology, 10(6):1519–1527, December 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Silva:1994:ASM

[SD94] Marina Silva and John A. Downing. Allometric scaling of minimal mammal densities. Escala alométrica de la densidad mínima en mamiferos. Conservation Biology, 8(3):732–743, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Snyder:1996:LCB

[SDB^{+}96] Noel F. R. Snyder, Scott R. Derrickson, Steven R. Beissinger, James W. Wiley, Thomas B. Smith, William D. Toone, and Brian Miller. Limitations of captive breeding in endangered species recovery. Las limitaciones de la cría en cautiverio en la recuperación de especies en peligro de extinción. Conservation Biology, 10(2):338–348, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Snyder:1997:LCB

[SDB^{+}97] Noel F. R. Snyder, Scott R. Derrickson, Steven R. Beissinger, James W. Wiley, Thomas B. Smith, William D. Toone, and Brian Miller. Limitations of captive breeding: Reply to Gippoliti and Carpaneto. Conservation Biology, 11(3):808–810, June 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stoms:1993:GAC

[SDC^{+}93] David M. Stoms, Frank W. Davis, Christopher B. Cogan, Marco O. Painho, Brean W. Duncan, Joseph Scepan, and J. Michael Scott. Geographic analysis of California condor sighting data. Análisis geográfico de datos de avistamiento del condor de California. Conservation Biology, 7(1):148–159, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Salafsky:1993:CER

[SDT93] Nick Salafsky, Barbara L. Dugelby, and John W. Terborgh. Can extractive reserves save the rain forest? An ecological and socioeconomic comparison of nontimber forest product extraction systems in Petén, Guatemala, and West Kalimantan, Indonesia. `?Pueden las reservas extractivas salvar la selva tropical lluviosa? Conservation Biology, 7(1):39–52, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schneider:1998:TMA

[SEC98] Daniel W. Schneider, Christopher D. Ellis, and Kevin S. Cummings. A transportation model assessment of the risk to native mussel communities from zebra mussel spread. Conservation Biology, 12(4):788–800, August 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Semlitsch:1998:BDT

[Sem98] Raymond D. Semlitsch. Biological delineation of terrestrial buffer zones for pond-breeding salamanders. Conservation Biology, 12(5):1113–1119, October 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shrader-Frechette:1996:TBP

[SF96] Kristin Shrader-Frechette. Throwing out the bathwater of positivism, keeping the baby of objectivity: Relativism and advocacy in conservation biology. Conservation Biology, 10(3):912– 914, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Swart:1997:CIG

[SF97] M. K. J. Swart and J. W. H. Ferguson. Conservation implications of genetic differentiation in Southern African populations of black rhinoceros (Diceros bicornis). Implicaciones de la diferenciacion genética en la conservación de poblaciones del rinoceronte Negro Diceros bicornis en el Sur de Africa. Conservation Biology, 11(1):79–83, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Simberloff:1992:MCC

[SFCM92] Daniel Simberloff, James A. Farr, James Cox, and David W. Mehlman. Movement corridors: Conservation bargains or poor investments? Corredores para el movimiento: `?Gangas de la conservación o malas inversiones? Conservation Biology, 6(4): 493–504, December 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Smith:1995:ECE

[SFLC95] Thomas B. Smith, Leonard A. Freed, Jaan Kaimanu Lepson, and John H. Carothers. Evolutionary consequences of extinctions in populations of a Hawaiian honeycreeper consecuencias evolutivas de las extinciones en poblaciones de un “honeycreeper” hawaiano. Conservation Biology, 9(1):107–113, February 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Silvertown:1996:IEM

[SFM96a] Jonathan Silvertown, Miguel Franco, and Eric Menges. Interpretation of elasticity matrices as an aid to the management of plant populations for conservation. La interpretación de matrices de elasticidad como una ayuda para el manejo de la conservación de poblaciones de plantas. Conservation Biology, 10 (2):591–597, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sutherst:1996:PGD

[SFM96b] Robert W. Sutherst, Robert B. Floyd, and Gunter F. Maywald. The potential geographical distribution of the cane toad, Bufo marinus L. in Australia. Conservation Biology, 10(1):294–299, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shrader-Frechette:1999:MSE

[SFM99] Kristin Shrader-Frechette and Earl D. McCoy. Molecular systematics, ethics, and biological decision making under uncertainty. Sistemática molecular, etica y toma de decisiones biológicas bajo incertidumbre. Conservation Biology, 13(5): 1008–1012, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stevens:1993:AER

[SG93] David R. Stevens and Nike J. Goodson. Assessing effects of removals for transplanting on a high-elevation bighorn sheep population. Conservation Biology, 7(4):908–915, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sanz:1998:SRC

[SG98] Virginia Sanz and Alejandro Grajal. Successful reintroduction of captive-raised yellow-shouldered Amazon parrots on Margarita Island, Venezuela. Conservation Biology, 12(2):430–441, April 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Soltis:1999:MSC

[SG99] Pamela S. Soltis and Matthew A. Gitzendanner. Molecular systematics and the conservation of rare species. Sistemática molecular y conservación de especies raras. Conservation Biology, 13(3):471–483, June 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Slade:1998:ARR

[SGA98] Norman A. Slade, Richard Gomulkiewicz, and Helen M. Alexander. Alternatives to Robinson and Redford’s method of assessing overharvest from incomplete demographic data. Conservation Biology, 12(1):148–155, February 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Saberwal:1994:LHC

[SGCJ94] Vasant K. Saberwal, James P. Gibbs, Ravi Chellam, and A. J. T. Johnsingh. Lion-human conflict in the gir forest, India conflicto entre leones y humanos en el bosque de gir, India. Conservation Biology, 8(2):501–507, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Snucins:1995:RAT

[SGK95] E. J. Snucins, J. M. Gunn, and W. Keller. Restoration of the Aurora trout to its acid-damaged native habitat. Conservation Biology, 9(5):1307–1311, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Seamans:1999:DTM

[SGMP99] Mark E. Seamans, R. J. Gutiérrez, Christopher A. May, and M. Zachariah Peery. Demography of two Mexican spotted owl populations. Demografía de dos poblaciones de la lechuza moteada Mexicana. Conservation Biology, 13(4):744–754, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schwinghamer:1998:EEO

[SGR^{+}98] Peter Schwinghamer, Donald C. Gordon, Jr., Terence W. Rowell, Jens Prena, David L. McKeown, G. Sonnichsen, and J. Y. Guigné. Effects of experimental otter trawling on surficial sediment properties of a sandy-bottom ecosystem on the Grand Banks of Newfoundland. Conservation Biology, 12(6):1215– 1222, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Steadman:1990:ABP

[SGW90] David W. Steadman, Ellis C. Greiner, and Christopher S. Wood. Absence of blood parasites in indigenous and introduced birds from the Cook Islands, South Pacific. Conservation Biology, 4(4):398–404, December 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shaffer:1990:PVA

[Sha90] Mark L. Shaffer. Population viability analysis. Conservation Biology, 4(1):39–40, March 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shankman:1993:CMV

[Sha93] David Shankman. Channel migration and vegetation patterns in the southeastern coastal plain. Migración de canales y patrones de vegetación en el sudeste de la planicie costera. Conservation Biology, 7(1):176–183, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Saunders:1991:BCE

[SHM91] Denis A. Saunders, Richard J. Hobbs, and Chris R. Margules. Biological consequences of ecosystem fragmentation: a review. Conservation Biology, 5(1):18–32, March 1991. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Smith:1997:RBP

[SHM97] A. P. Smith, N. Horning, and D. Moore. Regional biodiversity planning and lemur conservation with GIS in Western Madagascar. Planeacion de la biodiversidad regional y conservación de lemures con SIG en Madagascar Occidental. Conservation Biology, 11(2):498–512, April 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sheffer:1997:FEG

[SHMV97] Ruby J. Sheffer, Philip W. Hedrick, W. L. Minckley, and Anthony L. Velasco. Fitness in the endangered Gila topminnow. Adaptabilidad del poecílido de Gila en peligro. Conservation Biology, 11(1):162–171, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stanley:1997:EMP

[SHS^{+}97] Thomas R. Stanley, Joel T. Heinen, James R. Strittholt, Philip M. Fearnside, and Gillian Cooper-Driver. 619 ecosystem management projects. Conservation Biology, 11(4):1037– 1041, August 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sibbing:1994:CWH

[Sib94] Julie M. Sibbing. Conservation of western hemisphere shorebirds. Conservation Biology, 8(2):333, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sidle:1998:ACS

[Sid98] John G. Sidle. Arbitrary and capricious species conservation. Conservation Biology, 12(1):248–249, February 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Siebert:1993:RER

[Sie93] Stephen F. Siebert. Rattan and extractive reserves. Conservation Biology, 7(4):749–750, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Siegel:1996:SVA

[Sie96] James J. Siegel. “Subdivisions versus agriculture”: From false assumptions come false alternatives. Conservation Biology, 10 (5):1473–1474, October 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Silander:1994:YWU

[Sil94] John A. Silander, Jr. You want to use GIS? Conservation Biology, 8(4):1167–1168, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Simberloff:1994:CPP

[Sim94] Daniel Simberloff. Conservation planning for Papua New Guinea. Conservation Biology, 8(4):1166–1167, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Simonetti:1995:WCO

[Sim95] Javier A. Simonetti. Wildlife conservation outside parks is a disease-mediated task. Conservation Biology, 9(2):454–456, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Simmons:1996:PDV

[Sim96] Robert E. Simmons. Population declines, viable breeding areas, and management options for flamingos in Southern Africa. Declinaciones poblacionales, áreas reproductivas viables y opciones de manejo para los flamingos del sur de Africa. Conservation Biology, 10(2):504–514, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Safford:1998:SLB

[SJ98] R. J. Safford and C. G. Jones. Strategies for land-bird conservation on Mauritius. Conservation Biology, 12(1):169–176, February 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sweitzer:1997:NEP

[SJB97] Richard A. Sweitzer, Stephen H. Jenkins, and Joel Berger. Near-extinction of porcupines by mountain lions and consequences of ecosystem change in the Great Basin desert. Conservation Biology, 11(6):1407–1417, December 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Skagen:1993:TCM

[SK93] Susan K. Skagen and Fritz L. Knopf. Toward conservation of midcontinental shorebird migrations. Hacia la conservación de las migraciones de aves costeras del continente medio. Conservation Biology, 7(3):533–541, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Saberwal:1996:HDC

[SK96] Vasant K. Saberwal and Ashish Kothari. The human dimension in conservation biology curricula in developing countries. Las dimensiones humanas en el curriculum de la biología de la conservación en los paises en vias de desarrollo. Conservation Biology, 10(5):1328–1331, October 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Snow:1999:RSC

[SK99] Neil Snow and Philip L. Keating. Relevance of specimen citations to conservation. Conservation Biology, 13(4):943–944, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schindler:1998:ECA

[SKOO98] Daniel E. Schindler, James F. Kitchell, and Richard Ogutu-Ohwayo. Ecological consequences of alternative gill net fisheries for Nile perch in Lake Victoria. Conservation Biology, 12(1): 56–64, February 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

SoberonM:1993:USA

[SL93] Jorge Soberón M. and Jorge Llorente B. The use of species accumulation functions for the prediction of species richness. El uso de funciones dc acumulación de especies para la prediccion de la rigueza de especies. Conservation Biology, 7(3):480–488, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schieck:1995:EPS

[SLNP95] Jim Schieck, Ken Lertzman, Brian Nyberg, and Rick Page. Effects of patch size on birds in old-growth montane forests. Conservation Biology, 9(5):1072–1084, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stangel:1992:GVP

[SLS92] Peter W. Stangel, Michael R. Lennartz, and Michael H. Smith. Genetic variation and population structure of red-cockaded woodpeckers. Conservation Biology, 6(2):283–292, June 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Smith:1991:CVL

[SM91] James L. David Smith and Charles Mcdougal. The contribution of variance in lifetime reproduction to effective population size in tigers. Conservation Biology, 5(4):484–490, December 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Skelly:1997:RBM

[SM97] David K. Skelly and Eli Meir. Rule-based models for evaluating mechanisms of distributional change. Modelos basados en reglas para la evaluación de mecanismos de cambio distribucional. Conservation Biology, 11(2):531–538, April 1997. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Salafsky:1999:TRA

[SM99] Nick Salafsky and Richard Margoluis. Threat reduction assessment: a practical and cost-effective approach to evaluating conservation and development projects. Evaluación de amenazas de reducción: una estimación práctica y costo-efectiva para evaluar proyectos de conservación y desarrollo. Conservation Biology, 13(4):830–841, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sherwin:1991:MGV

[SMGB91] William B. Sherwin, Neil D. Murray, Jennifer A. Marshall Graves, and Peter R. Brown. Measurement of genetic variation in endangered populations: Bandicoots (Marsupialia: Peramelidae) as an example. Conservation Biology, 5(1):103–108, March 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Skagen:1998:CUR

[SMHK98] Susan K. Skagen, Cynthia P. Melcher, William H. Howe, and Fritz L. Knopf. Comparative use of riparian corridors and oases by migrating birds in Southeast Arizona. Conservation Biology, 12(4):896–909, August 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Suzan:1999:EEP

[SMPN99] Humberto Suzán, Guadalupe Malda, Duncan T. Patten, and Gary P. Nabhan. Effects of exploitation and park boundaries on legume trees in the Sonoran desert. Efectos de la explotación y los límites de parques en leguminosas del desierto de Sonora. Conservation Biology, 13(6):1497–1501, December 1999. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Salwasser:1997:PVC

[SMS97] Hal Salwasser, Douglas W. MacCleery, and Thomas A. Snellgrove. The pollyannas vs. the chicken littles — enough already! Conservation Biology, 11(1):283–286, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stockwell:1996:TPA

[SMV96] Craig A. Stockwell, Margaret Mulvey, and Gary L. Vinyard. Translocations and the preservation of allelic diversity. Translocaciones y la preservación de la diversidad alélica. Conservation Biology, 10(4):1133–1141, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Suzan:1994:NPF

[SNP94] Humberto Suzán, Gary P. Nabhan, and Duncan T. Patten. Nurse plant and floral biology of a rare night-blooming cereus, Peniocereus striatus (Brandegee) F. Buxbaum. Nodricismo y biología floral de una rara reina de la noche, Peniocereus striatus (Brandegee) F. Buxbaum. Conservation Biology, 8(2):461– 470, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Seidl:1994:UFB

[SO94] Amy L. Seidl and Paul A. Opler. Uncompahgre fritillary butterfly demographics: Response to Britten et al. Conservation Biology, 8(4):1156–1157, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [BBM94b] and reply [BBM94c].

Song:1996:PVA

[Son96] Yan-Ling Song. Population viability analysis for two isolated populations of haianan Eld’s deer. Conservation Biology, 10(5): 1467–1472, October 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Soule:1990:OAS

[Sou90] Michael E. Soulé. The onslaught of alien species, and other challenges in the coming decades. Conservation Biology, 4(3):

233–240, September 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Soule:1991:TPF

[Sou91] Michael E. Soulé. The “Two Point Five Society”. Conservation Biology, 5(2):255, June 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

South:1999:DSE

[Sou99] Andy South. Dispersal in spatially explicit population models. Dispersión en modelos poblacionales espacialmente explícitos. Conservation Biology, 13(5):1039–1046, October 1999. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Smith:1990:CAD

[SP90] Andrew T. Smith and Mary M. Peacock. Conspecific attraction and the determination of metapopulation colonization rates. Conservation Biology, 4(3):320–323, September 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sosa:1998:EPR

[SP98] Victoria Sosa and Teodoro Platas. Extinction and persistence of rare orchids in Veracruz, Mexico. Conservation Biology, 12 (2):451–455, April 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sinclair:1998:PEP

[SPD^{+}98] A. R. E. Sinclair, R. P. Pech, C. R. Dickman, D. Hik, P. Mahon, and A. E. Newsome. Predicting effects of predation on conservation of endangered prey. Conservation Biology, 12(3):564– 575, June 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Spellerberg:1995:PCB

[Spe95] Ian F. Spellerberg. Putting conservation biology into practice. Conservation Biology, 9(6):1663–1664, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sauer:1996:ECP

[SPP96] John R. Sauer, Grey W. Pendleton, and Bruce G. Peterjohn. Evaluating causes of population change in North American insectivorous songbirds. Evaluando las causas de los cambios poblacionales de las aves cantoras insectívoras de Norte América. Conservation Biology, 10(2):465–478, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Suarez:1997:NSD

[SPR97] Andrew V. Suarez, Karin S. Pfennig, and Scott K. Robinson. Nesting success of a disturbance-dependent songbird on different kinds of edges. Exito de nidación de un ave paserina dependiente de disturbaciones en diferentes tipos de bordes. Conservation Biology, 11(4):928–935, August 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Starfield:1995:MHM

[SRR95] Anthony M. Starfield, James D. Roth, and Katherine Ralls. “Mobbing” in Hawaiian monk seals (Monachus schauinslani): The value of simulation modeling in the absence of apparently crucial data. “Mobbing” en las focas Hawaianas: el valor de la modelación de simulación en la ausencia de datos aparentemente cruciales. Conservation Biology, 9(1):166–174, February 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Swart:1996:AGV

[SRV96] Jacobus A. A. Swart, Peter J. H. Reijnders, and Wilke Van Delden. Absence of genetic variation in harbor seals (Phoca vitulina) in the Dutch Wadden Sea and the British Wash. Conservation Biology, 10(1):289–293, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stokes:1994:DIC

[SS94] Michael K. Stokes and Norman A. Slade. Drought-induced cracks in the soil as refuges for small mammals: an unforeseen consequence of climatic change. Conservation Biology, 8 (2):577–580, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Su:1995:GDS

[SS95] Bing Su and Li-Ming Shi. Genetic diversity in the snub-nosed monkey (Rhinopithecus bieti) as estimated by protein electrophoresis. Conservation Biology, 9(4):947–951, August 1995.

CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stewart:1998:CDO

[SS98] David A. B. Stewart and Michael J. Samways. Conserving dragonfly (odonata) assemblages relative to river dynamics in an African savanna game reserve. Conservation Biology, 12 (3):683–692, June 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Seddon:1999:GSS

[SS99a] Philip J. Seddon and Pritpal S. Soorae. Guidelines for subspecific substitutions in wildlife restoration projects. Conservation Biology, 13(1):177–184, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stratford:1999:LET

[SS99b] Jeffrey A. Stratford and Philip C. Stouffer. Local extinctions of terrestrial insectivorous birds in a fragmented landscape near Manaus, Brazil. Extinciones locales de aves insectívoras terrestres en un paisaje fragmentado cercano a Manaos, Brazil. Conservation Biology, 13(6):1416–1423, December 1999. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sparrow:1994:TGM

[SSEM94] Helen R. Sparrow, Thomas D. Sisk, Paul R. Ehrlich, and Dennis D. Murphy. Techniques and guidelines for monitoring neotropical butterflies. Técnicas y lineamientos generales para el monitoreo de mariposas neotropicales. Conservation Biology, 8(3):800–809, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stone:1994:BDB

[SSS94] Paul A. Stone, Howard L. Snell, and Heidi M. Snell. Behavioral diversity as biological diversity: Introduced cats and Lava lizard wariness. Conservation Biology, 8(2):569–573, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Soltis:1992:AVA

[SSTL92] Pamela S. Soltis, Douglas E. Soltis, Tracy L. Tucker, and Frank A. Lang. Allozyme variability is absent in the narrow endemic Bensoniella oregona (Saxifragaceae). Conservation Biology, 6(1):131–134, March 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Short:1994:TVM

[ST94] Jeff Short and Bruce Turner. A test of the vegetation mosaic hypothesis: a hypothesis to explain the decline and extinction of Australian mammals. Test de la hipóstesis del mosaico de vegetación — una hipótesis para explicar la declinación y extinción de los mamíferos australianos. Conservation Biology, 8 (2):439–449, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sipes:1995:RBR

[ST95] Sedonia D. Sipes and Vincent J. Tepedino. Reproductive biology of the rare orchid, Spiranthes diluvialis: Breeding system, pollination, and implications for conservation. Biología reproductiva de una orquidea rara, Spiranthes diluvialis: sistema reproductivo, polinización, y consecuencias para la conservación. Conservation Biology, 9(4):929–938, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sutcliffe:1996:OCA

[ST96] Odette L. Sutcliffe and Chris D. Thomas. Open corridors appear to facilitate dispersal by ringlet butterflies (Aphantopus hyperantus) between woodland clearings. Corredores abiertos y dispersión de mariposas de ringlet (Aphantopus hyperantus) entre claros de bosque. Conservation Biology, 10(5):1359–1365, October 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Standley:1994:INF

[Sta94] Lisa A. Standley. Important new flora. Conservation Biology, 8(3):912–913, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stanley:1995:EMA

[Sta95] Thomas R. Stanley, Jr. Ecosystem management and the arrogance of humanism. El manejo de los ecosistemas y la arrogancia del humanismo. Conservation Biology, 9(2):255–262, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stanley:1996:ETC

[Sta96] Thomas R. Stanley, Jr. European tools to consider. Conservation Biology, 10(4):1300–1301, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stapp:1998:RRP

[Sta98] Paul Stapp. A reevaluation of the role of prairie dogs in Great Plains grasslands. Conservation Biology, 12(6):1253–1259, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shogren:1999:WEM

[STA^{+}99] Jason F. Shogren, John Tschirhart, Terry Anderson, Amy Whritenour Ando, Steven R. Beissinger, David Brookshire, Gardner M. Brown, Jr., Don Coursey, Robert Innes, Stephen M. Meyer, and Stephan Polasky. Why economics matters for endangered species protection. Porque es importante la economía para la protección de especies en peligro de extinción. Conservation Biology, 13(6):1257–1261, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stevenson:1993:KMB

[Ste93] R. D. Stevenson. “Kill not the moth or butterfly for the last judgment draweth nigh.”. Conservation Biology, 7(2):445–446, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Steinberg:1998:NKG

[Ste98] Michael K. Steinberg. Neotropical kitchen gardens as a potential research landscape for conservation biologists. Conservation Biology, 12(5):1150–1152, October 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stinner:1993:EAA

[Sti93] Benjamin R. Stinner. An ecological approach to agriculture. Conservation Biology, 7(3):741–742, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stier:1997:SEW

[Sti97] S. Cord Stier. The sound of T. S. Eliot’s whimper. Conservation Biology, 11(4):1035–1036, August 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Scott:1995:SRE

[STM95] J. Michael Scott, Timothy H. Tear, and L. Scott Mills. Socioeconomics and the recovery of endangered species: Biological assessment in a political world. Conservation Biology, 9(1): 214–216, February 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Storrow:1994:WG

[Sto94] Tim Storrow. Working with the government. Conservation Biology, 8(1):324–325, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Storfer:1995:SCB

[Sto95] Andrew Storfer. The Society for Conservation Biology: Progress or stasis? Conservation Biology, 9(5):982–983, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stoner:1996:PII

[Sto96] Kathryn E. Stoner. Prevalence and intensity of intestinal parasites in mantled howling monkeys (Alouatta palliata) in Northeastern Costa Rica: Implications for conservation biology. La ocurrencia e intensidad de parásitos intestinales en los monos Congo (Allouata paliata) en el noreste de Costa Rica: su relevancia para la implicaciones biologia de la conservación. Conservation Biology, 10(2):539–546, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stokland:1997:REB

[Sto97] Jogeir N. Stokland. Representativeness and efficiency of bird and insect conservation in Norwegian boreal forest reserves. Representatividad y eficiencia en la conservación de aves e insectos en las reservas de bosque boreal de Noruega. Conservation Biology, 11(1):101–111, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Strier:1991:DCE

[Str91] Karen B. Strier. Demography and conservation of an endangered primate, Brachyteles arachnoids. Conservation Biology, 5(2):214–218, June 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Strong:1994:VAB

[Str94] Paul I. V. Strong. Viability analysis in biological evaluations. Conservation Biology, 8(4):927–928, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Strauss:1995:ERW

[Str95] Ben Strauss. Education in the real world. Conservation Biology, 9(6):1346–1348, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stroich:1996:CCS

[Str96] Sam Stroich, Jr. Clearcut consolidation a status quo solution. Conservation Biology, 10(2):315, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Struhsaker:1998:BPR

[Str98] Thomas T. Struhsaker. A Biologist’s perspective on the role of sustainable harvest in conservation. Conservation Biology, 12 (4):930–932, August 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Strayer:1999:SPP

[Str99] David L. Strayer. Statistical power of presence-absence data to detect population declines. Poder estadístico de la presenciaausencia de datos para detectar disminuciones poblacionales. Conservation Biology, 13(5):1034–1038, October 1999. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sumardja:1995:L

[Sum95] Effendy A. Sumardja. Letter. Conservation Biology, 9(5):978– 979, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sun:1996:EPS

[Sun96] Mei Sun. Effects of population size, mating system, and evolutionary origin on genetic diversity in Spiranthes sinensis and S. hongkongensis. Efectos del tamaño poblacional, sistema de fecundación y origen evolutivo en la diversidad genética de Spiranthes sinensis y S. hongkongensis. Conservation Biology, 10 (3):785–795, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Scheel:1996:GWS

[SVC96] D. Scheel, T. L. S. Vincent, and Guy N. Cameron. Global warming and the species richness of bats in Texas. El calentamiento global y la riqueza de especies de los murciélagos en Texas. Conservation Biology, 10(2):452–464, April 1996. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sjogren:1994:CGD

[SW94] Per Sjögren and Per-Ivan Wyöni. Conservation genetics and detection of rare alleles in finite populations. Conservation Biology, 8(1):267–270, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schmidt:1999:EEL

[SW99a] Kenneth A. Schmidt and Christopher J. Whelan. Effects of exotic Lonicera and Rhamnus on songbird nest predation. Efectos de las especies exóticas Lonicera y Rhamnus en la depredación de nidos de aves cantoras. Conservation Biology, 13(6):1502– 1506, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schmidt:1999:RIN

[SW99b] Kenneth A. Schmidt and Christopher J. Whelan. The relative impacts of nest predation and brood parasitism on seasonal fecundity in songbirds. Conservation Biology, 13(1):46– 57, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Swanson:1998:ARC

[Swa98] Bradley J. Swanson. Autocorrelated rates of change in animal populations and their relationship to precipitation. Conservation Biology, 12(4):801–808, August 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Seehausen:1997:PRC

[SWK^{+}97] Ole Seehausen, Frans Witte, Egid F. Katunzi, Jan Smits, and Niels Bouton. Patterns of the remnant cichlid fauna in Southern Lake Victoria. Patrones de la fauna de cíclidos remanentes en el Sur del Lago Victoria. Conservation Biology, 11(4):890–904, August 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schneider:1994:EDA

[SY94] Richard R. Schneider and Peter Yodzis. Extinction dynamics in the American marten (Martes americana). Dinámica de las extinciones en la marta americana (Martes americana). Conservation Biology, 8(4):1058–1068, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Symanski:1996:DHL

[Sym96] Richard Symanski. Dances with horses: Lessons from the environmental fringe danzas con caballos: Lecciones del limite ambiental. Conservation Biology, 10(3):708–712, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tiebout:1997:CCI

[TA97] Harry M. Tiebout III and Roger A. Anderson. A comparison of corridors and intrinsic connectivity to promote dispersal in transient successional landscapes. Comparación de corredores y conectividad intrínseca para promover la dispersión en paisajes sucesionales transitorios. Conservation Biology, 11(3):620–627, June 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Taylor:1999:HHN

[TA99] Robert V. Taylor and Steven K. Albert. Human hunting of nongame birds at Zuni, New Mexico, U.S.A. Caza de aves no cinegéticas por humanos en Zuni, Neuvo México, U.S.A. Conservation Biology, 13(6):1398–1403, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Taylor:1995:RUP

[Tay95] Barbara L. Taylor. The reliability of using population viability analysis for risk classification of species. Confiabilidad de el uso del análisis de viabilidad poblacional para clasificacion de riesgo de especies. Conservation Biology, 9(3):551–558, June 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Trail:1993:IBH

[TB93] Pepper W. Trail and Luis F. Baptista. The impact of brownheaded cowbird parasitism on populations of the white-crowned sparrow. El impacto del parasitismo del “Brown-headed Cowbird” sobre poblaciones del “Nuttall’s white-crowned sparrow”. Conservation Biology, 7(2):309–315, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tiebout:1995:ERA

[TB95] Harry M. Tiebout III and Kristin E. Brugger. Ecological risk assessment of pesticides for terrestrial vertebrates: Evaluation and application of the U.S. Environmental Protection Agency’s quotient model. Evaluación del riesgo ecológico de los pesticidas para los vertebrados terrestres: evaluación y aplicación del modelo del cociente de USEPA. Conservation Biology, 9(6): 1605–1618, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tracy:1996:ISC

[TB96] C. Richard Tracy and Peter F. Brussard. The importance of science in conservation biology. Conservation Biology, 10(3):918– 919, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Todd:1998:ATC

[TB98] Charles R. Todd and Mark A. Burgman. Assessment of threat and conservation priorities under realistic levels of uncertainty and reliability. Conservation Biology, 12(5):966–974, October 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tomiuk:1997:GDL

[TBL^{+}97] Jürgen Tomiuk, Lutz Bachmann, Michael Leipoldt, Jörg U. Ganzhorn, Reiner Ries, Michaela Weis, and Volker Loeschcke. Genetic diversity of Lepilemur mustelinus ruficaudatus, a nocturnal lemur of Madagascar. Diversidad genética de Lepilemur mustelinus ruficaudatus, un lemur nocturno de Madagascar. Conservation Biology, 11(2):491–497, April 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Trombulak:1998:GU

[TBW^{+}98] Stephen C. Trombulak, Susan Bratton, Mark H. Waugh, David Pimentel, and James M. Peek. The Green University. Conservation Biology, 12(5):1158–1162, October 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tomback:1995:EBR

[TCK^{+}95] Diana F. Tomback, Jane Kees Clary, James Koehler, Raymond J. Hoff, and Stephen F. Arno. The effects of blister rust on post-fire regeneration of whitebark pine: The Sundance burn of Northern Idaho (U.S.A.). Los efectos de Cronartium ribicola en la regeneración de Pinus albicaulis posterior al incendio: El área quemada de “Sundance” en el Norte de Idaho. Conservation Biology, 9(3):654–664, June 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Turner:1996:CPS

[TCO^{+}96] I. M. Turner, K. S. Chua, J. S. Y. Ong, B. C. Soong, and H. T. W. Tan. A century of plant species loss from an isolated fragment of lowland tropical rain forest. Un siglo de pérdida de especies de plantas en un fragmento aislado de bosque tropical lluvioso en tierras bajas. Conservation Biology, 10(4):1229– 1244, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Touval:1994:PTC

[TD94] Jerome L. Touval and James M. Dietz. The problem of teaching conservation problem solving. Conservation Biology, 8(3): 902–904, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Taylor:1996:NEP

[TD96] Barbara L. Taylor and Andrew E. Dizon. The need to estimate power to link genetics and demography for conservation. Conservation Biology, 10(2):661–664, April 1996. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Temple:1990:NNE

[Tem90] Stanley A. Temple. The nasty necessity: Eradicating exotics. Conservation Biology, 4(2):113–115, June 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Temple:1991:NS

[Tem91] Stanley A. Temple. News of the Society. Conservation Biology, 5(4):434, December 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Temple:1992:CBW

[Tem92a] Stanley A. Temple. Conservation biologists and wildlife managers getting together. Conservation Biology, 6(1):4, March 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Temple:1992:HAM

[Tem92b] Stanley A. Temple. Highlights of the 1992 annual meeting. Conservation Biology, 6(3):312–313, September 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Temple:1992:SCB

[Tem92c] Stanley A. Temple. Is the Society for Conservation Biology having an impact? Conservation Biology, 6(2):160, June 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Temple:1992:SNE

[Tem92d] Stanley A. Temple. The search for a new Editor. Conservation Biology, 6(4):485, December 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Temple:1993:CFS

[Tem93a] Stanley A. Temple. A campaign to fund SCB’s public affairs programs. Conservation Biology, 7(1):4, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Temple:1993:FIT

[Tem93b] Stanley A. Temple. Finally, an introductory textbook! Conservation Biology, 7(4):961–962, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Temple:1993:RNE

[Tem93c] Stanley A. Temple. Reed Noss to edit Conservation Biology. Conservation Biology, 7(2):218, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Templeton:1994:BP

[Tem94] Alan R. Templeton. In and out of breeding problems. Conservation Biology, 8(2):608–610, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tella:1998:CBL

[TFHD98] José L. Tella, Manuela G. Forero, Fernando Hiraldo, and José A. Donázar. Conflicts between lesser kestrel conservation and European agricultural policies as identified by habitat use analyses. Conservation Biology, 12(3):593–604, June 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Taylor:1993:USP

[TG93] Barbara L. Taylor and Tim Gerrodette. The uses of statistical power in conservation biology: The vaquita and northern spotted owl. Los usos del poder estadístico en conservación biológica: la vaquita y el búho moteado del Norte. Conservation Biology, 7(3):489–500, September 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thiollay:1992:ISL

[Thi92] Jean-Marc Thiollay. Influence of selective logging on bird species diversity in a guianan rain forest. Conservation Biology, 6(1):47–63, March 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thiollay:1995:RTA

[Thi95] Jean-Marc Thiollay. The role of traditional agroforests in the conservation of rain forest bird diversity in Sumatra. El rol de los sistemas agroforestales tradicionales en la conservación de la diversidad de los pájaros de la selva tropical en Sumatra. Conservation Biology, 9(2):335–353, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thomas:1990:WDR

[Tho90] C. D. Thomas. What do real population dynamics tell us about minimum viable population sizes? Conservation Biology, 4(3): 324–327, September 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thomas:1991:CSB

[Tho91] Jack Ward Thomas. Comments on the Sustainable Biosphere Initiative. Conservation Biology, 5(4):552, December 1991. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See comments [GM91, Sch91, Wil91].

Thompson:1993:SRF

[Tho93] Frank R. Thompson III. Simulated responses of a forest-interior bird population to forest management options in central hardwood forests of the United States. Simulación de la respuesta de poblaciones de pájaros del interior de forestas a opciones en el manejo de forestas en bosques de maderas duras del centro de los Estados Unidos. Conservation Biology, 7(2):325–333, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thomas:1994:ECM

[Tho94a] C. D. Thomas. Extinction, colonization, and metapopulations: Environmental tracking by rare species. Extinción, colonización y metapoblaciones: adecuación ambiental de especies raras. Conservation Biology, 8(2):373–378, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thorne:1994:EUE

[Tho94b] K. Tom Thorne. Evaluating the U.S. Endangered Species Act. Conservation Biology, 8(3):914–916, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thorpe:1994:PTL

[Tho94c] John E. Thorpe. Performance thresholds and life-history flexibility in salmonids. Conservation Biology, 8(3):877–879, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thomas:1997:RPA

[Tho97] Len Thomas. Retrospective power analysis. Conservation Biology, 11(1):276–280, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thorbjarnarson:1999:CTS

[Tho99] John Thorbjarnarson. Crocodile tears and skins: International trade, economic constraints, and limits to the sustainable use of crocodilians. Conservation Biology, 13(3):465–470, June 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thompson:1999:HPD

[TJ99] Ken Thompson and Allan Jones. Human population density and prediction of local plant extinction in Britain. Conservation Biology, 13(1):185–189, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tomiuk:1994:ABD

[TL94] Jürgen Tomiuk and Volker Loeschcke. On the application of birth-death models in conservation biology. Conservation Biology, 8(2):574–576, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tyler:1998:IBI

[TLG^{+}98] Torrey Tyler, William J. Liss, Lisa M. Ganio, Gary L. Larson, Robert Hoffman, Elisabeth Deimling, and Gregg Lomnicky. Interaction between introduced trout and larval salamanders (Ambystoma macrodactylum) in high-elevation lakes. Conservation Biology, 12(1):94–105, February 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thomas:1996:IAM

[TM96] Len Thomas and Kathy Martin. The importance of analysis method for breeding bird survey population trend estimates. El efecto del método de análisis sobre los datos de las tendencias poblacionales obtenidos a partir de las evaluaciones de aves reproductoras. Conservation Biology, 10(2):479–490, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tilson:1998:L

[TN98] Ronald Tilson and Philip Nyhus. Letter. Conservation Biology, 12(2):261–262, April 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tewksbury:1999:SCW

[TNN^{+}99] Joshua Jordan Tewksbury, Gary Paul Nabhan, Donald Norman, Humberto Suzán, John Tuxill, and Jim Donovan. In situ conservation of wild chiles and their biotic associates. Conservation Biology, 13(1):98–107, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tomlinson:1992:DPR

[Tom92] P. B. Tomlinson. Deforestation provides a renewable resource. Conservation Biology, 6(2):306–307, June 1992. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tremaine:1993:VTR

[Tre93] Richard Tremaine. Valuing tropical rainforests. Conservation Biology, 7(1):7–8, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Trine:1998:WTP

[Tri98] Cheryl L. Trine. Wood thrush population sinks and implications for the scale of regional conservation strategies. Conservation Biology, 12(3):576–585, June 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Trombulak:1994:UEN

[Tro94] Stephen C. Trombulak. Undergraduate education and the next generation of conservation biologists. Conservation Biology, 8 (2):589–591, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Palazzo:1994:CV

[Tru94] José Truda Palazzo, Jr. Conservation values. Conservation Biology, 8(4):925–926, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See response [Vid94].

Tscharntke:1992:FPH

[Tsc92] Teja Tscharntke. Fragmentation of Phragmites habitats, minimum viable population size, habitat suitability, and local extinction of moths, midges, flies, aphids, and birds. Fragmentación de los ambientes de phragmites, mínimo tamaño viable de la población, adecuación del hábitat y extinción local de polillas, jejenes, moscas, áfidos y pájaros. Conservation Biology, 6(4):530–536, December 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tear:1995:RPE

[TSHG95] Timothy H. Tear, J. Michael Scott, Patricia H. Hayward, and Brad Griffith. Recovery plans and the Endangered Species Act: Are criticisms supported by data? Los planes de recuperación y el Acta de Especies en Peligro de Extinción: `?Están las críticas sustentadas por los datos? Conservation Biology, 9(1):182–195, February 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Timm:1997:HDE

[TSP97] Robert M. Timm, Rosa M. Salazar, and A. Townsend Peterson. Historical distribution of the extinct tropical seal, Monachus tropicalis (Carnivora: Phocidae). Conservation Biology, 11 (2):549–551, April 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Taberlet:1995:LCZ

[TSSB95] Pierre Taberlet, Jon E. Swenson, Finn Sandegren, and Anders Bjärvall. Localization of a contact zone between two highly divergent mitochondrial DNA lineages of the brown bear Ursus arctos in Scandinavia. Conservation Biology, 9(5):1255–1261, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tsubaki:1998:FAO

[Tsu98] Yoshitaka Tsubaki. Fluctuating asymmetry of the oriental fruit fly (Dacus dorsalis) during the process of its extinction from the Okinawa Islands. Conservation Biology, 12(4):926–929, August 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Turner:1999:TEF

[TTJ^{+}99] Andrew M. Turner, Joel C. Trexler, C. Frank Jordan, Sarah J. Slack, Pamela Geddes, John H. Chick, and William F. Loftus. Targeting ecosystem features for conservation: Standing crops in the Florida Everglades. La conservación enfocada a las características del ecosistema: Productividad en los Everglades de Florida. Conservation Biology, 13(4):898–911, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Turner:1994:SPS

[TTW^{+}94] I. M. Turner, H. T. W. Tan, Y. C. Wee, Ali Bin Ibrahim, P. T. Chew, and R. T. Corlett. A study of plant species extinction in Singapore: Lessons for the conservation of tropical biodiversity. Estudio de la extinción de especies vegetales en singapur: Lecciones para la conservación de la biodiversidad tropical. Conservation Biology, 8(3):705–712, September 1994.

CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Turner:1993:BCT

[Tur93] R. Eugene Turner. Brazilian crocodilian tears. Conservation Biology, 7(2):225, June 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thorbjarnarson:1999:EIM

[TV99] John Thorbjarnarson and Alvaro Velasco. Economic incentives for management of Venezuelan caiman. Incentivos económicos para el manejo del caimán de Venezuela. Conservation Biology, 13(2):397–406, April 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Townsend:1992:AER

[TW92a] Colin R. Townsend and Michael J. Winterbourn. Assessment of the environmental risk posed by an exotic fish: The proposed introduction of channel catfish (Ictalurus punctatus) to New Zealand. Conservation Biology, 6(2):273–282, June 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tyser:1992:AFG

[TW92b] Robin W. Tyser and Christopher A. Worley. Alien flora in grasslands adjacent to road and trail corridors in Glacier National Park, Montana (U.S.A.). Conservation Biology, 6(2): 253–262, June 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tutin:1997:URF

[TWMM97] C. E. G. Tutin, L. J. T. White, and A. Mackanga-Missandzou. The use by rain forest mammals of natural forest fragments in an equatorial African savanna. Utilización de fragmentos de bosque natural por mamíferos de selva lluviosa en una sabana ecuatorial africana. Conservation Biology, 11(5):1190–1203, October 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Utter:1994:PRA

[UA94] Fred M. Utter and Fred W. Allendorf. Phylogenetic relationships among species of Oncorhynchus: a consensus view.

Conservation Biology, 8(3):864–867, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Usher:1992:AIG

[UCB92] Michael B. Usher, Terence J. Crawford, and Jean L. Banwell. An American invasion of Great Britain: The case of the native and alien squirrel (Sciurus) species. Conservation Biology, 6 (1):108–115, March 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Uhl:1998:CBY

[Uhl98a] Christopher Uhl. Conservation biology in your own front yard. Conservation Biology, 12(6):1175–1177, December 1998. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Uhl:1998:PWH

[Uhl98b] Christopher Uhl. Perspectives on wildfire in the humid tropics. Conservation Biology, 12(5):942–943, October 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Uhl:1996:STC

[UKG^{+}96] Christopher Uhl, Dominik Kulakowski, Jeffrey Gerwing, Michelle Brown, and Mark Cochrane. Sustainability: a touchstone concept for university operations, education, and research. Conservation Biology, 10(5):1308–1311, October 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vitt:1998:IIT

[VAPCO98] Laurie J. Vitt, Teresa C. S. Avila-Pires, Janalee P. Caldwell, and Veronica R. L. Oliveira. The impact of individual tree harvesting on thermal environments of lizards in Amazonian rain forest. Conservation Biology, 12(3):654–664, June 1998. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

VanVuren:1999:RDI

[VB99] Dirk Van Vuren and Lizabeth Bowen. Reduced defenses in insular endemic plants: an evolutionary time frame. Conservation Biology, 13(1):211–212, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vucetich:1999:EIS

[VC99] John A. Vucetich and Scott Creel. Ecological interactions, social organization, and extinction risk in African wild dogs. Interacciones ecológicas, organización social, y riesgo de extinción de perros silvestres africanos. Conservation Biology, 13 (5):1172–1182, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vogler:1994:DUC

[VD94] Alfried P. Vogler and Rob Desalle. Diagnosing units of conservation management. Diagnosticando las unidades de manejo de conservación. Conservation Biology, 8(2):354–363, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vermeij:1993:BRE

[Ver93] Geerat J. Vermeij. Biogeography of recently extinct marine species: Implications for conservation. Biogeografia de especies marinas recientemente extinguidas: implicaciones para la conservación. Conservation Biology, 7(2):391–397, June 1993. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vickery:1994:EHA

[VHM94] Peter D. Vickery, Malcolm L. Hunter, Jr., and Scott M. Melvin. Effects of habitat area on the distribution of grassland birds in Maine. Los efectos de la superficie del hábitat en la distribución de los pájaros de pastizales en Maine. Conservation Biology, 8 (4):1087–1097, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Voris:1995:FAA

[VI95] Harold K. Voris and Robert F. Inger. Frog abundance along streams in bornean forests. Conservation Biology, 9(3):679– 683, June 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vidal:1993:AMC

[Vid93] Omar Vidal. Aquatic mammal conservation in Latin America: Problems and perspectives. Conservation Biology, 7(4): 788–795, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vidal:1994:AEI

[Vid94] Omar Vidal. Awareness and education instead of sanctions: a response to Truda Palazzo. Conservation Biology, 8(4):926– 927, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [Tru94].

Villasenor:1998:SCA

[VIO98] Jose Luis Villaseñor, Guillermo Ibarra, and Daniel Ocaña. Strategies for the conservation of Asteraceae in Mexico. Conservation Biology, 12(5):1066–1075, October 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Verissimo:1998:ZTE

[VJSU98] Adalberto Veríssimo, Carlos Souza Júnior, Steve Stone, and Christopher Uhl. Zoning of timber extraction in the Brazilian Amazon. Conservation Biology, 12(1):128–136, February 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vrijenhoek:1991:LTB

[VL91] Robert C. Vrijenhoek and Paul L. Leberg. Let’s not throw the baby out with the bathwater: a comment on management for MHC diversity in captive populations. Conservation Biology, 5(2):252–254, June 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [Hug91].

Valutis:1999:APR

[VM99] Laura L. Valutis and John M. Marzluff. The appropriateness of puppet-rearing birds for reintroduction. Conveniencia de la cría de aves con títeres para reintroducción. Conservation Biology, 13(3):584–591, June 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Veblen:1992:EII

[VMMK92] Thomas T. Veblen, Monica Mermoz, Carlos Martin, and Thomas Kitzberger. Ecological impacts of introduced animals in Nahuel Huapi National Park, Argentina. Conservation Biology, 6(1):71–83, March 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vogelmann:1995:AFF

[Vog95] James E. Vogelmann. Assessment of forest fragmentation in Southern New England using remote sensing and geographic information systems technology. Evaluación de la fragmentación del bosque en el sur de nueva inglaterra utilizando tecnologia de sensores remotos y sistemas de información geográficos. Conservation Biology, 9(2):439–449, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vandermeer:1997:ANC

[VP97] John Vandermeer and Ivette Perfecto. The agroecosystem: a need for the conservation Biologist’s lens. Conservation Biology, 11(3):591–592, June 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vucetich:1997:ESS

[VPW97] John A. Vucetich, Rolf O. Peterson, and Thomas A. Waite. Effects of social structure and prey dynamics on extinction risk in gray wolves. Efectos de la estructura social y dinámica de las presas sobre el riesgo de extinción de lobos grises. Conservation Biology, 11(4):957–965, August 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Virkkala:1994:SPA

[VRV^{+}94] Raimo Virkkala, Ari Rajasärkkä, Risto A. Väisänen, Markku Vickholm, and Erkki Virolainen. The significance of protected areas for the land birds of Southern Finland. La importancia de las áreas protegidas para las aves terrestres del sur de Finlandia. Conservation Biology, 8(2):532–544, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

VanRooy:1995:EIE

[VS95] Peter T. J. C. Van Rooy and Anton H. P. Stumpel. Ecological impact of economic development on sardinian herpetofauna. Impacto ecológico del desarrollo económico sobre la herpetofauna de Cerdeña. Conservation Biology, 9(2):263–269, April 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vaughn:1999:IDF

[VT99] Caryn C. Vaughn and Christopher M. Taylor. Impoundments and the decline of freshwater mussels: a case study of an extinction gradient. Represamiento y declive de mejillones de agua dulce: El caso de estudio de un gradiente de extinción. Conservation Biology, 13(4):912–920, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Villard:1999:FEF

[VTM99] Marc-André Villard, M. Kurtis Trzcinski, and Gray Merriam. Fragmentation effects on forest birds: Relative influence of woodland cover and configuration on landscape occupancy. Efectos de la fragmentación sobre aves de bosque: Importancia relativa de la cobertura y configuración forestal. Conservation Biology, 13(4):774–783, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

VanDierendonck:1996:URE

[VW96] Machteld C. Van Dierendonck and Michiel F. Wallis de Vries. Ungulate reintroductions: Experiences with the takhi or Przewalski horse (Equus ferus przewalskii) in Mongolia. Reintroducción de especies: Experiencias con el takhi (Equus ferus przewalskii) en Mongolia. Conservation Biology, 10(3):728– 740, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vucetich:1998:NCR

[VW98] John A. Vucetich and Thomas A. Waite. Number of censuses required for demographic estimation of effective population size. Conservation Biology, 12(5):1023–1030, October 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vila:1999:HBW

[VW99a] Carles Vilà and Robert K. Wayne. Hybridization between wolves and dogs. Conservation Biology, 13(1):195–198, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vucetich:1999:EHF

[VW99b] John A. Vucetich and Thomas A. Waite. Erosion of heterozygosity in fluctuating populations. Erosión de la heterocigosidad en poblacione fluctuantes. Conservation Biology, 13(4):860– 868, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Winston:1995:ACV

[WA95] Matthew R. Winston and Paul L. Angermeier. Assessing conservation value using centers of population density evaluando el valor conservacional utilizando centros de densidad de poblacion. Conservation Biology, 9(6):1518–1527, December 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Walsh:1995:PTP

[WABH95] Peter D. Walsh, H. Resit Akçakaya, Mark Burgham, and Alexander S. Harcourt. PVA in theory and practice. Conservation Biology, 9(4):704–708, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Walter:1990:SVP

[Wal90] Hartmut S. Walter. Small viable population: The red-tailed hawk of Socorro Island. Conservation Biology, 4(4):441–443, December 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Walker:1992:BER

[Wal92] Brian H. Walker. Biodiversity and ecological redundancy. Conservation Biology, 6(1):18–23, March 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Walker:1995:CBD

[Wal95] Brian Walker. Conserving biological diversity through ecosystem resilience. La conservación de la diversidad biológica a través de la resiliencia de los ecosistemas. Conservation Biology, 9(4):747–752, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Walsh:1996:ACU

[Wal96] Stephen J. Walsh. Aquatic conservation units. Conservation Biology, 10(6):1566–1567, December 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Walker:1999:EAC

[Wal99a] Brian Walker. The ecosystem approach to conservation: Reply to Goldstein. Conservation Biology, 13(2):436–437, April 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [Gol99b, Gol99a].

Wallace:1999:EET

[Wal99b] S. Scott Wallace. Evaluating the effects of three forms of marine reserve on northern abalone populations in British Columbia, Canada. Evaluación del efecto de tres tipos de reservas marinas sobre poblaciones de abulón del norte en Columbia Británica, Canadá. Conservation Biology, 13(4):882–887, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Waples:1994:GCR

[Wap94] Robin S. Waples. Genetic considerations in recovery efforts for Pacific salmon. Conservation Biology, 8(3):884–886, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Waples:1998:ESU

[Wap98] Robin S. Waples. Evolutionarily significant units, distinct population segments, and the Endangered Species Act: Reply to Pennock and Dimmick. Conservation Biology, 12(3):718–721, June 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [PD97].

Warner:1994:ALU

[War94a] Richard E. Warner. Agricultural land use and grassland habitat in Illinois: Future shock for Midwestern birds? Uso agícola de la tierra y hábitat del pastizal en Illinois: Futura conmoción para las aves del Oeste Medio? Conservation Biology, 8(1):147– 156, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Warren:1994:DMC

[War94b] Peter Warren. Drawing management conclusions in conservation biology. Conservation Biology, 8(4):928, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Warshall:1994:BMG

[War94c] Peter Warshall. The biopolitics of the Mt. Graham red squirrel (Tamiasciuris hudsonicus grahamensis). La biopolítica de la ardilla roja (Tamiasciuris hudsonicus grahamensis) del Monte Graham. Conservation Biology, 8(4):977–988, December 1994.

CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Warshall:1996:AAM

[War96] Peter Warshall. Astronomy and animals on Mt. Graham. Conservation Biology, 10(5):1479–1480, October 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Washitani:1996:PGC

[Was96] Izumi Washitani. Predicted genetic consequences of strong fertility selection due to pollinator loss in an isolated population of Primula sieboldii. Las concecuencias genéticas predecibles de la fuerte selección de fertilidad debido a la pérdida del polinizador en una población aislada de Primula sieboldii, un especie heterostilica en peligro. Conservation Biology, 10(1):59– 64, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wayne:1992:UMM

[Way92] Robert K. Wayne. On the use of morphologic and molecular genetic characters to investigate species status. Conservation Biology, 6(4):590–592, December 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wasser:1997:NPM

[WBKH97] Samuel K. Wasser, Kenneth Bevis, Gina King, and Eric Hanson. Noninvasive physiological measures of disturbance in the northern spotted owl. Conservation Biology, 11(4):1019–1022, August 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilcove:1998:MCE

[WC98] David S. Wilcove and Linus Y. Chen. Management costs for endangered species. Conservation Biology, 12(6):1405–1407, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Whitehead:1997:PDW

[WCD97] Hal Whitehead, Jenny Christal, and Susan Dufault. Past and distant whaling and the rapid decline of sperm whales off the Galápagos Islands. Conservation Biology, 11(6):1387–1396, December 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilkie:1998:MSS

[WCTM98] David S. Wilkie, Bryan Curran, Richard Tshombe, and Gilda A. Morelli. Modeling the sustainability of subsistence farming and hunting in the Ituri Forest of Zaire. Conservation Biology, 12(1):137–147, February 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wikramanayake:1998:EBM

[WDR^{+}98] Eric D. Wikramanayake, Eric Dinerstein, John G. Robinson, Ullas Karanth, Alan Rabinowitz, David Olson, Thomas Mathew, Prashant Hedao, Melissa Conner, Ginette Hemley, and Dorene Bolze. An ecology-based method for defining priorities for large mammal conservation: The tiger as case study. Conservation Biology, 12(4):865–878, August 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Willson:1994:ACF

[WDSA94] Mary F. Willson, Toni L. De Santo, Carlos Sabag, and Juan J. Armesto. Avian communities of fragmented south-temperate rainforests in Chile comunidades de aves en bosques fragmentados templados lluviosos del sur de Chile. Conservation Biology, 8(2):508–520, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Walters:1994:RCW

[WDWG94] Terrence W. Walters, Deena S. Decker-Walters, and Doria R. Gordon. Restoration considerations for wiregrass (Aristida stricta): Allozymic diversity of populations. Conservation Biology, 8(2):581–585, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Weaver:1995:ISS

[Wea95] Jan C. Weaver. Indicator species and scale of observation. Conservation Biology, 9(4):939–942, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Webb:1999:EDH

[Web99] Campbell O. Webb. Environment as destiny, history as science. Conservation Biology, 13(6):1520–1521, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wehausen:1999:REM

[Weh99] John D. Wehausen. Rapid extinction of mountain sheep populations revisited. Revisión de la extinción rápida de poblaciones le la oveja de montaña. Conservation Biology, 13(2):378–384, April 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Weir:1992:SRR

[Wei92] Jack Weir. The Sweetwater Rattlesnake Round-Up: a case study in environmental ethics. Conservation Biology, 6(1):116– 127, March 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Weiss:1999:CCC

[Wei99] Stuart B. Weiss. Cars, cows, and checkerspot butterflies: Nitrogen deposition and management of nutrient-poor grasslands for a threatened species. Carros, vacas, y mariposas: Deposición de nitrógeno y manejo de pastisales pobres en nitrógeno para una especie amenazada. Conservation Biology, 13(6):1476–1486, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Welsh:1990:RAO

[Wel90] Hartwell H. Welsh, Jr. Relictual amphibians and old-growth forests. Conservation Biology, 4(3):309–319, September 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Winker:1997:PML

[WER^{+}97] Kevin Winker, Patricia Escalante, John H. Rappole, Mario A. Ramos, Richard J. Oehlenschlager, and Dwain W. Warner. Periodic migration and lowland forest refugia in a “Sedentary” neotropical bird, Wetmore’s bush-tanager. Migración esporádica y refugios en bosques de llanuras en una especie sedentaria de ave neotropical, el chinchinero comú. Conservation Biology, 11(3):692–697, June 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Westman:1990:PME

[Wes90] Walter E. Westman. Park management of exotic plant species: Problems and issues. Conservation Biology, 4(3):251–260, September 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Western:1991:BCM

[Wes91] David Western. Biology and conservation: Making the relevant connection. Conservation Biology, 5(4):431–433, December 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Winter:1999:PAS

[WF99a] Maiken Winter and John Faaborg. Patterns of area sensitivity in grassland-nesting birds patrones de sensitividad de area en aves que nidan en pastizales. Conservation Biology, 13(6): 1424–1436, December 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Woinarski:1999:AES

[WF99b] J. C. Z. Woinarski and Alaric Fisher. The Australian Endangered Species Protection Act 1992. Conservation Biology, 13 (5):959–962, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Williams:1996:CRH

[WGM^{+}96] Paul Williams, David Gibbons, Chris Margules, Anthony Rebelo, Chris Humphries, and Robert Pressey. A comparison of richness hotspots, rarity hotspots, and complementary areas for conserving diversity of British birds. Comparación de los puntos máximos de riqueza de especies, máxima concentración de especies raras y areas complementarias para conservar la diversidad de aves Británicas. Conservation Biology, 10(1):155–174, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Warkentin:1995:SUG

[WGO95] Ian G. Warkentin, Russel Greenberg, and Javier Salgado Ortiz. Songbird use of gallery woodlands in recently cleared and older settled landscapes of the selva lacandona, Chiapas, Mexico. Conservation Biology, 9(5):1095–1106, October 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wolf:1996:AMT

[WGRT96] C. Magdalena Wolf, Brad Griffith, Christine Reed, and Stanley A. Temple. Avian and mammalian translocations: Update and reanalysis of 1987 survey data. Translocaciones de aves y mamíferos: Actualizaceón y re-análisis de datos generados. Conservation Biology, 10(4):1142–1154, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Willson:1995:AFK

[WH95] Mary F. Willson and Karl C. Halupka. Anadromous fish as Keystone species in vertebrate communities. Peces anádromos como especies claves en las comunidades de vertebrados. Conservation Biology, 9(3):489–497, June 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wiersma:1999:LNF

[Wie99] Yolanda Wiersma. Lessons from northwest forests. Conservation Biology, 13(5):1232–1233, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wikramanayake:1990:CER

[Wik90] Eric D. Wikramanayake. Conservation of endemic rain forest fishes of Sri Lanka: Results of a translocation experiment. Conservation Biology, 4(1):32–37, March 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilcove:1991:CSB

[Wil91] David S. Wilcove. Comments on the Sustainable Biosphere Initiative. Conservation Biology, 5(4):553, December 1991. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See comments [GM91, Sch91, Tho91].

Willers:1992:TSL

[Wil92] Bill Willers. Toward a science of letting things be. Conservation Biology, 6(4):605–607, December 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilkes:1993:VPP

[Wil93a] Garrison H. Wilkes. N. I. Vavilov — patriarch of plant genetic resources. Conservation Biology, 7(4):962–963, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Willers:1993:EVM

[Wil93b] Bill Willers. Ecocentrism versus management: a reply to Blumstein. Conservation Biology, 7(3):452–453, September 1993.

CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Willers:1994:SDN

[Wil94] Bill Willers. Sustainable development: a new world deception. Conservation Biology, 8(4):1146–1148, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Willers:1996:WEE

[Wil96] Bill Willers. War on ecology and environmentalism. Conservation Biology, 10(6):1563–1564, December 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Windsor:1995:ERP

[Win95] Donald A. Windsor. Equal rights for parasites. Conservation Biology, 9(1):1–2, February 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Winker:1996:CIB

[Win96a] Kevin Winker. The crumbling infrastructure of biodiversity: The avian example. El desmoronamiento de la infraestructura de la biodiversidad: El ejemplo de las aves. Conservation Biology, 10(3):703–707, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Winker:1996:SSV

[Win96b] Kevin Winker. Specimen shrinkage versus evolution: I’iwi morphology. Conservation Biology, 10(2):657–658, April 1996. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See rejoinder [FSCL96].

Windsor:1998:ESA

[Win98] Donald A. Windsor. The Endangered Species Act is analogous to Schindler’s list. Conservation Biology, 12(2):485–486, April 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

With:1997:ANL

[Wit97a] Kimberly A. With. The application of neutral landscape models in conservation biology. Aplicación de modelos de paisaje neutros en la biología de la conservación. Conservation Biology, 11(5):1069–1080, October 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

With:1997:TCB

[Wit97b] Kimberly A. With. The theory of conservation biology. Conservation Biology, 11(6):1436–1440, December 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

With:1999:ETS

[WK99] Kimberly A. With and Anthony W. King. Extinction thresholds for species in fractal landscapes. Umbrales de extinción para especies en paisajes fraccionados. Conservation Biology, 13(2):314–326, April 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilson:1994:PRP

[WKS^{+}94] Marcia H. Wilson, Cameron B. Kepler, Noel F. R. Snyder, Scott R. Derrickson, F. Josh Dein, James W. Wiley, Joseph M. Wunderle, Jr., Ariel E. Lugo, David L. Graham, and William D. Toone. Puerto Rican parrots and potential limitations of the metapopulation approach to species conservation. Las cotorras portorniqueñas y las limitaciones potenciales de la estrategia de metapoblaciones aplicada a la conservación de especies. Conservation Biology, 8(1):114–123, March 1994. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Winterhalder:1997:FRP

[WL97] Bruce Winterhalder and Flora Lu. A forager-resource population ecology model and implications for indigenous conservation. Conservation Biology, 11(6):1354–1364, December 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wayne:1992:MDV

[WLAH92] Robert K. Wayne, Niles Lehman, Marc W. Allard, and Rodney L. Honeycutt. Mitochondrial DNA variability of the gray wolf: Genetic consequences of population decline and habitat fragmentation. Variabilidad del ADN mitocondrial del lobo gris: concecuencias genéticas de la declinación poblacional y fragmentación del hábitat. Conservation Biology, 6(4):559–569, December 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Williams-Linera:1998:MFD

[WLDGGZ98] G. Williams-Linera, V. Domínguez-Gastelú, and M. E. García-Zurita. Microenvironment and floristics of different edges in a fragmented tropical rainforest. Conservation Biology, 12(5): 1091–1102, October 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wayne:1991:CGE

[WLG^{+}91] R. K. Wayne, N. Lehman, D. Girman, P. J. P. Gogan, D. A. Gilbert, K. Hansen, R. O. Peterson, U. S. Seal, A. Eisenhawer, L. D. Mech, and R. J. Krumenaker. Conservation genetics of the endangered Isle Royale gray wolf. Conservation Biology, 5(1):41–51, March 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilcove:1991:SOC

[WM91] David Wilcove and Dennis Murphy. The spotted owl controversy and Conservation Biology. Conservation Biology, 5(3): 261–262, September 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

White:1997:ARB

[WMB^{+}97] Denis White, Priscilla G. Minotti, Mary J. Barczak, Jean C. Sifneos, Kathryn E. Freemark, Mary V. Santelmann, Carl F. Steinitz, A. Ross Kiester, and Eric M. Preston. Assessing risks to biodiversity from future landscape change. Evaluación de riesgos para la biodiversidad debido a cambios futuros en el paisaje. Conservation Biology, 11(2):349–360, April 1997. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wright:1994:EEP

[WMH94] R. Gerald Wright, James G. MacCracken, and Joel Hall. An ecological evaluation of proposed new conservation areas in Idaho: Evaluating proposed Idaho national parks. Una evaluación ecológica de nuevas áreas de conservación propuestas en Idaho: Evaluando parques nacionales propuestos en Idaho. Conservation Biology, 8(1):207–216, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ward:1998:RBP

[WMKF98] Paul I. Ward, Nicole Mosberger, Claudia Kistler, and Oliver Fischer. The relationship between popularity and body size in zoo animals. Conservation Biology, 12(6):1408–1411, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilcove:1993:WEE

[WMW93] David S. Wilcove, Margaret McMillan, and Keith C. Winston. What exactly is an endangered species? An analysis of the U.S. Endangered Species List: 1985–1991. `?Que es exactamente una especie en peligro de extinción? Un análisis de la lista de especies en peligro de extinción de U.S.A. Conservation Biology, 7(1):87–93, March 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Watling:1998:DSM

[WN98a] Les Watling and Elliott A. Norse. Disturbance of the seabed by mobile fishing gear: a comparison to forest clearcutting. Conservation Biology, 12(6):1180–1197, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Watling:1998:SSE

[WN98b] Les Watling and Elliott A. Norse. Special section: Effects of mobile fishing gear on marine benthos. Conservation Biology, 12(6):1178–1179, December 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Walker:1998:PDT

[WOA98] Deette Walker, Guillermo Ortí, and John C. Avise. Phylogenetic distinctiveness of a threatened aquatic turtle (Sternotherus depressus). Conservation Biology, 12(3):639–645, June 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Woodworth:1999:MPD

[Woo99] Bethany L. Woodworth. Modeling population dynamics of a songbird exposed to parasitism and predation and evaluating management options. Conservation Biology, 13(1):67–76, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Waite:1996:DLH

[WP96] Thomas A. Waite and Patricia G. Parker. Dimensionless life histories and effective population size. Las histórias de vida sin dimensiones y tamaño efectivo poblacional. Conservation Biology, 10(5):1456–1462, October 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Weeks:1997:ASM

[WP97] Priscilla Weeks and Jane M. Packard. Acceptance of scientific management by natural resource dependent communities. Aceptación del manejo científico por comunidades dependientes de recursos naturales. Conservation Biology, 11(1):236–245, February 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Weaver:1996:RCL

[WPR96] John L. Weaver, Paul C. Paquet, and Leonard F. Ruggiero. Resilience and conservation of large carnivores in the Rocky Mountains. Resistencia y conservación de carnívoros mayores en las Montañas Rocallosas. Conservation Biology, 10(4):964– 976, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Warren:1996:HUM

[WPSZ96] C. D. Warren, J. M. Peek, G. L. Servheen, and P. Zager. Habitat use and movements of two ecotypes of translocated caribou in Idaho and British Columbia. La utilización del hábitat y la movilidad de dos ecotipos de caribues translocados en el norte de Idaho y en la Columbia Británica. Conservation Biology, 10 (2):547–553, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Weber:1996:GPL

[WR96] William Weber and Alan Rabinowitz. A global perspective on large carnivore conservation. Conservación de carnívoros mayores. Conservation Biology, 10(4):1046–1054, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wayne:1998:ORW

[WRG98] Robert K. Wayne, Michael S. Roy, and John L. Gittleman. Origin of the red wolf: Response to Nowak and Federoff and Gardener. Conservation Biology, 12(3):726–729, June 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [NF98].

Weiss:1991:FCS

[WRM^{+}91] Stuart B. Weiss, Paul M. Rich, Dennis D. Murphy, William H. Calvert, and Paul R. Ehrlich. Forest canopy structure at overwintering Monarch butterfly sites: Measurements with hemispherical photography. Conservation Biology, 5(2):165–175, June 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilson:1997:ELU

[WRSP97] Cathleen J. Wilson, Robin S. Reid, Nancy L. Stanton, and Brian D. Perry. Effects of land-use and Tsetse fly control on bird species richness in Southwestern Ethiopia. Efectos del uso del suelo y el control de la mosca Tsetse en la riqueza de especies del suroeste de Etiopía. Conservation Biology, 11(2):435–447, April 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wolf:1997:HDP

[WS97] Paul G. Wolf and Robert B. Sinclair. Highly differentiated populations of the narrow endemic plant Maguire primrose (Primula maguirei). Poblaciones altamente diferenciadas de la prímula Maguire, una planta endémica. Conservation Biology, 11(2):375–381, April 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilkie:1992:MLM

[WSB92] David S. Wilkie, John G. Sidle, and Georges C. Boundzanga. Mechanized logging, market hunting, and a bank loan in Congo. Explotación forestal mecanizada, cacería comercial y préstamos bancarios en el Congo. Conservation Biology, 6(4):570–580, December 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wolff:1997:EHL

[WSE97] Jerry O. Wolff, Eric M. Schauber, and W. Daniel Edge. Effects of habitat loss and fragmentation on the behavior and demography of gray-tailed voles. Efectos de la pérdida y fragmentación de hábitat sobre el comportamiento y la demografía de ratones de campo de cola gris. Conservation Biology, 11 (4):945–956, August 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Waples:1990:CGP

[WT90] Robin S. Waples and David J. Teel. Conservation genetics of Pacific salmon I. Temporal changes in allele frequency. Conservation Biology, 4(2):144–156, June 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Waits:1998:MDP

[WTWS98] Lisette P. Waits, Sandra L. Talbot, R. H. Ward, and G. F. Shields. Mitochondrial DNA phylogeography of the North American brown bear and implications for conservation. Conservation Biology, 12(2):408–417, April 1998. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wuerthner:1994:SVA

[Wue94] George Wuerthner. Subdivisions versus agriculture. Conservation Biology, 8(3):905–908, September 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wainwright:1998:PPS

[WW98a] Thomas C. Wainwright and Robin S. Waples. Prioritizing Pacific salmon stocks for conservation: Response to Allendorf et al. Conservation Biology, 12(5):1144–1147, October 1998. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [ABB^{+}97].

Waldman:1998:SRO

[WW98b] John R. Waldman and Isaac I. Wirgin. Status and restoration options for Atlantic sturgeon in North America. Conservation Biology, 12(3):631–638, June 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Walsh:1999:WIW

[WW99] Peter D. Walsh and Lee J. T. White. What it will take to monitor forest elephant populations monitoreo de poblaciones de elefantes en bosques. Conservation Biology, 13(5):1194–1202, October 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wyman:1990:WHA

[Wym90] Richard L. Wyman. What’s happening to the amphibians? Conservation Biology, 4(4):350–352, December 1990. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yaffee:1997:WEP

[Yaf97] Steven L. Yaffee. Why environmental policy nightmares recur. Por que son recurrentes las pesadillas en políticas ambientales. Conservation Biology, 11(2):328–337, April 1997. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yaffee:1999:TFE

[Yaf99] Steven L. Yaffee. Three faces of ecosystem management tres caras del manejo de ecosistemas. Conservation Biology, 13(4): 713–725, August 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yahner:1995:FDC

[Yah95] Richard H. Yahner. Forest-dividing corridors and neotropical migrant birds. Conservation Biology, 9(3):476–477, June 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yahner:1996:FFA

[Yah96] Richard H. Yahner. Forest fragmentation, artificial nest studies, and predator abundance. Conservation Biology, 10(2):672–673, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Young:1998:DFV

[YAM^{+}98] Deshea L. Young, Marc W. Allard, Jorge A. Moreno, Michael M. Miyamoto, Carlos R. Ruiz, and Raul A. Pérez-Rivera. DNA fingerprint variation and reproductive fitness in the plain pigeon. Conservation Biology, 12(1):225–227, February 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Young:1996:CPG

[YB96] A. G. Young and A. H. D. Brown. Comparative population genetic structure of the rare woodland shrub Daviesia suaveolens and its common congener D. mimosoides. Comparación de la estructura genética poblacional del arbusto raro de bosque Daviesia suaveolens y su congénere comú D. mimosoides. Conservation Biology, 10(4):1220–1228, August 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Young:1999:GSF

[YBZ99] A. G. Young, A. H. D. Brown, and F. A. Zich. Genetic structure of fragmented populations of the endangered daisy Rutidosis leptorrhynchoides. Estructura genética de poblaciones fragmentadas de la margarita Rutidosis leptorrhynchoides, amenazada de extinción. Conservation Biology, 13(2):256–265, April 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yeager:1997:ORT

[Yea97] Carey P. Yeager. Orangutan rehabilitation in Tanjung Puting National Park, Indonesia. Conservation Biology, 11(3):802– 805, June 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yosef:1992:TSI

[YG92] Reuven Yosef and Thomas C. Grubb, Jr. Territory size influences nutritional condition in nonbreeding loggerhead shrikes (Lanius ludovicianus): a ptilochronology approach. Conservation Biology, 6(3):447–449, September 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yonzon:1991:CTR

[YH91] Pralad B. Yonzon and Malcolm L. Hunter, Jr. Cheese, tourists, and red pandas in the Nepal Himalayas. Conservation Biology, 5(2):196–202, June 1991. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yahnke:1996:DFD

[YJG^{+}96] Christopher J. Yahnke, Warren E. Johnson, Eli Geffen, Deborah Smith, Fritz Hertel, Michael S. Roy, Cristian F. Bonacic, Todd K. Fuller, Blaire Van Valkenburgh, and Robert K. Wayne. Darwin’s fox: a distinct endangered species in a vanishing habitat. El zorro de Darwin: una especie particular en peligro en un hábitat que está desapareciendo. Conservation Biology, 10 (2):366–375, April 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yahner:1996:DAG

[YM96] Richard H. Yahner and Carolyn G. Mahan. Depredation of artificial ground nests in a managed, forested landscape. Conservation Biology, 10(1):285–288, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yahner:1997:BCF

[YM97] Richard H. Yahner and Carolyn G. Mahan. Behavioral considerations in fragmented landscapes. Conservation Biology, 11 (2):569–570, April 1997. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yosef:1994:EFR

[Yos94] Reuven Yosef. The effects of fencelines on the reproductive success of loggerhead shrikes. Conservation Biology, 8(1):281– 285, March 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Young:1993:DCM

[You93] Truman P. Young. Development and conservation: More on Caring. Conservation Biology, 7(4):750–751, December 1993. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Young:1994:RED

[You94a] Kenneth R. Young. Roads and the environmental degradation of tropical montane forests. Las carreteras y la degradación ambiental de los bosques tropicales montanos. Conservation Biology, 8(4):972–976, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Young:1994:NOL

[You94b] Truman P. Young. Natural die-offs of large mammals: Implications for conservation. Declinaciones naturales repentinas y pronunciadas en mamíferos grandes: implicaciones para la conservación. Conservation Biology, 8(2):410–418, June 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Young:1995:WML

[You95] Truman P. Young. Wildlife management links up with conservation biology. Conservation Biology, 9(4):969–971, August 1995. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Young:1999:CSI

[You99] Truman P. Young. Catastrophes are still interesting. Conservation Biology, 13(1):202, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yanes:1996:INP

[YS96] Miguel Yanes and Francisco Suárez. Incidental nest predation and lark conservation in an Iberian semiarid shrubsteppe. Depredación incidental de los nidos y la conservación de las alondras ibericas en los arbustos esteparios semiaridos. Conservation Biology, 10(3):881–887, June 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yu:1994:FTG

[Yu94] Douglas Yu. Free trade is green, protectionism is not. El libre comercio es verde, no así el proteccionismo. Conservation Biology, 8(4):989–996, December 1994. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yu:1996:NFF

[Yu96] Douglas W. Yu. New factor in free trade: Reply to Jenkins. Conservation Biology, 10(1):303–304, February 1996. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zhu:1998:EDS

[ZDM98] Danqing Zhu, Sandie Degnan, and Craig Moritz. Evolutionary distinctiveness and status of the endangered Lake Eacham rainbowfish (Melanotaenia eachamensis). Conservation Biology, 12 (1):80–93, February 1998. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zeide:1992:PGD

[Zei92] Boris Zeide. Has pine growth declined in the Southeastern United States? Conservation Biology, 6(2):185–195, June 1992. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zielinski:1999:BUR

[ZG99] William J. Zielinski and Steven T. Gellman. Bat use of remnant old-growth redwood stands. Conservation Biology, 13(1): 160–167, February 1999. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).